A Bibliography of Papers in Lecture Notes in Computer Science (1999), Part 2 of 2

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

29 July 2017 Version 1.08

Title word cross-reference

(111)7x7 [501]. (k − 1) [1300]. 0/1 [237]. 012 [248]. 1 [1300]. 128 [1102]. 2 [1402, 1750, 1755, 493, 546, 1307, 910, 212, 239, 1113, 244, 1270, 873, 320, 874, 1602]. 2R [1782]. 3 [1620, 1537, 1291, 184, 388, 1048, 379, 371, 1753, 493, 369, 235, 495, 405, 1676, 410, 1742, 254, 384, 368, 320, 406, 359, 363, 874, 361, 362, 1234, 1603]. 4 [242, 383, 376, 254]. 5 [886]. 6 [886]. A∗ [283]. b [639]. L̄R^{2} [972]. C [762]. N P [896]. T OY [969]. d [475, 1368]. ∃xx∀xx [957]. ILU [1711]. K [1808, 995, 1728, 885, 233, 475, 241, 1348, 1840, 1300, 1297]. κ [1374]. Kd [1254]. L [706]. λσ [977]. LU [1711]. m [245]. μ [1367]. n [1108]. N = p^{r}q [1784]. n × m [650]. O(1) [1371]. O(log |G|) [1308]. O(N) [1614]. ± [1522]. R [657, 1530, 1539, 1538, 1784]. S [914]. T [1730]. T (A) = T (B) [1378]. V [837].

-Average [762]. -Bit [1102]. -Calculus [1367]. -Categories [1374]. -Center [1755]. -complete [896]. -Cost [245]. -Coverings [1307]. -Cube [237]. -D [1753, 1270, 362]. -Edge [239]. -Edge-Connected [241, 1300, 873]. -Factor [244]. -join [1730]. -key [1113]. -matching [639]. -median [1297]. -Modell [837]. -Packings [1307]. -Partitioning [885]. -Point [1750]. -regular [914]. -SAT [184, 242]. -Server [1291]. -set [1728]. -Systems [706]. -Terminal [235]. -Tree [1530]. -Trees [1348, 1840, 1539, 1538, 1254, 657]. -Valued [1048]. -Vectors [248]. -Vertex-Connected [1300]. -way [886, 254].

0.1 [999].

1 [1205, 1204]. 10 [1521]. 10th [1175]. 16 [1016]. 1BC [1079]. 1Higher [967]. 1Higher-Order [967].

2 [48, 555, 556]. 2K [1884]. 2R [1783].

3 [142, 1096]. 3.0 [825]. 3D [360].

4 [520].

5.0 [998].

7 [432].

802.6 [140].

’97 [1780, 97]. ’98 [1869, 89, 91, 1870, 124, 110, 1861, 81, 113, 100, 136]. ’99 [742, 578]. 9th [1962].

Abdominal [667]. Abduction [324, 285, 1542]. Abilities [102, 103]. above [810]. ABox [580]. ABR [1034]. Abrupt [446]. Abstract [1326, 1275, 1327, 1291, 1328, 1785, 1330, 1790, 1795, 1810, 1335, 1269, 1303, 1342, 1345, 1801, 1770, 1347, 1283, 1352, 875, 24, 1140, 1357, 1358, 1365, 1070, 1289, 1766, 1008, 1294, 1373, 1071, 1376, 1061, 1799, 1010, 1060, 872, 1382, 604]. Abstracting [689]. Abstraction [1038, 1032, 1039, 1004, 1014, 1646, 1644]. Abstraction-Based [1014]. Abstractions [687, 62, 506]. Abstracts [1257]. Abuse [1792]. Abuse-Free [1792]. AC [975, 957, 960]. AC-Rewriting [975]. AC-RPO [960]. Accelerated [1106]. Accept [887]. Acceptance [1372]. Access [67, 68, 850, 64, 77, 1467, 73, 430, 1892, 79, 72, 1911, 1877, 794, 1439, 1285, 1989, 1982, 942, 851, 1997, 1991]. Access-Rights [67]. Accesses [1671, 543]. Accessibility [872]. Accessing [645, 1377]. Accommodating [1736]. ACCORD [629]. Accuracy [379]. Accurate [1221, 1676]. Achieving [1686]. Achromatic [1299]. Acquiring [1076, 714, 725]. Acquisition [727, 713, 730, 1489, 733, 1492, 724, 299, 105, 264, 704, 199, 191]. Action [628]. Actions [720]. Activating [1477]. Activation [1196]. Activation-Based [1196]. Active [78, 1581, 1597, 1473, 1586, 1582, 380, 1580, 870, 1458, 1589, 365, 708, 1600, 1598, 1587, 366, 1948, 1577, 1574, 1590, 1194, 1591, 1593, 1585, 1579, 1595, 1991, 1596]. Activities [1466]. Activity [498, 625]. Actors [1186, 21]. Acute [667]. Ad [1981, 1925]. Ad-Hoc [1981, 1925]. Ada [742]. Ada-Europe [742]. Adaptability [563]. Adaptation [1507, 1480, 1479, 1484, 1485, 937, 45]. Adapting [475, 99]. Adaptive [655, 1769, 808, 271, 1612, 762, 547, 859, 1613, 54, 1569, 1800, 367, 1264, 1561, 524, 1991, 1619]. Adaptively [1266]. Adaptively-Secure [1266]. Adaptiveness [1287]. Add [794]. Add-On [794]. Adding [1087]. Addition [531]. Additive [230, 231]. Address [1271, 1319, 517]. Addressless [1924]. Adjusted [1247]. Administration [672]. Admission [271]. Adopted [1170]. Advanced [309, 1880]. Advances [1447]. Advice [1250, 1569, 1515]. Aerodynamics [1702]. Affect [1471]. Affect-Driven [1471]. Affective [10, 1888]. Affine [479]. Affinely [470]. Affirmative [267, 172]. Against [1108, 1445]. Age [1464]. Agency [1565, 128]. Agency-Based [1565]. Agent [314, 1581, 1412, 1550, 1419, 1425, 1998, 115, 1558, 1000, 1555, 262, 1410, 1551, 1600, 1562, 415, 1554, 305, 1564, 16, 306, 1418, 1219, 102, 103, 1547, 1566, 80, 670, 1420, 309, 312, 1414, 1413, 1263, 1869, 1416, 310, 1417, 308, 705, 1411, 1870, 1567, 158]. Agent-Based [1000, 1555, 415, 1554, 16, 1550]. Agent-Mediated [1566]. Agent-Oriented [1412, 262, 1410]. Agents [3, 311, 1710, 1415, 1586, 9, 315, 10, 1559, 63, 100, 11, 74, 307, 99, 1570, 95, 14, 306, 118, 1557, 71, 1552, 96, 738, 1421, 1862, 106, 1568, 1560, 1561, 1563, 70]. Agglomerations [1656]. Aggregated [336]. Aggregated-Temporal [336]. Aggregation [637]. Aggressive [1671]. Agilo [127]. Agreeably [1382]. Agreement [1350, 905]. AGV [785]. AH [1645]. AH-graphs [1645]. Ahead [511]. AI [121, 90, 292]. AI-Lab [121]. Aid [300, 207]. Aided [1981]. ail [803]. Air [324]. Air-Crew [324]. Aircraft [1260]. Alarm [298]. Alarm-Correlation [298]. ALC [580]. Alcoholics [397]. Aleph [53]. Alex [300]. Alexithymia [300]. Algebra [30, 1453, 1686, 1180, 37, 1344, 1352, 339, 1041, 957, 1049, 1174]. Algebraic [976, 944, 1179, 1660]. Algebras [1212, 38, 1361, 1644]. Algorithm [331, 284, 1032, 1825, 1615, 49, 478, 1854, 1751, 1616, 315, 350, 1805, 919, 278, 1711, 1309, 912, 270, 246, 1808, 1812, 811, 920, 374, 1934, 883, 1707, 760, 1304, 254, 886, 1856, 367, 283, 1235, 320, 1815, 987, 344, 530, 1301, 878, 321, 1142, 638, 1149, 217, 196, 654, 198, 1622]. Algorithmic [229, 26]. Algorithms [227, 1305, 1290, 1717, 280, 1275, 1540, 1725, 1292, 1830, 764, 1753, 493, 1999, 234, 1855, 1539, 649, 813, 1842, 746, 756, 277, 1816, 791, 1207, 1760, 242, 1827, 1675, 38, 1281, 971, 250, 706, 251, 807, 895, 1299, 754, 1733, 1678, 1289, 1858, 1806, 25, 252, 1260, 1635, 1274, 1268, 276, 279, 747, 1264, 907, 655, 1123, 220, 640, 225, 222, 639, 210]. Aliasing [1963]. ALIEN [1572]. Alignment [401]. Alignments [1392]. All-Optical [1825]. All-or-Nothing [1768, 1796]. Allen [339]. Allocation [141, 1202, 925, 1201]. Allocentric [1668]. allowing [1150]. ALM [1221]. Almost [1111]. Alternating [1337, 1338]. Alternative [1024, 1145]. Amalgamation [1523, 1534]. Ambients [1331]. American [1184]. among [1580, 1552]. Amplifying [1895]. Analogical [1212, 1091]. Analyses [1005]. Analysing [1820, 793]. Analysis [1038, 303, 1238, 1169, 501, 1053, 687, 1280, 622, 1064, 399, 1170, 1607, 1789, 1098, 1221, 400, 918, 788, 387, 350, 801, 1231, 1813, 389, 381, 843, 1065, 427, 1241, 1242, 1281, 1427, 787, 1788, 1166, 1733, 562, 1723, 694, 1210, 1193, 529, 1057, 1724, 1162, 290, 1803, 786, 107, 1680, 318, 1931, 1639, 360, 439, 1068, 421, 361, 1201, 1168, 1682, 1012, 140, 838, 699, 1701, 695, 1661, 1152, 1638, 196]. Analytic [584]. Analytical [1669, 356]. Analyzing [1944]. Anatomical [369, 392]. Anchoring [1131]. Andhill [110, 1861]. Andhill-98 [110, 1861]. Anesthesia [824]. Angiograms [379]. Angle [484]. ANIMAL [372]. Animating [707]. Animation [1619]. Anisotropic [1363]. Ankles [391]. ANMAC [1590]. ANN [817, 821, 330]. ANN-Based [817, 821, 330]. Annotated [1475, 1476]. annotating [169]. Annotation [723, 457, 481]. Announcement [1990]. Annuli [1735]. ANON [1585]. Anonymity [1431]. Anonymous [63, 1430, 1443, 1924, 1799, 1441]. Answering [1124, 819]. Antibiotics [1515]. Anticipation [1570]. Antidictionaries [1334, 1387]. Anywhere [1806]. AP [1886]. AP3000 [545]. Aperios [568]. API [1200]. Apoptosis [69]. Appearance [380]. Applets [65]. appliance [1893]. Appliances [1902, 1885]. Applicable [1500]. Application [778, 1501, 1669, 280, 858, 1212, 767, 56, 1004, 1082, 1270, 293, 620, 616, 296, 392, 288, 1186, 1260, 368, 1702, 664, 828, 532, 1772, 363, 108, 102, 907, 330, 1701, 196]. Application-Specific [56]. Applications [533, 1835, 355, 1730, 496, 713, 1970, 1909, 1330, 518, 1889, 882, 825, 1580, 697, 1138, 1350, 292, 539, 1302, 1108, 155, 1693, 1953, 483, 555, 556, 433, 1699, 259, 1301, 1919, 204, 156]. Applied [376, 1491, 318, 439, 1952]. Applying [636, 607, 813, 580, 510, 620, 634, 734, 841, 1403]. Approach [1624, 452, 563, 679, 1930, 26, 836, 940, 771, 852, 1689, 676, 1263, 1466, 1869, 678, 115, 918, 850, 984, 1612, 1705, 346, 1265, 351, 1484, 308, 1485, 1675, 842, 1030, 1385, 629, 281, 757, 830, 473, 853, 1723, 142, 317, 1191, 529, 1386, 780, 1469, 635, 1870, 118, 786, 107, 1591, 851, 1431, 1567, 1601, 447, 1168, 907, 39, 1683, 1487, 726, 1468, 266, 1486, 193, 1134, 197, 188, 1662, 213, 1152, 1645, 211]. Approaches [1974, 1169, 809, 1789, 1024, 1697, 698]. Approximate [1857, 804, 1283, 1740, 1085, 1404, 1398, 407, 213, 1834, 1394, 1149]. Approximating [899, 1836, 876, 920, 375]. Approximation [227, 809, 1753, 1854, 234, 1837, 1278, 919, 242, 1349, 1851, 1838, 250, 251, 1297, 895, 1299, 252, 873, 1274, 1268, 878]. Approximations [897, 1360, 877, 1845]. AQ18 [197]. Arabic [760, 317]. Arbitration [1568]. Archaeology [1891]. Architectural [1959, 337, 1590]. Architecture [1703, 1572, 561, 805, 1589, 430, 1551, 1707, 415, 846, 1513, 1862, 50, 1579, 1700, 1575]. Architectures [1685, 1691, 1676, 1966]. Arcs [1407]. Area [1312, 1359, 879, 548]. areas [1664]. Argument [730, 1883]. Argumentation [1151, 17, 1421]. Arithmetic [1323, 1318, 1719]. Arity [1216]. Arm [771, 453]. Arms [770]. Arrangements [1817]. Array [1453]. Arrays [1731, 1284, 1748, 916, 1707, 1724, 1401]. Arrhythmias [676]. Arrow [1373]. Arrows [287]. ART [773, 129]. ART-based [773]. Artery [682, 382]. Artificial [767, 353, 820, 12, 274, 659, 207, 749, 826, 699].

ARTMAP [761, 766]. Arts [319]. Ascendant [437]. ASIUM [730]. ASL [1945]. Aspect [1972, 574, 566, 575]. Aspect-Oriented [1972, 574, 575]. Aspects [1173, 393]. assembly [194]. assessed [219]. Assessing [1134, 1982]. Assessment [474, 1433, 396, 745, 1096]. Assignment [1279, 1730, 478, 1853, 906, 326]. Assignments [1849]. Assistance [42]. Assistant [602, 1571, 1905]. Assisted [360, 499, 361, 1257, 1096]. Associated [979, 248]. Association [278, 1223, 216]. Associative [1316, 1690, 1317, 1680, 1377, 590]. Assume [1043]. Assume-Guarantee [1043]. Assumptions [1778, 1802]. Assuring [1219]. Asymmetric [1798, 1813, 1985]. Asymmetry [1286]. Asynchronous [141, 1004, 944, 34, 1679, 1177]. Atlas [380]. ATM [534, 145, 803, 143]. Atomicity [1654]. ATPG [1059]. ATSP [245]. Attached [535]. Attack [1101, 1109]. Attacks [1107, 1115, 1789, 1774, 1117]. Attention [503]. Attitudes [1425, 1191]. Attribute [1206]. Attributes [375]. Auctionlike [1566]. Auctions [1444]. audio [192]. Audiovisual [1903]. Audit [1432]. Auditable [1799]. Augmented [1876, 1915, 1881, 1906]. Augmenting [1300]. Australian [1253]. Authentication [1778, 1776, 1777, 1441, 1068, 1997]. Author [1978, 23, 137, 576, 946, 983, 1097, 1120, 1159, 1384, 1804, 1832, 1926, 1955, 2001]. authoring [1921]. Authorisation [674]. Authorization [61]. autoepistemic [160]. Automata [1026, 887, 1685, 1708, 1709, 1329, 1046, 1337, 1338, 1045, 1719, 1712, 1354, 1399, 1400, 1058, 1368, 326, 1372, 1376, 39]. Automata-Theoretic [1337, 1338, 39]. Automated [1034, 388, 8, 338, 1000, 593, 392, 582, 1724, 363, 1146]. Automatic [773, 1056, 387, 354, 441, 504, 454, 541, 397, 523, 28, 483, 1023, 83, 914, 566, 368, 320, 938, 1068, 323, 1060, 459, 1246, 330, 1683, 1012, 704, 199, 170]. Automatically [438]. Automating [615]. Automation [1170, 614, 154]. Autonomous [1477, 1549, 817, 1995, 780, 1167, 18, 1421, 21, 1933]. Autonomy [1586]. Autowave [1681]. Auxiliary [1610]. Avatar [1950]. Average [900, 1280, 762, 1356, 1373, 1]. Average-Case [1280, 1356, 1]. Avoiding [1129, 1130, 654]. Aware [561, 1918, 1922, 1908, 1890, 45, 1919]. Awareness [1899, 1923, 1878, 1900, 1888, 1886, 1667]. Axiomatization [31]. axioms [1133]. azo [207]. Azzurra [129].

B [383]. B-Spline [383]. BabyTigers [136]. BabyTigers-98 [136]. Background [1080]. Backpropagation [1085]. backtrack [185]. Backward [1987]. Bacteraemia [679]. Balance [299]. Balancing [1272, 1280, 530]. Ball [93, 94, 1830]. Ball-Receiving [93, 94]. Balls [1743, 1293]. Bandwidth [141]. BargainFinder [1876]. Barrier [1779]. Base [1475, 1476]. Based [303, 778, 1232, 284, 816, 1501, 4, 1205, 1685, 1708, 1164, 304, 1797, 1502, 141, 355, 712, 1472, 112, 1059, 864, 798, 1473, 1474, 68, 353, 588, 676, 472, 291, 1998, 1607, 282, 1263, 423, 1181, 808, 271, 1690, 822, 1608, 425, 474, 715, 77, 667, 349, 603, 1433, 437, 1506, 269, 680, 1881, 1000, 504, 1014, 1480, 1479, 1197, 348, 1481, 1482, 1555, 1748, 683, 718, 817, 821, 144, 38, 1190, 1943, 1892, 860, 861, 789, 688, 422, 827, 732, 11, 471, 1020, 1774, 1427, 1877, 1949, 1196, 794, 1713, 1600, 1720, 285, 1565, 1245, 787, 1465]. Based [436, 766, 744, 807, 1613, 376, 476, 1727, 485, 927, 1240, 677, 1511, 281, 473, 415, 1512, 288, 1980, 1494, 1495, 1008, 109, 1011, 674, 1554, 1621, 305, 267, 16, 1803, 1496, 802, 1514, 724, 662, 835, 1094, 322, 1576, 295, 489, 517, 96, 1515, 669, 481, 1497, 1699, 1450, 1499, 1516, 477, 459, 108, 1517, 299, 470, 1246, 1571, 289, 488, 1579, 1518, 75, 1487, 402, 726, 838, 1936, 1619, 133, 321, 1603, 120, 1351, 1486, 165, 773, 200, 1550, 204, 168, 199, 1664, 1198, 1848, 1149, 171, 330, 458]. Based-Systems [295]. Bases [1625, 1498, 907, 1137, 166]. Basic [1122, 1425, 87, 88, 1636, 537]. Basis [226, 1311, 1316, 1317, 1673, 1589, 406, 747]. Batch [893, 894]. Bayesian [1606, 1614, 1133, 1612, 1079, 1139, 1142, 1253, 1236, 22]. BDD [38]. BDD-Based [38]. BDDs [1031]. Be [1072, 1029, 1793, 554, 1239]. Beautiful [702]. Bedside [672]. Behavior [925, 820, 329, 1190, 1552, 96, 105]. Behavioral [554, 1252]. Behaviour [806, 1682]. Behavioural [3, 1164, 719, 1411]. Behaviours [1181, 1559]. Belief [1125, 160, 1137, 206, 1145, 1149, 1152]. Believable [14]. Benchmarks [1704, 925]. Benefits [61, 348, 59, 929]. Bent [1111]. Bernstein [809]. Best [609, 15]. Beta [1754]. Beta-Skeleton [1754]. Better [1899, 920, 862, 176]. Between [486, 1950, 1365, 1944, 819, 1797, 1043, 114, 1408, 1358, 1434, 1165, 1364, 409, 978, 1286, 1409]. Beyond [1453, 1779, 868]. bias [180]. Biases [1225]. biconnected [641]. bigraph [642]. Bin [1857, 652]. Bin-Packing [1857]. Binary [778, 358, 1287, 285, 909, 643, 328, 901, 902]. Bindings [1434]. Binomial [1230]. Biochemical [1830, 293]. Biodegradability [1078]. Bioelasticity [356]. Biological [533, 759, 1927]. Biological- [533]. Biomechanical [384]. Biomedical [678]. Biomorph [807]. Bipartite [244, 1302]. Birth [605, 219]. Birthday [1779]. Bisectors [1280]. Bisimulation [38, 1352, 1055]. Bit [1267, 1102, 491]. Black [1827]. Blackboard [1683]. Blacklisting [1802]. BLAST [1697]. Blinding [1429]. Blobworld [472]. Block [1459, 1102, 1103, 1117, 1112, 343, 492]. Block-Constrained [492]. Blood [685, 706, 382, 824]. Blue [919]. Body [1934, 1935, 402]. Body-Face [1935]. Boltzmann [768]. Boolean [882, 1144, 1027, 2, 1383]. Boolean-like [1144]. Boomerang [1109]. BOOSTER*Process [621]. Bootstrapping [1475, 1476]. Both [609, 1996, 1521, 1522, 1239]. bottom [1146]. bottom-up [1146]. Bound [906, 1754, 1]. Boundary [1613, 1858, 545]. Bounded [1323, 1395, 589, 1345, 895, 1633, 1089]. Boundedness [1340]. Bounds [1326, 230, 231, 885, 237, 1393, 1277, 1844, 1369, 1286]. Box [1180]. BPP [1846]. Brain [665, 820, 666, 705, 397, 405, 359, 386, 370]. BRAINS [118]. Branch [260]. Branch-and-Cut [260]. Branching [1326]. Breadth [1015]. Breadth-First [1015]. Break [439]. Breakdown [332]. Breakdowns [1859]. Breaking [1945]. Brick [1881]. Brick-Based [1881]. Broadband [1581, 1982]. Broadcast [1438, 1787, 1987, 1984]. Broadcasting [1269]. Broker [571]. Brokering [711, 307, 305, 1567]. Browse [429]. Browsing [428]. BSP [1806, 1701]. Buffer [1030]. Build [371]. Building [712, 495, 1080, 13, 1498, 659, 1246, 266, 214, 1154]. built [1663]. Bulk [1718, 1530, 2, 1383, 657]. Bulk-Insertion [1530]. Bulk-Synchronous [2, 1383]. Bureaucrat [1741]. Bus [342]. Business [619, 348, 313, 831, 1565, 621, 1564, 828, 833, 868]. Busy [703]. Bytecode [1066].

C [1819, 1818]. C0YNTHIA [1019]. Cabinet [1914]. Cache [56, 548, 1820, 1994, 655]. Cache-Coherence [56]. cache-efficient [655]. cached [656]. Caches [1377]. Caching [1459, 844, 1579]. Cadastral [1545]. CADE [1016]. CADE-16 [1016]. Calculations [530]. Calculi [584, 592, 591, 583, 1632, 1649, 600].

Calculus [1013, 976, 923, 590, 1626, 301, 597, 1367, 598, 19, 1648, 1154]. Call [309, 271]. Callosum [389]. Cambridge [123]. Camera [441]. Can [1072, 1029, 1793, 554, 1281, 1567]. Canceling [246]. Cancer [699, 695]. Candidate [1506, 1134]. Cannot [1375]. Capabilities [275]. Capability [109, 75]. Capability-Based [75]. Capacitated [234]. Capacities [238]. Capacity [1437, 326, 1274]. CAPP [1685]. Capture [567, 453]. Capturing [619, 1928, 1531]. carcinogenic [207]. Card [1433, 1429]. Cardiac [676, 376]. cardinal [1652]. Cardiography [358]. Care [698, 659]. Carlo [1149]. Carriers [1917]. Case [303, 1501, 304, 1502, 355, 1280, 1472, 1828, 1474, 1475, 1476, 1586, 1216, 1626, 350, 1506, 1435, 1480, 1479, 1197, 454, 1484, 1481, 1482, 1485, 683, 789, 1356, 1510, 1491, 1511, 1492, 1512, 1980, 1494, 1495, 529, 1371, 1496, 1514, 1515, 669, 1815, 1497, 1498, 1516, 1517, 1518, 1700, 1487, 1637, 695, 1486, 212, 1, 647]. Case-Based [1501, 304, 1502, 355, 1472, 1474, 1506, 1480, 1479, 1197, 1481, 1482, 683, 789, 1511, 1512, 1495, 1496, 1514, 1515, 669, 1497, 1516, 1517, 1518, 1487, 1486]. Case-Bases [1498]. Cases [635, 302, 1969]. Cash [1433, 1799, 1431, 1432]. CAT [119]. cataloguing [169]. Categorial [984]. Categories [1374]. Categorization [1657]. Categorizing [455]. Category [421]. Causal [1126, 1148]. Causation [1127]. Cayenne [153]. Cayley [1308, 1684]. CBR [1500, 1471, 1477, 1478, 1507, 1483, 1509, 1489, 1490, 1513]. CDL [1712]. Cells [820]. Cellular [1685, 1708, 1709, 1719, 298, 1712, 1707, 1368, 1680, 1681, 779]. Cellular-Automata [1709]. Cellular-Neural [1680]. Cenju [520]. Cenju-4 [520]. Center [786, 1755]. Centered [805, 1935, 625]. Centering [93, 94]. Centers [309, 295]. Central [393]. Centralized [1925]. Centric [14, 849].

Cerebellar [269]. Cerebellum [397]. Certain [243, 248]. Certificate [1435]. Certificates [67, 1346]. Certification [1434]. Cessation [701]. CG [1190, 1200, 1201]. CG-Based [1190]. CGs [1198]. Chain [764, 1283]. Chains [1332]. Challenge [91]. Challenges [1597, 2000, 1979]. Challenging [90]. Chameleon [1910]. Champion [85, 86]. Change [1445, 446, 1552]. Changing [1490, 208]. Channel [284, 26, 1437, 1178]. Character [82, 1085, 1948, 760]. Characterisation [962]. Characteristic [804]. Characteristics [802]. Characterization [777, 1797, 696, 1111, 1335, 24, 1336, 772, 1661]. Charging [1576]. ChatterBox [1916]. Cheap [109]. Check [1061]. Checker [40]. Checking [1066, 1032, 1031, 1044, 1059, 1048, 1688, 1337, 1338, 1435, 29, 1035, 721, 1040, 1051, 1055, 39, 1052, 868]. Checks [1183]. Chemical [533, 699]. Chemical-Parallel [533]. Child [824]. Children [1875]. Chip [548]. Choice [30]. Cholesky [1669]. CHPS [122]. Chromosome [340]. Chronic [397, 396]. Chvátal [243, 237]. CINDI [169]. Cipher [1104, 1102, 1103, 1117, 1116]. Ciphers [1114, 1110, 1112]. Circuit [890, 891]. Circuits [1323, 1056, 889]. Circulant [908]. Circular [641]. City [1548]. clamped [195]. Class [924, 612, 995, 992, 618, 160]. Classes [776, 923, 1182, 756, 939, 929, 1246]. Classical [820, 578]. Classification [1537, 858, 1472, 751, 1244, 437, 1261, 1262, 681, 1239, 766, 744, 1240, 1235, 468, 499, 328, 319, 1257, 219, 175, 178, 192]. Classifier [782, 1079, 281]. Classifiers [682, 519, 1247, 217]. Classify [940]. Clausal [1089]. CLDS [586]. Cleaning [1292]. Click [1901]. Client [557, 297, 846, 1997]. Client-Server [557, 297, 846]. Clients [1987]. Climate [538]. Climbing [1616]. Clinical [703, 687, 691, 658, 663, 686, 190]. Clock [1053]. Clocks [1679]. Close [1887]. Closed [1140, 1374]. Closure [1323, 965]. Closures [964]. Cluster [533, 1704, 755, 543, 221]. Clustered [394]. Clustering [387, 755, 1245, 1243, 1236, 1532, 1226, 1237, 1246]. Clusters [57, 1223, 1697, 1234]. CMOS [826]. CMTrio [135]. CMTrio-98 [135]. CMUnited [85, 86]. CMUnited-98 [85, 86]. CNC [354, 349]. CNF [1123]. Co [1211, 1894, 1921, 366, 71, 105]. Co-authoring [1921]. Co-dimension [366]. Co-evolution [105]. Co-identity [1211]. Co-location [1894]. Co-operating [71]. Coalitions [1424]. Code [77, 527, 1774, 74, 1724, 1715]. Codes [754, 469, 638]. Coding [1802, 491, 492, 1619]. cognition [1662]. Cognitive [1425, 1869, 115, 1659, 1927, 1657, 1870]. Coherence [56]. Coherency [1995, 1994]. Coin [1430]. Coinduction [1376]. COIRS [458]. Colabra [1463]. Colineal [273]. Collaborating [1425]. Collaboration [1463, 615, 1887, 849]. Collaborative [1950, 1554, 1516]. Collection [811]. Collections [1233, 750]. Color [284, 422, 467, 439, 319, 1837]. Colored [1182, 899]. Colored-Graph [899]. Coloring [1829, 746, 1298, 898]. Colors [419]. Coloured [141, 147, 1988, 1165, 140]. Combination [924, 564, 1247]. Combinational [1056]. combinations [1150]. Combinator [151]. Combinatorial [1703, 918, 883, 253]. Combine [698]. Combined [272]. Combining [1038, 687, 1073, 1074, 1644, 1083, 971, 54, 1489, 12, 1450, 929, 344, 1149]. Combustion [758]. Commentary [82, 83]. Commentator [84]. Commerce [1876, 1428, 1564, 1566, 1563]. Commercials [439]. Committees [1222]. Common [1753, 1407, 1194, 1405, 1406]. Communicating [1325]. Communication [1624, 1767, 49, 1419, 1950, 516, 1000, 1555, 144, 114, 520, 1918, 1562, 1165, 1677, 1230, 1980, 1696, 1724, 306, 1418, 1862, 1596, 133]. Communications [1580, 1692, 1924, 1219]. Communicative [1951]. Commutative [1174]. Compact [431, 1345, 702, 1498]. Compaction [249, 1734]. Comparative [1709, 1680]. Comparing [346, 454, 34]. Comparison [771, 379, 788, 682, 1244, 1344, 1808, 681, 523, 578, 1123]. Compensation [359]. Competence [1496, 726]. Competent [1498]. Competition [102]. Competitive [1736, 912]. Compilation [541, 525, 35]. Compile [567]. Compile-Time [567]. Compiler [523, 1008]. Complementary [1697]. Complements [1348]. Complete [1075, 1081, 1041, 1946, 909, 896]. Completely [1739]. Completeness [1833]. Completion [1030]. Complex [498, 1503, 1504, 349, 1719, 817, 821, 688]. complexes [1524, 1525]. Complexity [890, 1391, 899, 1760, 1356, 1734, 1373, 802, 994, 872, 1380, 989, 161, 1]. Compliance [663]. Component [353, 827, 866, 937, 1465, 830, 1260, 318, 1965]. Component-Based [353, 827, 1465]. Component-Oriented [1965]. Components [1467, 320]. Componentware [837]. Composition [1329, 1881, 829]. Compositional [1411, 1065, 293, 1025]. Comprehension [1571]. Compressed [443, 1385, 1386, 1387]. Compressing [1801]. Compression [493, 1334, 1399, 1400, 1289, 1401, 196]. Computable [1381]. Computation [1768, 1312, 1718, 917, 769, 1672, 1054, 397, 1721, 201, 1224, 1149, 1158]. Computational [1686, 1558, 1802, 562, 792, 538, 872, 1322, 52, 1659, 639]. Computations [62, 1315, 904, 45, 1864]. Computer [495, 1319, 1903, 520, 1267, 1934, 1512, 1924, 320, 360, 521, 300, 1219, 361, 528, 1257, 218].

Computer-Assisted [1257].

Computerized [272]. Computers [1897, 1875, 1910, 1990]. Computing [1898, 1882, 1891, 1669, 486, 1894, 1909, 1923, 825, 182, 1705, 10, 1296, 516, 1346, 1912, 1808, 1408, 1893, 1277, 1968, 254, 886, 823, 1890, 793, 1320, 650, 1883, 1871, 1409, 210, 647]. Concept [165, 1210, 1208, 459, 458]. Concept-Based [459, 458]. Concepts [614, 1508, 1439, 724, 268]. Conceptual [314, 607, 1214, 1212, 1213, 1202, 1265, 1197, 1207, 1508, 1196, 1199, 1215, 1204, 1210, 1193, 1191, 1194, 1184, 725, 869, 189, 1664]. Concert [1362]. CONCERTO [177]. Concrete [1975]. Concurrency [1987, 1986, 1454]. Concurrent [1689, 1795, 1794, 1827, 1177]. Conditional [1439, 206]. Conditions [1037, 240, 1001, 1653, 1858, 1088, 507]. Confidence [95]. Confidential [1924]. Configurable [56, 72, 1908]. Configuration [1467, 1194, 330, 140]. Configurations [534]. Configuring [831, 1579]. Conflict [1172]. Conflicting [1568]. Conflicts [841]. Confluent [1013, 583]. Conformance [1034, 1029]. Congestion [909, 1811]. ConGolog [855]. Congruences [954, 955]. Conjecture [1020]. Conjugacy [776]. Conjunctive [1630, 1092]. Connected [239, 241, 1300, 873]. Connecting [1128]. Connection [1013, 1658, 1358]. Connectionist [799]. Connections [978]. Connectives [597, 775]. Connectivity [1889, 1816]. Conquer [1717, 1073, 1074]. Consciousness [1316, 1317]. Considerations [810, 627]. Consistency [1843, 1998, 580, 1050]. Consistent [373, 1679]. Constant [1860, 238]. Constrain [1495]. Constrained [1735, 1751, 522, 241, 527, 390, 1277, 1222, 492, 221]. Constraint [1275, 1970, 678, 617, 616, 448, 305, 1716, 726, 1052, 120, 1351]. Constraint-Based [305, 726, 120, 1351].

Constraint-Rich [1052]. Constraints [863, 1213, 848, 1849, 1083, 385, 1143, 301, 1677, 1204, 1185, 1011, 947, 618]. Construct [1867]. Constructing [1778, 1284, 845]. Construction [1817, 1847, 1114, 1828, 388, 1813, 1357, 1366, 1137, 638]. Constructionist [1689]. Constructions [976, 1802, 1367]. Constructive [1634, 225]. Constructors [963, 595]. Constructs [1626, 637]. Consultation [736]. Consumer [1914]. Containing [1452, 873]. Content [423, 714, 425, 474, 476, 424, 1450, 477, 460, 470, 174, 166]. Content-Based [423, 425, 476, 1450, 477, 470]. Contention [27]. Contents [455]. Contest [1016]. Context [1899, 1624, 622, 1923, 1889, 1878, 975, 974, 1207, 1918, 1896, 889, 1890, 1208, 333, 988, 1880, 1872, 1556, 1155]. Context-Aware [1918, 1890]. Context-Awareness [1899, 1878]. Context-Dependent [1624]. Context-Free [889]. Context-Sensitive [1889, 975, 974, 333]. Contexts [1209]. Contextual [1206]. contiguous [1724]. Contingency [1844]. Continuations [1375]. Continuous [1943, 270, 364, 326, 1167, 1618]. Continuous-Time [1618]. Contour [1612, 502, 485]. Contour-Based [485]. Contours [366]. Contract [1792, 570]. Contraction [589]. Contribution [812]. Control [795, 312, 759, 771, 767, 68, 49, 338, 808, 146, 77, 762, 1445, 784, 1467, 814, 73, 817, 1599, 1258, 72, 293, 1941, 785, 753, 815, 1987, 803, 1584, 783, 1590, 544, 942, 735, 786, 1993, 18, 1431, 295, 1872, 796, 140, 797, 824, 187]. Controlled [764, 1574]. Controller [751, 782, 271, 269, 810, 40, 785, 154, 758, 120, 1351]. Controllers [773, 752, 770]. Controlling [759]. Conventional [401]. Convergence [1614, 1616, 276, 196, 195]. Conversational [700, 1569]. Conversations [1193]. Conversion [1251, 913]. Conversions [950, 951]. Convex [228, 787, 1276]. Conveying [1641]. Convolution [1721]. COOL [1705]. Cooperating [1558, 168]. Cooperation [670, 128]. Cooperative [1555, 1423, 846, 1557, 1552, 105, 1571, 133]. coordinates [1652]. Coordination [1500, 1894, 1416, 1672, 1702]. Cope [1424]. Coping [521]. copyright [651]. CORBA [943, 1164, 68, 1998, 944, 1698, 844, 865]. Core [923]. Corneal [395]. Cornuejols [226]. Cornuejols-Dawande [226]. Coronary [682]. corpora [199]. Corporate [1481, 1482, 737]. Corpus [389]. Correcting [754, 317]. Correction [1621]. Correctness [1030]. Correlated [1224]. Correlation [298, 1774]. Cortical [369, 372, 509, 370]. Cost [228, 1459, 1996, 245, 1850, 285, 753, 1254, 213, 647, 201]. Cost-Based [285]. Cost-Effective [1459]. COSTPN [146]. Costs [1285, 1839]. Counteract [1789]. Counterfeit [1433]. Counting [1323, 1814]. Coupled [1469, 785]. Course [90]. Cover [1839]. Coverage [227]. Covering [882, 250]. Coverings [1307]. CPAM [829]. Crashes [1679]. Crawling [435, 432]. Creating [1710, 1207, 313]. Credit [1429]. Crew [324]. Criminal [304]. Criteria [866, 1372]. Critical [1970, 239, 697, 343]. Critique [1127]. Critiquing [736]. Crop [795]. Cross [859, 693]. Cross-Organizational [859]. crossing [642]. Crossover [1281, 296]. Cryptanalysis [1119, 1099, 1783, 1108, 1764, 1103, 1780, 1782]. Crypto [1780]. Cryptographic [1313, 1801, 1765]. Cryptography [1766]. Crypton [1100, 1101, 1100]. Cryptosystem [1764, 1780]. Cryptosystems [1769, 1445]. CS [87, 88, 1116]. CS-Cipher [1116]. CSG [644]. CTIVE [1886]. CTL [1051]. Cube [237]. Cubic [1283]. Cultural [693]. Currency [913]. Curve [1528, 1529]. Curvilinear [386]. Customised [708].

Customizable [1592, 849]. Cut [227, 999, 592, 246, 1277, 260, 1852]. Cut-Free [592]. CutRes [999]. Cuts [1306, 233, 235, 254, 886]. cutting [653]. Cycle [1483, 674]. Cyclic [1057, 334].

D [493, 320, 1620, 1402, 1537, 511, 388, 379, 371, 1753, 493, 546, 910, 369, 212, 495, 405, 383, 1270, 1676, 376, 410, 1742, 542, 384, 368, 320, 406, 359, 363, 361, 362, 1602, 1234, 1603]. D- [320]. D/ [493]. D02NOW [536]. Dairy [1253]. Dalang [553]. Dantzig [250]. Dantzig-Wolfe [250]. Darwin [111]. Data [1661, 1232, 1273, 450, 1453, 1540, 798, 863, 412, 1671, 691, 1131, 1221, 1524, 506, 1244, 509, 1265, 1004, 1231, 1525, 2000, 1674, 1347, 1252, 336, 429, 1241, 1242, 688, 875, 1258, 1822, 1613, 1995, 536, 376, 661, 12, 1088, 1722, 1723, 694, 1584, 689, 1090, 1294, 1469, 364, 432, 1256, 802, 1194, 1401, 1543, 328, 849, 508, 660, 421, 1622, 1254, 460, 1255, 840, 1447, 550, 1579, 839, 319, 1762, 1321, 838, 1468, 1991, 1396, 1248, 1234, 1397, 1534, 393, 209, 204, 645, 197, 1142, 646, 171, 202]. Data-Based [802]. Data-Centric [849]. Data-Driven [550, 1248]. data-intensive [204]. Data-Mining [1447]. Database [452, 1453, 498, 870, 416, 436, 1989, 410, 483, 1456, 674, 433, 618, 318, 499, 1520, 413, 488, 1455, 216]. Databases [1537, 1546, 412, 850, 2000, 1706, 1967, 845, 794, 12, 1535, 1070, 1544, 457, 851, 1454, 1543, 1536, 189, 190]. Dataflow [838]. Dataflow-Based [838]. Datasets [1229, 1533]. Dawande [226]. DAWN [1183]. DBMS [1469]. DCL [331]. Deadlock [55]. Deadlock-Free [55]. DEAL [1102]. DEALed [1726]. dealing [172]. Death [69]. Debate [1132]. debuggers [222]. Debugging [927, 286]. Decentralization [1586]. Decentralized [335]. Decidability [959, 973, 994]. Decidable [1214, 1333, 1352, 1353, 988]. Deciding [963, 958, 1343, 1062]. Decision [679, 1502, 1044, 665, 7, 577, 588, 672, 678, 666, 667, 1225, 986, 789, 992, 285, 694, 288, 1161, 658, 1222, 1151, 1235, 663, 747, 675, 1667, 218, 213, 159]. decision-making [213, 159]. Decision-Support [667, 675]. Declarative [969, 1559, 970, 1237, 180]. Decomposable [1343]. Decomposition [1329, 547, 145, 1302, 1713, 787, 545, 186]. Decompositions [956, 258]. Decorrelated [1104]. Dedicated [723]. DEDS [338]. Deduction [8, 984, 950, 951, 1011, 1542, 1158]. Deductive [1456, 997]. Deep [678]. Default [1121, 1124, 1135, 1140, 1651, 1146, 1650]. Defeasible [17, 1156, 1157]. Defined [956, 248]. Definite [991]. Definition [836, 830, 156]. Definitions [1216, 1057, 1635]. Deformable [483, 370]. Deformation [1604, 385, 376, 384]. Degree [1745]. Degrees [876, 806]. Delays [1677]. Delegation [67, 937, 1570]. Deliverables [626]. Demand [1854, 1599]. Demands [1854, 247]. Dementia [693]. Demonstration [815, 1570]. Dempster [1145, 1149]. Dempster-Shafer [1145, 1149]. Dendrochronology [1403]. Denied [267]. deny [1150]. Dependencies [1678]. Dependent [1624, 153, 1631, 93, 94, 1994]. Deployment [1518]. Depth [1323]. Derandomization [1846]. Derivatives [1867, 1376]. Derived [1035]. Descent [1139]. Described [1399, 1400]. Describing [735]. Description [125, 996, 998, 999, 665, 581, 379, 666, 1000, 1483, 126, 1001, 1003, 1712, 100, 1020, 127, 1007, 116, 1008, 997, 118, 123, 131, 132, 108, 1017, 726, 1019, 124, 163, 170]. Description-Processing [100]. Descriptions [114, 689, 1255]. Design [685, 1920, 1930, 1975, 280, 1412, 82, 1687, 561, 648, 782, 1413, 1076, 342, 614, 269, 1818, 1653, 1508, 1411, 626, 1016, 708, 510, 1971, 620, 570, 456, 753, 1510, 1733, 1511, 757, 578, 252, 846, 618, 306, 130, 1514, 1268, 844, 1461, 739, 1575, 725, 833, 1990, 538, 1871, 1160, 1518, 869, 458, 537, 187, 1659, 157, 646, 642, 207]. Designing [1221, 557, 1573, 640, 1896, 1418, 345, 1885, 151, 840]. Designs [1304, 1033, 286]. DESIRE [294]. Desk [1481, 1482]. Desktop [1915]. Destination [1569]. detail [1646]. Detecting [466, 446, 317]. Detection [462, 505, 1181, 394, 682, 1433, 680, 442, 1252, 688, 377, 1714, 444, 1613, 272, 780, 1256, 368, 439, 97, 1255, 393]. Determination [747, 790]. Deterministic [54, 1845, 11]. Detrimental [1494]. Developer [1413]. Developing [355, 1197, 606, 292, 731, 1890, 96, 828, 1200]. Development [1420, 837, 1414, 1726, 1475, 1476, 353, 351, 1481, 1482, 827, 621, 625, 1209]. Developments [1166]. Device [1876, 1893, 532]. Devices [1995]. DFC [1104]. diabetes [202]. Diabetic [1502, 668]. Diagnosis [1501, 288, 306, 191]. Diagnostic [1505, 697, 692, 736, 264]. Diagnostics [1518]. Diagram [622, 1751]. Diagrams [1750, 1053, 671, 612, 285, 615, 287, 1028, 1161, 618, 625, 1462, 623]. Diameter [1296, 1841]. Dictionaries [454, 1371]. Dictionary [1746]. Dienes [777]. Difference [1053]. Different [1043, 1082, 1285, 1152]. Differential [1788]. Differentials [1773]. Differentiating [1718]. Differently [333]. Diffusion [1709]. Digital [1791, 1437, 1892, 1548, 1430, 1888, 710, 1516, 447, 1872, 1468, 169, 175, 177]. Digitized [486]. Dijkstra [182, 1815]. Dimension [1526, 1527, 366]. Dimensional [464, 1296, 383, 1357, 812, 1678, 1368, 432, 1724, 367, 779, 1224, 1294, 1404, 205]. Direct [118, 532]. Directed [252, 1125]. Direction [780]. Directions [1442, 286, 1652, 1642]. Directives [523]. Directory [1373]. Disambiguation [1470, 1087]. Disc [815]. Disc-Motor [815]. Disciplined [633]. Discontinuities [1615].

Discourse [1191]. Discovering [179, 1090, 1236, 212]. Discovery [1138, 336, 215, 1322, 156, 190]. Discrepancy [1837, 1834]. Discrete [1755, 1064, 1655, 1171, 28, 1153, 195]. Discriminant [1238]. Discriminative [463, 1227]. Disease [682]. Diseases [396, 681, 675]. Disjoint [1306, 260]. Disjunctive [331]. Dispatch [934]. Displace [1732]. Displacement [934]. Display [592]. Dissemination [1920, 1983, 1532]. Dissimilarity [486, 478]. Distance [1759, 308, 876, 1245, 1388, 1617, 386, 1403]. Distances [1826]. Distortion [401]. Distortions [1402]. Distributed [1771, 1703, 856, 67, 312, 66, 1726, 1550, 805, 850, 315, 516, 1000, 1266, 1691, 1589, 1350, 53, 1002, 632, 456, 931, 757, 25, 531, 1696, 433, 723, 1373, 1866, 851, 1884, 45, 69, 865]. Distribution [411, 813, 377, 1230, 1722, 1723, 1803, 1820, 402]. Distributionally [888]. Distributionally-Hard [888]. Distributions [28, 1254]. Distributive [994]. Divide [1717, 1073, 1074]. Divide-and-Conquer [1073, 1074]. Divisibility [1654]. DLP [579]. DNA [210, 723, 1320]. Document [750, 487, 174, 164, 176, 166]. Documentation [1501]. Documents [1448, 1198, 169, 177]. Does [1267]. Domain [685, 712, 1474, 573, 1506, 547, 270, 1713, 281, 345, 620]. Domains [1047, 1216, 11, 1140, 855, 733, 1062]. DOME [1710]. Domination [1813]. DOMINO [1455]. Double [1113, 796]. Down [1222, 1945, 204]. Downward [1080]. Draft [1184]. Drama [14, 512]. Drawbacks [61]. Drawing [878]. Drawings [249, 641]. Drawn [486]. Drawn-Sketches [486]. Driven [1471, 703, 1460, 715, 541, 536, 410, 550, 1248]. Drug [1753, 672]. DSM [572]. Dual [228]. DUMBO [1512]. Duplication [1756].

During [824]. dyes [207]. Dynamic [563, 141, 1755, 1761, 1064, 1921, 1419, 55, 984, 1608, 1786, 1445, 1467, 1748, 718, 1309, 1215, 1465, 444, 436, 1166, 927, 855, 420, 1948, 1028, 1677, 906, 364, 1576, 1454, 1441, 1532, 398, 22, 1396, 1596, 1397, 657, 217]. Dynamical [1249, 1153]. Dynamically [524]. Dynamics [1121, 527, 1427, 1236]. Dynamo98 [120, 1351]. Dynastically [765].

E-Cash [1432]. e-Colabra [1463]. E-mail [1450]. E0.3 [1010]. E2 [1103]. Each [1375]. Ear [258]. Ear-Decompositions [258]. Earth [537]. Ease [1003]. Easy [1819]. ECG [682, 680]. Echocardiography [363]. Ecological [1930]. Economic [348]. Economics [1141]. Economy [1440]. Edge [241, 502, 1300, 247, 873, 884, 239]. Edge-Splitting [247]. Edit [1403]. EDM [354]. EEG [1231]. Effect [543]. Effective [1459, 1073, 1074, 1634, 491]. Effectiveness [474, 1433]. Effects [930, 527, 290, 1820]. Efficiency [1790, 34]. Efficient [1305, 1728, 657, 1275, 1053, 1044, 1785, 1073, 1074, 577, 493, 1999, 1757, 1672, 746, 603, 892, 1849, 1787, 1809, 1748, 1347, 339, 1299, 920, 916, 1366, 1289, 1897, 638, 1161, 1057, 1294, 1732, 1591, 1421, 480, 1534, 655, 640]. Efficiently [1535, 990, 1536, 1137, 650]. Efforts [283]. Egocentric [1668]. Egyptian [318]. Eigen [489]. Eigen-Eyes [489]. Eikonal [1601]. Elastic [373, 1758, 385, 390, 375]. Elderly [819]. Election [1368]. Electric [1170, 815]. Electrical [817]. Electronic [311, 672, 1428, 1433, 516, 313, 1565, 1564, 1799, 1431, 1772, 1563, 169, 1640]. Electrophoresis [1402]. Elementary [1049]. Elements [1228, 1642, 1185, 1071]. Elicitation [741]. Elimination [999, 1014, 1760, 604]. Elliptic [545]. ELSI [1505]. EM-algorithm [1622]. Embedded [1900, 1172]. Embedding [1198, 909, 1021]. Embeddings [1171, 253, 1684]. Emergence [1681]. Emergency [1902, 310]. Emerging [1584]. Emission [1621]. Emotional [1950]. Emphasis [1975]. Empirical [679, 1657, 1815, 1556]. Empty [256, 257]. Emulation [550]. Enabling [1581, 505, 1878, 1989, 297]. Enacting [834]. Enclosing [1293]. Encoded [1992]. Encodings [1310]. Encounters [1887]. Encrypt [1105]. Encryption [1438, 1797, 1798, 1787, 1106]. End [1883]. Endotheliums [395]. Energy [1615, 1607, 1999, 1604, 1603]. Energy-Based [1607]. Enforcement [1467]. Engine [795, 433, 1451]. Engineer [836]. Engineering [1974, 713, 1828, 1909, 1417, 352, 719, 262, 1410, 1510, 1188, 1964, 828]. Engines [1560, 758, 204]. English [198]. Enhance [670, 631]. Enhanced [134, 382, 494, 210]. Enhancement [762, 835]. Enhancing [663]. Enjoyable [90]. Enough [316]. Enrichment [715]. ensemble [217]. entailment [1128]. Entailments [1122]. Enterprise [1463, 1445]. Entities [1434]. Entropy [1147, 407, 1128]. Enumerability [901, 902]. Enumeration [1728, 907]. Environment [1670, 1414, 557, 1463, 516, 1549, 442, 1691, 1481, 1482, 731, 1693, 1953, 1985, 1908, 723, 283, 1514, 1461, 1699, 549, 105, 838, 1994, 1919, 177]. Environment-Aware [1919]. Environments [1669, 1901, 1894, 496, 1415, 1550, 1950, 1490, 1987, 1984, 1983, 1695, 1986, 1960, 1571, 1643, 1642, 1663]. EPI [401]. Epileptic [793]. Epistatic [765]. Epistemology [1666]. Epsilon [593]. Epsilon-Terms [593]. Equal [1172]. Equality [1063, 577, 1635, 1062]. Equation [1601]. Equational [963, 971, 990]. Equations [988]. Equipment [1505, 313]. Equivalence [1797, 1688, 1054, 1352]. Equivalences [1344]. Erika [116]. Erratum [1383]. Error [1746, 416, 721, 754, 402]. Error-Tolerant [416]. Errors [317, 375].

Essential [1185]. Established [1588]. Establishing [26]. Esterel [1033]. Estimate [755]. Estimating [1826, 409]. Estimation [394, 1612, 1609, 1247, 161]. Euclidean [1296, 1297, 386, 654]. Europe [742]. Europeanwide [813]. Evaluating [1259, 749]. Evaluation [1779, 1474, 804, 48, 291, 1042, 1816, 1805, 144, 1911, 1812, 251, 1822, 54, 1086, 1370, 1732, 1541, 1538, 130, 1557, 1990, 675, 530, 1905]. Evaluative [1191]. Event [1064, 1171, 688, 1051, 19, 1154]. Events [793]. Everglades [44]. Everyday [1900]. Everything [660]. Everywhere [1873]. evidence [206, 1153]. Evolution [1959, 633, 839, 553, 105, 221]. Evolutionary [809, 769, 746, 756, 1484, 1485, 876, 755, 1281, 1302, 706, 1366, 222, 1224, 1286, 211]. Evolving [111, 208, 1839, 368]. Exact [1303, 1318, 654]. Examining [1490]. Example [1852, 302]. Examples [1703, 1975, 387, 418, 181]. Exceptions [1375]. Exchange [1545]. Excitement [945]. Excluding [419]. Executing [119]. Execution [519, 63, 628, 1185, 1561, 46]. Existential [965, 959]. Expanders [1308]. Expansion [1451, 459]. Expansions [357]. Expectation [1517]. Expected [1614]. Experience [685, 622, 1039, 1490, 610, 1575]. Experiences [1717, 1964, 1228, 1427, 1544, 669]. Experienceware [1487, 1486]. Experiment [1096, 190]. Experimental [713, 1825, 1831, 1913, 1816, 1808, 251, 1822, 642]. Experimentation [1814]. Experimenting [1428]. Experiments [193, 1810, 1078, 826]. Expert [683, 731, 297, 1723, 674, 725, 187]. Expertise [740]. Experts [724]. Explained [19]. Explanation [297, 1942]. Explanations [1126]. Explicit [1326, 1006, 1810, 101]. Exploitation [1722, 737]. Exploiting [670, 1063, 1261, 1262, 1241, 1242, 1553, 1254, 428]. Exploration [1473, 1449, 171]. Exploratory [10]. Exploring [1881]. Exponential [891]. Express [1375, 609]. EXPRESS-G [609]. expressed [172]. Expression [977, 1821]. Expressions [1180, 1452]. Expressive [1471, 1928]. Extended [1326, 1275, 1327, 1291, 1328, 1785, 1330, 1790, 1795, 1810, 1335, 1269, 1303, 1342, 298, 1345, 1801, 1770, 1347, 1283, 1352, 875, 1140, 1357, 1358, 766, 1365, 1289, 1766, 1294, 1373, 1071, 468, 1376, 1799, 872, 1382, 1619]. Extending [163, 1823, 1416, 1424, 1191, 736, 637]. Extensible [73, 1633, 1544]. Extension [1528, 1529, 325, 623, 553, 1150]. Extensional [1022]. Extensions [634, 1027, 183]. Exterior [1908]. External [1831, 1284, 875, 299, 1762, 1321]. Externalization [170]. Externalized [935, 936]. Extracted [587]. Extraction [331, 440, 437, 1506, 504, 1190, 461, 469]. Extreme [239]. Extremely [750]. Eye [1259, 706]. Eyes [489].

F [1005]. Face [505, 1935, 318]. Facial [500]. Facilitate [634]. Facilities [1672, 297]. Facility [1759, 234, 895, 550]. FaCT [579]. Factor [399, 398, 530, 244]. Factoring [1784]. Factorization [1669]. Factors [1757, 1912, 1969]. Failing [884]. Failure [442, 551, 1260, 290, 1255]. Fair [1440, 1444]. Fake [1315]. False [543]. familiar [1642]. Families [1307, 1207, 866, 735, 1834]. Fan [1629]. FAQ [1958]. Fast [1857, 1776, 49, 1829, 505, 917, 1271, 1777, 1293, 980, 981, 554, 919, 1310, 1774, 1104, 403, 254, 1404, 1821, 656, 404]. Faster [1303, 886]. Fault [501, 1849, 817, 1002, 1270, 1747, 1512]. Fault-Tolerance [1849]. Fault-Tolerant [1002, 1270]. Faults [1445]. Feasibility [1500, 525]. Feasible [229]. Feature [440, 464, 437, 376, 461, 469, 447, 1226, 203]. Feature-Map [447]. Features [463, 284, 1673, 1934, 461, 468, 1227, 370, 224]. Federated [1461]. Feedback [232, 423, 425, 477]. Feedforward [818]. FEM [539]. Fertility [360, 361]. Fetal [745]. Few [1742]. FFT [1056]. FHFC [753]. FHFC-Tool [753]. Fiduccia [648]. Fiduccia-Mattheyses [648]. Field [1891, 138, 509, 739, 1622]. Fieldbus [141]. Fields [1610]. FIIT [1693]. FIL [1778]. FIL-MACs [1778]. File [1459]. File-Block [1459]. Files [1692]. Filling [1528, 1529]. Filter [762, 320, 1560, 428, 505]. Filtering [687, 1132, 1554, 1450, 1264]. Financial [303, 1435, 1256]. Find [283]. Finding [1306, 1395, 1606, 1751, 1407, 335, 653, 873, 1856, 802, 1405, 1811, 1406]. Fine [1725, 942]. Fine-Grain [1725]. Fine-Grained [942]. Finite [887, 1341, 1354, 904, 1399, 1400, 1358, 1041, 905, 1372]. Finland [119]. Fire [44]. First [1015, 985, 977, 1079, 1343, 595, 1007, 1651, 1012, 601, 161, 1650, 180]. First-Order [985, 977, 1079, 1343, 595, 1007, 1012, 601, 180]. Fitness [819, 176, 222]. Fitting [1605]. Five [1182]. Fixed [1835, 389, 916, 1756]. Fixed-Parameter [1756]. Fixed-Size [916]. Flexibility [47, 869]. Flexible [554, 270, 992, 834, 1877, 347, 740, 1593]. Flexibly [1493]. Floods [310]. FLORES [488]. Flow [228, 49, 238, 270, 246, 251, 806, 1431]. Flow-Shop [270]. Flows [1676, 735]. Fluid [1676, 374]. Fluke [502]. Fly [1714]. FMR [733]. fMRI [1622]. Focus [503]. Focusing [819]. Fold [325]. Folding [1332, 1283, 325]. Follicle [360, 361]. Football [123]. Footprint [1499]. Footprint-Based [1499]. Footwear [1511]. Forecasts [913]. Forest [1856]. Form [588, 530, 1603]. Formal [1006, 943, 1029, 630, 1004, 351, 40, 352, 1961, 1427, 721, 562, 35, 1210, 608, 1457, 635, 978, 345, 1636, 1663]. Formalisations [261]. Formalism [1204]. Formality [616]. Formalization [789, 1005]. Formalizing [1367, 1645]. Formally [718]. Formation [799, 1552]. Formats [1344]. Forms [1092, 947, 172, 1152]. Formulae [958]. Formulas [956, 1062, 1027, 1123]. Formulation [482]. Forward [1791, 385, 1987, 323]. Forward-Secure [1791]. Foundation [1042, 168]. Foundations [928, 1491]. Founded [1055]. Four [383]. Four-Dimensional [383]. FPTAS [1382]. Fractal [807, 469, 492, 195]. Fragment [588]. Fragments [1214, 1333]. Frames [1668]. Framework [1920, 636, 607, 858, 388, 696, 1170, 380, 1189, 341, 1353, 992, 471, 1565, 1953, 1888, 95, 1574, 1590, 263, 432, 997, 306, 1576, 1421, 740, 1997, 1593, 1864, 1158]. Frameworks [311, 346]. Free [1847, 900, 958, 55, 592, 1792, 595, 244, 1858, 889, 779, 601, 1603]. Free-Form [1603]. Freedom [806]. Freehand [406, 362]. Freiburg [88, 87]. French [1942]. Frequency [689]. Freyd [1374]. Freyd- [1374]. Friendly [1692]. Frontal [505]. FT [1170]. Full [601]. Fully [1309, 199, 24, 960]. Function [1290, 1736, 809, 478, 680, 1508, 377, 1094, 406, 206, 222]. Functional [400, 1081, 378, 529, 526, 675, 212, 1644, 558]. Functions [1326, 773, 1779, 1759, 1063, 577, 1111, 1604, 1777, 594, 1030, 248, 1775, 1732, 962, 1145]. Fusion [509, 1613, 1155]. Future [568]. Fuzziness [743]. Fuzzy [795, 773, 776, 759, 771, 751, 761, 782, 1131, 788, 269, 762, 752, 810, 784, 791, 270, 770, 490, 794, 1245, 787, 785, 744, 753, 392, 733, 760, 1243, 823, 783, 267, 367, 786, 1543, 268, 1227, 745, 774, 772, 796, 264, 797, 265, 824, 266, 200, 183, 1152]. Fuzzy-Control [753]. Fuzzy-Logic-Based [794].

G [609]. GA [275]. GAL [1819, 1818]. Gale [259]. Galois [337]. Games [93, 94]. Gap [1395]. Gated [381]. Gathering [1477]. GAUSS [145]. Gaussian [1611]. Gaze [503]. GBI [1530]. GCD [903]. GDT [220]. GDT-RS [220]. Gene [1756]. Gene-Duplication [1756]. General [795, 1327, 1310, 1740, 1386, 413]. General-Purpose [795]. Generalisation [1179]. Generalization [816, 1736]. Generalized [911, 1530, 540, 913, 1372, 1611, 163, 215]. Generalizing [1092, 1523]. Generate [1471]. Generated [775]. Generating [1311, 506, 821, 1742, 599, 1516]. Generation [1118, 773, 751, 388, 1046, 346, 275, 454, 1770, 1508, 1035, 1946, 390, 1240, 1023, 1584, 288, 1161, 1724, 276, 1533, 487, 330]. Generator [810, 1846, 1007]. Generators [1508, 1308]. Generic [149, 1790, 150, 346, 420, 457]. Genericity [929]. Genetic [280, 764, 478, 1616, 825, 1244, 277, 278, 1675, 225, 754, 811, 281, 276, 279, 747, 210]. Genome [1390, 877]. Genomic [1392]. Genus [1345]. Geodesic [366]. GeoFEM [539]. Geographic [1657, 1652]. Geologic [1542]. Geometric [1295, 365, 1747, 647]. Geometry [881, 371]. Geophysical [539]. German [1553]. GESTALT [1392]. Gestural [1954]. Gesture [1940, 1939, 1938, 1951, 1936]. Gestures [1928, 1941, 1953, 1946, 1952, 1929]. GETESS [1553]. Getting [117]. Gibbs [1608]. Gigahertz [1032]. Given [227, 176]. Global [1037, 1364, 1365, 451, 117, 1461, 538, 840]. Glory [122]. GMD [131, 132]. Go [1749]. Go-with-the-Winner [1749]. Goal [868]. Goal-Satisfaction [868]. Goals [1204, 1156, 1157]. Goldreich [1780]. Goldreich-Goldwasser-Halevi [1780]. Goldwasser [1780]. Good [1273, 97, 1642]. GP [275]. Graded [987]. Gradient [282, 1139]. Gradient-Based [282].

Gradual [798]. Grain [1725]. Grained [942]. Grammars [956, 1943]. Grand [1234]. Granularity [858]. Granulation [1665]. Graph [452, 1205, 379, 1829, 899, 1855, 956, 1616, 1738, 1339, 746, 1265, 1816, 1342, 1196, 1199, 871, 1204, 1090, 1193, 1191, 253, 451, 841, 279, 1217, 872, 972, 878, 1662, 646]. Graph-Based [1205]. Graph-Theoretic [452]. grapheme [198]. Graphic [1076, 319]. Graphical [506]. Graphics [541]. Graphs [1306, 1214, 1211, 1298, 1730, 885, 1212, 1195, 1761, 908, 1202, 880, 1343, 1197, 1345, 1207, 1840, 244, 1310, 1309, 1215, 879, 1210, 886, 905, 1308, 1841, 1208, 1194, 1184, 1665, 1523, 258, 1684, 1645, 641]. Grasping [1933]. Grasshopper [1581]. Greedier [920]. GREFIT [1932]. Grey [282]. Grey-Gradient [282]. Grid [249, 879]. Grids [896, 741, 1652]. Gröbner [1625]. Ground [622, 957, 1129, 1130]. Group [1442, 1790, 400, 1721, 812, 1557]. Group-Theoretic [1721]. Grouped [543]. Grouping [1606]. Groups [1679, 1112, 1441]. Growing [973]. GSM [1980]. Guarantee [1043]. Guarantees [1737, 1848]. Guarded [986]. Guidance [836]. Guided [134]. Guideline [663]. Guidelines [658, 1495, 662, 686]. Guidelines-Based [662].

Hair [699]. Halevi [1780]. Half [891]. Half-Exponential [891]. Hamiltonian [1601]. Hand [486, 1953, 1935, 1932, 1910, 191]. Hand-Held [1910]. Handheld [1882, 1876, 1911, 1897, 1875, 1871]. Handling [1037, 112, 1137, 1152, 333, 18]. Hands [703]. Hands-Busy [703]. Hard [1311, 1326, 1415, 888, 1570, 757, 1303]. Hardness [1853, 1765]. Hardware [56, 1024, 1706, 1587, 286]. Harmonic [1603]. Hash [1777, 1732, 1777]. Haskell [154]. Head [694]. Headed [769]. Health [700, 659]. Healthcare [1914]. Heap [1828]. Heaps [874]. Heart [392, 745]. Hebrew [1470]. Held [1910]. Help [1175, 1481, 1482, 1281, 1567]. Help-Desk [1481, 1482]. Hemodynamic [378]. HEPASCORE [675]. Herbrand [599]. Hereditary [897]. Heterogeneous [1669, 850, 851, 158]. Heuristic [1478, 813, 336, 1749, 648, 652]. Heuristics [1813, 862, 642]. HFE [1764]. Hidden [818, 1240, 1765]. Hide [1436]. Hiding [1443]. Hierarchical [1325, 1272, 622, 1607, 956, 437, 744, 376, 323, 1226, 1237, 370, 1647]. Hierarchically [879]. hierarchies [1646]. Hierarchy [891, 268, 1645]. High [1118, 209, 450, 1032, 138, 1582, 1296, 1691, 1992, 539, 834, 812, 689, 521, 50, 572, 1224, 460]. High-Dimensional [1224]. High-Frequency [689]. High-Level [138, 1691, 834, 460]. High-Performance [1032, 539, 1968, 572, 209]. High-Speed [1118]. Higher [1745, 1022, 977, 1610, 966, 989]. Higher-Order [1022, 977, 966]. Highly [463, 1676, 1775]. Higman [962]. Hilbert [1175, 593]. Hill [1616]. Hints [1446]. Hippie [1907]. Histograms [1537]. History [77]. History-Based [77]. Hitting [1846]. HMM [1943]. HMM-Based [1943]. Hoarding [1922]. HOBE [828]. Hoc [1981, 1925]. HOL [977]. HOL- [977]. Holism [727]. Holographic [1379]. Homogeneity [395]. Homogeneous [448]. Hospitals [670]. Hot [792]. HotHelmet [1902]. House [828]. HP [1283]. HPF [515]. HTGviz [541]. HTMT [46]. HTN [1495]. Huge [1401]. Human [1937, 1187, 805, 1231, 1912, 993, 1934, 1954, 360, 1969, 361, 170]. Human-Oriented [993]. Human-Robot [1937]. Humboldt [113]. Hungarian [1082]. Hybrid [1459, 365, 341, 1199, 1822, 1166, 54, 1511, 102, 103, 600, 32, 1160]. Hybridization [1391, 1390]. Hydrodynamic [528]. Hyper [585]. Hypercubes [896]. Hypergraph [1389, 1812]. Hypergraphs [1835]. Hyperinference [815]. Hypermaps [1542]. Hypermedia [1501, 456, 854]. Hyperparameter [1609]. Hypertextual [454]. Hypotheses [1075].

I/O [1258, 543, 1683]. IADL [819]. IBM [1032]. ICARUS [1518]. Iconicity [1942]. ICU [696]. idea [170, 1107]. Ideas [1320]. Identical [1278]. Identification [1753, 817, 821, 363, 675]. Identified [502]. Identify [676]. Identifying [1934, 841]. Identities [1443]. Identity [1653, 1211]. IDL [845]. IEEE [142, 27, 140]. II [1354]. III [734]. IK [218]. ILF [1183]. Illegal [1181]. Illumination [444]. ILP [1075, 1085, 1096]. ILP-Assisted [1096]. ILU [545]. Image [1620, 778, 452, 463, 419, 440, 462, 472, 1607, 414, 373, 423, 1608, 425, 474, 464, 465, 762, 1507, 651, 429, 405, 484, 422, 1609, 490, 471, 436, 476, 410, 483, 424, 467, 451, 433, 453, 384, 1513, 1194, 468, 322, 1617, 1611, 499, 481, 1699, 477, 413, 401, 459, 470, 370, 488, 428, 480, 492, 458, 189, 196]. Image-Based [481]. Image-Retrieval [410]. Images [1402, 486, 1437, 805, 493, 435, 416, 365, 430, 1399, 1400, 766, 392, 409, 529, 781, 1621, 407, 367, 368, 323, 1602, 1603]. ImageScape [426]. Imaging [399, 509, 357, 364, 1621]. Imitation [1933]. Impact [1269, 1588]. Imperative [1628]. Implementation [1920, 1581, 1825, 1819, 1816, 1818, 44, 1809, 1559, 38, 1457, 1008, 823, 1698, 1716, 982, 572, 1160, 1177, 458, 648, 638, 1645]. Implementing [1685, 56, 1688, 1192, 639, 151, 162]. Implicant [1380]. Implication [777]. Implications [776, 1765]. Implicit [1004, 392, 601]. Importance [273]. Impredicative [594]. Imprimitive [1112]. Improve [423, 1162]. Improved [1292, 234, 372, 1098, 1046, 1842, 1846, 1446, 1722, 1844]. Improvement [444]. Improvements [585, 1003, 523]. Improving [185, 917, 1087, 468, 1288]. Improvised [283]. In-vehicle [1941]. Inapproximability [1328, 1380]. Incomplete [1355, 216]. Incompressibility [1330]. Incompressible [1676]. inconsistency [1137]. inconsistent [197]. Incorporating [401]. Increase [1986]. Increased [697]. Incremental [727, 291, 139, 733, 1137]. independence [1133, 1148]. Independent [1835, 1833, 1752, 897, 920, 1834, 1862, 1301]. Index [1978, 23, 137, 576, 946, 983, 1097, 1120, 1159, 1384, 1804, 1832, 1926, 1955, 2001, 459, 1155]. Indexation [462]. Indexing [1746, 284, 440, 728, 472, 434, 437, 442, 1526, 1527, 1398, 467, 432, 489, 460, 177]. Indistinguishability [1797]. Indistinguishability-Based [1797]. Individual [502, 114]. Individualized [1497]. Individuals [1921, 1742]. Indoor [321]. inducing [175]. Induction [1073, 1074, 1028, 1222, 601, 220, 218, 225, 203, 190]. Inductive [816, 1127, 1076, 98, 1136, 1080, 1081, 1016, 1070, 1367, 1090, 325, 181]. Industrial [355]. Industry [1253, 302]. Inequalities [230, 231, 243]. Inequality [1737, 1775]. Infants [673]. Inference [1620, 22]. Inferences [1607]. Inferring [1126, 1258]. Infinite [1038, 1654]. Infinite-State [1038]. Influence [671, 1664, 1148]. Informal [616, 1886, 608]. Information [1920, 856, 1767, 1902, 703, 1440, 1477, 1263, 1231, 516, 1261, 1262, 1549, 716, 1555, 1310, 430, 427, 429, 1892, 1548, 866, 307, 1877, 750, 1510, 1983, 410, 1925, 1577, 1766, 403, 1651, 1980, 1908, 1907, 117, 1803, 1209, 1552, 828, 1532, 1567, 1155, 1885, 559, 725, 867, 1990, 1545, 1264, 1571, 159, 1994, 404, 1650, 205, 172, 167]. Information-Theoretic [1766]. Information-Visualization [410]. Informational [1286]. InfoStick [1893]. Infrared [323]. Infrastructure [1548, 1600, 20]. Ingest [411]. Inhabitants [1950]. Initial [1715, 1575, 1556]. Initialization [1238]. Initiative [1205, 1727]. Injury [694]. Ink [1446]. inka [998]. Inlining [932]. Innovations [1687]. Innovative [107]. Input [1114, 1249, 1897, 1376, 1227, 330]. Input-Output [1249, 330]. INRECA [355]. Insect [1637, 372]. Insensitive [1036]. Insertion [1530]. Insights [639]. Inspection [320]. Inspired [698, 210]. Instance [1473, 1094, 200]. Instance-Based [1473, 200]. Instances [226, 1311]. Instantaneous [1696]. Instantiations [1062]. Instruction [518, 1810, 194]. Instruction-Level [518]. instructions [1643]. Instrumental [1929]. Insurance [1551]. Integer [228, 1275, 917]. Integers [903, 1707]. Integral [991, 381, 248, 1144]. Integrated [1624, 1670, 443, 291, 843, 1589, 106]. Integrates [684]. Integrating [1501, 1502, 496, 1999, 1478, 678, 1787, 621, 1510, 1469, 1884, 1969, 1237]. Integration [864, 1460, 1318, 1798, 716, 87, 88, 1711, 719, 992, 128, 12, 375, 347, 323, 1952, 217]. Integrity [535, 1143, 301, 618]. Intellectual [1440]. Intelligence [122, 353, 12, 792, 1879, 659, 699]. Intelligent [1581, 687, 138, 1221, 346, 1231, 308, 821, 329, 341, 307, 1551, 436, 1230, 1554, 391, 345, 1532, 21, 1661, 164, 172, 173, 177]. Intensity [409, 407]. Intensive [698, 204]. Intention [1556]. Intentional [977]. Intentions [15]. Inter [852, 1555, 620, 1893, 1692, 853]. Inter-Agent [1555]. Inter-appliance [1893]. Inter-domain [620]. Inter-model [852, 853]. Inter-Process [1692]. Interaction [1771, 1901, 1937, 1417, 1939, 422, 1893, 1896, 1888, 420, 1954, 1880, 428].

Interactions [1886, 1610]. Interactive [496, 602, 442, 482, 1971, 794, 42, 1489, 410, 14, 473, 1495, 391, 1451, 501, 1335, 1336]. Intercession [557]. Interchange [876, 1887]. Interdomain [1596]. Interest [462]. Interests [1568]. Interface [1718, 1583, 74, 436, 624, 536, 1953, 1569, 1866, 50]. Interfaces [1573, 1913, 79]. Interference [1580]. Interleaving [1493]. Internal [935, 936, 299, 645]. Internet [711, 864, 1416, 1867, 434, 667, 348, 1464, 1997, 20, 1566]. Internet-Based [667, 348]. Interoperability [1956]. Interoperable [611]. Interoperative [731]. Interpolate [744]. Interpolation [1117, 1370, 406]. Interposing [74]. Interposition [74]. Interpretation [1624, 504, 1629, 1940, 1660, 267, 322, 206, 1144]. Intersections [783]. Interval [678, 1309]. Intervals [334]. Interworking [1168]. Intracerebral [379]. Intractable [915]. Intranet [1571]. Intranet-Based [1571]. Introducing [1975, 80]. Introduction [149, 514, 512, 334, 256, 257, 513, 515]. Intrusion [1252]. Intuition [1183]. Intuitionisitic [587]. Intuitionistic [586]. Invariant [463, 466, 1182, 465, 1045, 444, 470, 22]. Invariant-Preserving [1182]. Inverse [1095]. Inverted [796]. Inverter [767]. Investigation [1556]. Investing [1443]. Invited [986, 1733, 1742, 1011, 1021, 1749, 1762]. Involving [588, 663]. IP [1583, 1271, 1924]. Irregular [1825, 55, 47]. Irregularity [407]. Irrelevance [1133]. Isabelle [1631]. Ischemic [680]. ISocRob [125]. Isolda [13]. isotope [399]. issuance [1435]. Issues [1442, 1825, 461, 1890, 1520, 259, 216]. Iterated [1112, 198]. Iteration [1626, 935, 936]. Iterations [250]. Iterative [525, 1607].

J [75]. J-Kernel [75]. Jaundice [671].

JAVA [436, 488, 569, 1066, 926, 925, 1961, 570, 1165, 942, 938, 1200, 75, 553, 46]. JAVA-Based [436]. Java-Processes [1165]. jContractor [570]. JDSL [645]. Jeopardy [948]. Jet [779]. Jitter [534]. Job [1851]. Join [1540, 1541, 1730]. Joining [1539]. Joins [1538]. Joint [391]. Judge [1239]. Jukebox [1459]. Jumping [554]. Junction [782, 22]. Junctions [1606]. Jury [1239].

KADS [346]. Kava [553]. Kawasaki [1608]. KBDistAl} [1248]. Kernel [75]. Key [1785, 1266, 1770, 1764, 1434, 1803, 1113]. Keyed [1106]. Keynote [1319]. Keys [1347, 1434, 1436]. Keywords [455, 176]. Khufu [1107]. Kimba [1007]. Kinetic [1674, 1349, 529]. KK [1018]. Knapsack [1838]. know [181]. Knowledge [685, 1232, 619, 4, 711, 727, 713, 7, 1076, 337, 1503, 1504, 1795, 714, 316, 1335, 1794, 1269, 715, 730, 1507, 1136, 1080, 1480, 1793, 718, 1138, 1251, 336, 707, 352, 719, 688, 732, 294, 339, 156, 1489, 1087, 1198, 733, 1722, 1188, 1090, 709, 267, 710, 736, 333, 724, 737, 701, 1336, 295, 664, 738, 669, 739, 725, 1237, 686, 299, 190, 289, 264, 704, 160, 168, 199, 191, 167]. Knowledge-Based [1232, 715, 688, 732, 267, 669, 289]. Knowledge-Level [1480]. Knowledges [816]. KoMoD [739]. KtSeqC [581]. KULRoT [98]. KUMP [542]. KUMP/ [542]. Kyoto [1548].

L0M [774]. Lab [121]. Labeled [319]. Labeling [1618]. Labelling [1301]. Labels [1758]. Laboratory [972]. ladder [1644]. Lambda [1626]. lambdaProlog [1008]. Lambek [590]. Laminar [1307]. Lamport [1679]. Landmark [388, 375]. landmarks [1640]. Landscape [44]. Language [60, 450, 153, 703, 1419, 617, 1943, 632, 1196, 1199, 1720, 456, 616, 1946, 628, 1456, 978, 566, 1462, 306, 1866, 1401, 702, 1944, 152, 1201, 1942, 704, 199, 180]. Languages [1420, 887, 1314, 1045, 892, 1342, 888, 965, 1198, 889, 1021, 1418, 345, 982, 151, 1246]. LANs [143]. Lapis [1083]. LARA [1582]. Large [1847, 340, 1784, 1233, 497, 1826, 1713, 750, 832, 1637, 199]. Large-Scale [497, 1637]. Latencies [521]. Latency [1036, 58]. Lateral [785]. Lattice [337, 880, 527, 1208, 256, 257, 178]. Lattices [1305, 1283, 994, 774]. Law [304]. Laws [797]. Layer [1000, 779]. Layered [99, 878]. Layout [1730, 880]. Lazy [1741, 1809, 1148, 1367]. LEAD [563]. Leader [1368]. League [121, 1869, 115, 1870, 118, 124]. Leaks [926]. Learn [1092, 181]. Learnability [1088, 1220]. Learnable [1220]. Learner [304, 308]. Learner-Model [308]. Learning [331, 1195, 426, 696, 1689, 676, 1076, 1478, 820, 387, 1135, 269, 730, 892, 1228, 1136, 308, 329, 1082, 197, 1083, 1084, 128, 692, 661, 327, 1253, 693, 101, 507, 109, 326, 551, 1680, 1071, 1094, 698, 659, 104, 108, 1933, 826, 1322, 198, 695, 200, 214, 181, 164, 1142, 175]. LEDA [1823, 1816, 1807]. LEDA-SM [1823]. Left [973]. Left-Linear [973]. Legacy [831]. Legal [1257]. Legged [91, 136, 135]. LEGO [1632, 92]. Lemma [962]. Lemmata [604]. Lempel [1401, 1386]. Length [1114, 468, 491, 638]. length-restricted [638]. Lesions [393]. less [181]. Let [1132]. Letters [701]. Level [450, 518, 138, 1878, 1314, 714, 1480, 1691, 554, 352, 471, 834, 1862, 460, 120, 1351]. Levels [686, 1646]. levelwise [180]. Leveraging [346]. Levinthal [1332]. lexicographic [1128]. Lexicographically [1396, 1397]. Lexicon [1192]. Libraries [1687, 411, 557, 710, 1468, 169]. Library [1830, 49, 570, 1491, 684, 645, 175]. Life [1483, 674, 223]. Life-Cycle [674]. Lifecycle [1698]. Lifelike [10]. Lightweight [1863]. like [1144]. Likelihood [1620]. Limb [1931]. Limited [890, 1767, 199]. Limited-syntax [199]. Line [1817, 1279, 1272, 910, 1751, 350, 497, 1966, 1934, 1289, 1729, 793, 1815, 1444, 161]. Line-Pattern [497]. Line-Scan [1934]. Linear [1238, 1843, 1686, 1044, 1099, 373, 1819, 1539, 235, 1046, 682, 892, 1740, 1851, 1713, 1357, 1297, 596, 973, 39, 374]. Linear-Elastic [373]. Linear-Time [1357, 1297, 39]. Linearity [589]. Lines [1817, 350]. Linguistic [563, 1221, 822, 794, 1087, 1150]. Linguistically [267]. Linguistics [712]. Linguistics-Based [712]. Lingware [351]. Linked [1347, 1288]. Linking [1055, 833]. Links [1981, 454, 1188, 722]. LINSOL [1711]. linTAP [596]. List [1290, 1810, 1506, 364]. Listings [1516]. Lists [1355, 1288]. Literate [619]. Live [1267]. Liver [396, 675]. LLC [142]. Load [1272, 1280, 517, 530]. Local [122, 1829, 371, 377, 1364, 1365, 1982, 905, 117, 407, 1697, 276, 439, 344, 470, 1123, 1158]. Localization [1937, 87, 392, 107, 88]. Locally [1343]. Locating [1996]. Location [1759, 1666, 911, 1996, 234, 1878, 817, 1877, 895, 1922, 1886, 1908, 551, 1862, 1994, 1894]. Location-Aware [1922, 1908]. Location-Based [1877]. Location-Dependent [1994]. Location-Independent [1862]. Lock [1986]. Locomotive [1518]. LOGCFL [1346]. Logfile [97]. LOGFLOW [542]. Logic [1324, 998, 4, 1124, 1044, 586, 1063, 577, 782, 1131, 1690, 1046, 574, 985, 977, 1341, 1045, 1080, 298, 1206, 791, 595, 580, 1081, 1258, 794, 1023, 1147, 596, 1088, 597, 25, 775, 1090, 783, 1095, 268, 20, 173, 987, 575, 1012, 824, 204, 160, 161, 183, 162]. Logic-Based [4]. Logical [956, 438, 997, 1158, 167]. Logically [1696]. Logics [996, 584, 1048, 592, 591, 985, 1483, 986, 1353, 1058, 1007, 1379, 600, 163]. Long [1267, 1488, 1707, 102, 103]. Long-Term [1488]. Longitudinal [785]. Look [991]. Look-Up [991]. Lookahead [1671].

Looking [511, 1521, 1522, 1987, 323]. Lookup [1271, 1371]. Loop [526]. Loops [1982]. Lossless [493]. Lossy [26]. Love [13]. Low [1314, 1335, 58, 753, 1371, 1373, 1834, 1336]. Low-Cost [753]. Low-Level [1314]. Low-Randomness [1335, 1336]. Lower [1326, 1277, 1286, 1]. LP [1848]. LP-based [1848]. LR [1083]. Luby [1110]. Luby-Rackoff [1110]. Lucent [494]. Lungs [392]. LV [383]. LZW [1385].

M [1886]. M6 [1108]. Machine [1859, 1909, 696, 214, 164, 730, 1850, 692, 661, 327, 693, 1848, 1008, 1954, 1071, 698, 1382, 46, 695, 178]. Machines [1860, 1325, 1278, 1875, 1696, 255, 1060]. Machining [330]. Macro [1375]. Macro-Express [1375]. MACs [1778]. MAD [288]. Made [1271, 1793, 1294]. Magic [1446, 1914]. MagiCol [126]. Magnetic [1621, 367]. mail [1450]. Maintain [869]. Maintainability [870]. Maintaining [1347, 1872, 1560]. Maintenance [1366, 660]. MAINZ [118]. Majority [905]. Make [90]. Making [1438, 1502, 672, 789, 1151, 1110, 611, 213, 159]. malleable [1797]. Mammograms [377]. Manage [1693, 664]. Management [1273, 450, 67, 858, 1502, 671, 66, 1550, 1996, 310, 2000, 1589, 860, 861, 620, 847, 1512, 1464, 668, 1590, 709, 16, 737, 1591, 1986, 560, 738, 1543, 1461, 1593, 1517, 832, 865, 1432]. Managing [837, 1594, 1503, 1504]. Mandala [430]. Manhattan [1836]. Manifold [786]. Manipulation [1489]. Manipulators [797]. manual [194]. Manufactured [1743]. Manufacturing [350, 349, 348, 347]. Many [903, 1007, 775, 1379]. Many-Valued [1007, 1379]. Map [1540, 1738, 1758, 544, 447, 1301, 266, 1664, 1154, 1646]. Mapping [609, 1389, 522, 409]. Maps [400, 277, 680, 1818, 1233, 476]. Market [226, 311]. Marketing [1506]. Markov [764, 1332, 1240, 1610, 1622]. Markovian [1065]. Marriage [1355, 259]. MARS [418]. MAS [1425]. Mask [1730]. Maslov [995]. Mass [393]. Masses [377]. Massive [1540]. Match [502]. Matching [1402, 463, 1746, 373, 369, 380, 1616, 915, 1385, 455, 1535, 1404, 1386, 1398, 1617, 1387, 370, 949, 1403, 989, 1536, 161, 219, 639]. Matchings [1302]. Matchmaking [1568]. Material [385]. Materials [804]. Mathematical [1687, 1042, 272]. MathWeb [1000]. Mating [1015]. Matrices [882, 1370, 2, 1383]. Matrix [892]. Matroid [907]. Mattheyses [648]. Maude [968]. Max [227, 232, 1852, 242]. Maxima [465]. Maximal [1395]. Maximally [233]. Maximum [227, 1306, 1620, 229, 910, 238, 1302, 920, 1147, 1301, 1128]. MaxSat [1369]. May [24]. MCMCML [1609]. MCS [1020]. Mean [893, 1622, 894]. Meaningful [732]. Means [395, 758, 185]. Measure [407, 1617, 200, 176]. Measurement [363, 386, 362, 1487, 1486]. Measurements [871]. Measures [36, 394, 405, 1245, 1086, 1220, 698, 224]. Measuring [1324, 395]. MEC [1688]. Mechanical [673, 739, 194]. Mechanism [423, 1922, 1990]. Mechanisms [1671, 1508, 59, 1563]. Mechanization [591]. Mechanizing [1054, 301]. Media [560]. MediaCup [1900]. Medial [363]. median [1297]. Mediated [1939, 1562, 1566]. Medical [805, 414, 672, 678, 1505, 509, 697, 683, 365, 707, 690, 281, 684, 368, 702, 664, 669, 698, 704]. Medicine [399, 661, 660, 1914]. Meet [1445]. MEG [357]. Membership [773, 1441]. Memory [1118, 1831, 1316, 1671, 1284, 1823, 1690, 918, 1317, 926, 1691, 1050, 1903, 912, 875, 58, 57, 1488, 1733, 507, 1696, 1680, 102, 103, 1762, 1321, 656]. Mental [1425]. Mereotopology [1655]. Merge [585]. Mergesort [1288]. Merging [890, 1143, 1366]. Mesh [546, 547, 1619]. Meshes [1270]. Message [1778, 1776, 1718, 1777, 58, 520, 57, 1702]. Message-Passing [58]. Messages [49, 1450]. Messaging [944, 1873]. Meta [574, 554, 1451, 997, 736, 1201, 575]. Meta-Logical [997]. Meta-Representation [1201]. Meta-search [1451]. Metadata [864, 1198]. Metadata-Based [864]. Metaheuristic [263]. MetaML [152]. Metamodel [616]. Metaprogramming [573]. Metaprograms [573]. Metatheory [1004, 1632]. Metathesaurus^{T M} [684]. Method [996, 751, 1330, 804, 282, 924, 275, 547, 442, 381, 1251, 1258, 1609, 787, 409, 272, 1897, 1230, 403, 1398, 408, 391, 802, 862, 835, 1944, 491, 1027, 867, 404, 214, 219, 934]. Methodologies [729]. Methodology [355, 1221, 1310, 1538, 1497, 220]. Methods [1686, 728, 1029, 236, 932, 1787, 1082, 681, 871, 1721, 1676, 766, 812, 1162, 735, 1680, 532, 669, 1969, 1534]. Metric [444]. Metrical [1353]. Metrics [1472, 870]. Metropolis [1608]. Microarchitecture [1049]. Microcalcifications [394, 272]. Microcontrollers [753]. Micrographs [501]. Micronet [274]. Microprocessor [1031, 577, 518]. Microprocessors [519]. Microsoft [794]. Microstructure [768]. Middle [1107]. Middleware [1898, 561, 565]. Migration [546, 1863]. MIMS [414]. Min [1833, 232, 1834]. Min-Max [232]. Min-Wise [1833, 1834]. Mind [353]. Mindstorms [92]. Minimal [585, 1732, 599, 165, 784, 201]. Minimization [1615, 879, 1603, 642]. Minimized [283]. Minimizing [893, 894]. Minimum [1287, 1836, 909, 1277, 254, 886, 1856, 1754, 1811]. Mining [1232, 1223, 661, 1469, 1256, 421, 1447, 1534, 209, 155, 202]. Miss [1107]. Mixing [1332, 102, 103]. Mixins [923]. Mixture [1605]. Miya2 [114]. Mobile [1891, 1981, 1581, 1710, 1909, 1921, 1586, 62, 1331, 1923, 1999, 315, 77, 1467, 1559, 1814, 63, 1562, 1887, 1896, 1995, 1983, 1985, 1980, 109, 306, 283, 1986, 1862, 1532, 1874, 1990, 1560, 105, 1561, 1864, 80, 1991, 1994, 70]. Mobile-Agent [1562]. Mobility [1331, 1698, 1997]. Mod [1108, 233]. mod- [233]. Modal [996, 985, 1626, 580, 578, 447, 987, 1913, 1912, 403, 668, 1496, 404]. Mode [821, 364, 290]. Model [1324, 856, 1767, 1708, 1066, 1032, 585, 312, 1180, 1031, 1044, 1059, 388, 68, 1316, 1048, 1688, 676, 700, 414, 1466, 1067, 1855, 1317, 1558, 1337, 1338, 1672, 277, 269, 1014, 1480, 308, 683, 1674, 29, 40, 352, 1283, 921, 383, 292, 527, 1035, 1020, 293, 627, 1007, 621, 1040, 376, 35, 390, 1240, 1051, 95, 677, 1055, 101, 288, 555, 556, 1457, 1243, 451, 846, 322, 457, 800, 1576, 1611, 1061, 347, 828, 268, 1227, 1885, 398, 413, 401, 102, 103, 39, 796, 1700, 52, 604, 1052, 1619, 393, 852, 214, 1664, 853, 1150, 1663, 44]. Model-Based [676, 1020, 677, 288]. Model-Checking [1035, 1055]. Modeled [632, 1201]. Modeling [314, 1669, 605, 1473, 518, 1249, 349, 630, 275, 716, 748, 632, 834, 624, 1166, 629, 1983, 392, 855, 799, 608, 1186, 332, 768, 779, 265]. Modell [837]. Modelling [1624, 619, 141, 1415, 43, 146, 1483, 1988, 843, 842, 706, 139, 1938, 205, 862, 302, 833, 1622, 867, 790, 140, 838]. Models [1685, 1709, 37, 371, 1606, 1718, 1607, 380, 1221, 714, 678, 1176, 557, 1608, 387, 1605, 495, 147, 365, 1675, 145, 1065, 708, 1971, 1427, 871, 1069, 378, 613, 1649, 877, 1230, 633, 384, 143, 841, 599, 739, 1096, 45, 611, 59, 538, 330, 1052, 46, 167]. Modes [1658, 1113]. Modifications [1465]. Modified [761, 1245, 196]. Modifiers [822]. Modular [1631, 916, 327, 933]. Modulation [767, 1437]. Modules [134]. Modulo [954, 955, 1627, 957]. Molecular [516]. Molecules [1753]. Moment [465]. Monadic [1341]. Monitoring [1231, 1171, 1230, 1552, 1561]. Monocular [453]. Monoids [1358]. Monotonicity [1842]. Monte [1149]. Morphisms [148, 1327]. Morphological [1470, 704]. Morphology [272]. Morphometry [386]. Morphosyntactic [1077]. Mosaics [431]. Motif [1282]. Motion [437, 381, 1927, 1949, 448, 1668, 453, 1639, 1638]. Motion-Based [437]. Motor [815]. Movement [1996, 551]. Movements [1933]. Movies [438]. Moving [1712, 1531, 1455]. Moving-Object [1531]. MPEG [445, 48, 432]. MPEG-2 [48]. MPEG-7 [432]. MPI [1704]. MPMA [182]. MR [665, 666, 405, 392, 407]. MRA [379, 366, 382]. MRI [705, 383, 378, 409, 408, 401]. Mult [1413]. Mult-agent [1413]. Multi [778, 670, 1420, 1298, 309, 312, 1414, 399, 1419, 1263, 342, 1869, 1416, 310, 115, 1810, 1524, 1837, 1913, 756, 1417, 1525, 308, 705, 898, 1912, 1411, 842, 1551, 904, 403, 668, 1294, 1404, 783, 305, 934, 1538, 1870, 1418, 1496, 1567, 447, 1960, 550, 1700, 528, 404, 120, 1351, 205]. Multi-Agent [115, 1551, 305, 670, 1420, 309, 312, 1414, 1419, 1263, 1869, 1416, 310, 1417, 308, 705, 1411, 1870, 1418, 1567]. Multi-color [1837]. Multi-coloring [1298, 898]. Multi-complexes [1524, 1525]. Multi-dimensional [1294, 1404, 205]. Multi-isotope [399]. Multi-level [120, 1351]. Multi-method [934]. Multi-Modal [447, 1913, 1912, 403, 668, 1496, 404]. Multi-objective [756]. Multi-paradigm [1420]. Multi-party [904]. Multi-phase [783, 528]. Multi-processing [550]. Multi-Threaded [1810, 1700]. Multi-tiered [342]. Multi-user [1960]. Multi-valued [778]. Multi-variant [842]. Multi-way [1538]. Multiagent [128]. Multicast [1803]. Multicoloring [1840]. Multicomputers [546]. Multicriteria [788, 213]. Multidatabases [1457]. Multidimensional [419, 1857, 1347, 838, 1234]. Multifeature [323]. Multigraph [1300]. Multilayer [319]. Multilevel [277, 847]. Multimedia [48, 412, 271, 434, 1706, 1980, 507, 1910]. Multimethods [933]. Multimodal [1930, 1953, 1952]. Multipacket [1270]. Multiparadigm [969, 970]. Multiperspective [1201]. Multiple [1474, 1424, 335, 1838, 615, 390, 1243, 934, 418, 333, 724, 368, 1377, 768, 105]. Multiple-Domain [1474]. Multiple-Platform [335]. Multiplex [1457]. Multiplication [1771, 916, 1707, 2, 1383]. Multiply [962]. Multipolar [357]. Multiprocessor [918, 1050, 1740, 548, 1715]. Multiprocessors [1733, 640]. MultiPurpose [823]. Multirobot [87, 88]. Multiscale [484]. Multisensor [109]. Multisensor-Based [109]. Multispectral [781, 1611]. Multithreaded [542]. Multithreading [526]. Music [1471, 1929]. Mutual [403, 71, 404]. My [1468]. Myocardial [381].

NAIST [109]. Names [732]. Nano [790]. Nano-Scale [790]. NANOS [514]. NAS [1704]. NASA [1547]. National [1184]. Natural [703, 1503, 1504, 950, 951, 1462, 1401, 199]. Nature [1768, 1319, 1640]. Navigating [1233]. Navigation [1881, 780, 710, 1639, 1637, 1638, 1154]. Nearest [1472, 876, 1254]. Nearest-Neighbor [1254]. Nearly [1297]. Need [1999]. Needs [1445]. Negation [890, 1150]. Negation-Limited [890]. negative [172]. Negotiating [1422, 1421]. Negotiation [1566]. Neighbor [1472, 876, 1254]. Neighborhood [343]. Neighborhoods [919]. Neonatal [671, 673]. Nested [1211]. Net [1164, 1176, 147, 1171, 1165, 142, 143, 768, 1174]. netlist [648]. Netlists [1032]. Nets [1169, 856, 1327, 141, 1180, 1688, 138, 1170, 1182, 1175, 954, 955, 1340, 1178, 1988, 1173, 1179, 807, 139, 1162, 800, 806, 1172, 1682, 140, 1153]. Network [228, 1581, 1597, 767, 1263, 1878, 822, 1580, 535, 683, 1589, 519, 1868, 1588, 681, 1600, 1085, 536, 1925, 446, 1253, 272, 757, 803, 252, 1574, 1982, 823, 1590, 1268, 1591, 1681, 779, 1575, 1276, 45, 1593, 749, 1585, 76, 941, 22, 1596, 649, 1142]. Network-Aware [45]. Network-Level [1878]. Networked [710]. Networking [1582, 1598, 564, 1579]. Networks [778, 78, 1981, 1273, 818, 759, 1825, 1586, 1594, 138, 55, 1853, 808, 271, 1269, 813, 147, 298, 44, 791, 1917, 814, 1836, 278, 748, 540, 1814, 63, 58, 1139, 766, 807, 1660, 1577, 273, 1512, 1584, 1590, 332, 1592, 1576, 47, 1578, 549, 559, 1585, 758, 1125, 209, 207]. Networks-Based [271]. Neural [778, 818, 759, 767, 1263, 808, 271, 822, 791, 683, 814, 748, 519, 681, 1085, 766, 807, 446, 1253, 272, 273, 757, 780, 781, 823, 1680, 800, 806, 768, 1681, 779, 749, 758, 207]. Neuro [809, 693]. Neuro-psychological [693]. Neuroanatomical [508]. Neuron [826]. Neurons [778]. Neutral [73]. Newborn [673]. News [715]. NIP [52]. NISZK [1793]. NIT [100]. NJR [538]. Noise [394, 1611]. Noisy [689, 212, 197]. Nomadic [1907]. Nominal [592, 591]. Non [1402, 1273, 1797, 1130, 590, 1335, 682, 1950, 784, 876, 11, 385, 374, 578, 1724, 1617, 558, 1336, 328, 401, 386, 1929, 1129, 207, 211]. Non-associative [590]. non-carcinogenic [207]. Non-classical [578]. Non-contiguous [1724]. Non-deterministic [11]. Non-Euclidean [386]. Non-functional [558]. Non-ground [1130, 1129]. Non-interactive [1335, 1336]. Non-linear [682, 374]. Non-malleable [1797]. Non-minimal [784]. Non-obvious [1929]. Non-rigid [1617, 401]. non-stationary [211]. Non-supervised [328]. Non-uniform [1402, 1273, 876, 385].

Non-verbal [1950]. Noninteractive [1793]. Nonlinear [814, 787, 1775, 407]. Nonmonotonic [1122, 1076]. Nonparametric [1238]. Normal [397, 1092, 1095]. Normalisation [393]. Normalization [964, 954, 950, 955, 1627, 951, 408, 952, 953]. Normed [1352]. Notation [605, 630]. Notational [1635]. Nothing [1768, 1796, 1694]. NOTIO [1200]. Notion [1187]. Notions [1797, 1664]. Novel [341, 317, 320]. NP [1335, 1303, 1336]. NP-Hard [1303]. nuanced [1150, 172]. Nuclear [399]. Nugget [749]. Number [1860, 1306, 1118, 1393, 755, 250, 1299, 642]. Numbers [900, 881, 1315, 1760, 1429, 901, 902]. Numerical [1169, 1318, 1676, 37, 221]. NuSMV [1067]. NVT [673].

O [1258, 543, 1683]. OAEP [1796]. OBDD [1001, 11]. OBDD-Based [11]. OBDDs [1360]. Oberon [555, 556]. Oberon-2 [555, 556]. Object [1974, 856, 854, 945, 1970, 1959, 1976, 1964, 1900, 482, 930, 1967, 617, 484, 1971, 471, 1199, 616, 621, 1693, 139, 1966, 1456, 1968, 1963, 618, 384, 1462, 1698, 1531, 646, 1962, 844, 1454, 347, 481, 1969, 323, 1960, 1956, 1246, 1863, 941, 52, 1957, 265, 174, 1664]. Object-Based [471, 481, 1246]. Object-Orientation [1957]. Object-Oriented [1974, 945, 1970, 1959, 1964, 930, 1967, 1199, 1968, 618, 646, 1454, 347, 1969, 1960, 941, 52, 265, 174]. Object-Process [1462]. objective [756]. Objects [940, 1528, 1529, 1999, 922, 53, 831, 921, 1712, 1229, 448, 1633, 931, 1657, 1973, 363, 76, 480, 865, 1455]. Objects’State [567]. Oblivious [1800]. OBOA [268]. Observation [110, 1861]. Observations [1319, 872, 212]. Observed [1258]. Obstacle [654]. Obstacle-avoiding [654]. obvious [1929]. OCEANS [513]. OCR [317]. Odense [1501]. Off [1279]. Off-Line [1279]. Offset [1735]. Offspring [276]. Oja [196]. Oja-RLS [196]. OLAP [1534]. Oligo [738]. Oligo-Agents [738]. OLOG [1456]. omega [1045]. omega-Automata [1045]. OML [631]. Omnidirectional [107, 108]. OMPC [572]. On-Chip [548]. On-Line [1817, 1272, 1289, 793, 1815, 1444, 161]. On-the-Fly [1714]. Once [1806]. One [1746, 958, 464, 1644, 959, 1204]. One-Dimensional [464]. One-Step [958, 959]. Online [1435, 912, 913, 1218, 436, 1762, 1321]. Onthology [1460]. Onthology-Driven [1460]. onto [896]. Ontological [717, 843]. Ontologies [316, 721, 722]. Ontology [685, 712, 1666, 715, 1653, 720, 1657, 684, 157]. Ontology-Driven [715]. OO [1975]. Opaque [104]. Opaque-Transition [104]. Open [670, 1442, 1414, 1573, 350, 1140, 571, 1883]. Open-End [1883]. OpenC [572]. OpenCorba [571]. OpenMP [1704]. Operant [820]. Operating [821, 75, 1957, 71]. Operation [1113, 142]. Operational [741]. Operations [1500, 1686, 348, 1683, 657]. Operator [1080, 33, 195]. Operators [1072, 1673, 1588, 1092, 994]. Opportunities [313, 1982, 1979]. Optical [1825, 147, 1982]. Optima [276]. Optimal [818, 1759, 1731, 1287, 765, 235, 1310, 1083, 249, 1744, 754, 1110, 1856, 283, 258]. optimality [763]. Optimisation [1301, 188]. Optimistic [1792, 1987]. Optimization [1703, 1238, 1728, 764, 522, 756, 277, 752, 1452, 250, 812, 327, 531, 417, 1683, 211, 221]. Optimizations [1706]. Optimized [1777, 1142, 1997]. Optimizing [1735, 791, 1775, 253, 1052]. Optimum [792]. Oracle8i [1544]. Order [1022, 985, 977, 1341, 1079, 1343, 595, 915, 1007, 1023, 1610, 966, 1012, 989, 601, 161, 180, 967]. Ordered [498, 1342]. Ordering [1528, 1360, 947, 1529]. Orderings [961]. Ordinal [1235]. Organisations [1595]. organization [645]. Organizational [1445, 859]. Organizations [659]. Organize [430]. Organizing [277, 680, 476, 652]. Orientation [240, 377, 107, 375, 1957]. Orientations [241, 1729]. Oriented [1974, 314, 431, 1412, 1972, 945, 1970, 1959, 574, 1964, 930, 1967, 993, 262, 1410, 1199, 1562, 690, 1693, 378, 830, 1968, 1193, 1963, 618, 290, 646, 1962, 1454, 96, 347, 1969, 1497, 421, 1965, 1960, 839, 575, 941, 52, 265, 174]. Originating [679]. Ornamentals [488]. ORTES [187]. Orthogonal [249, 952, 1734, 953]. Osaka [136, 108]. Osteoarthritic [391]. Other [681, 787, 1375]. OTHY [854]. OTM [844]. Otoneurological [681]. Outdoor [321]. Outliner [1911]. Outlines [481]. Output [1249, 330]. Outsourcing [859]. outward [1145]. Overlapping [1713]. Overlay [1540, 1585]. Overload [507]. Overview [81, 1687, 6, 39]. Oz [14]. Oz-Centric [14].

P [1340, 1167]. P/T [1340, 1167].

Pacemaker [677]. Package [1711]. Packet [1853, 1576]. Packet-Switched [1576]. Packing [1857, 250, 260, 652]. Packing-Covering [250]. Packings [1304, 1307]. Padding [1763]. Pages [289, 179]. Paging [912]. Pain [667]. Pairs [1395]. PalmORB [1884]. PalmPilot [1428]. Panel [1440, 1976]. Paraconsistent [1122]. Paradigm [329, 80, 1420, 174]. Paradigms [1717, 1320]. Paradox [1332]. Parallel [533, 314, 1169, 1669, 1717, 1670, 1685, 1708, 1687, 1718, 1751, 917, 1672, 1278, 44, 1674, 1675, 539, 520, 1713, 1720, 1714, 1721, 1676, 33, 1694, 531, 1968, 1696, 274, 544, 1856, 119, 1716, 532, 1699, 549, 538, 2, 1383, 530, 524, 1683, 52, 604, 528]. Parallel/High [1968]. Parallel/High-Performance [1968]. Parallelism [1810, 1693, 526].

Parallelization [1717, 1704, 522, 541, 1711, 527, 545, 1697]. Parallelizing [1673, 523]. Parameter [812, 327, 1938, 142, 1756, 790]. Parameterized [1737, 1028, 929]. Parametric [1324, 1034, 706, 1217, 1682]. Parametrized [1037]. Paramodulation [1022]. Pareto [763]. Pareto-optimality [763]. Paris [91]. Parity [240, 241]. Parsers [1083]. Parsing [443]. Part [1082, 1087, 957]. Part-of-Speech [1082, 1087]. Partial [1048, 594, 1348, 1840, 1432, 219, 167]. Partially [1342]. Parties [1444, 80]. Partition [1855]. Partitioned [104]. Partitioning [885, 1855, 279, 648]. Partitions [1724, 1647, 643]. Partner [1424]. Parts [227]. Party [1768, 1770, 904]. Pascal [1720]. PaSo [117]. PaSo-Team’98 [117]. Pass [114]. Passing [1718, 58, 520, 57, 1702]. Past [568]. Path [585, 1452, 875, 274, 947, 283, 961, 1596]. Paths [1835, 1808, 1363, 884, 650]. Patient [1502, 696, 668]. Pattern [771, 761, 614, 497, 1385, 1240, 1404, 1386, 1387, 949]. Patterns [926, 215]. PC [533]. PCPs [1857]. PCs [1704, 57]. PD [797]. PD/PID [797]. PDAs [1884]. PDE [545]. Peak [1686]. Pedagogical [10]. Peer [1993]. Peirce [1218]. Peircean [1195, 1189]. Pen [481]. Pen-Based [481]. Pendulum [796]. People [1917]. PEPA [36]. Perceptive [21]. Perceptron [319]. Perceptrons [801]. Perceptual [1879, 195]. Percolation [1708]. Perfect [1072, 1732]. Performance [1032, 1737, 1828, 48, 1704, 1582, 36, 271, 1690, 918, 535, 1827, 144, 1065, 1003, 539, 1984, 1694, 25, 1848, 1968, 529, 117, 1541, 1538, 50, 572, 1701, 209]. Performer [1929]. Performing [1540]. Perimeter [1758]. Periodic [1230]. Periodicity [1364, 1365]. Periods [1394]. Peripheral [382]. Permanent [1845]. Permutation [950, 951, 1112, 1834].

Permutation-Conversions [950, 951]. Permutations [1833]. Persian [760]. Persistence [557]. Persistent [931, 865]. Person [1937]. Personalise [701]. Personalised [1516]. Personalities [13]. Personalized [1468]. Perspective [1413, 1636, 1547]. PERT [344]. Perturbations [1819]. Pesticide [1708]. PET [385, 364]. Petaflop [521, 46]. Petaflops [1705]. Petri [856, 1164, 1327, 141, 1688, 138, 1182, 1175, 147, 1988, 1173, 1171, 1179, 1165, 139, 142, 1162, 143, 1153, 1174, 1682, 140]. Pfaffian [1858]. Pfaffians [883]. Phase [184, 1730, 350, 784, 806, 783, 528]. Phase-Space-Flow [806]. PhD [1962]. Phenomena [1305]. Phone [298, 1980]. Phonemes [1945]. Photographs [502]. Photography [1888]. Phylogenetic [1826]. Physical [1389, 819, 796]. Physicians [663]. Physics [473]. Physics-Based [473]. Physiologically [378]. Pictorial [489, 417, 1641]. Picture [1636, 487]. Pictures [1742]. PID [797]. Pieces [353, 1644]. Piecewise [238]. Piecewise-Constant [238]. PIL [1068]. PIL/SETHEO [1068]. Pile [1491]. Pipeline [1673, 544]. Pipelined [1056, 1033]. Pisot [979]. Pitfalls [606]. Pittsburgh [281]. Pixel [1613]. Pixel-Based [1613]. Place [9, 1562]. Placement [1727]. Plan [1423, 1868]. Planar [1818, 1840, 879, 253]. Plane [1817, 1291, 910, 1751, 654]. Planes [1584, 653]. Planning [6, 11, 690, 1495, 1274, 1497, 649]. Plans [690]. Plant [1170, 1258]. Planted [1855]. Plants [759, 817]. Plate [375]. Platform [1581, 335, 1919]. PLATINUM [1727]. Plausible [1122]. Play [114, 97, 133]. Players [114]. Playing [1436]. Plays [97]. Pluecker [643]. PML [1866]. PMLP [1687]. POBox [1897]. Pocket [1876]. PoG [1940]. Point [1402, 1901, 369, 1296, 339, 402, 644, 1750]. Point-Based [402]. point-in-polygon [644]. Pointers [921]. Pointing [1940]. Points [462, 239, 1256]. Policies [141, 64, 1848, 558]. Policy [73, 860, 861, 1494, 1578]. Policy-Based [860, 861]. Policy-Neutral [73]. Pollen [1917]. Polling [1350, 144]. Polling-Based [144]. Polygon [1534, 644]. Polygonal [1735, 1540]. polygons [644]. Polyhedra [248, 260]. Polyhedral [1304]. Polymorphism [1633]. Polynomial [590, 246, 1838, 891, 1370]. Polynomials [809, 980, 981]. Polytope [239]. Polytopes [1728, 237]. Pool [382]. Population [1675, 693]. Populations [1742, 363]. PORT [1218]. Portable [572]. Portal [1914]. Porting [610]. Portrait [489]. Position [1752, 815]. Position-Independent [1752]. Positional [1391]. Positioning [110, 1861, 1906]. Positive [1063, 1088]. Positron [1621]. Possibilistic [789, 1134]. Possibilities [61]. Poster [1977]. Posture [1937, 1935]. Postures [1932]. Potential [509]. Power [280, 1789, 1853, 1207, 817, 40, 1354, 1928, 1358, 1788, 1376, 295, 1227]. Power-Analysis [1789]. PowerPC [1031]. Practical [1438, 685, 636, 1273, 1029, 850, 1430, 529, 1386, 1110, 851, 644, 640]. Practice [663]. Practices [614]. Pre [505, 1795, 1794, 680, 1146, 1149]. pre-computation [1149]. Pre-filter [505]. Pre-processing [1795, 1794, 680, 1146]. Precedence [1677]. Precise [628]. Predicate [1039, 730, 33]. predication [1648]. Predicting [1078]. Prediction [1084, 1253, 1091, 529, 551, 1260, 500, 1701, 1155]. Predictions [1882]. Predictors [517]. Preference [1473]. Prefix [1041, 914, 638]. Prefix-Rewriting [914]. Prefixed [996]. Prefixes [1396, 1397]. Preliminary [1687, 1356]. Premature [276]. Prenex [1092]. Preorder [1361].

Prescribing [672]. Presence [1635]. Present [568]. Presentation [1992, 456]. Presentations [507]. Presenting [993]. Preserving [1182, 1992]. Pressure [824]. Preterm [219]. Preventing [276]. Preventive [817]. Preventive-Restoration [817]. Preview [1636]. Pricing [1576]. Primitive [119]. Primitives [1425, 363]. Principal [318]. Principles [1439]. Priorities [17]. Prioritized [1124, 162]. Priority [1831, 1824, 95, 1848, 656]. Priority/Confidence [95]. Private [1883]. Probabilistic [30, 26, 1847, 220, 29, 1350, 1233, 24, 1069, 1147, 25, 905, 31, 344, 39, 1845]. Probabilities [887, 1047]. Probability [1472, 28]. Probes [1389, 1871]. Problem [228, 1311, 229, 1741, 963, 311, 1389, 1291, 711, 1730, 728, 911, 1854, 1175, 1853, 236, 1337, 1338, 244, 912, 1241, 1242, 246, 1851, 1812, 1838, 1358, 1359, 1297, 1360, 296, 811, 1203, 906, 252, 1204, 545, 905, 1765, 326, 723, 735, 1557, 1268, 279, 1756, 872, 259, 878, 184, 170]. Problem-Solving [311, 711, 735]. Problems [1703, 230, 231, 340, 1857, 1755, 338, 234, 899, 765, 882, 880, 1303, 1805, 763, 580, 897, 915, 11, 247, 1239, 1744, 1802, 273, 1277, 1715, 990, 1174, 1276, 344, 1380, 155, 186, 223]. ProbVerus [29]. Procedure [588]. Procedures [1044, 577, 986, 242, 992]. Process [30, 837, 836, 1180, 37, 1344, 944, 541, 843, 352, 38, 842, 1352, 827, 293, 627, 1692, 621, 1361, 692, 1727, 1041, 830, 1679, 625, 1462, 802, 841, 302, 828, 532, 663, 330, 1663]. Process-Oriented [830]. Processes [7, 784, 1589, 834, 293, 1165, 33, 1493, 1185, 1973, 806, 368, 1618, 868]. Processing [778, 1583, 893, 752, 1507, 1231, 1674, 100, 1984, 1694, 1985, 1541, 1194, 702, 1699, 417, 894, 500, 704, 1795, 1794, 164, 680, 1146, 550]. Processor [1032, 1030, 1033]. Processors [1686]. Procrustes [387]. Produce [275]. Product [809, 590, 1966]. Product-Line [1966]. Production [343, 1643]. Products [862]. Profile [792, 517]. Profile-Based [517]. Profiles [1887]. prognosis [191]. Progol [1077]. Program [1034, 150, 506, 308, 541, 1711, 673, 35, 25]. Programmable [1573, 1575, 1578]. Programmed [69]. Programming [60, 1975, 1670, 149, 1972, 945, 1970, 1689, 1583, 1314, 235, 236, 984, 574, 1867, 1244, 354, 801, 1080, 298, 1691, 1081, 1868, 1720, 1147, 906, 1090, 1021, 566, 1699, 152, 1276, 549, 1965, 20, 575, 941, 170, 162]. Programming-Language [60]. Programs [1326, 1275, 1704, 1852, 1628, 1673, 242, 1961, 1714, 1088, 942, 1095, 938, 524, 160, 173]. Progress [41]. Project [1685, 1687, 1949, 514, 512, 1277, 16, 513, 1468, 515]. Projective [584]. Promotion [700]. Pronoun [598]. Proof [1313, 954, 955, 1627, 1628, 583, 1030, 1281, 42, 1023, 1367, 325]. Proofs [1183, 1054, 993, 301, 1379]. Propagating [924]. Propagation [1454, 1681, 1145, 1148]. Propagator [1381]. Proper [1309, 142]. Properties [229, 26, 1031, 1044, 1796, 1833, 1339, 1343, 718, 1040, 65, 267, 1095, 1163, 166]. Property [1440, 557, 1358, 1150]. Proportionate [1350]. Proposal [628]. Proposals [631]. Proposed [1184]. Proposition [1942]. Propositional [586]. Prospects [360, 361]. Prostate [699, 695]. Prosthesis [1903, 708]. Protean [1575]. Protection [60, 71]. protections [651]. Protein [1753, 1332, 1283]. Protocol [1981, 141, 1034, 1417, 1429, 1995, 139, 829, 142, 51, 1863, 1566, 140]. Protocols [56, 1036, 1313, 582, 664, 1068, 1012, 1563]. Prototype [864, 1582, 1243]. Prototypes [669]. Prototyping [943, 46]. Provability [985]. Provably [1273]. Prove [1061]. Prover [603, 1016, 42, 596, 1018]. Provers [992, 1021]. Provided [47]. Providing [73, 931, 942]. Proving [1006, 5, 602, 1000, 593, 1002, 1055]. Provisioning [1422]. Proximity [1535, 1536, 179]. Proxy [1998]. Prune [1222]. Pruning [1225, 1226]. PSA [1675]. Pseudo [1326, 1230]. Pseudo-Periodic [1230]. Pseudo-Random [1326]. Pseudoknots [1407]. Pseudorandom [1118, 1779]. PSpace [987]. psychological [693]. PTE [1096]. PTE-3 [1096]. Public [1785, 1266, 1764, 1434, 1815]. Public-Key [1266, 1434]. Publicly [1772]. Pulse [767]. Pulse-Width [767]. Pump [1255]. Purchaser [1268]. Pure [1205]. Purpose [795].

QoS [1923, 1992]. Quadratic

[801, 979, 906, 1276]. Quadruped [91, 134]. Quadtrees [1539]. Qualitative [26, 1667, 689, 288, 1668, 1151, 1134, 1648, 1652]. Quality [1743, 1496, 862, 1517, 839, 214]. Quality-Oriented [839]. Quantification [588, 163]. Quantified [1027]. Quantifier [958]. Quantitative [1974, 25]. Quantization [800]. Quantization-Model [800]. Quantum [887]. Quartet [1292]. Quasi [1172, 1133]. quasi-Bayesian [1133]. Quasi-Static [1172]. Quaternion [1781]. Queries [419, 863, 1342, 1452, 875, 1449, 424, 1800, 1519, 481, 189]. Query [419, 1124, 425, 482, 927, 1451, 417, 459, 188]. Query-Answering [1124]. Query-Based [927]. Querying [1448, 412, 794, 166, 177]. Questionnaire [819]. Queue [1824]. Queues [1831, 656]. Queuing [1700]. Quotients [1634].

R [794]. Race [1714]. Rackoff [1110]. Radial [107, 406]. Radio [1853, 1269]. Radiology [665, 666]. Radiosity [530]. RAID [543]. Railroad [1815]. Rainbow [147]. RAMSES [1891]. Random [1305, 1238, 1326, 1610, 1308, 276, 1622, 184]. Randomization [395]. Randomized [1852, 1856]. Randomness [1335, 1336]. Range [1853]. Rank [882, 237, 1748, 243]. Rank-Based [1748]. Ranked [333]. Ranking [1810, 1289]. Rapid [972]. Rapidly [1332, 1908]. RASCAL [560]. Rate [745]. Rates [1686, 1614]. Ratio [1736]. Rational [1370]. Rationals [1295]. ray [379]. RC5 [1099]. RC5P [1108]. RC6 [1099, 1098]. Rcane [1574]. Re [1964, 1435]. Re-engineering [1964]. Re-issuance [1435]. Reachability [1038, 1053, 1175, 1333, 1161]. Reaction [1948]. Reactiva’Motion [1949]. Reactive [1949, 106]. Real [1034, 1726, 1415, 1478, 1315, 1318, 126, 40, 42, 28, 629, 35, 33, 297, 757, 288, 806, 623, 1699, 359, 552, 363, 108, 550, 901, 1432, 902, 187, 223, 1640]. real-life [223]. Real-Time [1034, 1726, 1415, 1478, 40, 42, 28, 629, 35, 33, 757, 806, 623, 1699, 552, 550, 1432, 187]. Realistic [1619]. Reality [1895, 1881, 1906]. Realization [1414, 1719]. Really [1281, 159]. Rearrangement [1390, 877]. Reasoner [1511]. Reasoning [303, 670, 1121, 1122, 1501, 304, 355, 1472, 1212, 1474, 804, 1466, 261, 1424, 1249, 1506, 1480, 1479, 789, 1179, 1434, 744, 1652, 287, 582, 1512, 191, 1651, 668, 334, 1496, 1514, 1515, 669, 660, 1518, 868, 1659, 189, 1146, 1650, 205]. Receiving [93, 94]. Recognition [778, 498, 1937, 761, 441, 1943, 1939, 497, 681, 1085, 1662, 1935, 1938, 317, 1932, 544, 318, 1944, 439, 323, 97, 1945, 1936, 193, 1153]. Recognition-triggered [1662]. Recognizing [1309]. Recombination [765]. Recommendation [1516]. Reconciliation [1556]. Reconfigurability [1910]. Reconfiguration [55, 744]. Reconstruction [1826, 1621, 398]. Recording [703]. Recordings [1231]. Recovery [365, 384, 1093]. Rectangular [328]. Rectilinear [1755, 1739]. Recurrence [695]. Recurrent [748]. Recursive [1238, 1507, 1173, 1707, 962].

Red [919, 1827]. Red-Black [1827]. Red-Blue [919]. Redescribing [1228]. Reduce [543, 1254]. Reduced [1103]. Reducing [1996]. Reduction [226, 1305, 1491, 531, 802, 841]. reducts [201]. Redundancy [1371, 185]. Redundant [1849]. Reengineering [717, 709]. Reference [926, 1668]. Refinement [1072, 1080, 547, 1043, 693, 1092, 1248]. Refining [1075]. Reflecting [569]. Reflection [563, 1006, 554, 292, 567, 562, 564, 555, 556, 572, 552]. Reflections [617]. Reflective [61, 561, 557, 58, 570, 571, 566, 560, 559, 565, 553]. Refrigeration [748]. Regaining [945]. Region [472, 1539, 442, 422, 424, 142, 321]. Region-Based [472, 422]. Regional [284]. Regions [1858, 470, 1645]. Registration [369, 405, 374, 403, 451, 375, 370, 402, 404]. Regression [273, 1152]. Regular [466, 1342, 1452, 965, 1821, 1123, 184, 914]. regularities [180]. Reification [857]. Reinforcement [104]. Reinforces [110, 1861]. Related [1278, 720]. Relating [1623]. Relation [819]. Relational [1207, 845, 1069, 1469]. Relationally [1208]. Relations [1122, 1216, 1364, 1365, 1647]. Relationship [1793, 857]. Relationships [1535, 635, 1944, 1454, 1536, 212]. Relaxation [1618]. Relaxations [1276]. Release [255]. Relevance [423, 425, 477]. Relevancy [1014]. Relevant [750, 802, 203]. Reliability [697]. Reliable [87, 88, 742, 1863]. Relinearization [1764]. REMEX [1487, 1486]. Remotely [766, 1106]. Reorder [1030]. RePART [761]. Repeated [1389]. Replication [1993, 1885]. Report [622, 610]. Repositories [1460]. Repository [1552]. Represent [1191, 447]. Representation [4, 431, 340, 7, 1047, 337, 1612, 464, 1265, 1136, 1659, 484, 294, 1949, 339, 879, 1492, 1668, 461, 1194, 1217, 1201, 264, 868, 1603, 1667, 199]. Representational [268]. Representations [1761, 1228, 1531, 1648, 644]. representative [216]. Represented [1210]. Representing [174, 1745, 189, 1508, 1309, 806, 686]. Reprogramming [677]. Required [701]. Requirements [1546, 352, 719, 1411, 1839]. Rescue [1500]. Research [510, 128, 1954, 1520, 155]. Reset [1340]. Resilient [1775]. Resizable [1731]. Resolution [996, 999, 598, 1095, 204, 1022]. Resonance [1621, 367]. Resoponse [894]. Resource [561, 1550, 1202, 860, 861, 99, 1218, 1574, 531, 1277, 1885]. Resource-Adapting [99]. Resource-Aware [561]. Resource-Constrained [1277]. Resources [1801, 1510]. Response [893, 378, 1866, 1662]. Responsibilities [738]. Restoration [817, 1609, 551, 1274, 649]. Restricted [588, 1707, 1729, 638]. Restrictions [1058]. Restructuring [1714]. Result [1085, 959]. Results [1328, 1687, 1853, 126, 1349, 1747, 578, 382, 994]. Retail [1566]. Retention [1494]. Retrieval [452, 284, 431, 440, 472, 805, 478, 414, 466, 1263, 423, 425, 474, 464, 465, 715, 475, 437, 1265, 1507, 484, 490, 471, 476, 410, 424, 1492, 479, 1494, 433, 432, 1499, 477, 470, 487, 428, 480, 492, 458, 165, 167, 177]. Retrieval-Oriented [431]. Retrieval/Adaptation [1507]. Retrieve [429]. Retrospective [1226]. Reusability [1514]. Reuse [837, 728, 1976, 1463, 1507, 717, 1489, 634, 633]. Reusing [1478, 1508, 1487, 1486]. Reversals [877]. Reverse [1396, 1397]. Review [14, 1936]. Reviewing [97]. Revised [1100]. Revision [1132]. Revisited [995, 259]. Reward [1169]. Rewrite [964, 979, 974, 1011]. Rewrite-Based [1011]. Rewriting [958, 975, 980, 981, 959, 973, 952, 978, 914, 966, 953, 967, 972]. Rhythm [1951]. Rich [1052]. right [653]. Rights [67, 79]. Rigid [1333, 402, 1617, 401]. Rigid-Body [402]. Ring [1403]. Rings [1854]. RISC [1686]. Rise [605]. Risk [729, 801, 832, 1432, 219]. RLS [196]. RNA [1408, 1405, 1409, 1406]. Road [1614, 780]. RoboCup [132, 110, 1861, 81, 121, 90, 1869, 98, 91, 113, 99, 127, 101, 83, 109, 1870, 131, 108, 97, 102, 89, 84, 124]. RoboCup-98 [81, 113]. RoboCup-NAIST [109]. RoboCup-Team [132, 131]. RoboCup’98 [115, 130, 118]. RoboCuppers [127]. RoboRoos [124]. Robot [1501, 1937, 771, 1738, 1228, 91, 134, 92, 136, 109, 129, 118, 123, 347, 1154, 108, 1933, 105, 86, 135, 120, 1351]. Robotic [87, 88, 797]. Robotics [773, 772]. Robots [122, 15, 283, 18]. Robust [1771, 369, 1445, 504, 1293, 785, 448, 321]. Robustness [1833]. Rocco [84]. Role [66, 1889, 1653, 1927, 625, 1564]. Roll [792]. ROLLING [118, 792]. Room [1201]. Root [27]. Rooted [1803]. Rooted-Tree [1803]. ROSE [171]. ROSES [610]. Rotation [398]. Rough [1666, 1138, 803, 1260, 171, 202, 192, 193, 220, 213]. Round [1681]. Rounding [1852, 242]. Rounds [1101]. Route [335, 1638, 1639, 1643, 1642]. Routed [546]. Router [48, 54]. Routers [1573]. Routes [1641]. Routing [1835, 1981, 1825, 908, 55, 1854, 1345, 791, 1270, 54, 1925, 1595, 1684]. Routings [1858]. Row [934]. RPC [50]. RPO [960]. RS [220]. RSA [1763, 1770]. RT [1688]. RT-MEC [1688]. RUE [1022]. RUE-Resolution [1022]. Rule [331, 1502, 1073, 1074, 291, 1086, 1991, 188, 190]. Rule-Based [1502, 291]. Rules [1211, 798, 697, 612, 1458, 278, 1084, 1085, 1223, 1087, 747, 216, 175]. Run [937, 1806, 468, 1865]. Run-Time [937, 1865]. Running [1710]. Runtime [526].

Safe [63, 937]. Safety [1031, 707, 1040].

Salesman [236, 296]. Sample [1254]. Sample-Data [1254]. Sampling [1858, 1844]. Sanctuary [70]. SAPIA [322]. SAT [184, 185, 580, 242]. Satellite [622, 1609, 781, 322]. Satisfaction [868]. Satisfiability [958, 756]. Scalability [1945]. Scalable [1414, 49, 53, 1877, 1576, 1997]. Scale [340, 1526, 1527, 497, 377, 408, 779, 790, 1637]. Scaled [1735, 1208]. Scales [1043]. Scan [1934]. Scenario [835]. Scene [440, 446, 439]. Scenes [321]. Scent [1928]. SCG [1205, 1204]. SCG-1 [1205, 1204]. Scheduled [51]. Scheduling [1860, 1859, 1741, 1478, 1278, 763, 1850, 1740, 1851, 324, 1678, 1677, 1277, 1848, 255, 1172, 1276, 896, 344, 343]. Schema [1081, 1661]. Schemas [869]. Scheme [1791, 1785, 1851, 1838, 1297, 1222, 1772, 492, 264, 569]. Schemes [1790, 1783, 1278, 1798, 1803, 477, 1782]. Schwarz [464]. Science [1319, 1322]. Sciences [1503, 1504]. Scientific [518, 1460, 1968]. Sclerosis [368]. Scope [1416]. Scramble [1105]. Screening [1259, 693, 699]. Seal [1864]. Search [1500, 1537, 284, 1015, 1829, 1470, 282, 1287, 1478, 1614, 1549, 482, 754, 483, 1023, 1089, 1821, 433, 283, 418, 1224, 1560, 344, 655, 1123, 185, 224, 1451, 173, 180]. Searches [436, 1254]. Searching [1752, 1739, 1295, 1757, 1020, 1822, 1285, 1294, 1553, 169]. Second [1341, 915, 1023]. Second-Order [1341, 915, 1023]. Secondary [1823, 1405, 1406]. Secret [1772]. Sectoring [811]. Secure [1767, 1768, 1791, 1776, 1266, 1798, 847, 1803, 76, 1864, 1566]. Security [78, 1779, 66, 1796, 1769, 1763, 1416, 1113, 1600, 65, 1102, 582, 1116, 1012]. Seeing [1927, 1229]. Seek [1436]. Segmentation [445, 440, 372, 1608, 504, 705, 422, 1084, 483, 366, 424, 473, 760, 317, 781, 391, 1513, 367, 368, 1611, 321, 198]. Segmented [379]. Segmenting [438]. Segments [1751].

Selecting [1805]. Selection [277, 1423, 336, 444, 203, 792, 517, 1523, 1226, 188]. Selective [1993, 219]. Self [277, 680, 87, 88, 296, 476, 1905, 1579, 652]. Self-Configuring [1579]. Self-Crossover [296]. Self-Localization [87, 88]. Self-Organizing [680, 476]. Semantic [1192, 471, 1083, 1660, 157, 1188, 722, 1567, 299, 188, 179, 1645]. Semantically [1448]. Semantics [426, 1958, 1215, 613, 33, 287, 420, 628, 1973, 724, 702, 1454, 182]. Semi [501, 940, 37, 863, 1607]. Semi-interactive [501]. Semi-iterative [1607]. Semi-numerical [37]. Semi-Structured [940, 863]. Semiconductor [532]. Semidefinite [236, 1852, 242]. Sense [1874]. Sensed [766]. Sensible [311]. Sensitive [1889, 975, 974, 333]. Sensitivity [900, 497]. Sensor [87, 88]. Sensorimotor [128]. Separability [1790]. Separate [1073, 1074]. Separate-and-Conquer [1073, 1074]. Separation [233, 919]. Sequence [453, 723, 623]. Sequences [1620, 445, 1748, 384, 1377, 447, 1405, 1403, 1406]. Sequencing [1391]. Sequent [591, 595]. Sequential [1059, 1671, 1673, 1050, 1258, 362]. Sequents [601]. Serial [145, 405]. Series [687, 696, 688, 1256, 1376, 1224, 1646]. Seriously [1376]. Server [1272, 557, 715, 667, 1987, 297, 846, 551, 175, 1291]. Service [1422, 1998, 944, 1988, 620, 1697, 1698, 1552, 849, 175]. Services [1902, 1877, 1598, 1574, 69, 1994]. SeSKA [734]. Session [1977, 1440, 1976]. Set [1623, 588, 1296, 1846, 897, 920, 803, 1839, 1161, 1377, 1301, 1728, 176, 171, 202, 1429]. SETHEO [1068]. Sets [1835, 1326, 1847, 1540, 232, 821, 965, 1260, 1834, 1163, 192, 265, 165, 193]. Setup [1596]. Several [183]. Seymour [1306]. Shading [1602]. Shafer [1145, 1149]. Shape [1537, 387, 504, 389, 365, 1928, 708, 485, 483, 1602, 1603]. Shape-Extraction [504].

Shape-from-Shading [1602]. Shapes [484, 479, 772, 1659]. Shaping [1509]. Shapley [259]. Shared [918, 1050, 53, 57, 1692, 1733, 1694, 738]. Shared-Memory [918, 1050, 1733]. Shared-Nothing [1694]. Sharing [963, 316, 543, 1925, 1772, 1532]. Shelf [912]. Shellsort [1356, 1]. Shift [1385, 359]. Shift-And [1385]. Shipyard [1501]. Shooter [911]. Shop [270, 1851]. Shopbot [1141]. Short [1311, 1581, 102, 103]. Short-Term [1581]. Shortest [875, 1808, 1363, 884, 1380, 650]. Shot [444]. Shunting [344]. SICC [1545]. Side [1283]. Sidon [1847]. Siegenthaler [1775]. Sigmoid [680]. Sign [1943, 1946, 1947, 1944, 1942]. Signal [275, 752, 407, 192]. Signals [783]. Signature [1791, 1790, 107, 393]. Signatures [1442, 1781, 1446]. Signed [877]. Signing [1792]. Signs [1945]. Silicon [501]. Similar [479, 735]. Similarity [1537, 405, 1241, 1242, 1408, 744, 485, 418, 1557, 1409, 200]. Simple [1211, 426, 1829, 1271, 594, 595, 1430, 1361, 1649, 1294, 1772]. Simplices [256, 257]. Simplicial [1524, 1525]. Simplification [1729]. Simplified [1098]. Simulate [820]. Simulating [527, 332]. Simulation [1211, 1708, 1725, 141, 90, 99, 1361, 1983, 93, 94, 118, 532, 1160]. Simulations [1292, 1165, 528]. Simulator [1869, 115, 1690, 1870, 85, 537]. Simultaneous [1333]. Simultaneously [1996]. Single [543, 251, 884, 1382]. Single-Source [251]. SIROCCO [1492]. Sisyphus [1203, 734, 1201]. Sisyphus-I [1203, 1201]. Site [1750, 1911]. Situated [1914]. Situation [301, 332]. situational [1667]. Situations [1426]. Six [1101]. Size [1604, 916, 1707, 891, 124]. Sizes [227, 749]. Skeletal [690]. Skeleton [1699, 1754]. Skeleton-Based [1699]. Skeletons [1670, 389]. Sketch [425]. Sketches [486].

Skill [93, 94]. Skills [87, 88]. SKInT [1630]. Skipjack [1773]. SKIPPER [1699]. Skolem [1001]. Slide [1115]. Sliding [269]. Slotting [1854]. Slovene [1077]. Slow [398]. SM [1823]. Small [1118, 1326, 1847, 121, 1105, 1062, 702, 86, 124]. Small-Robot [86]. Small-Size [124]. Smallest [1293, 873]. Smart [1901, 1433, 313]. SMiLE [57]. Smithsonian [562]. Smoking [701]. Smooth [1992]. Smoothing [407]. SMP [1704]. SNAKE [1117]. SOAP [1973]. SOBER [1119]. Soccer [82, 112, 122, 90, 87, 88, 100, 95, 92, 93, 94, 83, 118, 106, 84, 120, 1351]. Social [670, 1424, 1548, 1888]. Socio [833]. Socio-technical [833]. Sociological [1450]. Sockets [50]. Soft [825, 384, 793]. Soft-Computing-Systems [793]. Software [1974, 837, 836, 1725, 713, 1195, 729, 353, 1706, 742, 842, 262, 1410, 632, 621, 634, 830, 829, 1172, 1487, 1486, 214, 213]. Soil [1708]. Solid [349]. Solos [1362]. Solution [226, 1205, 309, 37, 1337, 1338, 145, 1713, 812]. Solutions [1478, 1303, 1740]. Solvability [988]. Solve [1202, 1204]. Solved [947, 1174]. Solving [1703, 228, 311, 711, 728, 338, 926, 1241, 1242, 11, 1277, 326, 723, 735, 1715, 990, 1557, 1716, 223, 279]. Some [229, 1442, 1328, 1098, 1303, 892, 606, 681, 1807, 1071, 1095, 994, 872, 1447]. Sonography [396]. Sony [135]. Sorted [1080, 1396, 1397]. Sorting [148, 917, 1820]. Sound [1789]. Source [251, 357, 884]. Sources [685, 864, 412, 716, 333]. Space [419, 1767, 1528, 1529, 1731, 1612, 1339, 350, 1296, 1655, 543, 1935, 1654, 1089, 391, 806, 560, 1702, 386, 1874, 194, 643, 198]. Space-Filling [1528, 1529]. Space-Time [1339]. Spaces [1048, 429, 1918, 812, 1217, 1640]. Spanners [1747]. Spanning [239, 873, 1856, 1811]. Sparrows [128]. Sparse [1730, 1761, 1713, 532]. Spass [1017]. Spatial [1232, 1537, 450, 1540, 1894, 1528, 1529, 1187, 1546, 395, 1526, 1527, 1229, 1928, 449, 377, 807, 1535, 1651, 1541, 1538, 1544, 1944, 1543, 508, 1532, 1536, 1534, 1667, 1648, 1664, 1662, 1650, 1645]. Spatio [1453, 1520]. Spatio-Temporal [1453, 1520]. Spatiotemporal [1533]. Spawned [1594]. SPD [1713]. Spearmint [836]. Special [1975, 1444]. Specialization [938]. Specialized [1560]. Species [1756]. Specific [685, 56, 573, 346, 627]. Specification [943, 1164, 1170, 944, 719, 1411, 1065, 1946, 1947, 305, 618, 1025, 1578, 186]. Specifications [64, 1559, 721]. Specified [419, 873]. Specify [79]. Specifying [856, 36, 11]. SPECjvm98 [925]. SPECT [381, 376, 398]. Spectral [399]. Speculative [1673, 1986]. Speech [703, 1082, 1953, 1087, 1952, 1951, 193]. Speed [1118, 1496]. Speed-Up [1496]. Speeding [284, 907]. Speedup [1690]. Sperm [502]. Spiculated [393]. SPJ [840]. Spline [383]. Splines [375]. Split [226, 1744]. Splitting [247]. SPN [145]. SPNs [144]. Sports [1760, 494]. Spot [749]. Spots [1402]. Spotting [750]. Spreadsheet [449]. Square [900, 1777, 244]. Square-Free [900, 244]. SSD [1521, 1522]. ST [51]. Staalmarck [1027]. Stabbing [910]. Stability [787, 786, 869]. Stabilizer [280]. Stable [1355, 259, 797]. Stage [680]. Staged [152]. Staging [675]. Standard [605, 1163, 698, 1184]. Standpoint [797]. Star [1358, 1359]. State [1038, 1325, 1048, 696, 1341, 1319, 1051, 1153, 1060]. State/Event [1051]. StateChart [615]. Stateless [1779]. Static [141, 1755, 1371, 1005, 1172, 530]. Statically [933]. Station [344]. stationary [211]. Statistical [1250, 388, 1793, 389, 1938, 1220, 188]. Statistics [1282, 190]. Status [1435]. Steady [689]. Steel [1501, 792]. Steels [768]. Steiner [1392, 260, 654]. Step [958, 591, 959, 1651, 599, 1644, 1650]. Stepwise [692, 1051]. Stereotypes [1450]. Still [1437]. Stimulating [1954]. Stochastic [1169, 37, 7, 147, 38, 1943, 1848, 196]. Stock [305, 1155]. Stockholders [1443]. Stones [100]. Storage [1767, 1459, 522, 535, 1161]. Stored [1436]. Stories [1904]. Story [715, 438]. Strategic [259, 949]. Strategies [1273, 5, 1996, 1225, 1927, 1285, 1023, 1541, 1450, 221]. Strategy [1015, 1530, 1507, 755, 604]. Stratified [1474, 160, 1137]. Street [1752]. Streets [1739]. Stress [682]. Stress-ECG [682]. String [1393, 1398]. Strings [1394]. Strip [792]. Strong [954, 950, 955, 951, 1081, 564]. Strongly [1847, 246, 1224]. Structural [229, 687, 954, 400, 955, 1261, 1262, 145, 497, 1162, 841, 1942]. Structure [1620, 501, 818, 372, 1314, 1189, 516, 875, 1388, 559, 797, 1602, 1396, 1397, 204, 166, 1152]. Structured [940, 1047, 416, 1370, 460, 863]. Structures [1725, 466, 369, 1213, 1524, 506, 730, 1525, 1347, 1408, 1631, 1822, 1210, 1193, 914, 1217, 343, 1762, 1321, 1405, 1409, 1406, 645, 646]. Structuring [714, 813, 626]. Stubborn [1163]. Students [1962]. Studies [381, 1694, 1659]. Study [1500, 1669, 1709, 1828, 1825, 1831, 1586, 925, 454, 389, 356, 525, 1510, 281, 1657, 273, 1980, 1815, 1700, 695, 639, 647]. Stutter [1045]. Stutter-Invariant [1045]. Styles [1896]. Subgraph [239, 873]. subgroup [155]. Subjective [1426]. Subjects [397]. Submodular [246, 248]. Subsegments [1399, 1400]. Subsequences [1407, 1393]. Subset [897, 1765, 203]. Substances [720]. Substitution [1367]. Substructure [1753]. subsumption [204]. subsumption-based [204]. Subsystem [229]. Sufficient [1088]. Suffix [1284, 1809, 1357, 1401]. Sugeno [1144]. Suggestions [627]. Suitable [547]. Suite [1726]. Sulcal [371]. Sum [1847, 1298, 1765]. Sum-Free [1847]. Summaries [798, 1475, 1476]. Summarization [794]. Summary [1440]. Super [891, 1684, 1720]. Super-Polynomial [891]. Supercomputing [511]. Superscalar [1686, 519]. Superstring [1389]. Supervised [1239, 327, 328]. Supervision [134]. Supply [302]. Support [1232, 836, 665, 678, 1913, 666, 667, 1481, 1482, 53, 721, 570, 1510, 1886, 658, 290, 1576, 526, 1997, 1255, 675, 1661, 1667, 159]. Supporting [1420, 607, 1582, 1889, 1583, 715, 1241, 1242, 57, 1887, 410, 1888, 633, 1514, 758]. Surface [1602]. SurfaceNets [390]. Surfaces [349, 392, 1603]. Surfing [1516]. Survey [8, 89]. Survivability [1598]. Survival [695]. Swap [884]. Switch [145]. Switched [1576]. SWN [1170]. Symbol [1289, 957]. Symbolic [991, 1031, 1042, 1067, 29, 1011, 461, 1162, 1952, 1052]. Symbols [1131]. Symmetric [236, 1798, 980, 981, 1165, 640]. Synchronized [11, 982]. Synchronous [1718, 810, 516, 2, 1383]. Synergetic [1695]. Synergies [1931]. Synergy [1199, 1196]. Syntactic [1189, 960]. Syntax [1004, 199]. Syntenic [1388]. Synthesis [816, 1725, 10, 541, 1508, 1949, 615, 420, 1947, 565, 797]. Synthesizing [922, 329]. Synthetic [112]. System [1037, 1891, 1930, 998, 280, 999, 665, 622, 581, 1195, 782, 1582, 472, 1413, 414, 342, 968, 90, 1505, 310, 1558, 893, 666, 354, 346, 730, 680, 969, 1000, 979, 308, 705, 1940, 1481, 1482, 595, 930, 1001, 814, 335, 748, 817, 821, 1411, 1190, 1003, 134, 100, 1971, 866, 1020, 74, 847, 1720, 456, 971, 1465, 1007, 744, 1488, 1946, 327, 970, 677, 297, 281, 733, 1723, 1512, 1367, 1008, 674, 1908, 1554, 305, 1907, 1513, 997, 494, 1514, 17, 1005, 1194, 322, 1865, 844, 295, 96, 828, 738, 1010, 745, 1952, 833, 108, 1517]. System [1545, 675, 1518, 84, 75, 1017, 1019, 604, 1991, 894, 500, 194, 187, 169, 164, 218, 199, 191, 172, 213, 171, 458]. Systematic [1976, 565, 1123]. Systems [1038, 314, 670, 61, 856, 284, 26, 943, 1164, 858, 1414, 66, 1726, 1064, 1470, 1477, 1419, 291, 589, 1416, 678, 808, 146, 474, 1249, 848, 1433, 1004, 1266, 1197, 974, 1555, 683, 73, 718, 1171, 144, 341, 1050, 831, 356, 731, 72, 732, 34, 1941, 1713, 787, 706, 42, 1166, 28, 1983, 1694, 1925, 1051, 931, 1028, 757, 578, 317, 973, 1963, 914, 267, 1906, 997, 1418, 345, 786, 736, 662, 1209, 1962, 1167, 793, 1884, 521, 1431, 295, 347, 669, 623, 1520, 1025, 552, 725, 867, 41, 32, 1160, 1177, 264, 797, 1957, 1432, 168, 1153, 158]. systems [159]. Systolic [916]. SZK [1793].

T [1340, 130, 1167]. T-Team [130]. Table [991]. Tableau [587, 588, 583, 596, 597, 600]. Tableau-Based [588]. Tableaux [585, 598, 601, 578]. Tableaux-99 [578]. Tables [1345, 1844]. tabling [204, 223]. Tackling [765]. Tactical [114]. Tag [1084]. Tagged [1448, 383]. Tagging [1077, 1082]. Taking [1376]. Talk [986, 1733, 1742, 1011, 1021, 1749, 1762]. Talking [1570]. Target [402]. Task [546, 1478, 1202, 346, 1480, 427, 1740, 1677, 1193, 625, 290, 1201, 186]. Task-Oriented [1193, 290]. Task-Specific [346]. Tasks [1279, 346, 1490, 119, 1237, 203, 211]. taxonomic [199]. Taxonomy [5]. TCTL [1337, 1338]. Teaching [673]. Team [110, 1861, 125, 121, 98, 126, 87, 88, 114, 127, 116, 136, 129, 118, 131, 132, 85, 104, 108, 86, 135, 124, 133, 120, 1351, 111, 1869, 114, 1870, 130, 123, 131, 115]. Team-Partitioned [104]. Team’98 [117]. Teams [1921, 97]. Teamwork [101, 1556]. Technical [730, 806, 833]. Technique [771, 337, 506, 1557, 325, 699]. Techniques [1170, 1313, 752, 1347, 1961, 1774, 1359, 690, 694, 1514, 835, 841, 1447, 1579, 265, 1146]. Technologies [742, 770]. Technology [1976, 1900, 621, 1966, 1564, 1547, 1455].

Telecommunication [146, 791, 332]. telecommunications [649]. Telematics [1546]. Temperatures [804]. Temporal [1324, 445, 1453, 687, 1048, 1419, 1466, 691, 1757, 1046, 1045, 1458, 1251, 336, 1240, 597, 857, 333, 660, 1952, 867, 868, 1520]. Temporally [498, 22]. Temporary [1279]. Ten [1807]. Tense [592, 591]. Tensor [809]. Tentacle [770]. Term [1581, 1488, 957, 799, 973, 477, 102, 103, 972, 165]. Terminae [712]. Terminal [235]. Terminological [1650, 1651]. Terms [593]. Terrains [1363]. Test [1259, 395, 965, 1035]. Testbed [1218]. Tester [580]. Testing [1843, 900, 1743, 1029, 1175, 271, 1842, 1014, 308, 24, 31, 1841]. Tests [379, 1287, 693, 819, 644]. Text [1334, 1261, 1262, 1233, 1385, 1897, 1386, 1387, 178, 176]. text-document [176]. Texts [730, 760, 1553, 199, 156]. Textual [294]. Texture [467]. Teyjus [1008]. Thai [1085]. Their [882, 38, 635, 1454, 330]. Theorem [1006, 5, 232, 881, 602, 603, 240, 1000, 1629, 593, 1002, 1016, 992, 1302, 42, 1055, 1021, 1018]. Theorem-Proving [5]. Theoretic [452, 679, 1337, 1338, 1721, 1766, 1803, 39]. Theoretical [1173, 1220, 168]. Theoretically [1767, 1310]. Theories [1623, 963, 1135, 1140, 1089]. Theory [778, 743, 1029, 588, 1625, 1176, 1628, 1629, 1479, 594, 1814, 1058, 1104, 959, 1634, 957, 803, 978, 1093, 786, 1174, 994, 961, 874, 1322, 1248, 1133, 206, 1153]. Therapeutic [664]. Therapy [690, 1515]. There [1927]. Thermal [779]. Thin [375]. Thin-Plate [375]. Thinking [1316, 1317]. Third [80]. Thoracic [392]. Thoughts [1807]. Threaded [1810, 1700]. Three [432, 1697, 367, 779]. Three-Dimensional [432, 367]. Threshold [1305, 1769, 444]. Thresholds [1760]. Tiered [1748, 342]. Ties [1355]. Tight [1326, 1280, 1754]. Tighter [1328]. Tightly [1469]. Tightly-Coupled [1469]. Time [687, 1034, 1044, 1970, 1726, 1415, 1731, 1688, 696, 1478, 590, 893, 1269, 1339, 848, 603, 40, 1043, 688, 1740, 1851, 1838, 567, 1357, 937, 42, 1297, 690, 28, 629, 35, 33, 297, 757, 1230, 1839, 689, 364, 1371, 1856, 1256, 806, 1865, 623, 1699, 779, 359, 552, 363, 1224, 1618, 550, 39, 1682, 894, 1432, 187, 213, 223]. Time-Oriented [690]. Time-Series [1224]. time-tabling [223]. Time-Work [1856]. Timed [1026, 1053, 1329, 1337, 1338, 143, 1025, 41]. Timelocks [43]. Timeouts [43]. Times [255]. Timetabling [342]. Timing [660]. TIPPtool [1065]. Tissue [383, 384]. TMN [1597]. Today [1904]. TOGA [849]. Token [1892]. Token-Based [1892]. Tokens [1430]. Tolerance [1849]. Tolerant [416, 1002, 1270, 1747]. Tomography [358, 1621]. Too [1875]. Tool [1725, 712, 1726, 1688, 825, 541, 753, 1727, 1695, 1886, 664, 1068, 192, 753]. Toolbox [1064, 820]. Toolkit [53]. Tools [614, 1033, 1200, 1641, 1701, 157]. Top [1696, 1222, 204]. Top-Down [1222, 204]. Topic [1233]. Topographic [1602]. Topological [743, 448, 1519, 1647]. Topologies [47]. Topology [1272, 1416, 389, 1636]. Tori [909]. Total [167]. Tour [1234]. Toxicology [1096]. TOY [970]. Trace [1178, 921, 1358]. Traceability [1787]. Tracing [1786, 1430]. Track [741]. Trackies [108]. Tracking [1878, 1614, 442, 383, 448, 1455]. Tract [679]. Tractable [1214, 985, 915, 1756]. Trademark [478]. Tradeoffs [1339]. Traders [1168]. Traditional [758]. Traffic [312, 1546, 48, 782, 1583, 271, 1992, 803, 1230, 783, 1592]. Trailing [502]. Train [319]. Training [304, 665, 666, 1139]. Traitor [1785, 1786]. Trajectories [1938]. Trajectory [1042]. Transaction [1433, 1988, 1984, 1694, 1985, 846]. Transactions [1435]. Transducers [1341]. Transfer [1188, 1800, 722, 51].

Transference [881]. Transform [1796, 1824, 479]. Transformation [150, 752, 612]. Transformations [1670, 852, 1182, 373, 985, 613, 853, 768]. Transformer [33]. Transforming [974]. Transforms [502]. Transfusion [685]. Transient [1169]. Transition [104, 1177]. transitions [184]. Translations [60]. Transparent [73]. Transparently [74]. Transport [1815]. Transpositions [877]. Transversal [1812]. Trapdoors [1112]. Traveling [236, 296, 1268]. Treating [300]. Treatment [679, 759]. Treatments [1592]. Tree [1530, 1614, 1337, 1338, 1225, 1343, 1827, 965, 1744, 694, 288, 873, 884, 1803, 1235, 982, 1403, 22]. Tree-Decomposable [1343]. Trees [1745, 1539, 475, 1809, 1348, 898, 1840, 876, 1302, 1357, 1366, 909, 1538, 417, 1756, 1254, 896, 1286, 260, 1811, 654, 657, 218]. Triangle [1737]. Triangulations [1754]. trie [655]. Tries [1401]. triggered [1662]. Triplane [358]. Triple [1113]. Triples [490]. Tristão [13]. Truncated [1773]. Trunk [1931]. Trust [66, 1423, 1427, 59]. Trusted [1444, 80]. Trustee [1430]. truth [183]. TSP [1737, 1813, 919, 1349, 653]. Tuebingen [130]. TUML [867]. Tuning [810]. Tuple [1702]. Turbine [1461]. Turbo [1774]. Turbulent [779]. Turing [1381]. Turning [484, 608, 1256]. Tutor [673]. Tutorial [1197, 1025]. Tutoring [346, 345, 1497]. TV [439, 1516]. Twelf [997]. Two [1703, 1275, 1768, 1797, 1287, 1341, 1770, 1354, 1357, 1359, 1678, 255, 1724, 988, 1862, 779, 774, 260, 878]. Two-Dimensional [1357, 1678, 1724]. Two-Layered [878]. Two-Level [1862]. Two-Party [1768]. Two-Way [1341, 1354]. Type [1623, 1625, 1314, 1628, 1629, 817, 63, 937, 1634, 142]. Type-Safe [63, 937]. Typed [65, 760, 933, 1375]. Types [153, 1331, 64, 594, 1630, 928, 1631, 701, 1543, 929, 774].

Ubidata [1920]. Ubiquitous [1898, 1894, 2000, 1912, 1877, 1989, 1897, 1874, 1871]. UBU [112]. UFSC [1869, 115, 1870]. UFSC-Team [1869, 1870, 115]. UIB [218]. UIB-IK [218]. Ulm [128]. Ultrametric [1286]. Ultrasound [356, 406, 360, 361, 362]. UMAC [1776]. UML [636, 619, 609, 607, 605, 622, 630, 612, 631, 606, 626, 620, 1958, 615, 627, 616, 621, 624, 613, 629, 634, 628, 633, 610, 618, 625, 637, 623, 611]. UML-Specific [627]. UMLS [685, 684]. Un-coupled [785]. Unbundled [1982]. Uncertain [261, 740]. Uncertainty [6, 112, 1531, 18, 1543, 819, 1152]. Unconventional [770]. Undecidability [1175, 957]. underground [1643]. Underlying [1508]. Understanding [1899, 703, 194]. Underwater [13]. Undoing [1250]. Unexpected [1859]. unfamiliar [1642]. Unfold [325]. Unfold/Fold [325]. Unfoldings [1181]. Unidirectional [1981]. Unification [1671, 1213, 971, 1635]. Unified [340, 380, 1479, 827, 632, 1488, 263, 321]. Unified-Metaheuristic [263]. Uniform [852, 1678, 853, 1402, 1273, 876, 385]. Uniformly [1278]. Unifying [1086, 929]. UniMoK [971]. Uninterpreted [1063, 577]. Union [963]. Unit [1854, 698, 343]. United [133, 111]. Units [979, 438, 1866]. Universal [778, 1777, 965, 11, 1922, 994]. Universally [238]. Universals [1187]. University [123, 108, 130]. Unknown [1549, 283]. Unlabeled [319]. Unnoticeable [534]. Unplugged [2000]. Unreliable [1287]. Unsplittable [251]. Unsupervised [755, 1240, 1611]. Update [1290, 1132]. updates [182]. Updating [1125]. UPPAAL [40]. Upper [230, 231, 1369, 1931, 1286]. UQ [124]. Urban [811, 1643]. URCP [1913]. Urinary [679]. Usage [1576, 839]. Usage-Based [1576]. Use [685, 1440, 1819, 1917, 683, 1003, 732, 721, 815, 1896, 635, 1969, 604]. Used [810]. User [541, 422, 74, 1692, 1887, 624, 1948, 1186, 625, 551, 290, 96, 1450, 175, 1960]. User-Driven [541]. User-Friendly [1692]. User-Stereotypes [1450]. Using [563, 419, 431, 1053, 1894, 498, 1031, 1280, 759, 577, 804, 809, 696, 782, 138, 1704, 371, 337, 423, 380, 90, 765, 1334, 1077, 1757, 1202, 425, 464, 465, 756, 387, 277, 350, 801, 1816, 730, 1004, 752, 298, 1344, 1846, 389, 1483, 1988, 1589, 1001, 814, 748, 1171, 40, 144, 430, 1868, 1943, 1903, 1030, 502, 79, 913, 1446, 490, 285, 567, 385, 706, 42, 444, 815, 1953, 476, 390, 1240, 446, 693, 101, 803, 1230, 694, 1512, 1161, 1090, 357, 1495, 364, 783, 792, 1590, 846, 326, 618, 934, 384, 1538, 407, 418, 107, 1194]. Using [318, 468, 1033, 806, 1401, 560, 1639, 1611, 489, 302, 481, 469, 1236, 1172, 152, 1387, 1885, 398, 1516, 1952, 1523, 1593, 1622, 1517, 749, 1255, 319, 1455, 265, 1602, 1248, 1619, 194, 182, 219, 1652, 171, 1638, 1153, 572, 644, 178, 1158, 198]. Utilities [788]. Utility [112]. Utility-Based [112]. Utilization [1481, 1482, 548]. UTTORI [133]. UXF [611].

V [825]. V1.0 [1100]. Valid [230, 231]. Validating [856, 632]. Validation [98, 1001, 1185, 1033]. Value [545, 1715, 517, 1134]. Valued [1007, 775, 1379, 778, 1048]. Values [1517, 183]. Vampire [1009]. Vantage [480]. Variable [230, 231, 1114, 1216, 1360, 374, 491, 797, 601]. Variable-Arity [1216]. Variable-Bit-Length [491]. Variable-Input-Length [1114]. Variables [1275, 1130, 273, 1610, 988, 1224, 1129]. Variance [1382]. Variant [1373, 842]. Variants [1098, 1244, 1349, 408, 1268]. Variations [1852]. Varieties [994]. Various [1759]. Varying [1230]. Vasculature [379, 382]. VBR [1992]. VBR-Encoded [1992]. Vector [284, 767, 1612, 1255].

Vectorization [527]. Vectors [1748, 248]. Vegas [1354]. vehicle [1941]. Vehicles [495]. Ventilation [673]. Verbal [1641, 1950]. Verbs [1217]. Veri [739]. Veri-KoMoD [739]. Verifiable [991, 1772]. Verification [1037, 1038, 1006, 1066, 1034, 1056, 577, 1738, 1024, 98, 1001, 40, 1632, 28, 1028, 582, 295, 739, 1060, 32, 289]. Verifier [1067]. Verifying [1183, 1031, 718, 1050, 41]. Version [1356, 1100, 1103, 1461, 1017, 198]. Versus [1198, 891]. Vertex [232, 1300, 260]. Vertex-Disjoint [260]. Vertigo [681]. very [521]. Vessels [706]. VI [50]. via [1835, 964, 1847, 174, 1459, 369, 522, 1777, 1231, 794, 276, 363, 343]. Video [445, 431, 443, 411, 440, 48, 437, 1265, 442, 1992, 448, 420, 433, 457, 447]. VIE [673]. VIE-NVT [673]. View [68, 1509, 1086, 857, 489, 1662, 1468]. View-Based [68, 489]. view-graph [1662]. Viewer [420]. Viewing [494]. Views [845, 840]. VIL [1778]. VIL-MACs [1778]. Violated [233]. Virtual [3, 496, 1594, 932, 1950, 1917, 1508, 928, 1692, 1953, 420, 1742, 1592, 50, 21, 929, 194]. Virus [210]. Virus-enhanced [210]. Viscosity [374]. Visibility [879]. Vision [1131, 495, 134, 1934, 107, 320, 108, 1936, 494]. Vision-Based [108, 1936]. Visual [450, 498, 426, 475, 294, 455, 507, 109, 1932, 433, 421, 460, 838, 500, 1661, 1659]. Visualisation [415]. Visualization [1232, 509, 427, 429, 510, 690, 410, 1886, 382, 508]. Visualizing [691, 926, 278]. Visually [1939]. Vitreous [527]. VLSI [1730, 648]. Voice [1866]. Voice-Response [1866]. VoIP [1918, 1979]. Volume [362]. Volumes [397, 385, 390]. Voronoi [1750, 1751]. Voting [905, 1772]. Voxel [405]. Voyager [998]. vs [1132, 1608, 1654, 1702]. VSDITLU [991]. VUB [121].

Waldmeister [1003]. Wall [381]. WAP [1988]. Warehouse [840, 839].

Warehousing [1468]. WARM [638]. WARM-UP [638]. Waste [811]. Wastewater [759]. Water [302]. watermarking [651]. Watermarks [1437]. Wave [1516, 1381]. Wavelength [1825]. Wavelet [465, 502]. Way [1341, 993, 1354, 1204, 886, 1538, 254]. wayfinding [1663]. Ways [1521, 1522]. WCL [1865]. Weakened [1778]. Weakly [1072, 952, 953]. Weakness [1781]. Wearable [1902, 1903, 1875, 1905]. Weaver [566]. Web [700, 1263, 9, 714, 1757, 425, 430, 1911, 871, 1198, 179, 432, 1552, 1553, 173, 289]. Web-Pages [289]. WebKB [1203]. WebMaster [289]. Website [1551]. Websites [307]. WebTwig [1911]. Weight [1302, 273, 1754]. Weighted [897, 1382, 639]. Weighting [477]. Weights [478]. Welding [1501]. Welds [749]. Well [1055]. Well-Founded [1055]. Went [1882]. We’re [769]. Whales [502]. Where [769]. Which [1861, 110]. while [1061]. Who [89]. whole [1644]. Wide [1312, 700, 9, 792, 939]. WideWeb [1448]. Width [767]. WinALT [1725]. Winner [1749]. Wire [354]. Wire-EDM [354]. Wired [1999]. Wireless [1982, 1979]. Wirelessly [1999]. WISDOM [164]. Wise [1833, 1834]. within [1819, 719, 1411]. without [1771, 1031, 43, 1802, 424, 1444]. Witness [1981]. Witness-Aided [1981]. Wolfe [250]. Women [819]. Word [963, 1084]. WordNet [1083]. Words [1364, 1365, 1944]. Work [1290, 792, 1856]. Workarounds [606]. Workflow [314, 858, 1466, 1176, 848, 847, 1465, 1464, 662, 865]. Workflows [859]. Working [799, 273, 507]. Workshop [1962]. Workspace [1915]. Workstation [534, 1697]. Workstations [44, 536, 47, 209]. World [1999, 1915, 1738, 288, 1872, 208, 1448, 700, 9]. World-WideWeb [1448]. Worlds [609]. Wormhole [546]. Wormhole-Routed [546]. Worst [1371]. Worth [47]. wound [191]. Write [1806]. Wrong [1882, 1490]. WWW [435, 691, 1261, 1262, 1449].

X [379]. X-ray [379]. XJ [1710]. XML [1509, 1989, 1577, 1198, 737]. XML-based [1198]. XMT [1810]. XZ [1528, 1529]. XZ-Ordering [1529, 1528].

Years [1807]. Yield [1326, 1834].

Z39.50 [139]. Zero [1795, 1335, 1794, 1793, 1336]. Zero-Knowledge [1795, 1335, 1794, 1336]. Ziv [1401, 1386]. Zone [1817]. Zooms [441].

References

Jiang:2000:LBA

[1] Tao Jiang, Ming Li, and Paul Vitányi. A lower bound on the average-case complexity of Shellsort. Journal of the ACM, 47(5): 905–911, 2000. CODEN JACOAH. ISSN 0004-5411. URL http://www.

                                              acm.org/pubs/citations/journals/
                                              jacm/2000-47-5/p905-jiang/.

Tiskin:1998:BSP

[2] A. Tiskin. Bulk-synchronous parallel multiplication of Boolean matrices. Lecture Notes in Computer Science, 1443:494–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). See erratum [1383].

Aylett:1999:BVA

[3] Ruth Aylett. Behavioural virtual agents. Lecture Notes in Computer Science, 1600:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1600/16000001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1600/16000001.pdf.

Baader:1999:LBK

[4] Franz Baader. Logic-based knowledge representation. Lecture Notes in Computer Science, 1600:13–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1600/16000013.
pdf.

Bonacina:1999:TTP

[5] Maria Paola Bonacina. A taxonomy of theorem-proving strategies. Lecture Notes in Computer Science, 1600:43– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1600/16000043.
pdf.

Blythe:1999:OPU

[6] Jim Blythe. An overview of planning under uncertainty. Lecture Notes in Computer Science, 1600:85– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000085.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1600/16000085.
                                        pdf.

Boutilier:1999:KRS

[7] Craig Boutilier. Knowledge representation for stochastic decision processes. Lecture Notes in Computer Science, 1600:111–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1600/16000111.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000111.
                                        pdf.

Bundy:1999:SAD

[8] Alan Bundy. A survey of automated deduction. Lecture Notes in Computer Science, 1600:153–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000153.
                                        pdf.

Ciancarini:1999:WWW

[9] Paolo Ciancarini, Robert Tolksdorf, and Fabio Vitali. The World Wide Web as a place for agents. Lecture Notes in Computer Science, 1600:175– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000175.
                                        pdf.

Elliott:1999:LPA

[10] Clark Elliott, Jeff Rickel, and James Lester. Lifelike pedagogical agents and affective computing: An exploratory synthesis. Lecture Notes in Computer Science, 1600:195–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1600/16000195.
pdf.

Jensen:1999:OBU

[11] Rune M. Jensen and Manuela M. Veloso. OBDD-based universal planning: Specifying and solving planning problems for synchronized agents in non-deterministic domains. Lecture Notes in Computer Science, 1600:213– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1600/16000213.
pdf.

Levy:1999:CAI

[12] Alon Y. Levy. Combining artificial intelligence and databases for data integration. Lecture Notes in Computer Science, 1600:249–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000249.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1600/16000249.
                                        pdf.

Martinho:1999:ULB

[13] Carlos Martinho and Ana Paiva. “underwater love”: Building Tristão and Isolda’s personalities. Lecture Notes in Computer Science, 1600:269–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000269.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000269.
                                        pdf.

Mateas:1999:OCR

[14] Michael Mateas. An Oz-centric review of interactive drama and believable agents. Lecture Notes in Computer Science, 1600:297–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000297.
                                        pdf.

Parsons:1999:RBI

[15] Simon Parsons, Ola Pettersson, Alessandro Saffiotti, and Michael Wooldridge. Robots with the best of intentions. Lecture Notes in Computer Science, 1600:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000329.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1600/16000329.
pdf.

Petrie:1999:ABP

[16] Charles Petrie, Sigrid Goldmann, and Andreas Raquet. Agent-based project management. Lecture Notes in Computer Science, 1600:339–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1600/16000339.
pdf.

Prakken:1999:SDA

[17] Henry Prakken and Giovanni Sartor. A system for defeasible argumentation, with defeasible priorities. Lecture Notes in Computer Science, 1600:365– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1600/16000365.
pdf.

Saffiotti:1999:HUC

[18] Alessandro Saffiotti. Handling uncertainty in control of autonomous robots. Lecture Notes in Computer Science, 1600:381–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1600/16000381.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1600/16000381.
pdf.

Shanahan:1999:ECE

[19] Murray Shanahan. The event calculus explained. Lecture Notes in Computer Science, 1600:409–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000409.
                                        pdf.

Tarau:1999:TLP

[20] Paul Tarau and Veronica Dahl. Towards a logic programming infrastructure for Internet programming. Lecture Notes in Computer Science, 1600: 431–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000431.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000431.
                                        pdf.

Thalmann:1999:TAP

[21] Daniel Thalmann and Hansrudi Noser. Towards autonomous, perceptive, and intelligent virtual actors. Lecture Notes in Computer Science, 1600:457– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1600/16000457.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1600/16000457.
                                        pdf.

Xiang:1999:TIJ

[22] Yang Xiang. Temporally invariant junction tree for inference in dynamic Bayesian network. Lecture Notes in Computer Science, 1600:473– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1600/16000473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1600/16000473.
pdf.

Anonymous:1999:AIa

[23] Anonymous. Author index. Lecture Notes in Computer Science, 1600:489– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1600/1600indx.pdf.

Jonsson:1999:FAC

[24] Bengt Jonsson and Wang Yi. Fully abstract characterization of probabilistic may testing. Lecture Notes in Computer Science, 1601:1–18, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010001.
pdf.

McIver:1999:QPL

[25] Annabelle K. McIver. Quantitative program logic and performance in probabilistic distributed algorithms. Lecture Notes in Computer Science, 1601:19–33, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1601/16010019.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1601/16010019.
pdf.

Baier:1999:EQP

[26] Christel Baier and Bettina Engelen. Establishing qualitative properties for probabilistic lossy channel systems: An algorithmic approach. Lecture Notes in Computer Science, 1601:34– 52, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010034.
pdf.

Stoelinga:1999:RCI

[27] Mariëlle Stoelinga and Frits Vaandrager. Root contention in IEEE 1394. Lecture Notes in Computer Science, 1601:53–74, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1601/16010053.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1601/16010053.
pdf.

Kwiatkowska:1999:AVR

[28] Marta Kwiatkowska, Gethin Norman, Roberto Segala, and Jeremy Sproston. Automatic verification of realtime systems with discrete probability distributions. Lecture Notes in Computer Science, 1601:75–95, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010075.
pdf.

Hartonas-Garmhausen:1999:PPS

[29] Vicky Hartonas-Garmhausen, Sergio Campos, and Ed Clarke. ProbVerus: Probabilistic symbolic model checking. Lecture Notes in Computer Science, 1601:96–110, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1601/16010096.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1601/16010096.
pdf.

Andova:1999:PAP

[30] Suzana Andova. Process algebra with probabilistic choice. Lecture Notes in Computer Science, 1601:111–129, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010111.
pdf.

Nunez:1999:APT

[31] Manuel Núñez. An axiomatization of probabilistic testing. Lecture Notes in Computer Science, 1601:130–150, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1601/16010130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1601/16010130.
                                        pdf.

Vaandrager:1999:VHS

[32] Frits Vaandrager. Verification of hybrid systems. Lecture Notes in Computer Science, 1601:151–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1601/16010151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1601/16010151.
                                        pdf.

Lermer:1999:POR

[33] Karl Lermer. A parallel operator for real-time processes with predicate transformer semantics. Lecture Notes in Computer Science, 1601:152– 171, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1601/16010152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1601/16010152.
                                        pdf.

Jenner:1999:CEA

[34] Lars Jenner and Walter Vogler. Comparing the efficiency of asynchronous systems. Lecture Notes in Computer Science, 1601:172–191, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1601/16010172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010172.
pdf.

Lermer:1999:FMR

[35] Karl Lermer and Colin Fidge. A formal model of real-time program compilation. Lecture Notes in Computer Science, 1601:192–210, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010192.
pdf.

Clark:1999:SPM

[36] Graham Clark, Stephen Gilmore, and Jane Hillston. Specifying performance measures for PEPA. Lecture Notes in Computer Science, 1601:211–227, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010211.
pdf.

Bohnenkamp:1999:SNS

[37] Henrik C. Bohnenkamp and Boudewijn R. Haverkort. Semi-numerical solution of stochastic process algebra models. Lecture Notes in Computer Science, 1601:228–243, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1601/16010228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1601/16010228.
                                        pdf.

Hermanns:1999:BAS

[38] Holger Hermanns and Markus Siegle. Bisimulation algorithms for stochastic process algebras and their BDDbased implementation. Lecture Notes in Computer Science, 1601:244–264, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1601/16010244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1601/16010244.
                                        pdf.

Vardi:1999:PLT

[39] Moshe Y. Vardi. Probabilistic lineartime model checking: An overview of the automata-theoretic approach. Lecture Notes in Computer Science, 1601: 265–276, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1601/16010265.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1601/16010265.
                                        pdf.

Havelund:1999:FVP

[40] Klaus Havelund, Kim Guldstrand Larsen, and Arne Skou. Formal verification of a power controller using the real-time model checker UPPAAL. Lecture Notes in Computer Science, 1601:277–298, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010277.
pdf.

Tripakis:1999:VPT

[41] Stavros Tripakis. Verifying progress in timed systems. Lecture Notes in Computer Science, 1601:299–314, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010299.
pdf.

Kolano:1999:PAR

[42] Paul Z. Kolano. Proof assistance for real-time systems using an interactive theorem prover. Lecture Notes in Computer Science, 1601:315–333, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1601/16010315.
pdf.

Bowman:1999:MTT

[43] Howard Bowman. Modelling timeouts without timelocks. Lecture Notes in Computer Science, 1601:334–354, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1601/16010334.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1601/16010334.
                                        pdf.

Fusen:1999:PIE

[44] H. Fusen and W. Jie. A parallel implementation of the Everglades Landscape Fire Model in networks of workstations. Lecture Notes in Computer Science, 1602:1–15, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steenkiste:1999:AMN

[45] P. Steenkiste. Adaptation models for network-aware distributed computations. Lecture Notes in Computer Science, 1602:16–31, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yerosheva:1999:PEM

[46] L. Yerosheva and P. M. Kogge. Prototyping execution models for HTMT petaflop machine in Java. Lecture Notes in Computer Science, 1602:32– 46, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silla:1999:IWF

[47] F. Silla and J. Duato. Is it worth the flexibility provided by irregular topologies in networks of workstations? Lecture Notes in Computer Science, 1602:47–61, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caminero:1999:PEM

[48] B. Caminero, F. J. Quiles, Y. Duato, and D. S. Love. Performance evaluation of the multimedia router with MPEG-2 video traffic. Lecture Notes in Computer Science, 1602:62–76, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canonico:1999:SFC

[49] R. Canonico, R. Cristaldi, and G. Iannello. A scalable flow control algorithm for the fast messages communication library. Lecture Notes in Computer Science, 1602:77–90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shah:1999:HPS

[50] H. V. Shah, C. Pu, and R. S. Madukkarumukumana. High performance sockets and RPC over virtual interface (VI) architecture. Lecture Notes in Computer Science, 1602:91–107, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philp:1999:STS

[51] I. R. Philp and Y.-L. Liong. The scheduled transfer (ST) protocol. Lecture Notes in Computer Science, 1602:108– 121, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1999:NPO

[52] P. Watson and S. Parastatidis. The NIP parallel object-oriented computational model. Lecture Notes in Computer Science, 1602:122–136, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herlihy:1999:ATS

[53] M. Herlihy. The Aleph Toolkit: Support for scalable distributed shared objects. Lecture Notes in Computer Science, 1602:137–149, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1999:CAD

[54] D. R. Kumar and W. A. Najjar. Combining adaptive and deterministic routing: Evaluation of a hybrid router. Lecture Notes in Computer Science, 1602:150–164, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casado:1999:DFR

[55] R. Casado, F. J. Quiles, J. L. Sanchez, and J. Duato. Deadlock-free routing in irregular networks with dynamic reconfiguration. Lecture Notes in Computer Science, 1602:165–180, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brooks:1999:IAS

[56] D. Brooks and M. Martonosi. Implementing application-specific cachecoherence protocols in configurable hardware. Lecture Notes in Computer Science, 1602:181–195, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karl:1999:SSM

[57] W. Karl, M. Leberecht, and M. Schulz. Supporting shared memory and message passing on clusters of PCs with a SMiLE. Lecture Notes in Computer Science, 1602:196–210, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacunski:1999:LLM

[58] M. Jacunski, V. Moorthy, P. P. Ware, and M. Pillai. Low latency messagepassing for reflective memory networks. Lecture Notes in Computer Science, 1602:211–224, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swarup:1999:TBM

[59] V. Swarup and J. T. Fabrega. Trust: Benefits, models, and mechanisms. Lecture Notes in Computer Science, 1603:3–18, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1999:PPL

[60] M. Abadi. Protection in programminglanguage translations. Lecture Notes in Computer Science, 1603:19–34, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:1999:RAS

[61] M. Ancona, W. Cazzola, and E. B. Fernandez. Reflective authorization systems: Possibilities, benefits, and drawbacks. Lecture Notes in Computer Science, 1603:35–50, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardelli:1999:AMC

[62] L. Cardelli. Abstractions for mobile computations. Lecture Notes in Computer Science, 1603:51–94, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennessey:1999:TSE

[63] M. Hennessey and J. Riely. Type-safe execution of mobile agents in anonymous networks. Lecture Notes in Computer Science, 1603:95–116, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1999:TSA

[64] R. De Nicola, G. Ferrari, and R. Pugliese. Types as specifications of access policies. Lecture Notes in Computer Science, 1603:117–146, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leroy:1999:SPT

[65] X. Leroy and F. Rouaix. Security properties of typed applets. Lecture Notes in Computer Science, 1603:147–184, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaze:1999:RTM

[66] M. Blaze, J. Feigenbaum, J. Ioannidis, and A. D. Keromytis. The role of trust management in distributed systems security. Lecture Notes in Computer Science, 1603:185–210, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aura:1999:DAR

[67] T. Aura. Distributed access-rights management with delegation certificates. Lecture Notes in Computer Science, 1603:211–236, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brose:1999:VBA

[68] G. Brose. A view-based access control model for CORBA. Lecture Notes in Computer Science, 1603:237–252, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tschudin:1999:APD

[69] C. Tschudin. Apoptosis — the programmed death of distributed services. Lecture Notes in Computer Science, 1603:253–260, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yee:1999:SMA

[70] B. S. Yee. A sanctuary for mobile agents. Lecture Notes in Computer Science, 1603:261–274, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roth:1999:MPC

[71] V. Roth. Mutual protection of cooperating agents. Lecture Notes in Computer Science, 1603:275–288, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaeger:1999:ACC

[72] T. Jaeger. Access control in configurable systems. Lecture Notes in Computer Science, 1603:289–316, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grimm:1999:PPN

[73] R. Grimm and B. N. Bershad. Providing policy-neutral and transparent access control in extensible systems. Lecture Notes in Computer Science, 1603:317–338, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1999:IAT

[74] M. B. Jones. Interposition agents: Transparently interposing user code at the system interface. Lecture Notes in Computer Science, 1603:339–368, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonEicken:1999:JKC

[75] T. Von Eicken, C.-C. Chang, G. Czajkowski, and C. Hawblitzel. J-Kernel: A capability-based operating system for Java. Lecture Notes in Computer Science, 1603:369–394, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDoorn:1999:SNO

[76] L. Van Doorn, M. Abadi, M. Burrows, and E. Wobber. Secure network objects. Lecture Notes in Computer Science, 1603:395–412, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edjlali:1999:HBA

[77] G. Edjlali, A. Acharya, and V. Chaudhary. History-based access control for mobile code. Lecture Notes in Computer Science, 1603:413–432, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexander:1999:SAN

[78] D. S. Alexander, W. A. Arbaugh, A. D. Keromytis, and J. M. Smith. Security in active networks. Lecture Notes in Computer Science, 1603:433–452, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hulaas:1999:UIS

[79] J. Hulaas, A. Villazon, and J. Harms. Using interfaces to specify access rights. Lecture Notes in Computer Science, 1603:453–468, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilhelm:1999:ITT

[80] U. G. Wilhelm, S. Staamann, and L. Buttyan. Introducing trusted third parties to the mobile agent paradigm. Lecture Notes in Computer Science, 1603:469–492, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asada:1999:OR

[81] Minoru Asada, Manuela Veloso, Milind Tambe, Itsuki Noda, Hiroaki Kitano, and Gerhard K. Kraetzschmar. Overview of RoboCup-98. Lecture Notes in Computer Science, 1604:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040001.
pdf.

Binsted:1999:CDS

[82] Kim Binsted and Sean Luke. Character design for soccer commentary. Lecture Notes in Computer Science, 1604: 22–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040022.
pdf.

Matsubara:1999:ASC

[83] Hitoshi Matsubara, Ian Frank, Kumiko Tanaka-Ishii, Itsuki Noda, Hideyuki Nakashima, and Kôiti Hasida. Automatic soccer commentary and RoboCup. Lecture Notes in Computer Science, 1604:34–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040034.
pdf.

Voelz:1999:RRS

[84] Dirk Voelz, Elisabeth André, Gerd Herzog, and Thomas Rist. Rocco: A RoboCup soccer commentator system. Lecture Notes in Computer Science, 1604:50–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1604/16040050.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1604/16040050.
pdf.

Stone:1999:CCS

[85] Peter Stone, Manuela Veloso, and Patrick Riley. The CMUnited-98 champion simulator team. Lecture Notes in Computer Science, 1604:61– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040061.
pdf.

Veloso:1999:CCS

[86] Manuela Veloso, Michael Bowling, Sorin Achim, Kwun Han, and Peter Stone. The CMUnited-98 champion small-robot team. Lecture Notes in Computer Science, 1604:77–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040077.
pdf.

Gulmann:1999:CFR

[87] J.-S. Gulmann, W. Hatzack, I. Herrmann, and B. Nebel. The CS Freiburg robotic soccer team: Reliable selflocalization, multirobot sensor integration, and basic soccer skills. Lecture Notes in Computer Science, 1604:93– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutmann:1999:CFR

[88] Jens-Steffen Gutmann, Wolfgang Hatzack, Immanuel Herrmann, Bernhard Nebel, Frank Rittinger, Augustinus Topor, Thilo Weigel, and Bruno Welsch. The CS freiburg robotic soccer team: Reliable selflocalization, multirobot sensor integration, and basic soccer skills. Lecture Notes in Computer Science, 1604:93– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040093.
pdf.

Verner:1999:SRW

[89] Igor M. Verner. The survey of RoboCup ’98: Who, how and why. Lecture Notes in Computer Science, 1604: 109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040109.
                                        pdf.

Coradeschi:1999:HMC

[90] Silvia Coradeschi and Jacek Malec. How to make a challenging AI course enjoyable using the RoboCup soccer simulation system. Lecture Notes in Computer Science, 1604:120–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040120.
                                        pdf.

Fujita:1999:QRR

[91] Masahiro Fujita, Stephane Zrehen, and Hiroaki Kitano. A quadruped robot for RoboCup legged robot challenge in Paris ’98. Lecture Notes in Computer Science, 1604:125–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040125.
                                        pdf.

Lund:1999:RSL

[92] Henrik Hautop Lund and Luigi Pagliarini. Robot soccer with LEGO mindstorms. Lecture Notes in Computer Science, 1604:141–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040141.
pdf.

Maeda:1999:BRS

[93] Kazuaki Maeda, Akinori Kohketsu, and Tomoichi Takahashi. Ballreceiving skill dependent on centering in soccer simulation games. Lecture Notes in Computer Science, 1604:152– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040152.
pdf.

Marda:1999:BRS

[94] K. Marda, A. Kohketsu, and T. Takahashi. Ball-receiving skill dependent on centering in soccer simulation games. Lecture Notes in Computer Science, 1604:152–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lubbers:1999:PCM

[95] Jan Lubbers and Rogier R. Spaans. The priority/confidence model as a framework for soccer agents. Lecture Notes in Computer Science, 1604:162– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040162.
pdf.

Scerri:1999:UOS

[96] Paul Scerri, Silvia Coradeschi, and Anders Törne. A user oriented system for developing behavior based agents. Lecture Notes in Computer Science, 1604: 173–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040173.
pdf.

Takahashi:1999:PRG

[97] Tomoichi Takahashi and Tadashi Naruse. From play recognition to good plays detection — reviewing RoboCup 97 teams from logfile. Lecture Notes in Computer Science, 1604: 187–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040187.
pdf.

Driessens:1999:IVV

[98] Kurt Driessens, Nico Jacobs, Nathalie Cossement, Patrick Monsieurs, and Luc De Raedt. Inductive verification and validation of the KUL-RoT RoboCup team. Lecture Notes in Computer Science, 1604:193–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040193.
pdf.

Jung:1999:LRA

[99] Christoph G. Jung. Layered and resource-adapting agents in the RoboCup simulation. Lecture Notes in Computer Science, 1604:207–220, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1604/16040207.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1604/16040207.pdf.

Ito:1999:DPS

[100] Nobuhiro Ito, Kouichi Nakagawa, Xiaoyong Du, and Naohiro Ishii. A description-processing system for soccer agents and NIT Stones 98. Lecture Notes in Computer Science, 1604: 221–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040221.
pdf.

Marsella:1999:UET

[101] Stacy Marsella, Jafar Adibi, Yaser Al-Onaizan, Ali Erdem, Randall Hill, Gal A. Kaminka, Zhun Qiu, and Milind Tambe. Using an explicit teamwork model and learning in RoboCup. Lecture Notes in Computer Science, 1604: 237–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040237.
                                        pdf.

Tortcrolo:1999:HAM

[102] F. Tortcrolo and C. Carbay. A hybrid agent model, mixing short term and long term memory abilities. an application to RoboCup competition. Lecture Notes in Computer Science, 1604: 246–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torterolo:1999:HAM

[103] Fausto Torterolo and Catherine Garbay. A hybrid agent model, mixing short term and long term memory abilities. Lecture Notes in Computer Science, 1604:246–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040246.
                                        pdf.

Stone:1999:TPO

[104] Peter Stone and Manuela Veloso. Team-partitioned, opaque-transition reinforcement learning. Lecture Notes in Computer Science, 1604:261–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040261.
pdf.

Uchibe:1999:CBA

[105] Eiji Uchibe, Masateru Nakamura, and Minoru Asada. Cooperative behavior acquisition in a multiple mobile robot environment by co-evolution. Lecture Notes in Computer Science, 1604:273– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040273.
pdf.

Shen:1999:IRS

[106] Wei-Min Shen, Jafar Adibi, Rogelio Adobbati, Srini Lanksham, Hadi Moradi, Behnam Salemi, and Sheila Tejada. Integrated reactive soccer agents. Lecture Notes in Computer Science, 1604:286–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040286.
pdf.

Price:1999:IAV

[107] Andrew R. Price and Trevor Jones. An innovative approach to vision, localization and orientation using omnidirectional radial signature analysis. Lecture Notes in Computer Science, 1604: 299–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040299.
pdf.

Suzuki:1999:AVB

[108] Sho’ji Suzuki, Tatsunori Kato, Hiroshi Ishizuka, Yasutake Takahashi, Eiji Uchibe, and Minoru Asada. An application of vision-based learning in RoboCup for a real robot with an omnidirectional vision system and the team description of Osaka University “Trackies”. Lecture Notes in Computer Science, 1604:316–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040316.
pdf.

Nakamura:1999:RNC

[109] Takayuki Nakamura, Kazunori Terada, Atsushi Shibata, and Hideaki Takeda. The RoboCup-NAIST: A cheap multisensor-based mobile robot with visual learning capability. Lecture Notes in Computer Science, 1604: 326–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040326.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1604/16040326.
pdf.

Andou:1999:ARTa

[110] T. Andou. Andhill-98: A RoboCup team which reinforces positioning with observation. Lecture Notes in Computer Science, 1604:338–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1999:ETD

[111] David Andre and Astro Teller. Evolving Team Darwin United. Lecture Notes in Computer Science, 1604:346– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040346.
pdf.

Boman:1999:UUB

[112] Magnus Boman, Helena Åberg, Åsa Åhman, Jens Andreasen, Mats Danielson, Carl-Gustaf Jansson, Johan Kummeneje, Harko Verhagen, and Johan Walter. UBU: Utility-based uncertainty handling in synthetic soccer. Lecture Notes in Computer Science, 1604:352–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1604/16040352.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1604/16040352.
pdf.

Gugenberger:1999:HR

[113] P. Gugenberger, J. Wendler, K. Schroeter, and H.-D. Burkhard. AT Humboldt in RoboCup-98. Lecture Notes in Computer Science, 1604:358–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040358.
                                        pdf.

Igarashi:1999:ITP

[114] Harukazu Igarashi, Shougo Kosue, and Masatoshi Miyahara. Individual tactical play and pass with communication between players — team descriptions of Team Miya2. Lecture Notes in Computer Science, 1604:364– 370, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040364.
                                        pdf.

DaCosta:1999:UTC

[115] A. C. P. L. Da Costa and G. Bittencourt. UFSC-team: A cognitive multiagent approach to the RoboCup’98 simulator league. Lecture Notes in Computer Science, 1604:371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumura:1999:DTE

[116] Takeshi Matsumura. Description of team erika. Lecture Notes in Computer Science, 1604:377–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040377.
pdf.

Pagello:1999:GGP

[117] E. Pagello, F. Montesello, F. Garelli, F. Candon, P. Chioetto, and S. Griggio. Getting global performance through local information in PaSo-Team’98. Lecture Notes in Computer Science, 1604: 384–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040384.
pdf.

Polani:1999:DAR

[118] Daniel Polani, Stefan Weber, and Thomas Uthmann. A direct approach to robot soccer agents: Description for the team MAINZ ROLLING BRAINS simulation league of RoboCup’98. Lecture Notes in Computer Science, 1604: 390–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040390.
pdf.

Riekki:1999:CFE

[119] Jukka Riekki, Jussi Pajala, Antti Tikanmäki, and Juha Röning. CAT finland: Executing primitive tasks in parallel. Lecture Notes in Computer Science, 1604:396–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1604/16040396.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1604/16040396.pdf.

Zhang:1999:MLCa

[120] Y. Zhang and A. K. Mackworth. A multi-level constraint-based controller for the Dynamo98 robot soccer team. Lecture Notes in Computer Science, 1604:402–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birk:1999:SLR

[121] Andreas Birk, Thomas Walle, Tony Belpaeme, Johan Parent, Tom De Vlaminck, and Holger Kenn. The small league RoboCup team of the VUB AI-lab. Lecture Notes in Computer Science, 1604:410–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040410.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040410.
                                        pdf.

Braunl:1999:CGS

[122] T. Braunl. CHPS glory — soccer robots with local intelligence. Lecture Notes in Computer Science, 1604:416– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowstron:1999:CUR

[123] A. Rowstron, B. Bradshaw, D. Crosby, T. Edmonds, S. Hodges, A. Hopper, S. Lloyd, J. Wang, and S. Wray. The Cambridge University Robot Football Team description. Lecture Notes in Computer Science, 1604:422–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040422.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040422.
pdf.

Wyeth:1999:URS

[124] Gordon Wyeth, Brett Browning, and Ashley Tews. The UQ RoboRoos small-size league team description for RoboCup ’98. Lecture Notes in Computer Science, 1604:428–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040428.
pdf.

Aparicio:1999:ITD

[125] Pedro Aparício, Rodrigo Ventura, Pedro Lima, and Carlos Pinto-Ferreira. ISocRob — team description. Lecture Notes in Computer Science, 1604:434– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1604/16040434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040434.
                                        pdf.

Gonzalez:1999:RMT

[126] E. González, H. Loaiza, A. Suárez, and C. Moreno. Real MagiCol 98: Team description and results. Lecture Notes in Computer Science, 1604:440– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040440.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040440.
                                        pdf.

Klupsch:1999:ARR

[127] Michael Klupsch, Maximilian Lückenhaus, Christoph Zierl, Ivan Laptev, Thorsten Bandlow, Marc Grimme, Ignaz Kellerer, and Fabian Schwarzer. Agilo RoboCuppers: RoboCup team description. Lecture Notes in Computer Science, 1604:446–451, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040446.
                                        pdf.

Kraetzschmar:1999:USR

[128] Gerhard K. Kraetzschmar, Stefan Enderle, Stefan Sablatnög, Thomas Boß, Mark Dettinger, Hans Braxmayer, Heiko Folkerts, Markus Klingler, Dominik Maschke, Gerd Mayer, Markus Müller, Alexander Neubeck, Marcus Ritter, Heiner Seidl, Robert Wörz, and Günther Palm. The Ulm Sparrows: Research into sensorimotor integration, agency, learning, and multiagent cooperation. Lecture Notes in Computer Science, 1604:452–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040452.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040452.
pdf.

Nardi:1999:AAR

[129] Daniele Nardi, Giorgio Clemente, and Enrico Pagello. ART azzurra robot team. Lecture Notes in Computer Science, 1604:458–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040458.
pdf.

Plagge:1999:DET

[130] Michael Plagge, Boris Diebold, Richard Günther, Jörn Ihlenburg, Dirk Jung, Keyan Zahedi, and Andreas Zell. Design and evaluation of the T-Team of the University of Tuebingen for RoboCup’98. Lecture Notes in Computer Science, 1604:464–472, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040464.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1604/16040464.
                                        pdf.

Sicgberg:1999:TDG

[131] A. Sicgberg, A. Bredenfeld, H. Guenther, and H. U. Kobialka. Team description of the GMD RoboCup-Team. Lecture Notes in Computer Science, 1604:473–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegberg:1999:TDG

[132] A. Siegberg, A. Bredenfeld, H. Guenther, H. U. Kobialka, B. Klaassen, U. Licht, K. L. Paap, P. G. Ploeger, H. Streich, J. Vollmer, J. Wilberg, R. Worst, and T. Christaller. Team description of the GMD robocup-team. Lecture Notes in Computer Science, 1604:473–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1604/16040473.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040473.
                                        pdf.

Yokota:1999:UUC

[133] K. Yokota, K. Ozaki, N. Watanabe, A. Matsumoto, D. Koyama, T. Ishikawa, K. Kawabata, H. Kaetsu, and H. Asama. UTTORI united: Cooperative team play based on communication. Lecture Notes in Computer Science, 1604:479–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040479.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1604/16040479.
pdf.

Hugel:1999:QRG

[134] Vincent Hugel, Patrick Bonnin, Jean Christophe Bouramoué, Didier Solheid, Pierre Blazevic, and Dominique Duhaut. Quadruped robot guided by enhanced vision system and supervision modules. Lecture Notes in Computer Science, 1604:485– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040485.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040485.
pdf.

Veloso:1999:CSL

[135] Manuela Veloso and William Uther. The CMTrio-98 sony legged robot team. Lecture Notes in Computer Science, 1604:491–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1604/16040491.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1604/16040491.pdf.

Mitsunaga:1999:BOL

[136] Noriaki Mitsunaga, Minoru Asada, and Chizuko Mishima. BabyTigers-98: Osaka legged robot team. Lecture Notes in Computer Science, 1604:498– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040498.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040498.
                                        pdf.

Anonymous:1999:AIb

[137] Anonymous. Author index. Lecture Notes in Computer Science, 1604:507– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1604/1604auth.pdf.

Capellmann:1999:UHL

[138] C. Capellmann, H. Dibold, and U. Herzog. Using high-level Petri nets in the field of intelligent networks. Lecture Notes in Computer Science, 1605:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakos:1999:IMZ

[139] C. Lakos and J. Lamp. The incremental modelling of the Z39.50 protocol with object Petri nets. Lecture Notes in Computer Science, 1605:37–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wheeler:1999:MAI

[140] G. Wheeler. The modelling and analysis of IEEE 802.6’s configuration control protocol with coloured Petri nets. Lecture Notes in Computer Science, 1605:69–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BenMnaouer:1999:CPN

[141] A. Ben Mnaouer, T. Sekiguchi, Y. Fujii, and T. Ito. Coloured Petri nets based modelling and simulation of the static and dynamic allocation policies of the asynchronous bandwidth in the fieldbus protocol. Lecture Notes in Computer Science, 1605:93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moon:1999:PRP

[142] H.-J. Moon, S. Y. Moon, and W. H. Kwon. Parameter region for the proper operation of the IEEE 802.2 LLC Type 3 protocol: A Petri net approach. Lecture Notes in Computer Science, 1605:131–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reid:1999:TPN

[143] M. Reid and W. M. Zuberek. Timed Petri net models of ATM LANs. Lecture Notes in Computer Science, 1605: 150–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haverkort:1999:PEP

[144] B. R. Haverkort. Performance evaluation of polling-based communication systems using SPNs. Lecture Notes in Computer Science, 1605:176–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haverkort:1999:SDS

[145] B. R. Haverkort and H. P. Idzenga. Structural decomposition and serial solution of SPN models of the ATM GAUSS switch. Lecture Notes in Computer Science, 1605:210–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMeer:1999:CMC

[146] H. De Meer, O.-R. Duesterhoeft, and S. Fischer. COSTPN for modelling and control of telecommunication systems. Lecture Notes in Computer Science, 1605:232–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franceschinis:1999:SCP

[147] G. Franceschinis, A. Fumagalli, and A. Silinguelli. Stochastic coloured Petri net models for rainbow optical networks. Lecture Notes in Computer Science, 1605:273–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Augusteijn:1999:SM

[148] L. Augusteijn. Sorting morphisms. Lecture Notes in Computer Science, 1608: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backhouse:1999:GPI

[149] R. Backhouse, P. Jansson, J. Jeuring, and L. Meertens. Generic programming — an introduction. Lecture Notes in Computer Science, 1608:28–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMoor:1999:GPT

[150] O. De Moor and G. Sittampalam. Generic program transformation. Lecture Notes in Computer Science, 1608: 116–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swierstra:1999:DIC

[151] S. D. Swierstra, P. R. A. Alcocer, and J. Saraiva. Designing and implementing combinator languages. Lecture Notes in Computer Science, 1608:150– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheard:1999:UMS

[152] T. Sheard. Using MetaML: A staged programming language. Lecture Notes in Computer Science, 1608:207–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Augustsson:1999:CLD

[153] L. Augustsson. Cayenne — A language with dependent types. Lecture Notes in Computer Science, 1608:240–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leijen:1999:HAC

[154] D. Leijen, E. Meijer, and J. Hook. Haskell as an automation controller. Lecture Notes in Computer Science, 1608:268–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloesgen:1999:ARP

[155] W. Kloesgen. Applications and research problems of subgroup mining. Lecture Notes in Computer Science, 1609:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kodratoff:1999:KDT

[156] Y. Kodratoff. Knowledge discovery in texts: A definition and applications. Lecture Notes in Computer Science, 1609:16–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meersman:1999:SOT

[157] R. A. Meersman. Semantic ontology tools in IS design. Lecture Notes in Computer Science, 1609:30–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subrahmanian:1999:HAS

[158] V. S. Subrahmanian. Heterogeneous agent systems. Lecture Notes in Computer Science, 1609:46–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiederhold:1999:ISR

[159] G. Wiederhold. Information systems that really support decision-making. Lecture Notes in Computer Science, 1609:56–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:CSP

[160] J. Chen. A class of stratified programs in autoepistemic logic of knowledge and belief. Lecture Notes in Computer Science, 1609:67–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordana:1999:LEM

[161] A. Giordana and L. Saitta. On-line estimation of matching complexity in first order logic. Lecture Notes in Computer Science, 1609:76–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1999:IPL

[162] C.-M. Wu and Y. Zhang. Implementing prioritized logic programming. Lecture Notes in Computer Science, 1609: 85–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badia:1999:EDL

[163] A. Badia. Extending description logics with generalized quantification. Lecture Notes in Computer Science, 1609: 94–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1999:MLI

[164] F. Esposito, D. Malerba, and F. A. Lisi. Machine learning for intelligent document processing: The WISDOM system. Lecture Notes in Computer Science, 1609:103–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alsaffar:1999:CBR

[165] A. H. Alsaffar, J. S. Deogun, V. V. Raghavan, and H. Sever. Concept based retrieval by minimal term sets. Lecture Notes in Computer Science, 1609:114–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambrix:1999:QDB

[166] P. Lambrix, N. Shahmehri, and S. Jacobsen. Querying document bases by content, structure and properties. Lecture Notes in Computer Science, 1609: 123–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebastiani:1999:TKP

[167] F. Sebastiani. Total knowledge and partial knowledge in logical models of information retrieval. Lecture Notes in Computer Science, 1609:133–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deen:1999:TTF

[168] S. M. Deen and C. A. Johnson. Towards a theoretical foundation for cooperating knowledge based systems. Lecture Notes in Computer Science, 1609:144–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desai:1999:CSC

[169] B. C. Desai, R. Shinghal, N. Shayan, and Y. Zhou. CINDI: A system for cataloguing, searching, and annotating electronic documents in digital libraries. Lecture Notes in Computer Science, 1609:154–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohsuga:1999:EHI

[170] S. Ohsuga and T. Aida. Externalization of human idea and problem description for automatic programming. Lecture Notes in Computer Science, 1609:163–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Predki:1999:RSB

[171] B. Predki and S. Wilk. Rough set based data exploration using ROSE system. Lecture Notes in Computer Science, 1609:172–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholczyk:1999:TIS

[172] D. Pacholczyk. Towards an intelligent system dealing with nuanced information expressed in affirmative or negative forms. Lecture Notes in Computer Science, 1609:181–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1999:LPI

[173] B. Thomas. Logic programs for intelligent Web search. Lecture Notes in Computer Science, 1609:190–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basili:1999:RDC

[174] R. Basili, M. Di Nanni, and M. T. Pazienza. Representing document content via an object-oriented paradigm. Lecture Notes in Computer Science, 1609:199–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semeraro:1999:LSI

[175] G. Semeraro, M. F. Costabile, F. Esposito, and N. Fanizzi. A learning server for inducing user classification rules in the digital library service. Lecture Notes in Computer Science, 1609:208–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kundu:1999:BFM

[176] S. Kundu. A better fitness measure of a text-document for a given set of keywords. Lecture Notes in Computer Science, 1609:217–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zarri:1999:CEI

[177] G. P. Zarri, E. Bertino, B. Black, and A. Brasher. CONCERTO, an environment for the ‘intelligent’ indexing, querying and retrieval of digital documents. Lecture Notes in Computer Science, 1609:226–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:TCU

[178] H. Wang and H. S. Nguyen. Text classification using lattice machine. Lecture Notes in Computer Science, 1609: 235–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merzbacher:1999:DSP

[179] M. Merzbacher. Discovering semantic proximity for Web pages. Lecture Notes in Computer Science, 1609:244– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1999:DLB

[180] I. Weber. A declarative language bias for levelwise search of first-order regularities. Lecture Notes in Computer Science, 1609:253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ejdys:1999:MWL

[181] P. Ejdys and G. Gora. The more we learn the less we know? on inductive learning from examples. Lecture Notes in Computer Science, 1609:262– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doherty:1999:CMU

[182] P. Doherty, W. Lukaszewicz, and E. Madalinska-Bugaj. Computing MPMA updates using Dijkstra’s semantics. Lecture Notes in Computer Science, 1609:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukaidono:1999:SET

[183] M. Mukaidono. Several extensions of truth values in fuzzy logic. Lecture Notes in Computer Science, 1609:282– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bejar:1999:PTR

[184] R. Bejar and F. Manya. Phase transitions in the regular random 3-SAT problem. Lecture Notes in Computer Science, 1609:292–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brisoux:1999:IBS

[185] L. Brisoux, E. Gregoire, and L. Saies. Improving backtrack search for SAT by means of redundancy. Lecture Notes in Computer Science, 1609:301–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1999:DTS

[186] H. S. Nguyen, S. H. Nguyen, and A. Skowron. Decomposition of task specification problems. Lecture Notes in Computer Science, 1609:310–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

An:1999:ODR

[187] A. An, N. J. Cercone, and C. Chan. ORTES: The design of a real-time control expert system. Lecture Notes in Computer Science, 1609:319–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lowden:1999:SAR

[188] B. G. T. Lowden and J. Robinson. A statistical approach to rule selection in semantic query optimisation. Lecture Notes in Computer Science, 1609:330– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hacid:1999:RRC

[189] M.-S. Hacid and C. Rigotti. Representing and reasoning on conceptual queries over image databases. Lecture Notes in Computer Science, 1609:340–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1999:KDC

[190] S. Tsumoto. Knowledge discovery in clinical databases — an experiment with rule induction and statistics. Lecture Notes in Computer Science, 1609:349–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michalewicz:1999:RAK

[191] M. Michalewicz, M. A. Klopotek, and S. T. Wierzchon. Reasoning and acquisition of knowledge in a system for hand wound diagnosis and prognosis. Lecture Notes in Computer Science, 1609:358–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wieczorkowska:1999:RST

[192] A. Wieczorkowska. Rough sets as a tool for audio signal classification. Lecture Notes in Computer Science, 1609: 367–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brindle:1999:ERS

[193] D. Brindle and W. Ziarko. Experiments with rough sets approach to speech recognition. Lecture Notes in Computer Science, 1609:376–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:1999:SUM

[194] N. Abe, K. Tanaka, and H. Taki. A system for understanding a mechanical assembly instruction manual using virtual space. Lecture Notes in Computer Science, 1609:385–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarbek:1999:PCD

[195] W. Skarbek. Perceptual convergence of discrete clamped fractal operator. Lecture Notes in Computer Science, 1609: 395–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarbek:1999:MOR

[196] W. Skarbek, A. Pietrowcew, and R. Sikora. Modified Oja-RLS algorithm — stochastic convergence analysis and application for image compression. Lecture Notes in Computer Science, 1609:402–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufman:1999:LIN

[197] K. A. Kaufman and R. S. Michalski. Learning from inconsistent and noisy data: The AQ18 approach. Lecture Notes in Computer Science, 1609:411– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1999:LEG

[198] J. J. Zhang, N. J. Cercone, and H. J. Hamilton. Learning English grapheme segmentation using the iterated version space algorithm. Lecture Notes in Computer Science, 1609:420–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwanska:1999:FAA

[199] L. Iwanska, N. Mata, and K. Kruger. Fully automatic acquisition of taxonomic knowledge from large corpora of texts: Limited-syntax knowledge representation system based on natural language. Lecture Notes in Computer Science, 1609:430–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botana:1999:FMS

[200] F. Botana. A fuzzy measure of similarity for instance-based learning. Lecture Notes in Computer Science, 1609: 439–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Susmaga:1999:CMC

[201] R. Susmaga. Computation of minimal cost reducts. Lecture Notes in Computer Science, 1609:448–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stepaniuk:1999:RSD

[202] J. Stepaniuk. Rough set data mining of diabetes data. Lecture Notes in Computer Science, 1609:457–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michaut:1999:SRF

[203] D. Michaut and P. Baptiste. Selection of a relevant feature subset for induction tasks. Lecture Notes in Computer Science, 1609:466–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmer:1999:DSS

[204] J. Calmer and P. Kullmann. A data structure for subsumption-based tabling in top-down resolution engines for data-intensive logic applications. Lecture Notes in Computer Science, 1609:475–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orgun:1999:MRA

[205] M. A. Orgun. Modelling and reasoning about multi-dimensional information. Lecture Notes in Computer Science, 1609:484–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klopotek:1999:ICB

[206] M. A. Klopotek and S. T. Wierzchon. An interpretation for the conditional belief function in the theory of evidence. Lecture Notes in Computer Science, 1609:494–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sztandera:1999:ANN

[207] L. M. Sztandera, C. Bock, M. Trachtman, and J. Velga. Artificial neural networks aid the design of noncarcinogenic azo dyes. Lecture Notes in Computer Science, 1609:503–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1999:ECWa

[208] K. De Jong. Evolving in a changing world. Lecture Notes in Computer Science, 1609:512–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anglano:1999:HPD

[209] C. Anglano, A. Giordana, and G. Lo Bello. High-performance data mining on networks of workstations. Lecture Notes in Computer Science, 1609:520– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulawka:1999:VEG

[210] J. J. Mulawka, P. Wasiewicz, and K. Pietak. Virus-enhanced genetic algorithms inspired by DNA computing. Lecture Notes in Computer Science, 1609:529–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trojanowski:1999:EAN

[211] K. Trojanowski and Z. Michalewicz. Evolutionary approach to non-stationary optimization tasks. Lecture Notes in Computer Science, 1609:538–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cwik:1999:DFR

[212] J. Cwik, J. Koronacki, and J. M. Zytkow. On discovering functional relationships when observations are noisy: 2D case. Lecture Notes in Computer Science, 1609:547–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peters:1999:ATR

[213] J. F. Peters, L. Han, and S. Ramanna. Approximate time rough software cost decision system: Multicriteria decisionmaking approach. Lecture Notes in Computer Science, 1609:556–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAlmeida:1999:MLM

[214] M. A. De Almeida and S. Matwin. Machine learning method for software quality model building. Lecture Notes in Computer Science, 1609:565–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1999:DGP

[215] S. H. Nguyen. Discovery of generalized patterns. Lecture Notes in Computer Science, 1609:574–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kryszkiewicz:1999:IDI

[216] M. Kryszkiewicz and H. Rybinski. Incomplete database issues for representative association rules. Lecture Notes in Computer Science, 1609:583–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puuronen:1999:DIA

[217] S. Puuronen, V. Terziyan, and A. Tsymbal. A dynamic integration algorithm for an ensemble of classifiers. Lecture Notes in Computer Science, 1609:592–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiol-Roig:1999:UIC

[218] G. Fiol-Roig. UIB-IK: A computer system for decision trees induction. Lecture Notes in Computer Science, 1609: 601–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grzymala-Busse:1999:PBR

[219] J. W. Grzymala-Busse, L. K. Goodwin, and X. Zhang. Preterm birth risk assessed by a new method of classification using selective partial matching. Lecture Notes in Computer Science, 1609:612–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dong:1999:PRI

[220] J. Dong, N. Zhong, and S. Ohsuga. Probabilistic rough induction: The GDT-RS methodology and algorithms. Lecture Notes in Computer Science, 1609:621–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanKemenade:1999:CES

[221] C. H. M. Van Kemenade and J. N. Kok. Cluster evolution strategies for constrained numerical optimization. Lecture Notes in Computer Science, 1609: 630–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mansanne:1999:EAF

[222] F. Mansanne, F. Carrere, A. Ehinger, and M. Schoenauer. Evolutionary algorithms as fitness function debuggers. Lecture Notes in Computer Science, 1609:639–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1999:SRL

[223] M. Schmidt. Solving real-life timetabling problems. Lecture Notes in Computer Science, 1609:648–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1999:MSF

[224] R. Jonsson. On measures of search features. Lecture Notes in Computer Science, 1609:657–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jelonek:1999:GAC

[225] J. Jelonek and M. Komosinski. Genetic algorithms in constructive induction. Lecture Notes in Computer Science, 1609:665–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aardal:1999:MSB

[226] Karen Aardal, Robert E. Bixby, Cor A. J. Hurkens, Arjen K. Lenstra, and Job W. Smeltink. Market split and basis reduction: Towards a solution of the Cornuejols-Dawande instances. Lecture Notes in Computer Science, 1610:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100001.
pdf.

Ageev:1999:AAM

[227] Alexander A. Ageev and Maxim I. Sviridenko. Approximation algorithms for maximum coverage and max cut with given sizes of parts. Lecture Notes in Computer Science, 1610:17– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100017.
pdf.

Ahuja:1999:SCC

[228] Ravindra K. Ahuja, Dorit S. Hochbaum, and James B. Orlin. Solving the convex cost integer dual network flow problem. Lecture Notes in Computer Science, 1610:31–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100031.
pdf.

Amaldi:1999:SSA

[229] Edoardo Amaldi, Marc E. Pfetsch, and Leslie E. Trotter, Jr. Some structural and algorithmic properties of the maximum feasible subsystem problem. Lecture Notes in Computer Science, 1610: 45–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100045.
pdf.

Atamtuerk:1999:VIP

[230] A. Atamtuerk, G. L. Nemhauser, and M. W. P. Savelsbergh. Valid inequalities for problems with additive variable upper bounds. Lecture Notes in Computer Science, 1610:60–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atamturk:1999:VIP

[231] Alper Atamtürk, George L. Nemhauser, and Martin W. P. Savelsbergh. Valid inequalities for problems with additive variable upper bounds. Lecture Notes in Computer Science, 1610:60– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100060.
pdf.

Cai:1999:MMT

[232] Mao cheng Cai, Xiaotie Deng, and Wenan Zang. A min-max theorem on feedback vertex sets. Lecture Notes in Computer Science, 1610:73– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1610/16100073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100073.
                                        pdf.

Caprara:1999:SMV

[233] Alberto Caprara, Matteo Fischetti, and Adam N. Letchford. On the separation of maximally violated mod- k cuts. Lecture Notes in Computer Science, 1610: 87–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100087.
                                        pdf.

Chudak:1999:IAA

[234] Fabián A. Chudak and David P. Williamson. Improved approximation algorithms for capacitated facility location problems. Lecture Notes in Computer Science, 1610:99– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100099.
                                        pdf.

Cunningham:1999:OTC

[235] William H. Cunningham and Lawrence Tang. Optimal 3-Terminal cuts and linear programming. Lecture Notes in Computer Science, 1610:114–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1610/16100114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100114.
pdf.

Cvetkovic:1999:SPM

[236] Dragos Cvetkovic, Mirjana Cangalovic, and Vera Kovacevic-Vujcic. Semidefinite programming methods for the symmetric traveling salesman problem. Lecture Notes in Computer Science, 1610:126–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1610/16100126.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1610/16100126.
pdf.

Eisenbrand:1999:BCR

[237] Friedrich Eisenbrand and Andreas S. Schulz. Bounds on the Chvatal rank of polytopes in the 0/1-cube. Lecture Notes in Computer Science, 1610:137– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100137.
pdf.

Fleischer:1999:UMF

[238] Lisa Fleischer. Universally maximum flow with piecewise-constant capacities. Lecture Notes in Computer Science, 1610:151–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1610/16100151.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100151.
                                        pdf.

Fonlupt:1999:CEP

[239] Jean Fonlupt and Ali Ridha Mahjoub. Critical extreme points of the 2-Edge connected spanning subgraph polytope. Lecture Notes in Computer Science, 1610:166–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1610/16100166.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1610/16100166.pdf.

Frank:1999:OTP

[240] András Frank, Tibor Jordán, and Zoltán Szigeti. An orientation theorem with parity conditions. Lecture Notes in Computer Science, 1610:183– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100183.
                                        pdf.

Frank:1999:PCE

[241] András Frank and Zoltán Király. Parity constrained k-edge-connected orientations. Lecture Notes in Computer Science, 1610:191–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1610/16100191.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1610/16100191.pdf.

Halperin:1999:AAM

[242] Eran Halperin and Uri Zwick. Approximation algorithms for MAX 4-SAT and rounding procedures for semidefinite programs. Lecture Notes in Computer Science, 1610:202–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100202.
pdf.

Hartmann:1999:CRC

[243] Mark Hartmann, Maurice Queyranne, and Yaoguang Wang. On the chvátal rank of certain inequalities. Lecture Notes in Computer Science, 1610:218– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100218.
pdf.

Hartvigsen:1999:SFF

[244] David Hartvigsen. The square-free 2-Factor problem in bipartite graphs. Lecture Notes in Computer Science, 1610:234–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1610/16100234.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/1610/16100234.
                                        pdf.

Helmberg:1999:CA

[245] Christoph Helmberg. The m-cost ATSP. Lecture Notes in Computer Science, 1610:242–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1610/16100242.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1610/16100242.pdf.

Iwata:1999:SPC

[246] Satoru Iwata, S. Thomas McCormick, and Maiko Shigeno. A strongly polynomial cut canceling algorithm for the submodular flow problem. Lecture Notes in Computer Science, 1610:259– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100259.
                                        pdf.

Jordan:1999:ESP

[247] Tibor Jordán. Edge-splitting problems with demands. Lecture Notes in Computer Science, 1610:273–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100273.
                                        pdf.

Kashiwabara:1999:IPA

[248] Kenji Kashiwabara, Masataka Nakamura, and Takashi Takabatake. Integral polyhedra associated with certain submodular functions defined on 012-vectors. Lecture Notes in Computer Science, 1610:289–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100289.
pdf.

Klau:1999:OCO

[249] Gunnar W. Klau and Petra Mutzel. Optimal compaction of orthogonal grid drawings. Lecture Notes in Computer Science, 1610:304–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100304.
pdf.

Klein:1999:NID

[250] Philip Klein and Neal Young. On the number of iterations for Dantzig-Wolfe optimization and packing-covering approximation algorithms. Lecture Notes in Computer Science, 1610:320–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100320.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1610/16100320.
                                        pdf.

Kolliopoulos:1999:EEA

[251] Stavros G. Kolliopoulos and Clifford Stein. Experimental evaluation of approximation algorithms for singlesource unsplittable flow. Lecture Notes in Computer Science, 1610:328– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100328.
                                        pdf.

Melkonian:1999:AAD

[252] Vardges Melkonian and Éva Tardos. Approximation algorithms for a directed network design problem. Lecture Notes in Computer Science, 1610: 345–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100345.
                                        pdf.

Mutzel:1999:OAC

[253] Petra Mutzel and René Weiskircher. Optimizing over all combinatorial embeddings of a planar graph. Lecture Notes in Computer Science, 1610:361– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100361.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1610/16100361.
pdf.

Nagamochi:1999:FACa

[254] Hiroshi Nagamochi and Toshihide Ibaraki. A fast algorithm for computing minimum 3-Way and 4-Way cuts. Lecture Notes in Computer Science, 1610: 377–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100377.
pdf.

Noga:1999:STM

[255] John Noga and Steve Seiden. Scheduling two machines with release times. Lecture Notes in Computer Science, 1610:391–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1610/16100391.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1610/16100391.
pdf.

Sebo:1999:IEL

[256] András Sebö. An introduction to empty lattice simplices. Lecture Notes in Computer Science, 1610:400–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100400.
pdf.

Seboe:1999:IEL

[257] A. Seboe. An introduction to empty lattice simplices. Lecture Notes in Computer Science, 1610:400–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szigeti:1999:OED

[258] Zoltán Szigeti. On optimal eardecompositions of graphs. Lecture Notes in Computer Science, 1610:415– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100415.
                                        pdf.

Teo:1999:GSS

[259] Chung-Piaw Teo, Jay Sethuraman, and Wee-Peng Tan. Gale–Shapley stable marriage problem revisited: Strategic issues and applications. Lecture Notes in Computer Science, 1610:429– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1610/16100429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1610/16100429.
                                        pdf.

Uchoa:1999:VDP

[260] Eduardo Uchoa and Marcus Poggi de Aragão. Vertex-disjoint packing of two Steiner trees: Polyhedra and branch-and-cut. Lecture Notes in Computer Science, 1610:439–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1610/16100439.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1610/16100439.
pdf.

DaSilva:1999:FUR

[261] F. S. C. Da Silva. Formalisations of uncertain reasoning. Lecture Notes in Computer Science, 1611:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jennings:1999:AOSa

[262] N. R. Jennings. Agent-oriented software engineering. Lecture Notes in Computer Science, 1611:4–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osman:1999:UMF

[263] I. H. Osman. A unified-metaheuristic framework. Lecture Notes in Computer Science, 1611:11–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:FKR

[264] S.-L. Wang and Y.-H. Wu. A fuzzy knowledge representation and acquisition scheme for diagnostic systems. Lecture Notes in Computer Science, 1611:13–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1999:MFS

[265] G. Y. C. Wong and H. W. Chun. Modeling fuzzy sets using object-oriented techniques. Lecture Notes in Computer Science, 1611:23–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zreak:1999:FAM

[266] H. Zreak, M. Alwan, and M. Khaddour. A fuzzy approach to map building. Lecture Notes in Computer Science, 1611:33–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholczyk:1999:TAI

[267] D. Pacholczyk. Towards an affirmative interpretation of linguistically denied fuzzy properties in knowledge-based systems. Lecture Notes in Computer Science, 1611:43–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sendelj:1999:RHF

[268] R. Sendelj and V. Devedzic. Representational hierarchy of fuzzy logic concepts in the OBOA model. Lecture Notes in Computer Science, 1611:54– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fan:1999:DFS

[269] H.-K. Fan, C. Chen, and C.-M. Hong. Design of fuzzy sliding controller based on cerebellar learning model. Lecture Notes in Computer Science, 1611:64– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1999:NFF

[270] T.-P. Hong and T.-T. Wang. A new fuzzy flexible flow-shop algorithm for continuous fuzzy domain. Lecture Notes in Computer Science, 1611:74– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dakroury:1999:TPN

[271] Y. Dakroury, A. A. El Al, and O. Badr. Testing the performance of a neural networks-based adaptive call admission controller with multimedia traffic. Lecture Notes in Computer Science, 1611:80–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manrique:1999:CNN

[272] D. Manrique, J. Rios, and A. Vilarrasa. A combined neural network and mathematical morphology method for computerized detection of microcalcifications. Lecture Notes in Computer Science, 1611:92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez:1999:SWI

[273] A. Martinez, J. Castellanos, C. Hernandez, and F. De Mingo. Study of weight importance in neural networks working with colineal variables in regression problems. Lecture Notes in Computer Science, 1611:101–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murray:1999:PPA

[274] G. Murray, T. Hendtlass, and J. Podlena. The parallel path artificial micronet. Lecture Notes in Computer Science, 1611:111–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ezzat:1999:GMP

[275] A. Ezzat, N. Inuzuka, and H. Itoh. A generation method to produce GA with GP capabilities for signal modeling. Lecture Notes in Computer Science, 1611:118–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocha:1999:PPC

[276] M. Rocha and J. Neves. Preventing premature convergence to local optima in genetic algorithms via random offspring generation. Lecture Notes in Computer Science, 1611:127–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Beltagy:1999:USO

[277] M. El-Beltagy and A. Keane. Using self organizing maps and genetic algorithms for model selection in multilevel optimization. Lecture Notes in Computer Science, 1611:137–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guillet:1999:GAV

[278] F. Guillet, P. Kuntz, and R. Lehn. A genetic algorithm for visualizing networks of association rules. Lecture Notes in Computer Science, 1611:145– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shazely:1999:GAS

[279] S. Shazely, H. Baraka, A. Abdel-Wahab, and H. Kamal. Genetic algorithms in solving graph partitioning problem. Lecture Notes in Computer Science, 1611:155–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaleh:1999:AGA

[280] A. I. Baaleh and A. F. Sakr. Application of genetic algorithms in power system stabilizer design. Lecture Notes in Computer Science, 1611:165–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mansilla:1999:SGC

[281] E. B. Mansilla, A. Mekaouche, and J. M. G. Guiu. A study of a genetic classifier system based on the pittsburgh approach on a medical domain. Lecture Notes in Computer Science, 1611:175–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:NGB

[282] C.-M. Chen, C.-M. Hong, and H. Fan. A new gradient-based search method: Grey-gradient search method. Lecture Notes in Computer Science, 1611:185– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piriyakumar:1999:IAM

[283] D. A. L. Piriyakumar and P. Levi. An improvised A∗ algorithm for mobile robots to find the optimal path in an unknown environment with minimized search efforts. Lecture Notes in Computer Science, 1611:195–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appas:1999:SVS

[284] A. R. Appas, A. M. Darwish, A. El-Dessouki, and S. I. Shaheen. Speeding the vector search algorithm for regional color channel features based indexing and retrieval systems. Lecture Notes in Computer Science, 1611:205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kato:1999:CBA

[285] S. Kato, S. Oono, H. Seki, and H. Itoh. Cost-based abduction using binary decision diagrams. Lecture Notes in Computer Science, 1611:215–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wotawa:1999:NDD

[286] F. Wotawa. New directions in debugging hardware designs. Lecture Notes in Computer Science, 1611:226– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ligozat:1999:RDS

[287] G. Ligozat. Reasoning with diagrams: The semantics of arrows. Lecture Notes in Computer Science, 1611:236– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milde:1999:MRW

[288] H. Milde, L. Hotz, J. Kahl, and B. Neumann. MAD: A real world application of qualitative model-based decision tree generation for diagnosis. Lecture Notes in Computer Science, 1611:246–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHarmelen:1999:WKB

[289] F. Van Harmelen and J. Van der Meer. WebMaster: Knowledge-based verification of Web-pages. Lecture Notes in Computer Science, 1611:256–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peter:1999:TTO

[290] G. Peter and D. Roesner. Towards task-oriented user support for failure mode and effects analysis. Lecture Notes in Computer Science, 1611:266– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chander:1999:IIE

[291] P. G. Chander, R. Shinghal, and T. Radhakrishnan. Incremental and integrated evaluation of rule-based systems. Lecture Notes in Computer Science, 1611:276–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibrahim:1999:MRD

[292] M. H. Ibrahim and F. A. Cummins. A model of reflection for developing AI applications. Lecture Notes in Computer Science, 1611:286–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1999:CPC

[293] C. M. Jonker and J. Treur. A compositional process control model and its application to biochemical processes. Lecture Notes in Computer Science, 1611:296–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1999:VTK

[294] C. M. Jonker, R. Kremer, P. Van Leeuwen, and D. Pan. Visual and textual knowledge representation in DE-SIRE. Lecture Notes in Computer Science, 1611:306–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1999:VKB

[295] J. Santos, L. Faria, C. Ramos, and Z. A. Vale. Verification of knowledge based-systems for power system control centers. Lecture Notes in Computer Science, 1611:316–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kundu:1999:SCA

[296] M. K. Kundu and N. R. Pal. Selfcrossover and its application to the traveling salesman problem. Lecture Notes in Computer Science, 1611:326– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malheiro:1999:ECS

[297] N. Malheiro, Z. A. Vale, C. Ramos, and J. Santos. Enabling client-server explanation facilities in a real time expert system. Lecture Notes in Computer Science, 1611:333–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froehlich:1999:UEL

[298] P. Froehlich, W. Nejdl, M. Schroeder, and C. Damasio. Using extended logic programming for alarm-correlation in cellular phone networks. Lecture Notes in Computer Science, 1611:343–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terziyan:1999:KAB

[299] V. Y. Terziyan and S. Puuronen. Knowledge acquisition based on semantic balance of internal and external knowledge. Lecture Notes in Computer Science, 1611:353–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1999:ACA

[300] R. N. Smith and W. J. Frawley. Alex: A computer aid for treating alexithymia. Lecture Notes in Computer Science, 1611:362–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kounalis:1999:MPI

[301] E. Kounalis and P. Urso. Mechanizing proofs of integrity constraints in the situation calculus. Lecture Notes in Computer Science, 1611:372–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saward:1999:UCP

[302] G. Saward. Using cases for process modelling: An example from the water supply industry. Lecture Notes in Computer Science, 1611:382–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiemeur:1999:FAC

[303] E. Aiemeur and K. Boudina. Financial analysis by case based reasoning. Lecture Notes in Computer Science, 1611:388–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belanger:1999:TLC

[304] S. Belanger, M. A. Thibodeau, and E. Aiemeur. Training of the learner in criminal law by case-based reasoning. Lecture Notes in Computer Science, 1611:398–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ooi:1999:CBA

[305] B. H. Ooi and A. K. Ghose. Constraintbased agent specification for a multiagent stock brokering system. Lecture Notes in Computer Science, 1611:409– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piccardi:1999:CLD

[306] C. Piccardi and F. Wotawa. A communication language and the design of a diagnosis agent — towards a framework for mobile diagnosis agents. Lecture Notes in Computer Science, 1611: 420–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1999:IBA

[307] C. M. Jonker and J. Treur. Information brokering agents in intelligent Websites. Lecture Notes in Computer Science, 1611:430–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavrilova:1999:LMA

[308] T. Gavrilova, A. Voinov, and I. Lescheva. Learner-model approach to multi-agent intelligent distance learning system for program testing. Lecture Notes in Computer Science, 1611:440–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1999:MAS

[309] B. Bauer and C. Klein. A multiagent solution for advanced call centers. Lecture Notes in Computer Science, 1611:450–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuena:1999:MAS

[310] J. Cuena and M. Molina. A multi-agent system for emergency management in floods. Lecture Notes in Computer Science, 1611:460–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barber:1999:PSF

[311] K. S. Barber, A. Goel, D. Han, and J. Kim. Problem-solving frameworks for sensible agents in an electronic market. Lecture Notes in Computer Science, 1611:470–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bel:1999:MDM

[312] C. Bel and W. Van Stokkum. A model for distributed multi-agent traffic control. Lecture Notes in Computer Science, 1611:480–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustavsson:1999:SEC

[313] R. Gustavsson. Smart equipment creating new electronic business opportunities. Lecture Notes in Computer Science, 1611:490–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aknine:1999:AOC

[314] S. Aknine and S. Pinson. Agent oriented conceptual modeling of parallel workflow systems. Lecture Notes in Computer Science, 1611:500–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deugo:1999:DAM

[315] D. Deugo. A distributed algorithm as mobile agents. Lecture Notes in Computer Science, 1611:510–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaSilva:1999:WOE

[316] F. S. C. Da Silva, W. Weber Vasconcelos, J. Agusti, and D. Robertson. Why ontologies are not enough for knowledge sharing. Lecture Notes in Computer Science, 1611:520–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mostafa:1999:NAD

[317] K. Mostafa, S. I. Shaheen, A. M. Darwish, and I. Farag. A novel approach for detecting and correcting segmentation and recognition errors in Arabic OCR systems. Lecture Notes in Computer Science, 1611:530–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ragab:1999:FRU

[318] M. E. Ragab, A. M. Darwish, E. M. Abed, and S. I. Shaheen. Face recognition using principal component analysis applied to an Egyptian face database. Lecture Notes in Computer Science, 1611:540–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verikas:1999:ULU

[319] A. Verikas, A. Gelzinis, and K. Malmqvist. Using labeled and unlabeled data to train a multilayer perceptron for color classification in graphic arts. Lecture Notes in Computer Science, 1611:550– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodrigues:1999:NCV

[320] M. A. Rodrigues and Y. Liu. A novel 3D-2D computer vision algorithm for automatic inspection of filter components. Lecture Notes in Computer Science, 1611:560–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zangrouba:1999:RUA

[321] E. Zangrouba, T. Hedidar, and A. Jaoua. A robust and unified algorithm for indoor and outdoor scenes based on region segmentation. Lecture Notes in Computer Science, 1611:570– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramparany:1999:SMB

[322] F. Ramparany and F. Sandt. SAPIA: A model based satellite image interpretation system. Lecture Notes in Computer Science, 1611:580–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shah:1999:HMI

[323] S. Shah and J. K. Aggarwal. Hierarchical multifeature integration for automatic object recognition in forward looking infrared images. Lecture Notes in Computer Science, 1611:590– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakas:1999:ACS

[324] A. C. Kakas and A. Michael. Air-crew scheduling through abduction. Lecture Notes in Computer Science, 1611:600– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salem:1999:UFI

[325] R. Salem and K. Bsaiees. Unfold/ fold inductive proof: An extension of a folding technique. Lecture Notes in Computer Science, 1611:612–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oommen:1999:SCA

[326] B. J. Oommen and T. D. Roberts. On solving the capacity assignment problem using continuous learning automata. Lecture Notes in Computer Science, 1611:622–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Link:1999:SPO

[327] M. Link and M. Ishitobi. Supervised parameter optimization of a modular machine learning system. Lecture Notes in Computer Science, 1611:632– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sellami:1999:NSR

[328] H. M. Sellami and A. Jaoua. Nonsupervised rectangular classification of binary data. Lecture Notes in Computer Science, 1611:642–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamdi:1999:SIB

[329] M. S. Hamdi and K. Kaiser. Synthesizing intelligent behavior: A learning paradigm. Lecture Notes in Computer Science, 1611:649–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viharos:1999:AIO

[330] Z. J. Viharos and L. Monostori. Automatic input-output configuration and generation of ANN-based process models and their application in machining. Lecture Notes in Computer Science, 1611:659–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abu-Soud:1999:DDL

[331] S. M. Abu-Soud and M. R. Tolun. DCL: A disjunctive learning algorithm for rule extraction. Lecture Notes in Computer Science, 1611:669–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osmani:1999:MSB

[332] A. Osmani. Modeling and simulating breakdown situation in telecommunication networks. Lecture Notes in Computer Science, 1611:679–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puuronen:1999:HCS

[333] S. Puuronen and H. Kaikova. Handling context-sensitive temporal knowledge from multiple differently ranked sources. Lecture Notes in Computer Science, 1611:688–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osmani:1999:IRA

[334] A. Osmani. Introduction to reasoning about cyclic intervals. Lecture Notes in Computer Science, 1611:698–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guesgen:1999:MPD

[335] H. W. Guesgen and D. Mitra. A multiple-platform decentralized route finding system. Lecture Notes in Computer Science, 1611:707–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamilton:1999:HSA

[336] H. Hamilton and D. J. Randall. Heuristic selection of aggregated-temporal data for knowledge discovery. Lecture Notes in Computer Science, 1611:714– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chitchian:1999:AKR

[337] D. Chitchian, S. Sariyildiz, and H. Koppelaar. Architectural knowledge representation using the Galois lattice technique. Lecture Notes in Computer Science, 1611:724–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capkovic:1999:ASD

[338] F. Capkovic. Automated solving of the DEDS control problems. Lecture Notes in Computer Science, 1611:735– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahl:1999:MEK

[339] J. Kahl, L. Hotz, H. Milde, and S. Wessel. A more efficient knowledge representation for Allen’s algebra and point algebra. Lecture Notes in Computer Science, 1611:747–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraka:1999:UCR

[340] H. H. Baraka, S. Eid, H. Kamal, and A. H. AbdelWahab. Unified chromosome representation for large scale problems. Lecture Notes in Computer Science, 1611:753–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hefny:1999:NFH

[341] H. A. Hefny, A. Abdel Wahab, A. A. Bahnasawi, and S. I. Shaheen. A novel framework for hybrid intelligent systems. Lecture Notes in Computer Science, 1611:761–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chun:1999:DMT

[342] H. W. Chun and S. H. C. Chan. The design of a multi-tiered bus timetabling system. Lecture Notes in Computer Science, 1611:771–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBael:1999:SPU

[343] P. Van Bael and M. Rijckaert. Scheduling of a production unit via critical block neighborhood structures. Lecture Notes in Computer Science, 1611:780– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomii:1999:ASS

[344] N. Tomii, L. J. Zhou, and N. Fukumura. An algorithm for station shunting scheduling problems combining probabilistic local search and PERT. Lecture Notes in Computer Science, 1611:788–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Popovic:1999:DIT

[345] D. Popovic and V. Devedzic. Designing an intelligent tutoring systems in the domain of formal languages. Lecture Notes in Computer Science, 1611:798–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Sheikh:1999:LTS

[346] E. El-Sheikh and J. Sticklen. Leveraging a task-specific approach for intelligent tutoring system generation: Comparing the generic tasks and KADS frameworks. Lecture Notes in Computer Science, 1611:809–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaefer:1999:OOR

[347] C. Schaefer and O. Lopez. An objectoriented robot model and its integration into flexible manufacturing systems. Lecture Notes in Computer Science, 1611:820–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gide:1999:EBI

[348] E. Gide and F. Soliman. The economic benefits of Internet-based business operations in manufacturing. Lecture Notes in Computer Science, 1611: 830–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elkeran:1999:CMC

[349] A. Elkeran and M. A. El-Baz. CNC manufacturing of complex surfaces based on solid modeling. Lecture Notes in Computer Science, 1611:841–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Fouly:1999:AML

[350] T. El-Fouly, N. Zerhoini, M. Ferney, and A. El Moundi. Analysis of manufacturing lines using a phase space algorithm: Open line case. Lecture Notes in Computer Science, 1611:849–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gargouri:1999:FAL

[351] B. Gargouri, M. Jmaiel, and A. Ben Hamadou. A formal approach to lingware development. Lecture Notes in Computer Science, 1611:859–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herlea:1999:FKL

[352] D. E. Herlea, C. M. Jonker, J. Treur, and N. J. E. Wijngaards. A formal knowledge level process model of requirements engineering. Lecture Notes in Computer Science, 1611:869–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cakic:1999:PMC

[353] J. Cakic and V. Devedzic. Pieces of mind: Component-based software development of artificial intelligence. Lecture Notes in Computer Science, 1611: 879–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Midany:1999:CAP

[354] T. T. El-Midany, A. Elkeran, and H. T. Radwan. CNC automatic programming system for Wire-EDM. Lecture Notes in Computer Science, 1611:889– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1999:DIC

[355] R. Bergmann, S. Breen, M. Goeker, and M. Manago. Developing industrial case-based reasoning applications: The INRECA methodology. Lecture Notes in Computer Science, 1612:i–xx, 1–888, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Insana:1999:ASB

[356] Michael F. Insana, Larry T. Cook, and Pawan Chaturvedi. Analytical study of bioelasticity ultrasound systems. Lecture Notes in Computer Science, 1613: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130001.
pdf.

Mosher:1999:MSI

[357] John C. Mosher, Richard M. Leahy, David W. Shattuck, and Sylvain Baillet. MEG source imaging using multipolar expansions. Lecture Notes in Computer Science, 1613:15– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130015.
pdf.

Carvalho:1999:BTT

[358] Bruno M. Carvalho, Gabor T. Herman, Samuel Matej, Claudia Salzberg, and Eilat Vardi. Binary tomography for triplane cardiography. Lecture Notes in Computer Science, 1613:29– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130029.
pdf.

Skrinjar:1999:RTB

[359] Oskar M. Skrinjar and James S. Duncan. Real time 3D brain shift compensation. Lecture Notes in Computer Science, 1613:42–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130042.
pdf.

Romeny:1999:CAH

[360] Bart M. ter Haar Romeny, Bart Titulaer, Stiliyan Kalitzin, Gabriëlle Scheffer, Frank Broekmans, Joes Staal, and Egbert te Velde. Computer assisted human follicle analysis for fertility prospects with 3D ultrasound. Lecture Notes in Computer Science, 1613: 56–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130056.
                                        pdf.

TerHaarRomeny:1999:CAH

[361] B. M. Ter Haar Romeny, B. Titulaer, S. Kalitzin, and G. Scheffer. Computer assisted human follicle analysis for fertility prospects with 3D ultrasound. Lecture Notes in Computer Science, 1613:56–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treece:1999:VMS

[362] Graham Treece, Richard Prager, Andrew Gee, and Laurence Berman. Volume measurement in sequential freehand 3-D ultrasound. Lecture Notes in Computer Science, 1613:70– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130070.
                                        pdf.

Stetten:1999:AIM

[363] George D. Stetten and Stephen M. Pizer. Automated identification and measurement of objects via populations of medial primitives, with application to real time 3D echocardiography. Lecture Notes in Computer Science, 1613: 84–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130084.
pdf.

Nichols:1999:CTD

[364] Thomas E. Nichols, Jinyi Qi, and Richard M. Leahy. Continuous time dynamic PET imaging using list mode data. Lecture Notes in Computer Science, 1613:98–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130098.
pdf.

Guo:1999:HGA

[365] Yanlin Guo and Baba C. Vemuri. Hybrid geometric active models for shape recovery in medical images. Lecture Notes in Computer Science, 1613:112– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130112.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1613/16130112.
                                        pdf.

Lorigo:1999:CDG

[366] Liana M. Lorigo, Olivier Faugeras, W. E. L. Grimson, Renaud Keriven, Ron Kikinis, and Carl-Fredrik Westin. Co-dimension 2 geodesic active contours for MRA segmentation. Lecture Notes in Computer Science, 1613:126– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130126.
                                        pdf.

Pham:1999:AFS

[367] Dzung L. Pham and Jerry L. Prince. An adaptive fuzzy segmentation algorithm for three-dimensional magnetic resonance images. Lecture Notes in Computer Science, 1613:140–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130140.
                                        pdf.

Rey:1999:ADS

[368] David Rey, Gérard Subsol, Hervé Delingette, and Nicholas Ayache. Automatic detection and segmentation of evolving processes in 3D medical images: Application to multiple sclerosis. Lecture Notes in Computer Science, 1613:154–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130154.
pdf.

Chui:1999:RCA

[369] Haili Chui, James Rambo, James Duncan, Robert Schultz, and Anand Rangarajan. Registration of cortical anatomical structures via robust 3D point matching. Lecture Notes in Computer Science, 1613:168–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130168.
pdf.

Vaillant:1999:HMC

[370] Marc Vaillant and Christos Davatzikos. Hierarchical matching of cortical features for deformable brain image registration. Lecture Notes in Computer Science, 1613:182–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130182.
pdf.

Caunce:1999:ULG

[371] A. Caunce and C. J. Taylor. Using local geometry to build 3D sulcal models. Lecture Notes in Computer Science, 1613:196–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1613/16130196.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130196.
                                        pdf.

Collins:1999:AIC

[372] D. Louis Collins, Alex P. Zijdenbos, Wim F. C. Baaré, and Alan C. Evans. ANIMAL+INSECT: Improved cortical structure segmentation. Lecture Notes in Computer Science, 1613:210– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130210.
                                        pdf.

Christensen:1999:CLE

[373] Gary E. Christensen. Consistent linear-elastic transformations for image matching. Lecture Notes in Computer Science, 1613:224–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130224.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130224.
                                        pdf.

Lester:1999:NLR

[374] Hava Lester, Simon R. Arridge, Kalvis M. Jansons, Louis Lemieux, Joseph V. Hajnal, and Anjela Oatridge. Non-linear registration with the variable viscosity fluid algorithm. Lecture Notes in Computer Science, 1613:238– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130238.
pdf.

Rohr:1999:ATP

[375] Karl Rohr, Mike Fornefett, and H. Siegfried Stiehl. Approximating thin-plate splines for elastic registration: Integration of landmark errors and orientation attributes. Lecture Notes in Computer Science, 1613:252– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130252.
pdf.

Laading:1999:HFB

[376] Jacob K. Laading, Colin McCulloch, Valen E. Johnson, David R. Gilland, and Ronald J. Jaszczak. A hierarchical feature based deformation model applied to 4D cardiac SPECT data. Lecture Notes in Computer Science, 1613: 266–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130266.
pdf.

Karssemeijer:1999:LOD

[377] Nico Karssemeijer. Local orientation distribution as a function of spatial scale for detection of masses in mammograms. Lecture Notes in Computer Science, 1613:280–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130280.
                                        pdf.

Kruggel:1999:POM

[378] Frithjof Kruggel and D. Yves von Cramon. Physiologically oriented models of the hemodynamic response in functional MRI. Lecture Notes in Computer Science, 1613:294–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130294.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130294.
                                        pdf.

Bullitt:1999:GDI

[379] Elizabeth Bullitt, Stephen Aylward, Alan Liu, Jeffrey Stone, Suresh K. Mukherji, Chris Coffey, Guido Gerig, and Stephen M. Pizer. 3D graph description of the intracerebral vasculature from segmented MRA and tests of accuracy by comparison with X-ray angiograms. Lecture Notes in Computer Science, 1613:308–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1613/16130308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130308.
pdf.

Cootes:1999:UFA

[380] T. F. Cootes, C. Beeston, G. J. Edwards, and C. J. Taylor. A unified framework for atlas matching using active appearance models. Lecture Notes in Computer Science, 1613:322– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130322.
pdf.

Goris:1999:IMA

[381] Michael L. Goris and Robert L. Van Uitert, Jr. An integral method for the analysis of wall motion in gated myocardial SPECT studies. Lecture Notes in Computer Science, 1613:334– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130334.
pdf.

Niessen:1999:EAV

[382] Wiro Niessen, Alexander Montauban van Swijndregt, Bernard Elsman, Onno Wink, Max Viergever, and Willem Mali. Enhanced artery visualization in blood pool MRA: Results in the peripheral vasculature. Lecture Notes in Computer Science, 1613:340– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130340.
pdf.

Huang:1999:FDL

[383] Jiantao Huang and Amir A. Amini. Four-dimensional LV tissue tracking from tagged MRI with a 4D B-spline model. Lecture Notes in Computer Science, 1613:346–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130346.
pdf.

Papademetris:1999:RST

[384] Xenophon Papademetris, Pengcheng Shi, Donald P. Dione, Albert J. Sinusas, R. Todd Constable, and James S. Duncan. Recovery of soft tissue object deformation from 3D image sequences using biomechanical models. Lecture Notes in Computer Science, 1613:352–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130352.
pdf.

Klein:1999:FDP

[385] Gregory J. Klein. Forward deformation of PET volumes using non-uniform elastic material constraints. Lecture Notes in Computer Science, 1613:358– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130358.
pdf.

Styner:1999:BMD

[386] Martin Styner, Thomas Coradi, and Guido Gerig. Brain morphometry by distance measurement in a non-Euclidean, curvilinear space. Lecture Notes in Computer Science, 1613:364– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130364.
pdf.

Duta:1999:LSM

[387] Nicolae Duta, Milan Sonka, and Anil K. Jain. Learning shape models from examples using automatic shape clustering and Procrustes analysis. Lecture Notes in Computer Science, 1613: 370–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130370.
pdf.

Brett:1999:FAL

[388] A. D. Brett and C. J. Taylor. A framework for automated landmark generation for automated 3D statistical model construction. Lecture Notes in Computer Science, 1613:376– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130376.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130376.
                                        pdf.

Golland:1999:SSA

[389] Polina Golland, W. Eric L. Grimson, and Ron Kikinis. Statistical shape analysis using fixed topology skeletons: Corpus callosum study. Lecture Notes in Computer Science, 1613:382– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130382.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130382.
                                        pdf.

Leventon:1999:MGM

[390] Michael E. Leventon and Sarah F. F. Gibson. Model generation from multiple volumes using constrained elastic SurfaceNets. Lecture Notes in Computer Science, 1613:388–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130388.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1613/16130388.
pdf.

Olabarriaga:1999:IIS

[391] S. D. Olabarriaga, A. W. M. Smeulders, A. C. A. Marijnissen, and K. L. Vincken. An intelligent interactive segmentation method for the joint space in osteoarthritic ankles. Lecture Notes in Computer Science, 1613:394– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130394.
pdf.

Lelieveldt:1999:AMF

[392] Boudewijn P. F. Lelieveldt, Milan Sonka, Lizann Bolinger, Thomas D. Scholz, Hein W. M. Kayser, Rob J. van der Geest, and Johan H. C. Reiber. Anatomical modeling with fuzzy implicit surfaces: Application to automated localization of the heart and lungs in thoracic MR images. Lecture Notes in Computer Science, 1613:400– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130400.
pdf.

Zwiggelaar:1999:DCM

[393] Reyer Zwiggelaar, Christopher J. Taylor, and Caroline M. E. Rubin. Detection of the central mass of spiculated lesions — signature normalisation and model data aspects. Lecture Notes in Computer Science, 1613:406– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130406.
pdf.

Csapodi:1999:NEM

[394] Márton Csapodi, Ágota Petrányi, György Liszka, Ákos Zarándy, and Tamás Roska. Noise estimation and measures for detection of clustered microcalcifications. Lecture Notes in Computer Science, 1613:412–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130412.
pdf.

Diaz:1999:MSH

[395] M. E. Díaz and G. Ayala. Measuring the spatial homogeneity in corneal endotheliums by means of a randomization test. Lecture Notes in Computer Science, 1613:418–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130418.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130418.
pdf.

Horng:1999:ACL

[396] M.-H. Horng, Y.-N. Sun, and X.- Z. Lin. The assessment of chronic liver diseases by sonography. Lecture Notes in Computer Science, 1613:424– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130424.
pdf.

Hartmann:1999:ACB

[397] S. L. Hartmann, M. H. Parks, H. Schlack, W. Riddle, R. R. Price, P. R. Martin, and B. M. Dawant. Automatic computation of brain and cerebellum volumes in normal subjects and chronic alcoholics. Lecture Notes in Computer Science, 1613:430– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130430.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130430.
pdf.

Sitek:1999:RSR

[398] Arkadiusz Sitek, Edward V. R. Di Bella, and Grant T. Gullberg. Reconstruction from slow rotation dynamic SPECT using a factor model. Lecture Notes in Computer Science, 1613:436– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130436.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1613/16130436.
                                        pdf.

Buvat:1999:SFA

[399] I. Buvat, S. Hapdey, H. Benali, A. Todd-Pokropek, and R. Di Paola. Spectral factor analysis for multi-isotope imaging in nuclear medicine. Lecture Notes in Computer Science, 1613:442–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130442.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130442.
                                        pdf.

Coulon:1999:SGA

[400] Olivier Coulon, Jean-Francois Mangin, Jean-Baptiste Poline, Vincent Frouin, and Isabelle Bloch. Structural group analysis of functional maps. Lecture Notes in Computer Science, 1613:448– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130448.
                                        pdf.

Studholme:1999:IID

[401] Colin Studholme, R. Todd Constable, and James S. Duncan. Incorporating an image distortion model in non-rigid alignment of EPI with conventional MRI. Lecture Notes in Computer Science, 1613:454–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130454.
pdf.

West:1999:DTR

[402] Jay B. West and J. Michael Fitzpatrick. The distribution of target registration error in rigid-body, pointbased registration. Lecture Notes in Computer Science, 1613:460–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130460.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130460.
pdf.

Meihe:1999:FMI

[403] Xu Meihe, Rajagopalan Srinivasan, and Wieslaw L. Nowinski. A fast mutual information method for multimodal registration. Lecture Notes in Computer Science, 1613:466–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130466.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130466.
pdf.

Xu:1999:FMI

[404] M. Xu, R. Srinivasan, and W. L. Nowinski. A fast mutual information method for multi-modal registration. Lecture Notes in Computer Science, 1613:466–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holden:1999:VSM

[405] Mark Holden, Derek L. G. Hill, Erika R. E. Denton, Jo M. Jarosz, Tim C. S. Cox, and David J. Hawkes. Voxel similarity measures for 3D serial MR brain image registration. Lecture Notes in Computer Science, 1613:472– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130472.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130472.
                                        pdf.

Rohling:1999:RBF

[406] Robert Rohling, Andrew Gee, Laurence Berman, and Graham Treece. Radial basis function interpolation for freehand 3D ultrasound. Lecture Notes in Computer Science, 1613:478– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1613/16130478.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1613/16130478.
                                        pdf.

Parker:1999:NSM

[407] Geoffrey J. M. Parker, Julia A. Schnabel, and Gareth J. Barker. Nonlinear smoothing of MR images using approximate entropy — A local measure of signal intensity irregularity. Lecture Notes in Computer Science, 1613:484– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130484.
pdf.

Nyul:1999:NVM

[408] László G. Nyúl and Jayaram K. Udupa. New variants of a method of MRI scale normalization. Lecture Notes in Computer Science, 1613:490–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130490.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130490.
pdf.

Machado:1999:MEI

[409] Alexei M. C. Machado, Mario F. M. Campos, and James C. Gee. Method for estimating the intensity mapping between MRI images. Lecture Notes in Computer Science, 1613:496–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1613/16130496.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1613/16130496.
pdf.

Leissler:1999:SIR

[410] M. Leissler, M. Hemmje, and E. J. Neuhold. Supporting image-retrieval by database driven interactive 3D information-visualization. Lecture Notes in Computer Science, 1614:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolle:1999:VLI

[411] R. M. Bolle and A. Hampapur. Video libraries: From ingest to distribution. Lecture Notes in Computer Science, 1614:15–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1999:QMD

[412] S.-K. Chang, G. Costagliola, and E. Jungert. Querying multimedia data sources and databases. Lecture Notes in Computer Science, 1614:19–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stanchev:1999:GID

[413] P. L. Stanchev. General image database model. Lecture Notes in Computer Science, 1614:29–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chbeir:1999:SMI

[414] R. Chbeir, Y. Amghar, and A. Flory. System for medical image retrieval: The MIMS model. Lecture Notes in Computer Science, 1614:37–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meddes:1999:ABV

[415] J. Meddes and E. McKenzie. An agentbased visualisation architecture. Lecture Notes in Computer Science, 1614: 43–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferro:1999:ETD

[416] A. Ferro, G. Gallo, and R. Giugno. Error-tolerant database for structured images. Lecture Notes in Computer Science, 1614:51–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soffer:1999:QPO

[417] A. Soffer and H. Samet. Query processing and optimization for pictorial query trees. Lecture Notes in Computer Science, 1614:60–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Porkaew:1999:SSU

[418] K. Porkaew, S. Mehrotra, M. Ortega, and K. Chakrabarti. Similarity search using multiple examples in MARS. Lecture Notes in Computer Science, 1614:68–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Androutsos:1999:ESC

[419] D. Androutsos, K. N. Plataniotis, and A. N. Venetsanopoulos. Excluding specified colors from image queries using a multidimensional query space. Lecture Notes in Computer Science, 1614:76–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindley:1999:GVI

[420] C. A. Lindley and A.-M. Vercoustre. Generic viewer interaction semantics for dynamic virtual video synthesis. Lecture Notes in Computer Science, 1614:83–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shiohara:1999:COA

[421] H. Shiohara, Y. Iizuka, T. Maruyama, and S. Isobe. Category oriented analysis for visual data mining. Lecture Notes in Computer Science, 1614:91– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ikonomakis:1999:UIR

[422] N. Ikonomakis, K. N. Plataniotis, and A. N. Venetsanopoulos. User interaction in region-based color image segmentation. Lecture Notes in Computer Science, 1614:99–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciocca:1999:URF

[423] G. Ciocca and R. Schettini. Using a relevance feedback mechanism to improve content-based image retrieval. Lecture Notes in Computer Science, 1614:107–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malki:1999:RQS

[424] J. Malki, N. Boujemaa, C. Nastar, and A. Winter. Region queries without segmentation for image retrieval by content. Lecture Notes in Computer Science, 1614:115–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiSciascio:1999:CBI

[425] E. Di Sciascio, G. Mingolla, and M. Mongiello. Content-based image retrieval over the Web using query by sketch and relevance feedback. Lecture Notes in Computer Science, 1614:123– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buijs:1999:VLS

[426] J. M. Buijs and M. S. Lew. Visual learning of simple semantics in ImageScape. Lecture Notes in Computer Science, 1614:131–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hibino:1999:TAI

[427] S. L. Hibino. Task analysis for information visualization. Lecture Notes in Computer Science, 1614:139–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vendrig:1999:FIB

[428] J. Vendrig, M. Worring, and A. W. M. Smeulders. Filter image browsing: Exploiting interaction in image retrieval. Lecture Notes in Computer Science, 1614:147–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiroike:1999:VIS

[429] A. Hiroike, Y. Musha, A. Sugimoto, and Y. Mori. Visualization of information spaces to retrieve and browse image data. Lecture Notes in Computer Science, 1614:155–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helfman:1999:MAU

[430] J. I. Helfman. Mandala: An architecture for using images to access and organize Web information. Lecture Notes in Computer Science, 1614:163– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldi:1999:CRO

[431] C. Baldi, C. Colombo, and A. Del Bimbo. A compact and retrievaloriented video representation using mosaics. Lecture Notes in Computer Science, 1614:171–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paquet:1999:CIR

[432] E. Paquet and M. Rioux. Crawling, indexing and retrieval of threedimensional data on the Web in the framework of MPEG-7. Lecture Notes in Computer Science, 1614:179–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohm:1999:VSE

[433] J.-R. Ohm, F. Bunjamin, W. Liebsch, and B. Makai. A visual search engine for distributed image and video database retrieval applications. Lecture Notes in Computer Science, 1614:187– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberman:1999:IMI

[434] B. Eberman, B. Fidler, R. Iannucci, and C. Joerg. Indexing multimedia for the Internet. Lecture Notes in Computer Science, 1614:195–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1999:CIW

[435] J. Cho and S. Mukherjea. Crawling for images on the WWW. Lecture Notes in Computer Science, 1614:203–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konstantinou:1999:DJB

[436] V. Konstantinou and A. Psarrou. A dynamic JAVA-based intelligent interface for online image database searches. Lecture Notes in Computer Science, 1614:211–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fablet:1999:MBF

[437] R. Fablet and P. Bouthemy. Motionbased feature extraction and ascendant hierarchical classification for video indexing and retrieval. Lecture Notes in Computer Science, 1614:221–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanjalic:1999:ASM

[438] A. Hanjalic, R. L. Lagendijk, and J. Biemond. Automatically segmenting movies into logical story units. Lecture Notes in Computer Science, 1614: 229–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1999:LCA

[439] J. M. Sanchez, X. Binefa, J. Vitria, and P. Radeva. Local color analysis for scene break detection applied to TV commercials recognition. Lecture Notes in Computer Science, 1614:237– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouthemy:1999:SSI

[440] P. Bouthemy, C. Garcia, R. Ronfard, and G. Tziritas. Scene segmentation and image feature extraction for video indexing and retrieval. Lecture Notes in Computer Science, 1614:245–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1999:ARC

[441] S. Fischer, I. Rimac, and R. Steinmetz. Automatic recognition of camera zooms. Lecture Notes in Computer Science, 1614:253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelgon:1999:RTM

[442] M. Gelgon, P. Bouthemy, and T. Dubois. A region tracking method with failure detection for an interactive video indexing environment. Lecture Notes in Computer Science, 1614: 261–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhandarkar:1999:IPC

[443] S. M. Bhandarkar, Y. S. Warke, and A. A. Khombhadia. Integrated parsing of compressed video. Lecture Notes in Computer Science, 1614:269–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kong:1999:ISD

[444] W. Kong, X. Ding, H. Lu, and S. Ma. Improvement of shot detection using illumination invariant metric and dynamic threshold selection. Lecture Notes in Computer Science, 1614:277– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardizzone:1999:TSM

[445] E. Ardizzone, C. Lodato, and S. Lopes. Temporal segmentation of MPEG video sequences. Lecture Notes in Computer Science, 1614:283–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1999:DAS

[446] H. B. Lu and Y. J. Zhang. Detecting abrupt scene change using neural network. Lecture Notes in Computer Science, 1614:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasan:1999:MMF

[447] U. Srinivasan and C. Lindley. Multimodal feature-map: An approach to represent digital video sequences. Lecture Notes in Computer Science, 1614: 299–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liao:1999:RTV

[448] M. Liao, Y. Li, S. Ma, and H. Lu. Robust tracking of video objects through topological constraint on homogeneous motion. Lecture Notes in Computer Science, 1614:307–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwerks:1999:SS

[449] G. S. Iwerks and H. Samet. The spatial spreadsheet. Lecture Notes in Computer Science, 1614:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aufure-Portier:1999:HLV

[450] M.-A. Aufure-Portier and C. Bonhomme. A high level visual language for spatial data management. Lecture Notes in Computer Science, 1614:325– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolov:1999:GGM

[451] S. G. Nikolov, D. R. Bull, and C. N. Canagarajah. A global graph model of image registration. Lecture Notes in Computer Science, 1614:333–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aksoy:1999:GTA

[452] S. Aksoy and R. M. Haralick. A graphtheoretic approach to image database retrieval. Lecture Notes in Computer Science, 1614:341–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pan:1999:MCA

[453] C. Pan and S. Ma. Motion capture of arm from a monocular image sequence. Lecture Notes in Computer Science, 1614:349–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gagliardi:1999:CDA

[454] I. Gagliardi and B. Zonta. Comparing dictionaries for the automatic generation of hypertextual links: A case study. Lecture Notes in Computer Science, 1614:358–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1999:CVC

[455] J.-H. Lim. Categorizing visual contents by matching visual “keywords”. Lecture Notes in Computer Science, 1614: 367–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katsumoto:1999:DPL

[456] M. Katsumoto and S.-I. Iisaku. Design of the presentation language for distributed hypermedia system. Lecture Notes in Computer Science, 1614: 375–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rehatschek:1999:GAM

[457] H. Rehatschek and H. Mueller. A generic annotation model for video databases. Lecture Notes in Computer Science, 1614:383–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1999:DIC

[458] H. Yang, H. Kim, and J. Yang. Design and implementation of COIRS (a COncept-Based Image Retrieval System). Lecture Notes in Computer Science, 1614:391–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutanto:1999:AIE

[459] D. Sutanto and C. H. C. Leung. Automatic index expansion for conceptbased image query. Lecture Notes in Computer Science, 1614:399–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tam:1999:SHL

[460] A. M. Tam and C. H. C. Leung. Structured high-level indexing of visual data content. Lecture Notes in Computer Science, 1614:409–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palhang:1999:FEI

[461] M. Palhang and A. Sowmya. Feature extraction: Issues, new features, and symbolic representation. Lecture Notes in Computer Science, 1614:418– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bres:1999:DIP

[462] S. Bres and J.-M Jolion. Detection of interest points for image indexation. Lecture Notes in Computer Science, 1614:427–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferez:1999:HDI

[463] R. Alferez and Y.-F. Wang. Highly discriminative invariant features for image matching. Lecture Notes in Computer Science, 1614:435–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1999:IRU

[464] X. Ding, W. Kong, C. Hu, and S. Ma. Image retrieval using Schwarz representation of one-dimensional feature. Lecture Notes in Computer Science, 1614: 443–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Do:1999:IIR

[465] M. Do, S. Ayer, and M. Vetterli. Invariant image retrieval using wavelet maxima moment. Lecture Notes in Computer Science, 1614:451–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chetverikov:1999:DRS

[466] D. Chetverikov. Detecting regular structures for invariant retrieval. Lecture Notes in Computer Science, 1614: 459–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nes:1999:CIT

[467] N. Nes and M. C. D’Ornellas. Color image texture indexing. Lecture Notes in Computer Science, 1614:467–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahman:1999:IIC

[468] S. M. Rahman, G. C. Karmaker, and R. J. Bignall. Improving image classification using extended run length features. Lecture Notes in Computer Science, 1614:475–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schouten:1999:FEU

[469] B. A. M. Schouten and P. M. De Zeeuw. Feature extraction using fractal codes. Lecture Notes in Computer Science, 1614:483–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuytelaars:1999:CBI

[470] T. Tuytelaars and L. Van Gool. Content-based image retrieval based on local affinely invariant regions. Lecture Notes in Computer Science, 1614:493– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jia:1999:FOB

[471] L. Jia and L. Kitchen. A framework for object-based image retrieval at the semantic level. Lecture Notes in Computer Science, 1614:501–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carson:1999:BSR

[472] C. Carson, M. Thomas, S. Belongie, and J. M. Hellerstein. Blobworld: A system for region-based image indexing and retrieval. Lecture Notes in Computer Science, 1614:509–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maxwell:1999:PBA

[473] B. A. Maxwell. A physics-based approach to interactive segmentation. Lecture Notes in Computer Science, 1614:517–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimai:1999:AEC

[474] A. Dimai. Assessment of effectiveness of content based image retrieval systems. Lecture Notes in Computer Science, 1614:525–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egas:1999:ATV

[475] R. Egas, D. P. Huijsmans, M. Lew, and N. Sebe. Adapting k-d trees to visual retrieval. Lecture Notes in Computer Science, 1614:533–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laaksonen:1999:CBI

[476] J. Laaksonen, M. Koskela, and E. Oja. Content-based image retrieval using self-organizing maps. Lecture Notes in Computer Science, 1614:541–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Squire:1999:RFT

[477] D. Squire, W. Mueller, and H. Mueller. Relevance feedback and term weighting schemes for content-based image retrieval. Lecture Notes in Computer Science, 1614:549–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1999:GAW

[478] D. Y.-M. Chan and I. King. Genetic algorithm for weights assignment in dissimilarity function for trademark retrieval. Lecture Notes in Computer Science, 1614:557–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mokhtarian:1999:RSS

[479] F. Mokhtarian and S. Abbasi. Retrieval of similar shapes under affine transform. Lecture Notes in Computer Science, 1614:566–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vleugels:1999:EIR

[480] J. Vleugels and R. Veltkamp. Efficient image retrieval through vantage objects. Lecture Notes in Computer Science, 1614:575–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schomaker:1999:UPB

[481] L. Schomaker, E. De Leau, and L. Vuurpijl. Using pen-based outlines for object-based annotation and imagebased queries. Lecture Notes in Computer Science, 1614:585–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gevers:1999:IQF

[482] T. Gevers and A. W. M. Smeulders. Interactive query formulation for object search. Lecture Notes in Computer Science, 1614:593–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1999:ADS

[483] L. Liu and S. Sclaroff. Automatic deformable shape segmentation for image database search applications. Lecture Notes in Computer Science, 1614:601– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iannizzotto:1999:MTA

[484] G. Iannizzotto and L. Vita. A multiscale turning angle representation of object shapes for image retrieval. Lecture Notes in Computer Science, 1614: 609–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Latecki:1999:CBS

[485] L. J. Latecki and R. Lakaemper. Contour-based shape similarity. Lecture Notes in Computer Science, 1614: 617–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banfi:1999:CDB

[486] F. Banfi and R. Ingold. Computing dissimilarity between hand drawnsketches and digitized images. Lecture Notes in Computer Science, 1614:625– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeemter:1999:DGP

[487] K. Van Deemter. Document generation and picture retrieval. Lecture Notes in Computer Science, 1614:632–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderHeijden:1999:FJB

[488] G. Van der Heijden, G. Polder, and J. W. Van Eck. FLORES: A JAVA based image database for ornamentals. Lecture Notes in Computer Science, 1614:641–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saraceno:1999:PPI

[489] C. Saraceno, M. Reiter, P. Kammerer, and E. Zolda. Pictorial portrait indexing using view-based eigeneyes. Lecture Notes in Computer Science, 1614:649–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeong:1999:IRU

[490] S. H. Jeong, J. D. Yang, H. J. Yang, and J. H. Choi. Image retrieval using fuzzy triples. Lecture Notes in Computer Science, 1614:657–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sahni:1999:VBL

[491] S. Sahni, B. C. Vemuri, F. Chen, and C. Kapoor. Variable-bit-length coding: An effective coding method. Lecture Notes in Computer Science, 1614:665– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:BCF

[492] Z. Wang, Z. Chi, D. Deng, and Y. Yu. Block-constrained fractal coding scheme for image retrieval. Lecture Notes in Computer Science, 1614: 673–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:EAL

[493] F. Chen, S. Sahni, and B. C. Vemuri. Efficient algorithms for lossless compression of 2D/3D images. Lecture Notes in Computer Science, 1614:681– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pingali:1999:LVS

[494] C. S. Pingali, Y. Jean, and I. Carlbom. Lucent Vision: A system for enhanced sports viewing. Lecture Notes in Computer Science, 1614:689–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraile:1999:BMV

[495] R. Fraile and S. J. Maybank. Building 3D models of vehicles for computer vision. Lecture Notes in Computer Science, 1614:697–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biancardi:1999:IAI

[496] A. Biancardi and V. Moccia. Integrating applications into interactive virtual environments. Lecture Notes in Computer Science, 1614:703–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huet:1999:SSL

[497] B. Huet and E. R. Hancock. Structural sensitivity for large-scale linepattern recognition. Lecture Notes in Computer Science, 1614:711–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhonsle:1999:CVA

[498] S. Bhonsle, A. Gupta, S. Santini, and M. Worring. Complex visual activity recognition using a temporally ordered database. Lecture Notes in Computer Science, 1614:719–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santini:1999:IDA

[499] S. Santini, M. Worring, E. Hunter, and V. Kouznetsova. Image database assisted classification. Lecture Notes in Computer Science, 1614:727–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1999:VPS

[500] U. Xu, J. Wu, and S. Ma. Visual processing system for facial prediction. Lecture Notes in Computer Science, 1614:735–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Androutsos:1999:SIS

[501] P. Androutsos, H. E. Ruda, and A. N. Venetsanopoulos. Semi-interactive structure and fault analysis of (111)7x7 silicon micrographs. Lecture Notes in Computer Science, 1614:745–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huele:1999:UWT

[502] R. Huele and J. N. Ciano. Using wavelet transforms to match photographs of individual sperm whales identified by contour of the trailing edge of the fluke. Lecture Notes in Computer Science, 1614:753–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stiefelhagen:1999:GFA

[503] R. Stiefelhagen, M. Finke, J. Yang, and A. Waibel. From gaze to focus of attention. Lecture Notes in Computer Science, 1614:761–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frederix:1999:AIB

[504] G. Frederix and E. J. Pauwels. Automatic interpretation based on robust segmentation and shape-extraction. Lecture Notes in Computer Science, 1614:769–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1999:PFE

[505] S. C. Y. Chan and P. H. Lewis. A pre-filter enabling fast frontal face detection. Lecture Notes in Computer Science, 1614:777–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demetrescu:1999:TGG

[506] C. Demetrescu and I. Finocchi. A technique for generating graphical abstractions of program data structures. Lecture Notes in Computer Science, 1614: 785–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1999:VPM

[507] R. Moreno and R. E. Mayer. Visual presentations in multimedia learning: Conditions that overload visual working memory. Lecture Notes in Computer Science, 1614:793–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahabi:1999:VSN

[508] C. Shahabi, A. E. Dashti, C. Burns, and S. Ghandeharizadeh. Visualization of spatial neuroanatomical data. Lecture Notes in Computer Science, 1614: 801–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erie:1999:VCP

[509] M. C. Erie, C. H. Chu, and R. D. Sidman. Visualization of the cortical potential field by medical imaging data fusion. Lecture Notes in Computer Science, 1614:809–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janecek:1999:AVR

[510] P. Janecek. Applying visualization research towards design. Lecture Notes in Computer Science, 1614:817–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bell:1999:SDL

[511] G. Bell. Supercomputing !D looking ahead. Lecture Notes in Computer Science, 1615:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lonsdale:1999:IDP

[512] G. Lonsdale. An introduction to DRAMA project. Lecture Notes in Computer Science, 1615:3–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wijshoff:1999:IOP

[513] H. A. G. Wijshoff. An introduction to OCEANS project. Lecture Notes in Computer Science, 1615:5–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Labarta:1999:INP

[514] J. Labarta. An introduction to NANOS project. Lecture Notes in Computer Science, 1615:7–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zima:1999:IHP

[515] H. P. Zima. An introduction to HPF+ project. Lecture Notes in Computer Science, 1615:9–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fluekiger:1999:SCM

[516] P. F. Fluekiger, S. Portmann, and H. P. Luethi. Synchronous communication of molecular and electronic structure information in a distributed computing environment. Lecture Notes in Computer Science, 1615:11–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1999:PBS

[517] T. Sato. Profile-based selection of load value and address predictors. Lecture Notes in Computer Science, 1615:17– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cameron:1999:ILM

[518] K. W. Cameron, Y. Luo, and J. Scharzmeier. Instruction-level microprocessor modeling of scientific applications. Lecture Notes in Computer Science, 1615:29–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammami:1999:NNC

[519] O. Hammami. Neural network classifiers execution on superscalar microprocessors. Lecture Notes in Computer Science, 1615:41–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanoh:1999:MPC

[520] Y. Kanoh, M. Nakamura, T. Hirose, and T. Hosomi. Message passing communication in a parallel computer Cenju-4. Lecture Notes in Computer Science, 1615:55–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryan:1999:CVH

[521] S. Ryan, J. N. Amaral, G. Gao, and Z. Ruiz. Coping with very high latencies in petaflop computer systems. Lecture Notes in Computer Science, 1615:71–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1999:PCS

[522] A. Cohen. Parallelization via constrained storage mapping optimization. Lecture Notes in Computer Science, 1615:83–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusano:1999:CAP

[523] K. Kusano and M. Sato. A comparison of automatic parallelizing compiler and improvements by compiler directives. Lecture Notes in Computer Science, 1615:95–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voss:1999:DAP

[524] M. Voss and R. Eigenmann. Dynamically adaptive parallel programs. Lecture Notes in Computer Science, 1615: 109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kisuki:1999:FSI

[525] T. Kisuki, P. M. W. Knijnenburg, M. F. P. O’Boyle, and F. Bodin. A feasibility study in iterative compilation. Lecture Notes in Computer Science, 1615:121–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saito:1999:MRS

[526] H. Saito, N. Stavrakos, and C. Polychronopoulos. Multithreading runtime support for loop and functional parallelism. Lecture Notes in Computer Science, 1615:133–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imparato:1999:PVE

[527] A. Imparato, M. Giordano, and M. M. Furnari. Parallelization and vectorization effects on a code simulating a vitreous lattice model with constrained dynamics. Lecture Notes in Computer Science, 1615:145–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xiao:1999:MPH

[528] F. Xiao and T. Ebisuzaki. Multi-phase hydrodynamic simulations on parallel computer. Lecture Notes in Computer Science, 1615:157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munz:1999:KAF

[529] F. Munz, T. Ludwig, S. Ziegler, and P. Bartenstein. Kinetic analysis of functional images: The case for a practical approach to performance prediction. Lecture Notes in Computer Science, 1615:169–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uejima:1999:PRE

[530] A. Uejima and K. Yamazaki. Parallel radiosity: Evaluation of parallel form factor calculations and a static load balancing algorithm. Lecture Notes in Computer Science, 1615:181–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miki:1999:PDO

[531] M. Miki, T. Hiroyasu, and T. Ikeda. Parallel distributed optimization by resource addition and reduction. Lecture Notes in Computer Science, 1615:194– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schenk:1999:APS

[532] O. Schenk, K. Gaertner, and W. Fichtner. Application of parallel sparse direct methods in semiconductor device and process simulation. Lecture Notes in Computer Science, 1615:206– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akiyama:1999:BCP

[533] Y. Akiyama, K. Onizuka, T. Noguchi, and M. Ando. Biological- and chemicalparallel applications on a PC cluster. Lecture Notes in Computer Science, 1615:220–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chowdhury:1999:UJA

[534] A. Chowdhury and O. Frieder. Unnoticeable jitter in ATM workstation configurations. Lecture Notes in Computer Science, 1615:234–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gobioff:1999:IPN

[535] H. Gobioff, D. Nagle, and G. Gibson. Integrity and performance in network attached storage. Lecture Notes in Computer Science, 1615:244–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kyriacou:1999:NID

[536] C. Kyriacou and S. Evripidou. Network interface for a data driven network of workstations (D02NOW). Lecture Notes in Computer Science, 1615: 257–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokokawa:1999:BDE

[537] M. Yokokawa, S. Habata, S. Kawai, and H. Ito. Basic design of the Earth simulator. Lecture Notes in Computer Science, 1615:269–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1999:PCD

[538] Y. Tanaka, N. Goto, M. Kakei, and T. Inoue. Parallel computational design of NJR global climate models. Lecture Notes in Computer Science, 1615: 281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iizuka:1999:GHP

[539] M. Iizuka, H. Nakamura, K. Garatani, and K. Nkajima. GeoFEM: Highperformance parallel FEM for geophysical applications. Lecture Notes in Computer Science, 1615:292–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harwood:1999:GN

[540] A. Harwood and H. Shen. Generalized networks. Lecture Notes in Computer Science, 1615:304–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1999:HGT

[541] M. Giordano and M. M. Furnari. HT-Gviz: A graphics tool for the synthesis of automatic and user-driven program parallelization in the compilation process. Lecture Notes in Computer Science, 1615:312–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nemeth:1999:MLK

[542] Z. Nemeth, H. Tomiyasu, P. Kacsuk, and M. Amamiya. Multithreaded LOGFLOW on KUMP/ D. Lecture Notes in Computer Science, 1615:320– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jin:1999:GRA

[543] H. Jin, J. He, Q. Chen, and K. Hwang. Grouped RAID accesses to reduce false sharing effect in cluster with single I/O space. Lecture Notes in Computer Science, 1615:328–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oshitani:1999:PMR

[544] T. Oshitani and T. Watanabe. Parallel map recognition by pipeline control. Lecture Notes in Computer Science, 1615:336–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moriya:1999:PID

[545] K. Moriya and T. Nodera. Parallelization of ILU decomposition for elliptic boundary value problem of the PDE on AP3000. Lecture Notes in Computer Science, 1615:344–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:TMW

[546] T.-S. Chen. Task migration in 2D wormhole-routed mesh multicomputers. Lecture Notes in Computer Science, 1615:354–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furuyama:1999:SDD

[547] S. Furuyama and T. Matsuzawa. A suitable domain decomposition for the adaptive mesh refinement method. Lecture Notes in Computer Science, 1615: 363–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oi:1999:UCA

[548] H. Oi and N. Ranganathan. Utilization of cache area in on-chip multiprocessor. Lecture Notes in Computer Science, 1615:373–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stankovic:1999:PPE

[549] N. Stankovic and K. Zhang. A parallel programming environment for networks. Lecture Notes in Computer Science, 1615:381–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urata:1999:EFD

[550] T. Urata and H. Nishikawa. Emulation facility for data-driven real-time multiprocessing. Lecture Notes in Computer Science, 1615:391–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1999:FRL

[551] C. Y. Park, J.-M. Gil, Y.-H. Han, and C.-S. Hwang. Failure restoration for location server with user movement learning and prediction. Lecture Notes in Computer Science, 1615:399– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stankovic:1999:RRT

[552] John A. Stankovic. Reflection in real-time systems. Lecture Notes in Computer Science, 1616:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160001.
pdf.

Welch:1999:DKE

[553] Ian Welch and Robert Stroud. From Dalang to Kava — the evolution of a reflective Java extension. Lecture Notes in Computer Science, 1616:2– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160002.
pdf.

Golm:1999:JML

[554] Michael Golm and Jürgen Kleinöder. Jumping to the meta level: Behavioral reflection can be fast and flexible. Lecture Notes in Computer Science, 1616: 22–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160022.
pdf.

Moessenboek:1999:ORM

[555] H. Moessenboek and C. Steindl. The Oberon-2 reflection model and its applications. Lecture Notes in Computer Science, 1616:40–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossenbock:1999:ORM

[556] Hanspeter Mössenböck and Christoph Steindl. The Oberon-2 reflection model and its applications. Lecture Notes in Computer Science, 1616:40– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160040.
pdf.

Demphlous:1999:DPL

[557] Stéphane Demphlous and Franck Lebastard. Designing persistence libraries in reflective models with intercession property for a clientserver environment. Lecture Notes in Computer Science, 1616:54–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160054.
pdf.

Robben:1999:NFP

[558] Bert Robben, Bart Vanhaute, Wouter Joosen, and Pierre Verbaeten. Nonfunctional policies. Lecture Notes in Computer Science, 1616:74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160074.
pdf.

Stefani:1999:RSI

[559] Jean-Bernard Stefani. On the reflective structure of information networks. Lecture Notes in Computer Science, 1616: 93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160093.
pdf.

Robbins:1999:RMS

[560] Wayne Robbins and Nicolas D. Georganas. Reflective media space management using RASCAL. Lecture Notes in Computer Science, 1616:95– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160095.
pdf.

Blair:1999:DRA

[561] Gordon S. Blair, Fábio Costa, Geoff Coulson, Fabien Delpiano, Hector Duran, Bruno Dumant, François Horn, Nikos Parlavantzas, and Jean-Bernard Stefani. The design of a resource-aware reflective middleware architecture. Lecture Notes in Computer Science, 1616: 115–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160115.
pdf.

Lemmens:1999:FAS

[562] Inge M. C. Lemmens and Peter J. Braspenning. A formal analysis of Smithsonian computational reflection. Lecture Notes in Computer Science, 1616:135–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160135.
pdf.

Amano:1999:RDA

[563] Noriki Amano and Takuo Watanabe. Reflection for dynamic adaptability: A linguistic approach using LEAD++. Lecture Notes in Computer Science, 1616:138–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1616/16160138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1616/16160138.
                                        pdf.

Matthijs:1999:NRS

[564] Frank Matthijs, Peter Kenens, Wouter Joosen, and Pierre Verbaeten. Networking and reflection: A strong combination. Lecture Notes in Computer Science, 1616:141–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1616/16160141.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1616/16160141.pdf.

Tuma:1999:TSS

[565] Petr Tuma, Valerie Issarny, and Apostolos Zarras. Towards systematic synthesis of reflective middleware. Lecture Notes in Computer Science, 1616:144– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1616/16160144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1616/16160144.
                                        pdf.

Pawlak:1999:AAW

[566] Renaud Pawlak, Laurence Duchien, and Gérard Florin. An automatic aspect weaver with a reflective programming language. Lecture Notes in Computer Science, 1616:147–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160147.
pdf.

Killijian:1999:UCT

[567] Marc-Olivier Killijian, Juan-Carlos Ruiz-Garcia, and Jean-Charles Fabre. Using compile-time reflection for objects’state capture. Lecture Notes in Computer Science, 1616:150–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160150.
pdf.

Yokote:1999:PPF

[568] Yasuhiko Yokote. Past, present, and future of aperios. Lecture Notes in Computer Science, 1616:153–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160153.
pdf.

Anderson:1999:RJS

[569] Kenneth R. Anderson and Timothy J. Hickey. Reflecting Java into Scheme.

Lecture Notes in Computer Science, 1616:154–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1616/16160154.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1616/16160154.
                                        pdf.

Karaorman:1999:JRJ

[570] Murat Karaorman, Urs Hölzle, and John Bruno. jContractor: A reflective Java library to support design by contract. Lecture Notes in Computer Science, 1616:175–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1616/16160175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1616/16160175.
                                        pdf.

Ledoux:1999:ORO

[571] Thomas Ledoux. OpenCorba: A reflective open broker. Lecture Notes in Computer Science, 1616:197–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1616/16160197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1616/16160197.
                                        pdf.

Sohda:1999:OPH

[572] Yukihiko Sohda, Hirotaka Ogawa, and Satoshi Matsuoka. OMPC++ — A portable high-performance implementation of DSM using OpenC++ reflection. Lecture Notes in Computer Science, 1616:215–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1616/16160215.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1616/16160215.pdf.

Cazenave:1999:MDS

[573] Tristan Cazenave. Metaprogramming domain specific metaprograms. Lecture Notes in Computer Science, 1616:235– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1616/16160235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1616/16160235.
pdf.

DeVolder:1999:AOL

[574] K. De Volder and T. D’Hondt. Aspectoriented logic meta programming. Lecture Notes in Computer Science, 1616: 250–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volder:1999:AOL

[575] Kris De Volder and Theo D’Hondt. Aspect-oriented logic meta programming. Lecture Notes in Computer Science, 1616:250–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1616/16160250.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1616/16160250.pdf.

Anonymous:1999:AIc

[576] Anonymous. Author index. Lecture Notes in Computer Science, 1616:273– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1616/1616auth.pdf.

Bryant:1999:MVU

[577] R. E. Bryant. Microprocessor verification using efficient decision procedures for a logic of equality with uninterpreted functions. Lecture Notes in Computer Science, 1617:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massacci:1999:DRT

[578] F. Massacci. Design and results of the Tableaux-99 non-classical (modal) systems comparison. Lecture Notes in Computer Science, 1617:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel-Schneider:1999:DF

[579] P. F. Patel-Schneider and I. Horrocks. DLP and FaCT. Lecture Notes in Computer Science, 1617:19–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haarslev:1999:AAA

[580] V. Haarslev and R. Moeller. Applying an ALC ABox consistency tester to modal logic SAT problems. Lecture Notes in Computer Science, 1617:24– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyapati:1999:KSD

[581] V. Boyapati and R. Gore. KtSeqC: System description. Lecture Notes in Computer Science, 1617:29–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massacci:1999:ARV

[582] F. Massacci. Automated reasoning and the verification of security protocols. Lecture Notes in Computer Science, 1617:32–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haehnle:1999:PCT

[583] R. Haehnle and B. Beckert. Proof confluent tableau calculi. Lecture Notes in Computer Science, 1617:34–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1999:ACP

[584] M. Baaz and C. G. Femueller. Analytic calculi for projective logics. Lecture Notes in Computer Science, 1617: 36–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1999:MPI

[585] P. Baumgartner, J. D. Horton, and B. Spencer. Merge path improvements for minimal model hyper tableaux. Lecture Notes in Computer Science, 1617: 51–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broda:1999:CPI

[586] K. Broda and D. Gabbay. CLDS for propositional intuitionistic logic. Lecture Notes in Computer Science, 1617: 66–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caldwell:1999:ITE

[587] J. Caldwell. Intuitionisitic tableau extracted. Lecture Notes in Computer Science, 1617:82–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cantone:1999:TBD

[588] D. Cantone and C. G. Zarba. A tableau-based decision procedure for a fragment of set theory involving a restricted form of quantification. Lecture Notes in Computer Science, 1617:97– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciabattoni:1999:BCS

[589] A. Ciabattoni. Bounded contraction in systems with linearity. Lecture Notes in Computer Science, 1617:113–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGroote:1999:NAL

[590] P. De Groote. The non-associative lambek calculus with product in polynomial time. Lecture Notes in Computer Science, 1617:128–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1999:SCN

[591] S. Demri. Sequent calculi for nominal tense logics: A step towards mechanization? Lecture Notes in Computer Science, 1617:140–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1999:CFD

[592] S. Demri and R. Gore. Cut-free display calculi for nominal tense logics. Lecture Notes in Computer Science, 1617:155–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giese:1999:HET

[593] M. Giese and W. Ahrendt. Hilbert’s epsilon-terms in automated theorem proving. Lecture Notes in Computer Science, 1617:171–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilmore:1999:PFI

[594] P. C. Gilmore. Partial functions in an impredicative simple theory of types. Lecture Notes in Computer Science, 1617:186–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault-Larrecq:1999:SSS

[595] J. Goubault-Larrecq. A simple sequent system for first-order logic with free constructors. Lecture Notes in Computer Science, 1617:202–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantel:1999:LTP

[596] H. Mantel and J. Otten. linTAP: A tableau prover for linear logic. Lecture Notes in Computer Science, 1617: 217–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1999:TCT

[597] W. May. A tableau calculus for a temporal logic with temporal connectives. Lecture Notes in Computer Science, 1617:232–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monz:1999:TCP

[598] C. Monz and M. De Rijke. A tableaux calculus for pronoun resolution. Lecture Notes in Computer Science, 1617: 247–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuetz:1999:GMH

[599] H. Schuetz. Generating minimal Herbrand models step by step. Lecture Notes in Computer Science, 1617:263– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tzakova:1999:TCH

[600] M. Tzakova. Tableau calculi for hybrid logics. Lecture Notes in Computer Science, 1617:278–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirth:1999:FFO

[601] C.-P. Wirth. Full first-order free variable sequents and tableaux in implicit induction. Lecture Notes in Computer Science, 1617:293–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Endriss:1999:ITP

[602] U. Endriss. An interactive theorem proving assistant. Lecture Notes in Computer Science, 1617:308–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Endriss:1999:TEB

[603] U. Endriss. A time efficient KE based theorem prover. Lecture Notes in Computer Science, 1617:313–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1999:SPU

[604] A. Wolf and J. Draeger. Strategy parallel use of model elimination with lemmata — system abstract. Lecture Notes in Computer Science, 1617:319– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bezivin:1999:UBR

[605] J. Bezivin and P.-A Muller. UML: The birth and rise of a standard modeling notation. Lecture Notes in Computer Science, 1618:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hitz:1999:DUS

[606] M. Hitz and G. Kappel. Developing with UML — some pitfalls and workarounds. Lecture Notes in Computer Science, 1618:9–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1999:SAU

[607] C. Atkinson. Supporting and applying the UML conceptual framework. Lecture Notes in Computer Science, 1618:21–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morand:1999:MIT

[608] B. Morand. Modeling: Is it turning informal into formal? Lecture Notes in Computer Science, 1618:37–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:1999:BBW

[609] F. Arnold and G. Podehl. Best of both worlds — A mapping from EXPRESS-G to UML. Lecture Notes in Computer Science, 1618:49–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olive:1999:PRU

[610] A. Olive and M.-R Sancho. Porting ROSES to UML — an experience report. Lecture Notes in Computer Science, 1618:64–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1999:MUM

[611] J. Suzuki and Y. Yamamoto. Making I. UML models interoperable with UXF. Lecture Notes in Computer Science, 1618:78–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1999:TRU

[612] M. Gogolla and M. Richters. Transformation rules for UML class diagrams. Lecture Notes in Computer Science, 1618:92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lano:1999:STU

[613] K. Lano and J. Bicarregui. Semantics and transformations for UML models. Lecture Notes in Computer Science, 1618:107–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desfray:1999:ADP

[614] P. Desfray. Automation of design pattern: Concepts, tools and practices. Lecture Notes in Computer Science, 1618:120–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khriss:1999:ASU

[615] I. Khriss, M. Elkoutbi, and R. K. Keller. Automating the synthesis of UML StateChart diagrams from multiple collaboration diagrams. Lecture Notes in Computer Science, 1618:132– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleppe:1999:IFO

[616] A. Kleppe, J. Warmer, and S. Cook. Informal formality? the object constraint language and its application in the UML metamodel. Lecture Notes in Computer Science, 1618:148–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamie:1999:ROC

[617] A. Hamie, F. Civello, J. Howse, and S. Kent. Reflections on the object constraint language. Lecture Notes in Computer Science, 1618:162–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ou:1999:UUC

[618] Y. Ou. On using UML class diagrams for object-oriented database design specification of integrity constraints. Lecture Notes in Computer Science, 1618:173–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arlow:1999:LMC

[619] J. Arlow, W. Emmerich, and J. Quinn.

Literate modelling — capturing business knowledge with the UML. Lecture Notes in Computer Science, 1618: 189–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kande:1999:AUD

[620] M. Mancona Kande, S. Mazaher, O. Prnjat, and L. Sacks. Applying UML to design an inter-domain service management application. Lecture Notes in Computer Science, 1618:200– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korthaus:1999:BSD

[621] A. Korthaus and S. Kuhlins. BOOSTER*Process: A software development process model integrating business object technology and UML. Lecture Notes in Computer Science, 1618:215–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourdeau:1999:HCD

[622] E. Bourdeau, P. Lugagne, and P. Roques. Hierarchical context diagram with UML: An experience report on satellite ground system analysis. Lecture Notes in Computer Science, 1618:227–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seeman:1999:EUS

[623] J. Seeman and J. Wolff v. Gudenberg. Extension of UML sequence diagrams for real-time systems. Lecture Notes in Computer Science, 1618:240–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovacevic:1999:UUI

[624] S. Kovacevic. UML and user interface modeling. Lecture Notes in Computer Science, 1618:253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paech:1999:RAD

[625] B. Paech. On the role of activity diagrams in UML — A user task centered development process for UML. Lecture Notes in Computer Science, 1618:267– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hruby:1999:SUD

[626] P. Hruby. Structuring UML design deliverables. Lecture Notes in Computer Science, 1618:278–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kivisto:1999:CSU

[627] K. Kivisto. Considerations of and suggestions for a UML-Specific process model. Lecture Notes in Computer Science, 1618:294–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellor:1999:ALU

[628] S. J. Mellor, S. R. Tockey, R. Arthaud, and P. Leblanc. An action language for UML: Proposal for a precise execution semantics. Lecture Notes in Computer Science, 1618:307–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lanusse:1999:RTM

[629] A. Lanusse, S. Gerard, and F. Terrier. Real-time modeling with UML: The ACCORD approach. Lecture Notes in Computer Science, 1618:319–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1999:UFM

[630] A. Evans, R. France, K. Lano, and B. Rumpe. The UML as a formal modeling notation. Lecture Notes in Computer Science, 1618:336–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henderson-Sellers:1999:OPE

[631] B. Henderson-Sellers. OML: Proposals to enhance UML. Lecture Notes in Computer Science, 1618:349–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jezequel:1999:VDS

[632] J.-M. Jezequel, A. Le Guennec, and F. Pennaneac’h. Validating distributed software modeled with the unified modeling language. Lecture Notes in Computer Science, 1618:365–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mens:1999:SDR

[633] T. Mens, C. Lucas, and P. Steyaert. Supporting disciplined reuse and evolution of UML models. Lecture Notes in Computer Science, 1618:378–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lester:1999:AUE

[634] N. G. Lester, F. G. Wilkie, and D. W. Bustard. Applying UML extensions to facilitate software reuse. Lecture Notes in Computer Science, 1618:393– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oevergaard:1999:FAU

[635] G. Oevergaard and K. Palmkvist. A formal approach to use cases and their relationships. Lecture Notes in Computer Science, 1618:406–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1999:PFA

[636] P. Allen. A practical framework for applying UML. Lecture Notes in Computer Science, 1618:419–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saksena:1999:EAC

[637] M. Saksena, M. M. Larrondo-Petrie, R. B. France, and M. P. Evett. Extending aggregation constructs in UML. Lecture Notes in Computer Science, 1618:434–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milidiu:1999:EIW

[638] R. L. Milidiu, A. A. Pessoa, and E. S. Laber. Efficient implementation of the WARM-UP algorithm for the construction of length-restricted prefix codes. Lecture Notes in Computer Science, 1619:1–17, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller-Hannemann:1999:IWM

[639] M. Mueller-Hannemann and A. Schwartz. Implementing weighted b-matching algorithms: Insights from a computational study. Lecture Notes in Computer Science, 1619:18–36, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helman:1999:DPE

[640] D. R. Helman and J. Jaja. Designing practical efficient algorithms for symmetric multiprocessors. Lecture Notes in Computer Science, 1619:37– 56, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Six:1999:CDB

[641] J. M. Six and I. G. Tollis. Circular drawings of biconnected graphs. Lecture Notes in Computer Science, 1619: 57–73, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stallmann:1999:HED

[642] M. Stallmann, F. Brglez, and D. Gosh. Heuristics and experimental design for bigraph crossing number minimization. Lecture Notes in Computer Science, 1619:74–93, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mount:1999:BSP

[643] D. M. Mount and F.-T. Pu. Binary space partitions in Pluecker space. Lecture Notes in Computer Science, 1619: 94–113, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walker:1999:PPP

[644] R. J. Walker and J. Snoeyink. Practical point-in-polygon tests using CSG representations of polygons. Lecture Notes in Computer Science, 1619:114–123, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goodrich:1999:AIO

[645] M. T. Goodrich, M. Handy, B. Hudson, and R. Tamassia. Accessing the internal organization of data structures in the JDSL library. Lecture Notes in Computer Science, 1619:124–139, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pizzonia:1999:OOD

[646] M. Pizzonia and G. Di Battista. Object-oriented design of graph oriented data structures. Lecture Notes in Computer Science, 1619:140–155, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schirra:1999:CSC

[647] S. Schirra. A case study on the cost of geometric computing. Lecture Notes in Computer Science, 1619:156–176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caldwell:1999:DIF

[648] A. E. Caldwell, A. B. Kahng, and I. L. Markov. Design and implementation of the Fiduccia-Mattheyses heuristic for VLSI netlist partitioning. Lecture Notes in Computer Science, 1619:177–193, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cwilich:1999:ARP

[649] S. Cwilich, M. Deng, D. F. Lynch, and S. J. Phillips. Algorithms for restoration planning in a telecommunications network. Lecture Notes in Computer Science, 1619:194–209, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shibuya:1999:CSP

[650] T. Shibuya. Computing the n × m shortest paths efficiently. Lecture Notes in Computer Science, 1619:210–225, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heileman:1999:IWC

[651] G. L. Heileman, C. E. Pizano, and C. T. Abdallah. Image watermarking for copyright protections. Lecture Notes in Computer Science, 1619:226– 245, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Csirik:1999:SOP

[652] J. Csirik, D. S. Johnson, C. Kenyon, and P. W. Shor. A self organizing bin packing heuristic. Lecture Notes in Computer Science, 1619:246–265, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levine:1999:FRC

[653] M. S. Levine. Finding the right cutting planes for the TSP. Lecture Notes in Computer Science, 1619:266–281, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zachariasen:1999:OAE

[654] M. Zachariasen and P. Winter. Obstacle-avoiding Euclidean Steiner trees in the plane: An exact algorithm. Lecture Notes in Computer Science, 1619:282–295, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acharya:1999:AAC

[655] A. Acharya, H. Zhu, and K. Shen. Adaptive algorithms for cache-efficient trie search. Lecture Notes in Computer Science, 1619:296–311, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanders:1999:FPQ

[656] P. Sanders. Fast priority queues for cached memory. Lecture Notes in Computer Science, 1619:312–327, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arge:1999:EBO

[657] L. Arge, K. H. Hinrichs, J. Vahrenhold, and J. S. Vitter. Efficient bulk operations on dynamic R-trees. Lecture Notes in Computer Science, 1619:328– 348, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molino:1999:CGD

[658] Gianpaolo Molino. From clinical guidelines to decision support. Lecture Notes in Computer Science, 1620:3– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200003.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1620/16200003.
pdf.

Stefanelli:1999:AIB

[659] M. Stefanelli. Artificial intelligence for building learning health care organizations. Lecture Notes in Computer Science, 1620:13–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200013.
pdf.

Shahar:1999:TET

[660] Yuval Shahar. Timing is everything: Temporal reasoning and temporal data maintenance in medicine. Lecture Notes in Computer Science, 1620:30– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200030.
pdf.

Lavrac:1999:MLD

[661] Nada Lavrac. Machine learning for data mining in medicine. Lecture Notes in Computer Science, 1620:47– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200047.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1620/16200047.
pdf.

Quaglini:1999:GBW

[662] S. Quaglini, C. Mossa, C. Fassino, M. Stefanelli, A. Cavallini, and G. Micieli. Guidelines-based workflow systems. Lecture Notes in Computer Science, 1620:65–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1620/16200065.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1620/16200065.pdf.

Seroussi:1999:ECP

[663] Brigitte Séroussi, Jacques Bouaud, and Éric-Charles Antoine. Enhancing clinical practice guideline compliance by involving physicians in the decision process. Lecture Notes in Computer Science, 1620:76–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200076.
pdf.

Sauvagnac:1999:ATP

[664] C. Sauvagnac, J. Stines, A. Lesur, P. Falzon, and P. Bey. Application of therapeutic protocols: A tool to manage medical knowledge. Lecture Notes in Computer Science, 1620:86– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1620/16200086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200086.
                                        pdf.

Boulay:1999:DDT

[665] Benedict du Boulay, Briony Teather, George du Boulay, Nathan Jeffrey, Derek Teather, Mike Sharples, and Lisa Cuthbert. From description to decision: Towards a decision support training system for MR radiology of the brain. Lecture Notes in Computer Science, 1620:93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200093.
                                        pdf.

DuBoulay:1999:DDT

[666] B. Du Boulay, B. Teather, G. Du Boulay, and N. Jeffrey. From description to decision: Towards a decision support training system for MR radiology of the brain. Lecture Notes in Computer Science, 1620:93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eich:1999:IBD

[667] H. P. Eich and C. Ohmann. Internetbased decision-support server for acute abdominal pain. Lecture Notes in Computer Science, 1620:103–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200103.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1620/16200103.
pdf.

Montani:1999:MMR

[668] Stefania Montani, Riccardo Bellazzi, Luigi Portinale, Alberto Riva, and Mario Stefanelli. Multi-modal reasoning in diabetic patient management. Lecture Notes in Computer Science, 1620:113–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1620/16200113.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1620/16200113.
pdf.

Schmidt:1999:ECB

[669] Rainer Schmidt, Bernhard Pollwein, and Lothar Gierl. Experiences with case-based reasoning methods and prototypes for medical knowledgebased systems. Lecture Notes in Computer Science, 1620:124–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200124.
pdf.

Aknine:1999:ESR

[670] Samir Aknine. Exploiting social reasoning of open multi-agent systems to enhance cooperation in hospitals. Lecture Notes in Computer Science, 1620: 133–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200133.
                                        pdf.

Bielza:1999:IDN

[671] Concha Bielza, Sixto Ríos-Insua, and Manuel Gómez. Influence diagrams for neonatal jaundice management. Lecture Notes in Computer Science, 1620: 138–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200138.
                                        pdf.

Clark:1999:EDP

[672] I. R. Clark, B. A. McCauley, I. M. Young, P. G. Nightingale, M. Peters, N. T. Richards, and D. Adu. Electronic drug prescribing and administration — bedside medical decision making. Lecture Notes in Computer Science, 1620: 143–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200143.
                                        pdf.

Horn:1999:NVT

[673] Werner Horn, Christian Popow, Christoph Stocker, and Silvia Miksch. Neonatal ventilation tutor (VIE-NVT), a teaching program for the mechanical ventilation of newborn infants. Lecture Notes in Computer Science, 1620:

148–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200148.
pdf.

O:1999:LCB

[674] Ying-Lie O. A life-cycle based authorisation expert database system. Lecture Notes in Computer Science, 1620:153– 157, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200153.
pdf.

Torchio:1999:DSS

[675] Mauro Torchio, Stefania Battista, Fabrizio Bar, Cristina Pollet, Marina Marzuoli, Maria Cesira Bucchi, Roberto Pagni, and Gianpaolo Molino. A decision-support system for the identification, staging, and functional evaluation of liver diseases (HEPASCORE). Lecture Notes in Computer Science, 1620:158–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200158.
pdf.

Carrault:1999:MBA

[676] G. Carrault, M.-O. Cordier, René Quiniou, M. Garreau, J. J. Bellanger, and A. Bardou. A modelbased approach for learning to identify cardiac arrhythmias. Lecture Notes in Computer Science, 1620:165– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200165.
pdf.

Lucas:1999:MBS

[677] Peter Lucas, Astrid Tholen, and Geeske van Oort. An model-based system for pacemaker reprogramming. Lecture Notes in Computer Science, 1620:175–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1620/16200175.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1620/16200175.
pdf.

Cruz:1999:IDB

[678] Jorge Cruz, Pedro Barahona, and Frédéric Benhamou. Integrating deep biomedical models into medical decision support systems: An interval constraint approach. Lecture Notes in Computer Science, 1620:185–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200185.
pdf.

Andreassen:1999:DTA

[679] Steen Andreassen, Leonard Leibovici, Henrik C. Schønheyder, Brian Kristensen, Christian Riekehr, Anders Geill Kjær, and Kristian G. Olesen. A decision theoretic approach to empirical treatment of bacteraemia originating from the urinary tract. Lecture Notes in Computer Science, 1620:197– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200197.
pdf.

Fernandez:1999:EID

[680] E. A. Fernandez, J. Presedo, and S. Barro. An ECG ischemic detection system based on self-organizing maps and a sigmoid function preprocessing stage. Lecture Notes in Computer Science, 1620:207–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200207.
pdf.

Juhola:1999:NNR

[681] Martti Juhola, Jorma Laurikkala, Kati Viikki, Yrjö Auramo, Erna Kentala, and Ilmari Pyykkö. Neural network recognition of otoneurological vertigo diseases with comparison of some other classification methods. Lecture Notes in Computer Science, 1620:217– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200217.
                                        pdf.

Dorffner:1999:CLN

[682] Georg Dorffner, Ernst Leitgeb, and Heinz Koller. A comparison of linear and non-linear classifiers for the detection of coronary artery disease in Stress-ECG. Lecture Notes in Computer Science, 1620:227–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200227.
                                        pdf.

Goodridge:1999:CBN

[683] Wayne Goodridge, Hadrian Peter, and Akin Abayomi. The case-based neural network model and its use in medical expert systems. Lecture Notes in Computer Science, 1620:232–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200232.
                                        pdf.

Pisanelli:1999:MOL

[684] Domenico M. Pisanelli, Aldo Gangemi, and Geri Steve. A medical ontology library that integrates the UMLS Metathesaurus^{T M}. Lecture Notes in Computer Science, 1620:239–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200239.
pdf.

Achour:1999:UUK

[685] Soumeya Achour, Michel Dojat, Jean-Marc Brethon, Gil Blain, and Eric Lepage. The use of the UMLS knowledge sources for the design of a domain specific ontology: A practical experience in blood transfusion. Lecture Notes in Computer Science, 1620:249– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200249.
pdf.

Terenziani:1999:RKL

[686] Paolo Terenziani, Paolo Raviola, Oscar Bruschi, Mauro Torchio, Marina Marzuoli, and Gianpaolo Molino. Representing knowledge levels in clinical guidelines. Lecture Notes in Computer Science, 1620:254–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200254.
pdf.

Bellazzi:1999:IAC

[687] R. Bellazzi, C. Larizza, P. Magni, S. Montani, and G. De Nicolao. Intelligent analysis of clinical time series by combining structural filtering and temporal abstractions. Lecture Notes in Computer Science, 1620:261– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200261.
                                        pdf.

Hunter:1999:KBE

[688] Jim Hunter and Neil McIntosh. Knowledge-based event detection in complex time series data. Lecture Notes in Computer Science, 1620:271– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200271.
                                        pdf.

Miksch:1999:ASQ

[689] Silvia Miksch, Andreas Seyfang, Werner Horn, and Christian Popow. Abstracting steady qualitative descriptions over time from noisy, high-frequency data. Lecture Notes in Computer Science, 1620:281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200281.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1620/16200281.
pdf.

Kosara:1999:VTT

[690] Robert Kosara and Silvia Miksch. Visualization techniques for timeoriented, skeletal plans in medical therapy planning. Lecture Notes in Computer Science, 1620:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200291.
pdf.

Combi:1999:VTC

[691] Carlo Combi, Luisa Portoni, and Francesco Pinciroli. Visualizing temporal clinical data on the WWW. Lecture Notes in Computer Science, 1620: 301–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200301.
pdf.

Kukar:1999:MLS

[692] Matjaz Kukar and Ciril Groselj. Machine learning in stepwise diagnostic process. Lecture Notes in Computer Science, 1620:315–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200315.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1620/16200315.
                                        pdf.

Mani:1999:RNP

[693] Subramani Mani, Malcolm B. Dick, Michael J. Pazzani, Evelyn L. Teng, Daniel Kempler, and I. Maribell Taussig. Refinement of neuropsychological tests for dementia screening in a cross cultural population using machine learning. Lecture Notes in Computer Science, 1620:326–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200326.
                                        pdf.

McQuatt:1999:AHI

[694] A. McQuatt, P. J. D. Andrews, D. Sleeman, V. Corruble, and P. A. Jones. The analysis of head injury data using decision tree techniques. Lecture Notes in Computer Science, 1620:336– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200336.
                                        pdf.

Zupan:1999:MLS

[695] Blaz Zupan, Janez Demsar, Michael W. Kattan, J. Robert Beck, and I. Bratko. Machine learning for survival analysis: A case study on recurrence of prostate cancer. Lecture Notes in Computer Science, 1620:346–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200346.
pdf.

Calvelo:1999:IPS

[696] Daniel Calvelo, Marie-C. Chambrin, Denis Pomorski, and Pierre Ravaux. ICU patient state characterization using machine learning in a time series framework. Lecture Notes in Computer Science, 1620:356–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200356.
pdf.

Gamberger:1999:DRI

[697] Dragan Gamberger, Nada Lavrac, and Ciril Groselj. Diagnostic rules of increased reliability for critical medical applications. Lecture Notes in Computer Science, 1620:361–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200361.
pdf.

Sierra:1999:MLI

[698] Basilio Sierra, Nicolás Serrano, Pedro Larrañaga, Eliseo J. Plasencia, Iñaki Inza, Juan José Jiménez, Jose María De la Rosa, and María Luisa Mora. Machine learning inspired approaches to combine standard medical measures at an intensive care unit. Lecture Notes in Computer Science, 1620:366– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200366.
pdf.

Wu:1999:STP

[699] Ping Wu, Kok Liang Heng, Shuo Wang Yang, Yi Feng Chen, Ravuru Subramanyam Mohan, and Peter Huat Chye Lim. A screening technique for prostate cancer by hair chemical analysis and artificial intelligence. Lecture Notes in Computer Science, 1620:372–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200372.
pdf.

Cawsey:1999:CMH

[700] Alison Cawsey, Floriana Grasso, and Ray Jones. A conversational model for health promotion on the World Wide Web. Lecture Notes in Computer Science, 1620:379–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200379.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1620/16200379.
pdf.

Reiter:1999:TKR

[701] Ehud Reiter, Roma Robertson, and Liesl Osman. Types of knowledge required to personalise smoking cessation letters. Lecture Notes in Computer Science, 1620:389–399, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200389.
pdf.

Romacker:1999:SBC

[702] Martin Romacker, Stefan Schulz, and Udo Hahn. Small is beautiful — compact semantics for medical language processing. Lecture Notes in Computer Science, 1620:400–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200400.
pdf.

Barker:1999:SDN

[703] D. J. Barker, S. C. Lynch, D. S. Simpson, and W. A. Corbett. Speech driven natural language understanding for hands-busy recording of clinical information. Lecture Notes in Computer Science, 1620:411–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200411.
                                        pdf.

Zweigenbaum:1999:AAM

[704] Pierre Zweigenbaum and Natalia Grabar. Automatic acquisition of morphological knowledge for medical language processing. Lecture Notes in Computer Science, 1620:416–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200416.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200416.
                                        pdf.

Germond:1999:MAS

[705] Laurence Germond, Michel Dojat, Chris Taylor, and Catherine Garbay. A multi-agent system for MRI brain segmentation. Lecture Notes in Computer Science, 1620:423–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1620/16200423.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1620/16200423.
                                        pdf.

Kokai:1999:MBV

[706] Gabriella Kókai, Zoltán Tóth, and Róbert Ványi. Modelling blood vessels of the eye with parametric L-systems using evolutionary algorithms. Lecture Notes in Computer Science, 1620:433– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200433.
pdf.

Hammond:1999:AMS

[707] Peter Hammond, Paul Wells, and Sanjay Modgil. Animating medical and safety knowledge. Lecture Notes in Computer Science, 1620:443– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200443.
pdf.

Hutton:1999:ASM

[708] Tim J. Hutton, Peter Hammond, and John C. Davenport. Active shape models for customised prosthesis design. Lecture Notes in Computer Science, 1620:448–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1620/16200448.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1620/16200448.
pdf.

OLeary:1999:RKM

[709] D. E. O’Leary. Reengineering and knowledge management. Lecture Notes in Computer Science, 1621:1–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papazoglou:1999:KNN

[710] M. P. Papazoglou and J. Hoppenbrouwers. Knowledge navigation in networked digital libraries. Lecture Notes in Computer Science, 1621:13– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benjamins:1999:TBP

[711] V. R. Benjamins, B. Wielinga, J. Wielemaker, and D. Fensel. Towards brokering problem-solving knowledge on the Internet. Lecture Notes in Computer Science, 1621:33–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biebow:1999:TLB

[712] B. Biebow and S. Szulman. Terminae: A linguistics-based tool for the building of a domain ontology. Lecture Notes in Computer Science, 1621:49–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birk:1999:AKA

[713] A. Birk, D. Surmann, and K.-D. Althoff. Applications of knowledge acquisition in experimental software engineering. Lecture Notes in Computer Science, 1621:67–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crow:1999:ASW

[714] L. Crow and N. Shadbolt. Acquiring and structuring Web content with knowledge level models. Lecture Notes in Computer Science, 1621:85– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingue:1999:KBN

[715] J. Domingue and E. Motta. A knowledge-based news server supporting ontology-driven story enrichment and knowledge retrieval. Lecture Notes in Computer Science, 1621:103–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goasdoue:1999:MIS

[716] F. Goasdoue and C. Reynaud. Modeling information sources for information integration. Lecture Notes in Computer Science, 1621:121–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez-Perez:1999:ORR

[717] A. Gomez-Perez and D. Rojas-Amaya. Ontological reengineering for reuse. Lecture Notes in Computer Science, 1621:139–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groot:1999:FVD

[718] P. Groot, A. Ten Teije, and F. Van Harmelen. Formally verifying dynamic properties of knowledge based systems. Lecture Notes in Computer Science, 1621:157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herlea:1999:IBR

[719] D. E. Herlea, C. M. Jonker, J. Treur, and N. J. E. Wijngaards. Integration of behavioural requirements specification within knowledge engineering. Lecture Notes in Computer Science, 1621:173–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoefling:1999:TOS

[720] B. Hoefling, T. Liebig, D. Roesner, and L. Webel. Towards an ontology for substances and related actions. Lecture Notes in Computer Science, 1621:191–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalfoglou:1999:UFO

[721] Y. Kalfoglou and D. Robertson. Use of formal ontologies to support error checking in specifications. Lecture Notes in Computer Science, 1621:207– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oussalah:1999:OST

[722] M. Oussalah and K. Messaadia. The ontologies of semantic and transfer links. Lecture Notes in Computer Science, 1621:225–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parmentier:1999:DPS

[723] T. Parmentier and D. Ziebelin. Distributed problem solving environment dedicated to DNA sequence annotation. Lecture Notes in Computer Science, 1621:243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puuronen:1999:KAM

[724] S. Puuronen and V. Terziyan. Knowledge acquisition from multiple experts based on semantics of concepts. Lecture Notes in Computer Science, 1621: 259–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stumme:1999:AEK

[725] G. Stumme. Acquiring expert knowledge for the design of conceptual information systems. Lecture Notes in Computer Science, 1621:275–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

White:1999:CBA

[726] S. White and D. Sleeman. A constraintbased approach to the description of competence. Lecture Notes in Computer Science, 1621:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beydoun:1999:HIK

[727] G. Beydoun and A. Hoffmann. Holism and incremental knowledge acquisition. Lecture Notes in Computer Science, 1621:309–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuker:1999:IPS

[728] J. Breuker. Indexing problem solving methods for reuse. Lecture Notes in Computer Science, 1621:315–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cairo:1999:SMR

[729] O. Cairo, J. Barreiro, and F. Solsona. Software methodologies at risk. Lecture Notes in Computer Science, 1621: 323–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faure:1999:KAP

[730] D. Faure and C. Nedellec. Knowledge acquisition of predicate argument structures from technical texts using machine learning: The system ASIUM. Lecture Notes in Computer Science, 1621:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Izumi:1999:IED

[731] N. Izumi, A. Maruyama, A. Suzuki, and T. Yamaguchi. An interoperative environment for developing expert systems. Lecture Notes in Computer Science, 1621:335–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1999:UMN

[732] M. G. Jansen and P. Beys. On the use of meaningful names in knowledgebased systems. Lecture Notes in Computer Science, 1621:341–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez-Bejar:1999:FIK

[733] R. Martinez-Bejar, F. Ibañez-Cruz, L.- G. Thong, and T. M. Cao. FMR: An incremental knowledge acquisition system for fuzzy domains. Lecture Notes in Computer Science, 1621:349– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parpola:1999:ASS

[734] P. Parpola. Applying SeSKA to Sisyphus III. Lecture Notes in Computer Science, 1621:355–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perkuhn:1999:DSC

[735] R. Perkuhn. Describing similar control flows for families of problem-solving methods. Lecture Notes in Computer Science, 1621:361–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puppe:1999:MKE

[736] F. Puppe. Meta knowledge for extending diagnostic consultation to critiquing systems. Lecture Notes in Computer Science, 1621:367–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabarijoana:1999:EXC

[737] A. Rabarijoana, R. Dieng, and O. Corby. Exploitation of XML for corporate knowledge management. Lecture Notes in Computer Science, 1621: 373–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmalhofer:1999:OAS

[738] F. Schmalhofer and L. Van Elst. An oligo-agents system with shared responsibilities for knowledge management. Lecture Notes in Computer Science, 1621:379–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sellini:1999:VKV

[739] F. Sellini and P.-A. Yvars. Veri-KoMoD: Verification of knowledge models in the mechanical design field. Lecture Notes in Computer Science, 1621:385–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuckenschmidt:1999:FFU

[740] H. Stuckenschmidt and K. C. Ranze. A flexible framework for uncertain expertise. Lecture Notes in Computer Science, 1621:391–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1999:EOT

[741] R. P. Wolf. Elicitation of operational track grids. Lecture Notes in Computer Science, 1621:397–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harbour:1999:RST

[742] M. G. Harbour and J. A. De la Puente. Reliable software technologies Ada-Europe ’99. Lecture Notes in Computer Science, 1622:i–xiii, 1–449, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albrecht:1999:TTF

[743] R. F. Albrecht. Topological theory of fuzziness. Lecture Notes in Computer Science, 1625:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovacs:1999:SBS

[744] S. Kovacs. Similarity based system reconfiguration by fuzzy classification and hierarchical interpolate fuzzy reasoning. Lecture Notes in Computer Science, 1625:12–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skinner:1999:FSF

[745] J. F. Skinner, Y. M. Garibaldi, and E. C. Ifeachor. A fuzzy system for fetal heart rate assessment. Lecture Notes in Computer Science, 1625:20–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drechsler:1999:EGC

[746] N. Drechsler, W. Guenther, and R. Drechsler. Efficient graph coloring by evolutionary algorithms. Lecture Notes in Computer Science, 1625: 30–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1999:DDR

[747] A. Takahashi and A. Borisov. Determination of decision rules on the basis of genetic algorithms. Lecture Notes in Computer Science, 1625:40–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habtom:1999:MRS

[748] R. Habtom. Modeling a refrigeration system using recurrent neural networks. Lecture Notes in Computer Science, 1625:47–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1999:ENS

[749] Y. Tan, P. Fang, Y. Zhang, and S. Yang. Evaluating nugget sizes of spot welds by using artificial neural network. Lecture Notes in Computer Science, 1625:53–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohonen:1999:SRI

[750] T. Kohonen. Spotting relevant information in extremely large document collections. Lecture Notes in Computer Science, 1625:59–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgulya:1999:FCG

[751] I. Borgulya. Fuzzy controller generation with a fuzzy classification method. Lecture Notes in Computer Science, 1625:62–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1999:TOF

[752] F. Fernandez and J. Gutierrez. Transformation and optimization of fuzzy controllers using signal processing techniques. Lecture Notes in Computer Science, 1625:75–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1999:FCD

[753] K.-D. Kramer, J. Kirschner, and S. Woehlbier. Fuzzy-control design tool for low-cost microcontrollers (FHFC-Tool). Lecture Notes in Computer Science, 1625:88–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lacan:1999:SOE

[754] J. Lacan and P. Chatonnay. Search of optimal error correcting codes with genetic algorithms. Lecture Notes in Computer Science, 1625:93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imai:1999:UCE

[755] K. Imai, N. Kamiura, and Y. Hata. An unsupervised clustering with evolutionary strategy to estimate the cluster number. Lecture Notes in Computer Science, 1625:99–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drechsler:1999:MOO

[756] N. Drechsler, R. Drechsler, and B. Becker. Multi-objective optimization in evolutionary algorithms using satisfiability classes. Lecture Notes in Computer Science, 1625:108–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martyna:1999:NNA

[757] J. Martyna. Neural network approach to design of distributed hard real-time systems. Lecture Notes in Computer Science, 1625:118–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ungerer:1999:STC

[758] C. Ungerer, D. Stuebener, C. Kirchmair, and M. Sturm. Supporting traditional controller of combustion engines by means of neural networks. Lecture Notes in Computer Science, 1625:132– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bongards:1999:CBW

[759] M. Bongards. Controlling biological wastewater treatment plants using fuzzy control and neural networks. Lecture Notes in Computer Science, 1625: 142–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menhaj:1999:NFC

[760] M. B. Menhaj and F. Razzazi. A new fuzzy character segmentation algorithm for Persian / Arabic typed texts. Lecture Notes in Computer Science, 1625:151–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canuto:1999:RMF

[761] A. Canuto, G. Howells, and M. Fairhurst. RePART: A modified fuzzy ARTMAP for pattern recognition. Lecture Notes in Computer Science, 1625:159–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farbiz:1999:AAF

[762] F. Farbiz, M. B. Menhaj, and S. A. Motamedi. An adaptive C-average fuzzy control filter for image enhancement. Lecture Notes in Computer Science, 1625:169–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez-Skarmeta:1999:POS

[763] A. F. Gomez-Skarmeta, F. Jimenez, and J. Ibañez. Pareto-optimality in scheduling problems. Lecture Notes in Computer Science, 1625:177–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cao:1999:CMC

[764] Y. Cao and L. Cao. Controlled Markov chain optimization of genetic algorithms. Lecture Notes in Computer Science, 1625:186–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotta:1999:TEP

[765] C. Cotta and J. M. Troya. Tackling epistatic problems using dynastically optimal recombination. Lecture Notes in Computer Science, 1625:197– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopco:1999:EMC

[766] N. Kopco, P. Sincak, and H. Veregin. Extended methods for classification of remotely sensed images based on ARTMAP neural networks. Lecture Notes in Computer Science, 1625:206– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandstetter:1999:AAN

[767] P. Brandstetter and M. Skotnica. Application of artificial neural network in control of vector pulse-width modulation inverter. Lecture Notes in Computer Science, 1625:220–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitter:1999:MMM

[768] E. D. Schmitter. Modeling multiple microstructure transformations in steels with a Boltzmann neural net. Lecture Notes in Computer Science, 1625:225– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1999:ECWb

[769] K. De Jong. Evolutionary computation: Where we are and where we’re headed. Lecture Notes in Computer Science, 1625:230–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivanescu:1999:FCU

[770] M. Ivanescu and V. Stoian. Fuzzy controllers by unconventional technologies for tentacle arms. Lecture Notes in Computer Science, 1625:232–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonkovic:1999:CRA

[771] M. Bonkovic, D. Stipanicev, and M. Stula. Control of robot arm approach by fuzzy pattern comparison technique. Lecture Notes in Computer Science, 1625:246–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varachiu:1999:FSC

[772] N. Varachiu. A fuzzy shapes characterization for robotics. Lecture Notes in Computer Science, 1625:253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attolico:1999:ABA

[773] G. Attolico, A. Itta, G. Cicirelli, and T. D’Orazio. ART-based automatic generation of membership functions for fuzzy controllers in robotics. Lecture Notes in Computer Science, 1625:259– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trajkovski:1999:TTL

[774] G. Trajkovski and B. Cukic. On two types of L0M fuzzy lattices. Lecture Notes in Computer Science, 1625:272– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mesiar:1999:GCM

[775] R. Mesiar. Generated connectives in many valued logic. Lecture Notes in Computer Science, 1625:282–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baczynski:1999:CCF

[776] M. Baczynski and J. Drewniak. Conjugacy classes of fuzzy implications. Lecture Notes in Computer Science, 1625: 287–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baczynski:1999:CDI

[777] M. Baczynski. Characterization of dienes implication. Lecture Notes in Computer Science, 1625:299–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aizenberg:1999:NNB

[778] I. N. Aizenberg. Neural networks based on multi-valued and universal binary neurons: Theory, application to image processing and recognition. Lecture Notes in Computer Science, 1625:306–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shabani:1999:MTT

[779] A. Shabani, M. B. Menhaj, and H. B. Tabrizi. Modeling of thermal two dimensional free turbulent jet by a three layer two time scale cellular neural network. Lecture Notes in Computer Science, 1625:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neagoe:1999:NAD

[780] V. Neagoe, M. Valcu, and B. Sabac. A neural approach for detection of road direction in autonomous navigation. Lecture Notes in Computer Science, 1625:324–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neagoe:1999:NSM

[781] V. Neagoe and I. Fratila. A neural segmentation of multispectral satellite images. Lecture Notes in Computer Science, 1625:334–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cao:1999:DTJ

[782] Y. J. Cao, N. Ireson, L. Bull, and R. Miles. Design of a traffic junction controller using classifier system and fuzzy logic. Lecture Notes in Computer Science, 1625:342–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niittymaeki:1999:UFL

[783] J. Niittymaeki. Using fuzzy logic to control traffic signals at multi-phase intersections. Lecture Notes in Computer Science, 1625:354–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gharieb:1999:FCN

[784] W. Gharieb. Fuzzy control to nonminimal phase processes. Lecture Notes in Computer Science, 1625:363–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kodagoda:1999:RCF

[785] K. R. S. Kodagoda, W. S. Wijesoma, and E. K. Teoh. Robust un-coupled fuzzy controller for longitudinal and lateral control of an AGV. Lecture Notes in Computer Science, 1625:370– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Precup:1999:CMT

[786] R.-E. Precup, S. Preitl, and S. Solyom. Center manifold theory approach to the stability analysis of fuzzy control systems. Lecture Notes in Computer Science, 1625:382–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knicker:1999:SAF

[787] R. Knicker. Stability analysis of fuzzy and other nonlinear systems based on the method of convex decomposition. Lecture Notes in Computer Science, 1625:391–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:1999:FUC

[788] H. Deng and C.-H Yeh. Fuzzy utilities comparison in multicriteria analysis. Lecture Notes in Computer Science, 1625:401–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huellermeier:1999:PFC

[789] E. Huellermeier. A possibilistic formalization of case-based reasoning and decision making. Lecture Notes in Computer Science, 1625:411–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1999:PDN

[790] M. Thomas, C. Pacha, and K. Goser. Parameter determination for nanoscale modelling. Lecture Notes in Computer Science, 1625:421–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilsdorf:1999:ORA

[791] I. Gilsdorf and W. Brauer. Optimizing routing algorithms in telecommunication networks with neural networks and fuzzy logic. Lecture Notes in Computer Science, 1625:427–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolle:1999:OWR

[792] L. Nolle, A. Armstrong, A. Hopgood, and A. Ware. Optimum work roll profile selection in the hot rolling of wide steel strip using computational intelligence. Lecture Notes in Computer Science, 1625:435–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reuter:1999:AEE

[793] M. Reuter and C. Zemke. Analysing epileptic events on-line by softcomputing-systems. Lecture Notes in Computer Science, 1625:453–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacprzyk:1999:ILS

[794] J. Kacprzyk and S. Zadrozny. On interactive linguistic summarization of databases via a fuzzy-logic-based querying add-on to Microsoft Access[R]. Lecture Notes in Computer Science, 1625:462–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahmed:1999:GPF

[795] M. Ahmed, E. Damiani, and A. G. B. Tettamanzi. A general-purpose fuzzy engine for crop control. Lecture Notes in Computer Science, 1625:473–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vascak:1999:FCP

[796] J. Vascak. Fuzzy control of a physical double inverted pendulum model. Lecture Notes in Computer Science, 1625: 482–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wijesoma:1999:SSF

[797] W. S. Wijesoma and K. R. S. Kodagoda. Synthesis of stable fuzzy PD/PID control laws for robotic manipulators from a variable structure systems standpoint. Lecture Notes in Computer Science, 1625:495–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosc:1999:DSB

[798] P. Bosc, O. Pivert, and L. Ughetto. On data summaries based on gradual rules. Lecture Notes in Computer Science, 1625:512–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marshall:1999:WTC

[799] P. Marshall and Z. Bandar. Working towards connectionist modeling of term formation. Lecture Notes in Computer Science, 1625:522–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reuter:1999:QMN

[800] M. Reuter. A quantization-model of the neural nets. Lecture Notes in Computer Science, 1625:530–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falkowski:1999:RAU

[801] B.-J. Falkowski. Risk analysis using perceptrons and quadratic programming. Lecture Notes in Computer Science, 1625:543–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Praczyk:1999:FRP

[802] J. Praczyk, H. Kiendl, and T. Slawinski. Finding relevant process characteristics with a method for data-based complexity reduction. Lecture Notes in Computer Science, 1625:548–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martyna:1999:TCA

[803] J. Martyna. Traffic control in ail ATM network using rough set theory. Lecture Notes in Computer Science, 1625:556–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butkiewicz:1999:ECT

[804] B. S. Butkiewicz, T. Mroczek, and M. W. Grzybek. Evaluation of characteristic temperatures of materials using an approximate reasoning method. Lecture Notes in Computer Science, 1625: 569–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelletti:1999:HCA

[805] R. Castelletti, E. Damiani, G. Righini, and R. Khosla. A human centered architecture for distributed retrieval of medical images. Lecture Notes in Computer Science, 1625:578–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reuter:1999:RRT

[806] M. Reuter and D. P. F. Moeller. Representing the real-time behaviour of technical processes in neural nets by using the phase-space-flow of the degrees of freedom of the nets. Lecture Notes in Computer Science, 1625:589–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kromer:1999:SNN

[807] T. Kromer. Spatial neural networks based on fractal algorithms biomorph nets of nets of. Lecture Notes in Computer Science, 1625:603–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dafinca:1999:ACS

[808] L. Dafinca. Adaptive control systems based on neural networks. Lecture Notes in Computer Science, 1625:615– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buzoianu:1999:FAU

[809] M. Buzoianu, F. Oltean, and A. Agapie. Function approximation using tensor product Bernstein polynomials — neuro and evolutionary approaches. Lecture Notes in Computer Science, 1625:625–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filip:1999:TCA

[810] I. Filip, O. Prostean, and D. Curiac. Tuning considerations above a fuzzy controller used for the synchronous generator. Lecture Notes in Computer Science, 1625:632–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamata:1999:SGA

[811] M. T. Lamata, J. I. Pelaez, J. C. Sierra, and J. M. Bravo. A sectoring genetic algorithm for the urban waste collection problem. Lecture Notes in Computer Science, 1625:642–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kraeht:1999:POG

[812] C. Kraeht, H. Geyer, P. Ulbig, and S. Schulz. Parameter optimization of group contribution methods in high dimensional solution spaces. Lecture Notes in Computer Science, 1625:647– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djamschidi:1999:AHA

[813] R. Djamschidi and A. Bruckner. Applying heuristic algorithms on structuring europeanwide distribution networks. Lecture Notes in Computer Science, 1625:658–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigore:1999:CNS

[814] Ovidiu Grigore and Octavian Grigore. The control of a nonlinear system using neural networks. Lecture Notes in Computer Science, 1625:667–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krause:1999:DUH

[815] P. Krause. Demonstration of the use of hyperinference: Position control using an electric disc-motor. Lecture Notes in Computer Science, 1625:671–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astanin:1999:SKB

[816] S. V. Astanin and T. G. Kalashnikova. Synthesis of knowledges based on inductive generalization. Lecture Notes in Computer Science, 1625:676–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halinka:1999:ABF

[817] A. Halinka, M. Szewczyk, and B. Witek. ANN-based fault type and location identification system for autonomous preventive-restoration control of complex electrical power system plants. Lecture Notes in Computer Science, 1625:682–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachiller:1999:OHS

[818] P. Bachiller, R. M. Perez, P. Martinez, and P. L. Aguilar. Optimal hidden structure for feedforward neural networks. Lecture Notes in Computer Science, 1625:684–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uchida:1999:RBI

[819] H. Uchida, Y. Hata, K. Suei, and H. Nakagawa. Relation between the IADL and physical fitness tests focusing on uncertainty of answering questionnaire in elderly women. Lecture Notes in Computer Science, 1625:686– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doeben-Henisch:1999:TAB

[820] C. Doeben-Henisch and J. P. Hasebrook. A toolbox of artificial brain cells to simulate classical and operant learning behavior. Lecture Notes in Computer Science, 1625:689–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halinka:1999:IAB

[821] A. Halinka, M. Szewczyk, and B. Witek. Intelligent, ANN-based identification system of complex generating sets operating mode. Lecture Notes in Computer Science, 1625:692–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCock:1999:NNB

[822] M. De Cock, F. Vynckier, and E. E. Kerre. A neural network based on linguistic modifiers. Lecture Notes in Computer Science, 1625:694–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neagu:1999:FCM

[823] C.-D. Neagu and V. Palade. Fuzzy computing in a MultiPurpose neural network implementation. Lecture Notes in Computer Science, 1625:697–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yardimci:1999:FLC

[824] A. Yardimci and A. S. Onural. Fuzzy logic control of child blood pressure during anesthesia. Lecture Notes in Computer Science, 1625:701–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dediu:1999:SCG

[825] A. H. Dediu, A. Agapie, and N. Varachiu. Soft computing genetic tool V 3.0 — applications. Lecture Notes in Computer Science, 1625:704– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varshavsky:1999:LEC

[826] V. Varshavsky and V. Marakhovsky. Learning experiments with CMOS artificial neuron. Lecture Notes in Computer Science, 1625:706–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobsen:1999:UPC

[827] Ivar Jacobsen. The unified process for component-based development. Lecture Notes in Computer Science, 1626: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260001.
pdf.

Scheer:1999:BPM

[828] August-Wilhelm Scheer and Michael Hoffmann. From business process model to application system — developing an information system with the House of Business Engineering (HOBE). Lecture Notes in Computer Science, 1626:2–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260002.
pdf.

Melloul:1999:CPS

[829] Laurence Melloul, Dorothea Beringer, Neal Sample, and Gio Wiederhold. CPAM, A protocol for software composition. Lecture Notes in Computer Science, 1626:11–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260011.
pdf.

Matthes:1999:POA

[830] Florian Matthes, Holm Wegner, and Patrick Hupe. A processoriented approach to software component definition. Lecture Notes in Computer Science, 1626:26–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260026.
pdf.

Heuvel:1999:CBO

[831] Willem-Jan van den Heuvel, Mike Papazoglou, and Manfred A. Jeusfeld. Configuring business objects from legacy systems. Lecture Notes in Computer Science, 1626:41–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260041.
pdf.

Verhoef:1999:RMI

[832] Denis Verhoef and Marcel Franckson. Risk management for IT in the large. Lecture Notes in Computer Science, 1626:57–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1626/16260057.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1626/16260057.
pdf.

Sutcliffe:1999:LBM

[833] Alistair G. Sutcliffe and Shailey Minocha. Linking business modelling to socio-technical system design. Lecture Notes in Computer Science, 1626: 73–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260073.
                                        pdf.

Joeris:1999:TFH

[834] Gregor Joeris and Otthein Herzog. Towards flexible and high-level modeling and enacting of processes. Lecture Notes in Computer Science, 1626:88– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260088.
                                        pdf.

Ralyte:1999:MES

[835] Jolita Ralyté, Colette Rolland, and Véronique Plihon. Method enhancement with scenario based techniques. Lecture Notes in Computer Science, 1626:103–118, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1626/16260103.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1626/16260103.pdf.

Becker-Kornstaedt:1999:SPE

[836] Ulrike Becker-Kornstaedt, Dirk Hamann, Ralf Kempkens, Peter Rösch, Martin Verlage, Richard Webby, and Jörg Zettel. Support for the process engineer: The spearmint approach to software process definition and process guidance. Lecture Notes in Computer Science, 1626:119– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260119.
pdf.

Ansorge:1999:MCD

[837] Dirk Ansorge, Klaus Bergner, Bernd Deifel, Nicolas Hawlitzky, Christoph Maier, Barbara Paech, Andreas Rausch, Marc Sihling, Veronika Thurner, and Sascha Vogel. Managing componentware development — software reuse and the V - Modell process. Lecture Notes in Computer Science, 1626:134–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260134.
pdf.

Wietek:1999:MMD

[838] Frank Wietek. Modelling multidimensional data in a dataflow-based visual data analysis environment. Lecture Notes in Computer Science, 1626:149– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260149.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260149.
                                        pdf.

Vassiliadis:1999:TQO

[839] Panos Vassiliadis, Mokrane Bouzeghoub, and Christoph Quix. Towards quality-oriented data warehouse usage and evolution. Lecture Notes in Computer Science, 1626:164–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260164.
                                        pdf.

Theodoratos:1999:DGD

[840] Dimitri Theodoratos, Spyros Ligoudistianos, and Timos Sellis. Designing the global data warehouse with SPJ views. Lecture Notes in Computer Science, 1626:180–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260180.
                                        pdf.

Sadiq:1999:AGR

[841] Wasim Sadiq and Maria E. Orlowska. Applying graph reduction techniques for identifying structural conflicts in process models. Lecture Notes in Computer Science, 1626:195– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1626/16260195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260195.
pdf.

Hesse:1999:MVA

[842] Wolfgang Hesse and Jörg Noack. A multi-variant approach to software process modelling. Lecture Notes in Computer Science, 1626:210–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260210.
pdf.

Green:1999:OAI

[843] Peter Green and Michael Rosemann. An ontological analysis of integrated process modelling. Lecture Notes in Computer Science, 1626:225–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260225.
pdf.

Sandholm:1999:DOC

[844] Thomas Sandholm, Stefan Tai, Dirk Slama, and Eamon Walshe. Design of object caching in a CORBA OTM system. Lecture Notes in Computer Science, 1626:241–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/1626/16260241.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1626/16260241.pdf.

Jungfer:1999:CIV

[845] Kim Jungfer, Ulf Leser, and Patricia Rodriguez-Tomé. Constructing IDL views on relational databases. Lecture Notes in Computer Science, 1626:255– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260255.
                                        pdf.

Oh:1999:DCT

[846] Am suk Oh. The design of cooperative transaction model by using client-server architecture. Lecture Notes in Computer Science, 1626:269– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260269.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260269.
                                        pdf.

Kang:1999:MSW

[847] Myong H. Kang, Judith N. Froscher, Amit P. Sheth, Krys J. Kochut, and John A. Miller. A multilevel secure workflow management system. Lecture Notes in Computer Science, 1626:271– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1626/16260271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260271.
pdf.

Eder:1999:TCW

[848] Johann Eder, Euthimios Panagos, and Michael Rabinovich. Time constraints in workflow systems. Lecture Notes in Computer Science, 1626:286– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260286.
pdf.

Sellentin:1999:TCS

[849] Jürgen Sellentin, Aiko Frank, and Bernhard Mitschang. TOGA — A customizable service for datacentric collaboration. Lecture Notes in Computer Science, 1626:301–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260301.
pdf.

DeFerreiraRezende:1999:PAA

[850] F. De Ferreira Rezende, U. Hewnsen, G. De Sa Oliveira, and R. C. G. Pereira. A practical approach to access heterogeneous and distributed databases. Lecture Notes in Computer Science, 1626:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rezende:1999:PAA

[851] Fernando de Ferreira Rezende, Ulrich Hermsen, Georgiane de Sá Oliveira, Renata Costa Guedes Pereira, and Jochen Rütschlin. A practical approach to access heterogeneous and distributed databases. Lecture Notes in Computer Science, 1626:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260317.
                                        pdf.

Brien:1999:UAI

[852] Peter Mc. Brien and Alexandra Poulovassilis. A uniform approach to inter-model transformations. Lecture Notes in Computer Science, 1626:333– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260333.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260333.
                                        pdf.

McBrien:1999:UAI

[853] P. McBrien and A. Poulovassilis. A uniform approach to inter-model transformations. Lecture Notes in Computer Science, 1626:333–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbeau:1999:OOH

[854] Franck Barbeau and José Martinez. OTHY: Object to HYpermedia. Lecture Notes in Computer Science, 1626:

349–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260349.
pdf.

Lesperance:1999:MDD

[855] Yves Lespérance, Todd G. Kelley, John Mylopoulos, and Eric S. K. Yu. Modeling dynamic domains with ConGolog. Lecture Notes in Computer Science, 1626:365–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260365.
pdf.

Aoumeur:1999:TOP

[856] Nasreddine Aoumeur and Gunter Saake. Towards an object Petri nets model for specifying and validating distributed information systems. Lecture Notes in Computer Science, 1626:381– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260381.
pdf.

Olive:1999:RRT

[857] Antoni Olivé. Relationship reification: A temporal view. Lecture Notes in Computer Science, 1626:396– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260396.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260396.
                                        pdf.

Becker:1999:TCF

[858] Jörg Becker and Michael zur Mühlen. Towards a classification framework for application granularity in workflow management systems. Lecture Notes in Computer Science, 1626:411– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260411.
                                        pdf.

Klingemann:1999:AOC

[859] Justus Klingemann, Jürgen Wäsch, and Karl Aberer. Adaptive outsourcing in cross-organizational workflows. Lecture Notes in Computer Science, 1626: 417–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1626/16260417.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1626/16260417.
                                        pdf.

Huan:1999:PBR

[860] Yan-Nong Huan and Ming-Chien Shan. Policy-based resource management. Lecture Notes in Computer Science, 1626:422–??, 1999. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1626/16260422.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1626/16260422.
pdf.

Huang:1999:PBR

[861] Y.-N. Huang and M.-C. Shan. Policybased resource management. Lecture Notes in Computer Science, 1626:422– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prakash:1999:MMH

[862] Naveen Prakash and Ritu Sibal. Modelling method heuristics for better quality products. Lecture Notes in Computer Science, 1626:429–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260429.
pdf.

Calvanese:1999:QCS

[863] Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. Queries and constraints on semistructured data. Lecture Notes in Computer Science, 1626:434–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260434.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1626/16260434.
pdf.

Bornhovd:1999:PMB

[864] Christof Bornhövd and Alejandro P. Buchmann. A prototype for metadatabased integration of Internet sources. Lecture Notes in Computer Science, 1626:439–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1626/16260439.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1626/16260439.
pdf.

Weske:1999:WMT

[865] Mathias Weske. Workflow management through distributed and persistent CORBA workflow objects. Lecture Notes in Computer Science, 1626: 446–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260446.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260446.
pdf.

Jarzabek:1999:CCI

[866] Stan Jarzabek. Component criteria for information system families. Lecture Notes in Computer Science, 1626:451– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260451.
pdf.

Svinterikou:1999:TMM

[867] Marianthi Svinterikou and Babis Theodoulidis. TUML: A method for modelling temporal information systems. Lecture Notes in Computer Science, 1626:456–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1626/16260456.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1626/16260456.pdf.

Yi:1999:BGR

[868] Choong ho Yi and Paul Johannesson. Beyond goal representation: Checking goal-satisfaction by temporal reasoning with business processes. Lecture Notes in Computer Science, 1626:462– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260462.
pdf.

Wedemeijer:1999:DFM

[869] Lex Wedemeijer. Design the flexibility, maintain the stability of conceptual schemas. Lecture Notes in Computer Science, 1626:467–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1626/16260467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1626/16260467.
pdf.

Diaz:1999:MAD

[870] Oscar Díaz and Mario Piattini. Metrics for active database maintainability. Lecture Notes in Computer Science, 1626:472–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1626/16260472.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1626/16260472.
pdf.

Kleinberg:1999:WGM

[871] Jon M. Kleinberg, Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, and Andrew S. Tomkins. The Web as a graph: Measurements, models and methods. Lecture Notes in Computer Science, 1627:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270001.
pdf.

Tarui:1999:SOC

[872] Jun Tarui and Seinosuke Toda. Some observations on the computational complexity of graph accessibility problem (extended abstract). Lecture Notes in Computer Science, 1627:18– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270018.
pdf.

Nagamochi:1999:AFS

[873] Hiroshi Nagamochi and Toshihide Ibaraki. An approximation for finding a smallest 2-Edge-Connected subgraph containing a specified spanning tree. Lecture Notes in Computer Science, 1627:31–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1627/16270031.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1627/16270031.
pdf.

Takaoka:1999:TH

[874] Tadao Takaoka. Theory of 2-3 heaps. Lecture Notes in Computer Science, 1627:41–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1627/16270041.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1627/16270041.
pdf.

Hutchinson:1999:EMD

[875] David Hutchinson, Anil Maheshwari, and Norbert Zeh. An external memory data structure for shortest path queries (extended abstract). Lecture Notes in Computer Science, 1627:51– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270051.
pdf.

Hon:1999:ANN

[876] Wing-Kai Hon and Tak-Wah Lam. Approximating the nearest neighbor interchange distance for evolutionary trees with non-uniform degrees. Lecture Notes in Computer Science, 1627:61– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270061.
pdf.

Lin:1999:SGR

[877] Guo-Hui Lin and Guoliang Xue. Signed genome rearrangement by reversals and transpositions: Models and approximations. Lecture Notes in Computer Science, 1627:71–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270071.
pdf.

Yamaguchi:1999:AAT

[878] Atsuko Yamaguchi and Akihiro Sugimoto. An approximation algorithm for the two-layered graph drawing problem. Lecture Notes in Computer Science, 1627:81–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270081.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1627/16270081.
pdf.

Lin:1999:AMG

[879] Xuemin Lin and Peter Eades. Area minimization for grid visibility representation of hierarchically planar graphs. Lecture Notes in Computer Science, 1627:92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1627/16270092.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1627/16270092.pdf.

Diaz:1999:LPL

[880] Josep Díaz, Mathew D. Penrose, Jordi Petit, and María Serna. Layout problems on lattice graphs. Lecture Notes in Computer Science, 1627:103– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270103.
pdf.

Cai:1999:NTT

[881] Jin-Yi Cai. A new transference theorem in the geometry of numbers. Lecture Notes in Computer Science, 1627: 113–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270113.
pdf.

Damm:1999:CRP

[882] Carsten Damm, Ki Hang Kim, and Fred Roush. On covering and rank problems for Boolean matrices and their applications. Lecture Notes in Computer Science, 1627:123–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270123.
pdf.

Mahajan:1999:CAP

[883] Meena Mahajan, P. R. Subramanya, and V. Vinay. A combinatorial algorithm for pfaffians. Lecture Notes in Computer Science, 1627:134–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270134.
pdf.

Nardelli:1999:HSF

[884] Enrico Nardelli, Guido Proietti, and Peter Widmayer. How to swap a failing edge of a single source shortest paths tree. Lecture Notes in Computer Science, 1627:144–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270144.
pdf.

Bezrukov:1999:BPG

[885] S. L. Bezrukov, R. Elsässer, and U.- P. Schroeder. On bounds for the k-partitioning of graphs. Lecture Notes in Computer Science, 1627:154– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270154.
pdf.

Nagamochi:1999:FACb

[886] Hiroshi Nagamochi, Shigeki Katayama, and Toshihide Ibaraki. A faster algorithm for computing minimum 5-way and 6-way cuts in graphs. Lecture Notes in Computer Science, 1627:164– 173, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270164.
pdf.

Ambainis:1999:PAL

[887] Andris Ambainis, Richard Bonner, Rusins Freivalds, and Arnolds Kikusts. Probabilities to accept languages by quantum finite automata. Lecture Notes in Computer Science, 1627:174– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270174.
pdf.

Fortnow:1999:DHL

[888] Lance Fortnow, A. Pavan, and Alan L. Selman. Distributionallyhard languages. Lecture Notes in Computer Science, 1627:184–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270184.
                                        pdf.

McKenzie:1999:CCF

[889] Pierre McKenzie, Klaus Reinhardt, and V. Vinay. Circuits and contextfree languages. Lecture Notes in Computer Science, 1627:194–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270194.
                                        pdf.

Amano:1999:NLC

[890] Kazuyuki Amano, Akira Maruoka, and Jun Tarui. On the negation-limited circuit complexity of merging. Lecture Notes in Computer Science, 1627:204– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270204.
                                        pdf.

Miltersen:1999:SPV

[891] Peter Bro Miltersen, N. V. Vinodchandran, and Osamu Watanabe. Super-polynomial versus halfexponential circuit size in the exponential hierarchy. Lecture Notes in Computer Science, 1627:210–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270210.
pdf.

Fernau:1999:ELS

[892] Henning Fernau. Efficient learning of some linear matrix languages. Lecture Notes in Computer Science, 1627:221– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270221.
pdf.

Deng:1999:MMR

[893] Xiaotie Deng and Yuzhong Zhang. Minimizing mean response time in batch processing system. Lecture Notes in Computer Science, 1627:231– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270231.
pdf.

Xiaotie:1999:MMR

[894] D. Xiaotie and Z. Yuzhong. Minimizing mean resoponse time in batch processing system. Lecture Notes in Computer Science, 1627:231–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krysta:1999:AAB

[895] Piotr Krysta and Roberto Solis-Oba. Approximation algorithms for bounded facility location. Lecture Notes in Computer Science, 1627:241– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270241.
                                        pdf.

Tayu:1999:STH

[896] Satoshi Tayu. Scheduling trees onto hypercubes and grids is N P-complete. Lecture Notes in Computer Science, 1627:251–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1627/16270251.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270251.
                                        pdf.

Halldorsson:1999:AWI

[897] Magnús M. Halldórsson. Approximations of weighted independent set and hereditary subset problems. Lecture Notes in Computer Science, 1627:261– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270261.
pdf.

Halldorsson:1999:MCT

[898] Magnús M. Halldórsson, Guy Kortsarz, and Andrzej Proskurowski. Multi-coloring trees. Lecture Notes in Computer Science, 1627:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270271.
pdf.

Clementi:1999:CAC

[899] Andrea E. F. Clementi, Pierluigi Crescenzi, and Gianluca Rossi. On the complexity of approximating coloredgraph problems. Lecture Notes in Computer Science, 1627:281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270281.
pdf.

Bernasconi:1999:AST

[900] Anna Bernasconi, Carsten Damm, and Igor E. Shparlinski. On the average sensitivity of testing square-free numbers. Lecture Notes in Computer Science, 1627:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270291.
                                        pdf.

Xizhong:1999:BER

[901] Z. Xizhong. Binary enumerability of real numbers. Lecture Notes in Computer Science, 1627:300–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1999:BER

[902] Xizhong Zheng. Binary enumerability of real numbers. Lecture Notes in Computer Science, 1627:300–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270300.
                                        pdf.

Cooperman:1999:GMI

[903] Gene Cooperman, Sandra Feisel, Joachim von zur Gathen, and George Havas. GCD of many integers. Lecture Notes in Computer Science, 1627: 310–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270310.
                                        pdf.

Jurdzinski:1999:MPF

[904] Tomasz Jurdzinski, Mirosław Kutyłowski, and Krzysztof Lorys. Multi-party finite computations. Lecture Notes in Computer Science, 1627:318–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270318.
pdf.

Nakata:1999:PLM

[905] Toshio Nakata, Hiroshi Imahayashi, and Masafumi Yamashita. Probabilistic local majority voting for the agreement problem on finite graphs. Lecture Notes in Computer Science, 1627: 330–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270330.
pdf.

Marzetta:1999:DPB

[906] Ambros Marzetta and Adrian Brüngger. A dynamic programming bound for the quadratic assignment problem. Lecture Notes in Computer Science, 1627:339– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270339.
pdf.

Uno:1999:NAS

[907] Takeaki Uno. A new approach for speeding up enumeration algorithms and its application for matroid bases. Lecture Notes in Computer Science, 1627:349–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270349.
pdf.

Cai:1999:RCG

[908] Jin-Yi Cai, George Havas, Bernard Mans, Ajay Nerurkar, Jean-Pierre Seifert, and Igor Shparlinski. On routing in circulant graphs. Lecture Notes in Computer Science, 1627:360– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270360.
pdf.

Matsubayashi:1999:MCE

[909] Akira Matsubayashi and Ryo Takasu. Minimum congestion embedding of complete binary trees into tori. Lecture Notes in Computer Science, 1627: 370–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270370.
pdf.

Chin:1999:MSL

[910] Francis Y. L. Chin, Cao An Wang, and Fu Lee Wang. Maximum stabbing line in 2D plane. Lecture Notes in Computer Science, 1627:379–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270379.
pdf.

Chaudhuri:1999:GSL

[911] Jeet Chaudhuri and Subhas C. Nandy. Generalized shooter location problem. Lecture Notes in Computer Science, 1627:389–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270389.
pdf.

Hong:1999:COA

[912] Sung-Pil Hong. A competitive online algorithm for the paging problem with “shelf” memory. Lecture Notes in Computer Science, 1627:400– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270400.
pdf.

Iwama:1999:UGF

[913] Kazuo Iwama and Kouki Yonezawa. Using generalized forecasts for online currency conversion. Lecture Notes in Computer Science, 1627:409– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270409.
                                        pdf.

Otto:1999:RPR

[914] Friedrich Otto. On S-regular prefix-rewriting systems and automatic structures. Lecture Notes in Computer Science, 1627:422–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270422.
                                        pdf.

Hirata:1999:TIS

[915] Kouichi Hirata, Keizo Yamada, and Masateru Harao. Tractable and intractable second-order matching problems. Lecture Notes in Computer Science, 1627:432–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1627/16270432.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1627/16270432.pdf.

Lee:1999:EFS

[916] Sung-Woo Lee, Hyun-Sung Kim, Jung-Joon Kim, Tae-Geun Kim, and Kee-Young Yoo. Efficient fixedsize systolic arrays for the modular multiplication. Lecture Notes in Computer Science, 1627:442–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270442.
pdf.

Chong:1999:IPC

[917] Ka Wong Chong, Yijie Han, Yoshihide Igarashi, and Tak Wah Lam. Improving parallel computation with fast integer sorting. Lecture Notes in Computer Science, 1627:452–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270452.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1627/16270452.
pdf.

Das:1999:CAP

[918] Sajal K. Das, Bhabani P. Sinha, and Rajarshi Chaudhuri. A combinatorial approach to performance analysis of a shared-memory multiprocessor. Lecture Notes in Computer Science, 1627: 462–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1627/16270462.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1627/16270462.
                                        pdf.

Gudmundsson:1999:FAA

[919] Joachim Gudmundsson and Christos Levcopoulos. A fast approximation algorithm for TSP with neighborhoods and red-blue separation. Lecture Notes in Computer Science, 1627: 473–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270473.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270473.
                                        pdf.

Lau:1999:GBE

[920] H. Y. Lau and H. F. Ting. The greedier the better: An efficient algorithm for approximating maximum independent set. Lecture Notes in Computer Science, 1627:483–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1627/16270483.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1627/16270483.
                                        pdf.

Hoare:1999:TMP

[921] C. A. R. Hoare and He Jifeng. A trace model for pointers and objects. Lecture Notes in Computer Science, 1628: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1628/16280001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1628/16280001.
pdf.

Czarnecki:1999:SO

[922] Krzysztof Czarnecki and Ulrich W. Eisenecker. Synthesizing objects. Lecture Notes in Computer Science, 1628: 18–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280018.
pdf.

Bono:1999:CCC

[923] Viviana Bono, Amit Patel, and Vitaly Shmatikov. A core calculus of classes and mixins. Lecture Notes in Computer Science, 1628:43–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280043.
pdf.

Ernst:1999:PCM

[924] Erik Ernst. Propagating class and method combination. Lecture Notes in Computer Science, 1628:67–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280067.
pdf.

Dieckman:1999:SAB

[925] Sylvia Dieckman and Urs Hölzle. A study of the allocation behavior of the SPECjvm98 Java benchmarks. Lecture Notes in Computer Science, 1628:92– 115, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1628/16280092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1628/16280092.
                                        pdf.

DePauw:1999:VRP

[926] Wim De Pauw and Gary Sevitsky. Visualizing reference patterns for solving memory leaks in Java. Lecture Notes in Computer Science, 1628:116– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1628/16280116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1628/16280116.
                                        pdf.

Lencevicius:1999:DQB

[927] Raimondas Lencevicius, Urs Hölzle, and Ambuj K. Singh. Dynamic query-based debugging. Lecture Notes in Computer Science, 1628:135–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1628/16280135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1628/16280135.
                                        pdf.

Igarashi:1999:FVT

[928] Atsushi Igarashi and Benjamin C. Pierce. Foundations for virtual types. Lecture Notes in Computer Science, 1628:161–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1628/16280161.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1628/16280161.
pdf.

Thorup:1999:UGC

[929] Kresten Krab Thorup and Mads Torgersen. Unifying genericity: Combining the benefits of virtual types and parameterized classes. Lecture Notes in Computer Science, 1628:186– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280186.
pdf.

Greenhouse:1999:OOE

[930] Aaron Greenhouse and John Boyland. An object-oriented effects system. Lecture Notes in Computer Science, 1628: 205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280205.
pdf.

Liskov:1999:PPO

[931] Barbara Liskov, Miguel Castro, Liuba Shrira, and Atul Adya. Providing persistent objects in distributed systems. Lecture Notes in Computer Science, 1628:230–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1628/16280230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1628/16280230.
                                        pdf.

Detlefs:1999:IVM

[932] David Detlefs and Ole Agesen. Inlining of virtual methods. Lecture Notes in Computer Science, 1628:258– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1628/16280258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1628/16280258.
                                        pdf.

Millstein:1999:MST

[933] Todd Millstein and Craig Chambers. Modular statically typed multimethods. Lecture Notes in Computer Science, 1628:279–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1628/16280279.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1628/16280279.pdf.

Pang:1999:MMD

[934] Candy Pang, Wade Holst, Yuri Leontiev, and Duane Szafron. Multimethod dispatch using multiple row displacement. Lecture Notes in Computer Science, 1628:304–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280304.
pdf.

Kuehne:1999:IIE

[935] T. Kuehne. Internal iteration externalized. Lecture Notes in Computer Science, 1628:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuhne:1999:IIE

[936] Thomas Kühne. Internal iteration externalized. Lecture Notes in Computer Science, 1628:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280329.
pdf.

Kniesel:1999:TSD

[937] Günther Kniesel. Type-safe delegation for run-time component adaptation. Lecture Notes in Computer Science, 1628:351–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1628/16280351.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1628/16280351.
pdf.

Schultz:1999:TAS

[938] Ulrik Pagh Schultz, Julia L. Lawall, Charles Consel, and Gilles Muller. Towards automatic specialization of Java programs. Lecture Notes in Computer Science, 1628:367–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280367.
pdf.

Serrano:1999:WC

[939] Manuel Serrano. Wide classes. Lecture Notes in Computer Science, 1628: 391–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280391.
pdf.

Bertino:1999:ACS

[940] Elisa Bertino, Giovanna Guerrini, Isabella Merlo, and Marco Mesiti. An approach to classify semi-structured objects. Lecture Notes in Computer Science, 1628:416–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1628/16280416.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1628/16280416.pdf.

Waldo:1999:OOP

[941] Jim Waldo. Object-oriented programming on the network. Lecture Notes in Computer Science, 1628:441– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280441.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280441.
pdf.

Pandey:1999:PFG

[942] Raju Pandey and Brant Hashii. Providing fine-grained access control for Java programs. Lecture Notes in Computer Science, 1628:449–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1628/16280449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1628/16280449.
pdf.

Bastide:1999:FSP

[943] Rémi Bastide, Ousmane Sy, and Philippe Palanque. Formal specification and prototyping of CORBA systems. Lecture Notes in Computer Science, 1628:474–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1628/16280474.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        1628/16280474.pdf.

Gaspari:1999:PAS

[944] Mauro Gaspari and Gianluigi Zavattaro. A process algebraic specification of the new asynchronous CORBA messaging service. Lecture Notes in Computer Science, 1628:495–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1628/16280495.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1628/16280495.
                                        pdf.

Black:1999:OOP

[945] Andrew P. Black. Object-oriented programming: Regaining the excitement. Lecture Notes in Computer Science, 1628:519–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1628/16280519.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1628/16280519.
                                        pdf.

Anonymous:1999:AId

[946] Anonymous. Author index. Lecture Notes in Computer Science, 1628:529– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1628/1628ind.pdf.

Nieuwenhuis:1999:SFP

[947] Robert Nieuwenhuis and José Miguel Rivero. Solved forms for path ordering constraints. Lecture Notes in Computer Science, 1631:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310001.
pdf.

Dershowitz:1999:J

[948] Nachum Dershowitz and Subrata Mitra. Jeopardy. Lecture Notes in Computer Science, 1631:16–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310016.
pdf.

Visser:1999:SPM

[949] Eelco Visser. Strategic pattern matching. Lecture Notes in Computer Science, 1631:30–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1631/16310030.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1631/16310030.pdf.

DeGroote:1999:SNN

[950] P. De Groote. On the strong normalization of natural deduction with permutation-conversions. Lecture Notes in Computer Science, 1631:45– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groote:1999:SNN

[951] Philippe de Groote. On the strong normalization of natural deduction with permutation-conversions. Lecture Notes in Computer Science, 1631:45– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310045.
                                        pdf.

Oostrom:1999:NWO

[952] Vincent van Oostrom. Normalization in weakly orthogonal rewriting. Lecture Notes in Computer Science, 1631: 60–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310060.
                                        pdf.

VanOostrom:1999:NWO

[953] V. Van Oostrom. Normalization in weakly orthogonal rewriting. Lecture Notes in Computer Science, 1631:60– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosmo:1999:SNP

[954] Roberto Di Cosmo and Stefano Guerrini. Strong normalization of proof nets modulo structural congruences. Lecture Notes in Computer Science, 1631:

75–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310075.
pdf.

DiCosmo:1999:SNP

[955] R. Di Cosmo and S. Guerrini. Strong normalization of proof nets modulo structural congruences. Lecture Notes in Computer Science, 1631:75– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1999:HGD

[956] Bruno Courcelle. Hierarchical graph decompositions defined by grammars and logical formulas. Lecture Notes in Computer Science, 1631:90–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310090.
pdf.

Marcinkowski:1999:UPT

[957] Jerzy Marcinkowski. Undecidability of the ∃xx∀xx part of the theory of ground term algebra modulo an AC symbol. Lecture Notes in Computer Science, 1631:92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310092.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310092.
                                        pdf.

Caron:1999:DSQ

[958] Anne-Cécile Caron, Franck Seynhaeve, Sophie Tison, and Marc Tommasi. Deciding the satisfiability of quantifier free formulae on one-step rewriting. Lecture Notes in Computer Science, 1631:103–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310103.
                                        pdf.

Limet:1999:NRA

[959] Sébastien Limet and Pierre Réty. A new result about the decidability of the existential one-step rewriting theory. Lecture Notes in Computer Science, 1631:118–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310118.
                                        pdf.

Rubio:1999:FSA

[960] Albert Rubio. A fully syntactic AC-RPO. Lecture Notes in Computer Science, 1631:133–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1631/16310133.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1631/16310133.pdf.

Stuber:1999:TPO

[961] Jürgen Stuber. Theory path orderings. Lecture Notes in Computer Science, 1631:148–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1631/16310148.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1631/16310148.
pdf.

Touzet:1999:CMR

[962] Hélène Touzet. A characterisation of multiply recursive functions with Higman’s lemma. Lecture Notes in Computer Science, 1631:163–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310163.
pdf.

Baader:1999:DWP

[963] Franz Baader and Cesare Tinelli. Deciding the word problem in the union of equational theories sharing constructors. Lecture Notes in Computer Science, 1631:175–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310175.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1631/16310175.
                                        pdf.

Bachmair:1999:NRC

[964] L. Bachmair, C. R. Ramakrishnan, I. V. Ramakrishnan, and A. Tiwari. Normalization via rewrite closures. Lecture Notes in Computer Science, 1631:190–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310190.
                                        pdf.

Hofbauer:1999:TSU

[965] Dieter Hofbauer and Maria Huber. Test sets for the universal and existential closure of regular tree languages. Lecture Notes in Computer Science, 1631: 205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310205.
                                        pdf.

Raamsdonk:1999:HOR

[966] Femke van Raamsdonk. Higherorder rewriting. Lecture Notes in Computer Science, 1631:220–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310220.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1631/16310220.
pdf.

VanRaamsdonk:1999:OR

[967] F. Van Raamsdonk. 1Higher-Order rewriting. Lecture Notes in Computer Science, 1631:220–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clavel:1999:MS

[968] M. Clavel, F. Durán, S. Eker, P. Lincoln, N. Martí-Oliet, J. Meseguer, and J. F. Quesada. The Maude system. Lecture Notes in Computer Science, 1631: 240–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310240.
pdf.

Fraguas:1999:MDS

[969] F. J. López Fraguas and J. Sánchez Hernández. T OY: A multiparadigm declarative system. Lecture Notes in Computer Science, 1631:244–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310244.
pdf.

Lopez-Fraguas:1999:TMD

[970] F. J. Lopez-Fraguas and J. Sanchez-Hernandez. TOY: A multiparadigm declarative system. Lecture Notes in Computer Science, 1631:244–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kepser:1999:USC

[971] Stephan Kepser and Jörn Richts. UniMoK: A system for combining equational unification algorithms. Lecture Notes in Computer Science, 1631:248– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310248.
pdf.

Verma:1999:LRT

[972] Rakesh Verma and Shalitha Senanayake. L̄R^{2}: A laboratory for rapid term graph rewriting. Lecture Notes in Computer Science, 1631:252–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310252.
pdf.

Nagaya:1999:DLL

[973] Takashi Nagaya and Yoshihito Toyama. Decidability for left-linear growing term rewriting systems. Lecture Notes in Computer Science, 1631:256–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310256.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1631/16310256.
pdf.

Giesl:1999:TCS

[974] Jürgen Giesl and Aart Middeldorp. Transforming context-sensitive rewrite systems. Lecture Notes in Computer Science, 1631:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310271.
pdf.

Ferreira:1999:CSA

[975] M. C. F. Ferreira and A. L. Ribeiro. Context-sensitive AC-rewriting. Lecture Notes in Computer Science, 1631: 286–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310286.
pdf.

Blanqui:1999:CAC

[976] Frédéric Blanqui, Jean-Pierre Jouannaud, and Mitsuhiro Okada. The calculus of algebraic constructions. Lecture Notes in Computer Science, 1631: 301–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310301.
pdf.

Dowek:1999:HIF

[977] Gilles Dowek, Thérèse Hardin, and Claude Kirchner. HOL-λσ: An intentional first-order expression of higher-order logic. Lecture Notes in Computer Science, 1631:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310317.
                                        pdf.

Otto:1999:CBR

[978] Friedrich Otto. On the connections between rewriting and formal language theory. Lecture Notes in Computer Science, 1631:332–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310332.
                                        pdf.

Frougny:1999:RSA

[979] Christiane Frougny and Jacques Sakarovitch. A rewrite system associated with quadratic Pisot units. Lecture Notes in Computer Science, 1631: 356–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1631/16310356.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1631/16310356.
                                        pdf.

Gobel:1999:FRS

[980] Manfred Göbel. Fast rewriting of symmetric polynomials. Lecture Notes in Computer Science, 1631:371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310371.
pdf.

Goebel:1999:FRS

[981] M. Goebel. Fast rewriting of symmetric polynomials. Lecture Notes in Computer Science, 1631:371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saubion:1999:ITS

[982] Frédéric Saubion and Igor Stéphan. On implementation of tree synchronized languages. Lecture Notes in Computer Science, 1631:382–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1631/16310382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1631/16310382.
pdf.

Anonymous:1999:AIe

[983] Anonymous. Author index. Lecture Notes in Computer Science, 1631:397– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1631/16310397.pdf.

DeGroote:1999:DPA

[984] P. De Groote. A dynamic programming approach to categorial deduction. Lecture Notes in Computer Science, 1632:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1999:TTM

[985] S. Demri and R. Gore. Tractable transformations from modal provability logics into first-order logic. Lecture Notes in Computer Science, 1632:16– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graedel:1999:ITD

[986] E. Graedel. Invited talk: Decision procedures for guarded logics. Lecture Notes in Computer Science, 1632:31– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tobies:1999:PAG

[987] S. Tobies. A PSpace algorithm for graded modal logic. Lecture Notes in Computer Science, 1632:52–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt-Schauss:1999:SCE

[988] M. Schmidt-Schauss and K. U. Schulz. Solvability of context equations with two context variables is decidable. Lecture Notes in Computer Science, 1632: 67–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wierzbicki:1999:CHO

[989] T. Wierzbicki. Complexity of the higher order matching. Lecture Notes in Computer Science, 1632:82–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pichler:1999:SEP

[990] R. Pichler. Solving equational problems efficiently. Lecture Notes in Computer Science, 1632:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adams:1999:VVS

[991] A. A. Adams, H. Gottliebsen, S. A. Linton, and U. Martin. VSDITLU: A verifiable symbolic definite integral table look-up. Lecture Notes in Computer Science, 1632:112–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janicic:1999:FFI

[992] P. Janicic, A. Bundy, and I. Green. A framework for the flexible integration of a class of decision procedures into theorem provers. Lecture Notes in Computer Science, 1632:127–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horacek:1999:PPH

[993] H. Horacek. Presenting proofs in a human-oriented way. Lecture Notes in Computer Science, 1632:142–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sofronie-Stokkermans:1999:UTV

[994] V. Sofronie-Stokkermans. On the universal theory of varieties of distributive lattices with operators: Some decidability and complexity results. Lecture Notes in Computer Science, 1632: 157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hustadt:1999:MCR

[995] U. Hustadt and R. A. Schmidt. Maslov’s class K revisited. Lecture Notes in Computer Science, 1632:172– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Areces:1999:PRR

[996] C. Areces, H. De Nivelle, and M. De Rijke. Prefixed resolution: A resolution method for modal and description logics. Lecture Notes in Computer Science, 1632:187–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfenning:1999:SDT

[997] F. Pfenning and C. Schuermann. System description: Twelf — A metalogical framework for deductive systems. Lecture Notes in Computer Science, 1632:202–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Autexier:1999:SDI

[998] S. Autexier, D. Hutter, H. Mantel, and A. Schairer. System description: inka 5.0 — A logic voyager. Lecture Notes in Computer Science, 1632:207–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1999:SDC

[999] M. Baaz, A. Leitsch, and G. Moser. System description: CutRes 0.1: Cut elimination by resolution. Lecture Notes in Computer Science, 1632:212– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franke:1999:SDM

[1000] A. Franke and M. Kohlhase. System description: MathWeb, an agent-based communication layer for distributed automated theorem proving. Lecture Notes in Computer Science, 1632:217– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gribomont:1999:SDU

[1001] E. Pascal Gribomont and N. Salloum.

System description: Using OBDD’s for the validation of Skolem verification conditions. Lecture Notes in Computer Science, 1632:222–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hickey:1999:FTD

[1002] J. Hickey. Fault-tolerant distributed theorem proving. Lecture Notes in Computer Science, 1632:227–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hillenbrand:1999:SDW

[1003] T. Hillenbrand, A. Jaeger, and B. Loechner. System description: Waldmeister — improvements in performance and ease of use. Lecture Notes in Computer Science, 1632:232–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felty:1999:FMU

[1004] A. P. Felty, D. J. Howe, and A. Roychoudhury. Formal metatheory using implicit syntax, and an application to data abstraction for asynchronous systems. Lecture Notes in Computer Science, 1632:237–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prost:1999:FSA

[1005] F. Prost. A formalization of static analyses in system F. Lecture Notes in Computer Science, 1632:252–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artemov:1999:ERT

[1006] S. N. Artemov. On explicit reflection in theorem proving and formal verification. Lecture Notes in Computer Science, 1632:267–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konrad:1999:SDK

[1007] K. Konrad and D. A. Wolfram. System description: Kimba, A model generator for many-valued first-order logics. Lecture Notes in Computer Science, 1632:282–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nadathur:1999:SDT

[1008] G. Nadathur and D. J. Mitchell. System description: Teyjus — A compiler and abstract machine based implementation of lambdaProlog. Lecture Notes in Computer Science, 1632:287– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riazanov:1999:V

[1009] A. Riazanov and A. Voronkov. Vampire. Lecture Notes in Computer Science, 1632:292–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1999:SAE

[1010] S. Schulz. System abstract: E0.3. Lecture Notes in Computer Science, 1632: 297–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieuwenhuis:1999:ITR

[1011] R. Nieuwenhuis. Invited talk: Rewritebased deduction and symbolic constraints. Lecture Notes in Computer Science, 1632:302–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weidenbach:1999:TAA

[1012] C. Weidenbach. Towards an automatic analysis of security protocols in firstorder logic. Lecture Notes in Computer Science, 1632:314–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1999:CCC

[1013] P. Baumgartner, N. Eisinger, and U. Furbach. A confluent connection calculus. Lecture Notes in Computer Science, 1632:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1999:ABR

[1014] M. Fuchs and D. Fuchs. Abstractionbased relevancy testing for model elimination. Lecture Notes in Computer Science, 1632:344–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bishop:1999:BFS

[1015] M. Bishop. A breadth-first strategy for mating search. Lecture Notes in Computer Science, 1632:359–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutter:1999:DCI

[1016] D. Hutter and A. Bundy. The design of the CADE-16 inductive theorem prover contest. Lecture Notes in Computer Science, 1632:374–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weidenbach:1999:SDS

[1017] C. Weidenbach, B. Afshordel, U. Brahm, and C. Cohrs. System description: Spass version 1.0.0. Lecture Notes in Computer Science, 1632:378–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1999:KTP

[1018] A. Voronkov. KK: A theorem prover for K. Lecture Notes in Computer Science, 1632:383–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whittle:1999:SDC

[1019] J. Whittle, A. Bundy, R. Boulton, and H. Lowe. System description: C0YNTHIA. Lecture Notes in Computer Science, 1632:388–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jian:1999:SDM

[1020] Z. Jian. System description: MCS:

Model-based conjecture searching. Lecture Notes in Computer Science, 1632: 393–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1999:ITE

[1021] T. Nipkow. Invited talk: Embedding programming languages in theorem provers. Lecture Notes in Computer Science, 1632:398–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benzmueller:1999:EHO

[1022] C. Benzmueller. Extensional higherorder paramodulation and RUE-Resolution. Lecture Notes in Computer Science, 1632:399–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopes:1999:AGP

[1023] R. H. C. Lopes. Automatic generation of proof search strategies for secondorder logic. Lecture Notes in Computer Science, 1632:414–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dill:1999:AAH

[1024] D. L. Dill. Alternative approaches to hardware verification. Lecture Notes in Computer Science, 1633:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sifakis:1999:CST

[1025] J. Sifakis. The compositional specification of timed systems — A tutorial. Lecture Notes in Computer Science, 1633:2–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1999:TA

[1026] R. Alur. Timed automata. Lecture Notes in Computer Science, 1633:8–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staalmarck:1999:SME

[1027] G. Staalmarck. Staalmarck’s method with extensions to quantified Boolean formulas. Lecture Notes in Computer Science, 1633:23–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1999:VPS

[1028] Z. Manna and H. B. Sipma. Verification of parameterized systems by dynamic induction on diagrams. Lecture Notes in Computer Science, 1633:25– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brinksma:1999:FMC

[1029] E. Brinksma. Formal methods for conformance testing: Theory can be practical! Lecture Notes in Computer Science, 1633:44–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosabettu:1999:PCP

[1030] R. Hosabettu, M. Srivas, and G. Gopalakrishnan. Proof of correctness of a processor with reorder buffer using the completion functions approach. Lecture Notes in Computer Science, 1633: 47–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biere:1999:VSP

[1031] A. Biere, E. Clarke, R. Raimi, and Y. Zhu. Verifying safety properties of a PowerPC microprocessor using symbolic model checking without BDDs. Lecture Notes in Computer Science, 1633:60–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1999:MCI

[1032] J. Baumgartner, T. Heyman, V. Singhal, and A. Aziz. Model checking the IBM gigahertz processor: An abstraction algorithm for high-performance netlists. Lecture Notes in Computer Science, 1633:72–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramesh:1999:VPP

[1033] S. Ramesh and P. Bhaduri. Validation of pipelined processor designs using Esterel tools. Lecture Notes in Computer Science, 1633:84–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berard:1999:AVP

[1034] B. Berard and L. Fribourg. Automated verification of a parametric realtime program: The ABR conformance protocol. Lecture Notes in Computer Science, 1633:96–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeron:1999:TGD

[1035] T. Jeron and P. Morel. Test generation derived from model-checking. Lecture Notes in Computer Science, 1633: 108–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carloni:1999:LIP

[1036] L. P. Carloni, K. L. McMillan, and A. L. Sangiovanni-Vincentelli. Latency insensitive protocols. Lecture Notes in Computer Science, 1633:123–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulla:1999:HGC

[1037] P. A. Abdulla, A. Bouajjani, B. Jonsson, and M. Nilsson. Handling global conditions in parametrized system verification. Lecture Notes in Computer Science, 1633:134–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulla:1999:VIS

[1038] P. A. Abdulla, A. Annichini, S. Bensalem, and A. Bouajjani. Verification of infinite-state systems by combining abstraction and reachability analysis. Lecture Notes in Computer Science, 1633:146–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1999:EPA

[1039] S. Das, D. L. Dill, and S. Park. Experience with predicate abstraction. Lecture Notes in Computer Science, 1633: 160–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kupferman:1999:MCS

[1040] O. Kupferman and M. Y. Vardi. Model checking of safety properties. Lecture Notes in Computer Science, 1633:172– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langerak:1999:CFP

[1041] R. Langerak and E. Brinksma. A complete finite prefix for process algebra. Lecture Notes in Computer Science, 1633:184–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1999:MFS

[1042] C.-T. Chou. The mathematical foundation of symbolic trajectory evaluation. Lecture Notes in Computer Science, 1633:196–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1999:AGR

[1043] T. A. Henzinger, S. Qadeer, and S. K.

Rajamani. Assume-guarantee refinement between different time scales. Lecture Notes in Computer Science, 1633:208–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloem:1999:EDP

[1044] R. Bloem, K. Ravi, and F. Somenzi.

Efficient decision procedures for model checking of linear time logic properties. Lecture Notes in Computer Science, 1633:222–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Etessami:1999:SIL

[1045] K. Etessami. Stutter-invariant languages, omega-automata, and temporal logic. Lecture Notes in Computer Science, 1633:236–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniele:1999:IAG

[1046] M. Daniele, F. Giunchiglia, and M. Y. Vardi. Improved automata generation for linear temporal logic. Lecture Notes in Computer Science, 1633:249– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozga:1999:RPS

[1047] M. Bozga and O. Maler. On the representation of probabilities over structured domains. Lecture Notes in Computer Science, 1633:261–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruns:1999:MCP

[1048] G. Bruns and P. Godefroid. Model checking partial state spaces with 3-Valued temporal logics. Lecture Notes in Computer Science, 1633:274–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthews:1999:EMA

[1049] J. Matthews and J. Launchbury. Elementary microarchitecture algebra. Lecture Notes in Computer Science, 1633:288–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1999:VSC

[1050] T. A. Henzinger, S. Qadeer, and S. K. Rajamani. Verifying sequential consistency on shared-memory multiprocessor systems. Lecture Notes in Computer Science, 1633:301–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lind-Nielsen:1999:SCM

[1051] J. Lind-Nielsen and H. R. Andersen. Stepwise CTL model checking of state/event systems. Lecture Notes in Computer Science, 1633:316–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1999:OSM

[1052] B. Yang, R. Simmons, R. E. Bryant, and D. R. O’Hallaron. Optimizing symbolic model checking for constraint-rich models. Lecture Notes in Computer Science, 1633:328–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Behrmann:1999:ETR

[1053] G. Behrmann, K. G. Larsen, J. Pearson, and C. Weise. Efficient timed reachability analysis using clock difference diagrams. Lecture Notes in Computer Science, 1633:341–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glusman:1999:MPC

[1054] M. Glusman and S. Katz. Mechanizing proofs of computation equivalence. Lecture Notes in Computer Science, 1633:354–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manolios:1999:LTP

[1055] P. Manolios, K. S. Namjoshi, and R. Sumners. Linking theorem proving and model-checking with wellfounded bisimulation. Lecture Notes in Computer Science, 1633:369–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjesse:1999:AVC

[1056] P. Bjesse. Automatic verification of combinational and pipelined FFT circuits. Lecture Notes in Computer Science, 1633:380–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Namjoshi:1999:EAC

[1057] K. S. Namjoshi and R. P. Kurshan. Efficient analysis of cyclic definitions. Lecture Notes in Computer Science, 1633:394–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1999:TRL

[1058] N. Klarlund. A theory of restrictions for logics and automata. Lecture Notes in Computer Science, 1633:406– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boppana:1999:MCB

[1059] V. Boppana, S. P. Rajan, K. Takayama, and M. Fujita. Model checking based on sequential ATPG. Lecture Notes in Computer Science, 1633:418–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spielmann:1999:AVA

[1060] M. Spielmann. Automatic verification of abstract state machines. Lecture Notes in Computer Science, 1633:431– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saiedi:1999:AMC

[1061] H. Saiedi and N. Shankar. Abstract and model check while you prove. Lecture Notes in Computer Science, 1633: 443–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pnueli:1999:DEF

[1062] A. Pnueli, Y. Rodeh, O. Shtrichman, and M. Siegel. Deciding equality formulas by small domains instantiations. Lecture Notes in Computer Science, 1633:455–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1999:EPE

[1063] R. E. Bryant, S. German, and M. N.

Velev. Exploiting positive equality in a logic of equality with uninterpreted functions. Lecture Notes in Computer Science, 1633:470–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchholz:1999:TAD

[1064] P. Buchholz and P. Kemper. A toolbox for the analysis of discrete event dynamic systems. Lecture Notes in Computer Science, 1633:483–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermanns:1999:TCS

[1065] H. Hermanns, V. Mertsiotakis, and M. Siegle. TIPPtool: Compositional specification and analysis of Markovian performance models. Lecture Notes in Computer Science, 1633:487–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1999:JBV

[1066] D. Basin, S. Friedrich, J. Posegga, and H. Vogt. Java bytecode verification by model checking. Lecture Notes in Computer Science, 1633:491–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cimatti:1999:NNS

[1067] A. Cimatti, E. Clarke, F. Giunchiglia, and M. Roveri. NuSMV: A new symbolic model verifier. Lecture Notes in Computer Science, 1633:495–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumann:1999:PST

[1068] J. Schumann. PIL/SETHEO: A tool for the automatic analysis of authentication protocols. Lecture Notes in Computer Science, 1633:500–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koller:1999:PRM

[1069] Daphne Koller. Probabilistic relational models. Lecture Notes in Computer Science, 1634:3–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1634/16340003.
                                        pdf.

Mannila:1999:IDA

[1070] Heikki Mannila. Inductive databases (abstract). Lecture Notes in Computer Science, 1634:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1634/16340014.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1634/16340014.pdf.

Quinlan:1999:SEM

[1071] J. R. Quinlan. Some elements of machine learning (extended abstract). Lecture Notes in Computer Science, 1634:15–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1634/16340015.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1634/16340015.
pdf.

Badea:1999:ROC

[1072] Liviu Badea and Monica Stanciu. Refinement operators can be (weakly) perfect. Lecture Notes in Computer Science, 1634:21–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340021.
pdf.

Bostroem:1999:CDC

[1073] H. Bostroem and L. Asker. Combining divide-and-conquer and separate-and-conquer for efficient and effective rule induction. Lecture Notes in Computer Science, 1634:33–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bostrom:1999:CDC

[1074] Henrik Boström and Lars Asker. Combining divide-and-conquer and separate-and-conquer for efficient and effective rule induction. Lecture Notes in Computer Science, 1634:33– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340033.
pdf.

Bratko:1999:RCH

[1075] Ivan Bratko. Refining complete hypotheses in ILP. Lecture Notes in Computer Science, 1634:44–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340044.
pdf.

Chiba:1999:AGD

[1076] Kazuya Chiba, Hayato Ohwada, and Fumio Mizoguchi. Acquiring graphic design knowledge with nonmonotonic inductive learning. Lecture Notes in Computer Science, 1634:56–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340056.
pdf.

Cussens:1999:MTS

[1077] James Cussens, Sašo Džeroski, and Tomaž Erjavec. Morphosyntactic tagging of slovene using progol. Lecture Notes in Computer Science, 1634:68– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340068.
pdf.

Dzeroski:1999:EPB

[1078] Sašo Džeroski, Hendrik Blockeel, Boris Kompare, Stefan Kramer, Bernhard Pfahringer, and Wim Van Laer. Experiments in predicting biodegradability. Lecture Notes in Computer Science, 1634:80–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340080.
pdf.

Flach:1999:FOB

[1079] Peter Flach and Nicolas Lachiche. 1BC:

A first-order Bayesian classifier. Lecture Notes in Computer Science, 1634: 92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340092.
pdf.

Frisch:1999:SDR

[1080] Alan M. Frisch. Sorted downward refinement: Building background knowledge into a refinement operator for inductive logic programming. Lecture Notes in Computer Science, 1634:104– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1634/16340104.
                                        pdf.

Hernandez-Orallo:1999:SCS

[1081] J. Hernández-Orallo and M. J. Ramírez-Quintana. A strong complete schema for inductive functional logic programming. Lecture Notes in Computer Science, 1634:116–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1634/16340116.
                                        pdf.

Horvath:1999:ADL

[1082] Tamás Horváth, Zoltán Alexin, Tibor Gyimóthy, and Stefan Wrobel. Application of different learning methods to Hungarian part-of-speech tagging. Lecture Notes in Computer Science, 1634: 128–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340128.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1634/16340128.
pdf.

Kazakov:1999:CAW

[1083] Dimitar Kazakov. Combining Lapis and WordNet for the learning of LR parsers with optimal semantic constraints. Lecture Notes in Computer Science, 1634:140–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340140.
pdf.

Kazakov:1999:LWS

[1084] Dimitar Kazakov, Suresh Manandhar, and Tomaž Erjavec. Learning word segmentation rules for tag prediction. Lecture Notes in Computer Science, 1634: 152–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340152.
pdf.

Kijsirikul:1999:AIR

[1085] Boonserm Kijsirikul and Sukree Sinthupinyo. Approximate ILP rules by backpropagation neural network: A result on Thai character recognition. Lecture Notes in Computer Science, 1634: 162–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340162.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1634/16340162.
                                        pdf.

Lavrac:1999:REM

[1086] Nada Lavrač, Peter Flach, and Blaz Zupan. Rule evaluation measures: A unifying view. Lecture Notes in Computer Science, 1634:174–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1634/16340174.
                                        pdf.

Lindberg:1999:IPS

[1087] Nikolaj Lindberg and Martin Eineborg. Improving part-of-speech disambiguation rules by adding linguistic knowledge. Lecture Notes in Computer Science, 1634:186–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1634/16340186.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1634/16340186.pdf.

Martin:1999:SCL

[1088] Eric Martin and Arun Sharma. On sufficient conditions for learnability of logic programs from positive data. Lecture Notes in Computer Science, 1634: 198–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340198.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1634/16340198.
pdf.

Midelfart:1999:BSS

[1089] Herman Midelfart. A bounded search space of clausal theories. Lecture Notes in Computer Science, 1634:210– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340210.
pdf.

Miyahara:1999:DNK

[1090] Tetsuhiro Miyahara, Takayoshi Shoudai, Tomoyuki Uchida, Tetsuji Kuboyama, Kenichi Takahashi, and Hiroaki Ueda. Discovering new knowledge from graph data using inductive logic programming. Lecture Notes in Computer Science, 1634:222–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340222.
pdf.

Muggleton:1999:AP

[1091] Stephen Muggleton and Michael Bain.

Analogical prediction. Lecture Notes in Computer Science, 1634:234–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340234.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1634/16340234.
                                        pdf.

Nienhuys-Cheng:1999:GRO

[1092] Shan-Hwei Nienhuys-Cheng, Wim Van Laer, Jan Ramon, and Luc De Raedt. Generalizing refinement operators to learn prenex conjunctive normal forms. Lecture Notes in Computer Science, 1634:245–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1634/16340245.
                                        pdf.

Parson:1999:TR

[1093] Rupert Parson, Khalid Khan, and Stephen Muggleton. Theory recovery. Lecture Notes in Computer Science, 1634:257–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1634/16340257.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1634/16340257.
                                        pdf.

Ramon:1999:IBF

[1094] Jan Ramon and Luc De Raedt. Instance based function learning. Lecture Notes in Computer Science, 1634: 268–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1634/16340268.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1634/16340268.
pdf.

Sakama:1999:SPI

[1095] Chiaki Sakama. Some properties of inverse resolution in normal logic programs. Lecture Notes in Computer Science, 1634:279–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340279.
pdf.

Srinivasan:1999:AIA

[1096] Ashwin Srinivasan, Ross D. King, and Douglas W. Bristol. An assessment of ILP-Assisted models for toxicology and the PTE-3 experiment. Lecture Notes in Computer Science, 1634:291– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1634/16340291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1634/16340291.
pdf.

Anonymous:1999:AIf

[1097] Anonymous. Author index. Lecture Notes in Computer Science, 1634:303– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1634/1634auth.pdf.

Contini:1999:IAS

[1098] Scott Contini, Ronald L. Rivest, M. J. B. Robshaw, and Yiqun Lisa Yin. Improved analysis of some simplified variants of RC6. Lecture Notes in Computer Science, 1636:1– 15, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360001.
pdf.

Borst:1999:LCR

[1099] Johan Borst, Bart Preneel, and Joos Vandewalle. Linear cryptanalysis of RC5 and RC6. Lecture Notes in Computer Science, 1636:16–30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360016.
pdf.

Lim:1999:RVC

[1100] Chae Hoon Lim. A revised version of Crypton — Crypton V1.0. Lecture Notes in Computer Science, 1636:31– 45, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360031.
pdf.

DHalluin:1999:ASR

[1101] Carl D’Halluin, Gert Bijnens, Vincent Rijmen, and Bart Preneel. Attack on six rounds of Crypton. Lecture Notes in Computer Science, 1636:46– 59, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360046.
pdf.

Lucks:1999:SBB

[1102] Stefan Lucks. On the security of the 128-bit block cipher DEAL. Lecture Notes in Computer Science, 1636:60– 70, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360060.
pdf.

Matsui:1999:CRV

[1103] Mitsuru Matsui and Toshio Tokita.

Cryptanalysis of a reduced version of the block cipher E2. Lecture Notes in Computer Science, 1636:71– 80, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360071.
pdf.

Knudsen:1999:DFC

[1104] Lars R. Knudsen and Vincent Rijmen. On the decorrelated fast cipher (DFC) and its theory. Lecture Notes in Computer Science, 1636:81– 94, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1636/16360081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1636/16360081.
                                        pdf.

Jakobsson:1999:SAE

[1105] Markus Jakobsson, Julien P. Stern, and Moti Yung. Scramble all, encrypt small. Lecture Notes in Computer Science, 1636:95–111, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1636/16360095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1636/16360095.
                                        pdf.

Lucks:1999:ARK

[1106] Stefan Lucks. Accelerated remotely keyed encryption. Lecture Notes in Computer Science, 1636:112–123, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1636/16360112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1636/16360112.
                                        pdf.

Biham:1999:MMA

[1107] Eli Biham, Alex Biryukov, and Adi Shamir. Miss in the middle attacks on IDEA and Khufu. Lecture Notes in Computer Science, 1636:124–138, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360124.
pdf.

Kelsey:1999:MCA

[1108] John Kelsey, Bruce Schneier, and David Wagner. Mod n cryptanalysis, with applications against RC5P and M6. Lecture Notes in Computer Science, 1636:139–155, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360139.
pdf.

Wagner:1999:BA

[1109] David Wagner. The boomerang attack. Lecture Notes in Computer Science, 1636:156–170, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1636/16360156.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1636/16360156.pdf.

Patel:1999:TML

[1110] Sarvar Patel, Zulfikar Ramzan, and Ganapathy S. Sundaram. Towards making Luby-Rackoff ciphers optimal and practical. Lecture Notes in Computer Science, 1636:171–185, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360171.
pdf.

Canteaut:1999:NCA

[1111] Anne Canteaut, Pascale Charpin, and Hans Dobbertin. A new characterization of almost bent functions. Lecture Notes in Computer Science, 1636:186– 200, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360186.
pdf.

Paterson:1999:IPG

[1112] Kenneth G. Paterson. Imprimitive permutation groups and trapdoors in iterated block ciphers. Lecture Notes in Computer Science, 1636:201–214, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360201.
pdf.

Handschuh:1999:SDK

[1113] Helena Handschuh and Bart Preneel.

On the security of double and 2-key triple modes of operation. Lecture Notes in Computer Science, 1636:215– 230, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360215.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360215.
pdf.

Bellare:1999:CVI

[1114] Mihir Bellare and Phillip Rogaway.

On the construction of variableinput-length ciphers. Lecture Notes in Computer Science, 1636:231–244, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360231.
pdf.

Biryukov:1999:SA

[1115] Alex Biryukov and David Wagner. Slide attacks. Lecture Notes in Computer Science, 1636:245–259, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360245.
pdf.

Vaudenay:1999:SCC

[1116] Serge Vaudenay. On the security of CS-cipher. Lecture Notes in Computer Science, 1636:260–274, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360260.
pdf.

Moriai:1999:IAB

[1117] Shiho Moriai, Takeshi Shimoyama, and Toshinobu Kaneko. Interpolation attacks of the block cipher: SNAKE. Lecture Notes in Computer Science, 1636:275–289, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360275.
pdf.

Aiello:1999:HSP

[1118] William Aiello, S. Rajagopalan, and Ramarathnam Venkatesan. Highspeed pseudorandom number generation with small memory. Lecture Notes in Computer Science, 1636:290– 304, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360290.
pdf.

Bleichenbacher:1999:SC

[1119] Daniel Bleichenbacher and Sarvar Patel. SOBER cryptanalysis. Lecture Notes in Computer Science, 1636:305– 316, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1636/16360305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1636/16360305.
pdf.

Anonymous:1999:AIg

[1120] Anonymous. Author index. Lecture Notes in Computer Science, 1636:317– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1636/1636auth.pdf.

Antoniou:1999:DDR

[1121] Grigoris Antoniou. On the dynamics of default reasoning. Lecture Notes in Computer Science, 1638:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380001.
pdf.

Arieli:1999:NPR

[1122] Ofer Arieli and Arnon Avron. Nonmonotonic and paraconsistent reasoning: From basic entailments to plausible relations. Lecture Notes in Computer Science, 1638:11–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380011.
pdf.

Bejar:1999:CSL

[1123] Ramón Béjar and Felip Manyà. A comparison of systematic and local search algorithms for regular CNF formulas. Lecture Notes in Computer Science, 1638:22–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380022.
pdf.

Benhammadi:1999:QAP

[1124] F. Benhammadi, P. Nicolas, and T. Schaub. Query-answering in prioritized default logic. Lecture Notes in Computer Science, 1638:32–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380032.
pdf.

Yaghlane:1999:UDB

[1125] Boutheina Ben Yaghlane and Khaled Mellouli. Updating directed belief networks. Lecture Notes in Computer Science, 1638:43–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1638/16380043.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1638/16380043.pdf.

Besnard:1999:ICE

[1126] Philippe Besnard and Marie-Odile Cordier. Inferring causal explanations. Lecture Notes in Computer Science, 1638:55–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1638/16380055.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1638/16380055.
pdf.

Borgelt:1999:CIC

[1127] Christian Borgelt and Rudolf Kruse. A critique of inductive causation. Lecture Notes in Computer Science, 1638: 68–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380068.
pdf.

Bourne:1999:CLM

[1128] Rachel A. Bourne and Simon Parsons. Connecting lexicographic with maximum entropy entailment. Lecture Notes in Computer Science, 1638:80– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380080.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380080.
                                        pdf.

Bruening:1999:ANG

[1129] S. Bruening and T. Schaub. Avoiding non-ground variables. Lecture Notes in Computer Science, 1638:92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruning:1999:ANG

[1130] Stefan Brüning and Torsten Schaub. Avoiding non-ground variables. Lecture Notes in Computer Science, 1638: 92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380092.
                                        pdf.

Coradeschi:1999:ASV

[1131] Silvia Coradeschi and Alessandro Saffiotti. Anchoring symbols to vision data by fuzzy logic. Lecture Notes in Computer Science, 1638:104– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380104.
                                        pdf.

Cossart:1999:FVR

[1132] Corine Cossart and Catherine Tessier. Filtering vs revision and update: Let us debate! Lecture Notes in Computer Science, 1638:116–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380116.
pdf.

Cozman:1999:IIA

[1133] Fabio Gagliardi Cozman. Irrelevance and independence axioms in quasi-Bayesian theory. Lecture Notes in Computer Science, 1638:128–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380128.
pdf.

Dubois:1999:AVC

[1134] Didier Dubois, Michel Grabisch, and Henri Prade. Assessing the value of a candidate: A qualitative possibilistic approach. Lecture Notes in Computer Science, 1638:137–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380137.
pdf.

Duval:1999:LDT

[1135] Béatrice Duval and Pascal Nicolas.

Learning default theories. Lecture Notes in Computer Science, 1638:148– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380148.
                                        pdf.

Flach:1999:KRI

[1136] Peter A. Flach. Knowledge representation for inductive learning. Lecture Notes in Computer Science, 1638:160– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380160.
                                        pdf.

Gregoire:1999:HIE

[1137] Eric Grégoire. Handling inconsistency efficiently in the incremental construction of stratified belief bases. Lecture Notes in Computer Science, 1638:168– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380168.
                                        pdf.

Guan:1999:RKD

[1138] J. W. Guan and D. A. Bell. Rough knowledge discovery and applications. Lecture Notes in Computer Science, 1638:179–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/1638/16380179.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1638/16380179.
pdf.

Jensen:1999:GDT

[1139] Finn Verner Jensen. Gradient descent training of Bayesian networks. Lecture Notes in Computer Science, 1638:190– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380190.
pdf.

Kaminski:1999:ODT

[1140] Michael Kaminski. Open default theories over closed domains (extended abstract). Lecture Notes in Computer Science, 1638:201–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380201.
pdf.

Kephart:1999:SE

[1141] Jeffrey O. Kephart and Amy R. Greenwald. Shopbot economics. Lecture Notes in Computer Science, 1638:208– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380208.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1638/16380208.
                                        pdf.

Khalfallah:1999:OAL

[1142] Fédia Khalfallah and Khaled Mellouli. Optimized algorithm for learning Bayesian network from data. Lecture Notes in Computer Science, 1638: 221–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380221.
                                        pdf.

Konieczny:1999:MIC

[1143] Sébastien Konieczny and Ramón Pino Pérez. Merging with integrity constraints. Lecture Notes in Computer Science, 1638:233–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1638/16380233.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1638/16380233.pdf.

Kramosil:1999:BLI

[1144] Ivan Kramosil. Boolean-like interpretation of Sugeno integral. Lecture Notes in Computer Science, 1638:245– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380245.
                                        pdf.

Lehmann:1999:AOP

[1145] Norbert Lehmann and Rolf Haenni. An alternative to outward propagation for Dempster-Shafer belief functions. Lecture Notes in Computer Science, 1638: 256–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380256.
pdf.

Linke:1999:BPP

[1146] Thomas Linke and Torsten Schaub. On bottom-up pre-processing techniques for automated default reasoning. Lecture Notes in Computer Science, 1638: 268–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380268.
pdf.

Lukasiewicz:1999:PLP

[1147] Thomas Lukasiewicz and Gabriele Kern-Isberner. Probabilistic logic programming under maximum entropy. Lecture Notes in Computer Science, 1638:279–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1638/16380279.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1638/16380279.
pdf.

Madsen:1999:LPI

[1148] Anders L. Madsen and Bruce D’Ambrosio. Lazy propagation and independence of causal influence. Lecture Notes in Computer Science, 1638:293–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380293.
pdf.

Moral:1999:MCA

[1149] Serafín Moral and Antonio Salmerón. A Monte Carlo algorithm for combining Dempster-Shafer belief based on approximate pre-computation. Lecture Notes in Computer Science, 1638:305– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380305.
pdf.

Pacholczyk:1999:ELN

[1150] Daniel Pacholczyk. An extension of a linguistic negation model allowing us to deny nuanced property combinations. Lecture Notes in Computer Science, 1638:316–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380316.
pdf.

Parsons:1999:AQD

[1151] Simon Parsons and Shaw Green.

Argumentation and qualitative decision making. Lecture Notes in Computer Science, 1638:328–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380328.
pdf.

Petit-Renaud:1999:HDF

[1152] Simon Petit-Renaud and Thierry Denœux. Handling different forms of uncertainty in regression analysis: A fuzzy belief structure approach. Lecture Notes in Computer Science, 1638: 340–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380340.
pdf.

Rombaut:1999:SRD

[1153] Michèle Rombaut, Iman Jarkass, and Thierry Denœux. State recognition in discrete dynamical systems using Petri nets and evidence theory. Lecture Notes in Computer Science, 1638: 352–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380352.
pdf.

Shanahan:1999:RNM

[1154] Murray Shanahan and Mark Witkowski. Robot navigation and map building with the event calculus. Lecture Notes in Computer Science, 1638:362– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380362.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380362.
                                        pdf.

Siekmann:1999:IFC

[1155] Stefan Siekmann, Jörg Gebhardt, and Rudolf Kruse. Information fusion in the context of stock index prediction. Lecture Notes in Computer Science, 1638: 363–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1638/16380363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1638/16380363.
                                        pdf.

Torre:1999:DG

[1156] Leendert van der Torre. Defeasible goals. Lecture Notes in Computer Science, 1638:374–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1638/16380374.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1638/16380374.pdf.

vanderTorre:1999:DG

[1157] Leendert van der Torre. Defeasible goals. Lecture Notes in Computer Science, 1638:374–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1638/16380374.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1638/16380374.pdf.

Wilson:1999:LDU

[1158] Nic Wilson and Jérôme Mengin. Logical deduction using the local computation framework. Lecture Notes in Computer Science, 1638:386–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1638/16380386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1638/16380386.
pdf.

Anonymous:1999:AIh

[1159] Anonymous. Author index. Lecture Notes in Computer Science, 1638:397– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1638/1638indx.pdf.

Varaiya:1999:DSI

[1160] P. Varaiya. Design, simulation, and implementation of hybrid systems. Lecture Notes in Computer Science, 1639: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miner:1999:ERS

[1161] A. S. Miner and G. Ciardo. Efficient reachability set generation and storage using decision diagrams. Lecture Notes in Computer Science, 1639:6–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pastor:1999:SMI

[1162] E. Pastor, J. Cortadella, and M. A. Peña. Structural methods to improve the symbolic analysis of Petri nets. Lecture Notes in Computer Science, 1639:26–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1999:SSS

[1163] K. Schmidt. Stubborn sets for standard properties. Lecture Notes in Computer Science, 1639:46–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bastide:1999:PNB

[1164] R. Bastide, P. Palanque, O. Sy, and D.-H. Le. Petri net based behavioural specification of CORBA systems. Lecture Notes in Computer Science, 1639: 66–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummer:1999:SCB

[1165] O. Kummer, D. Moldt, and F. Wienberg. Symmetric communication between coloured Petri net simulations and Java-processes. Lecture Notes in Computer Science, 1639:86–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krogh:1999:RDM

[1166] B. H. Krogh. Recent developments in modeling and analysis of hybrid dynamic systems. Lecture Notes in Computer Science, 1639:106–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Recalde:1999:ACP

[1167] L. Recalde, E. Teruel, and M. Silva.

Autonomous continuous P/T systems. Lecture Notes in Computer Science, 1639:107–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokmakoff:1999:AAI

[1168] A. Tokmakoff and J. Billington. An approach to the analysis of interworking traders. Lecture Notes in Computer Science, 1639:127–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allmaier:1999:PAN

[1169] S. Allmaier and D. Kreische. Parallel approaches to the numerical transient analysis of stochastic reward nets. Lecture Notes in Computer Science, 1639: 147–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capra:1999:SNF

[1170] L. Capra, R. Gaeta, and O. Botti. SWN nets as a framework for the specification and the analysis of FT techniques adopted in electric plant automation. Lecture Notes in Computer Science, 1639:168–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hadjicostis:1999:MDE

[1171] C. N. Hadjicostis and G. C. Verghese. Monitoring discrete event systems using Petri net embeddings. Lecture Notes in Computer Science, 1639:188– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sgroi:1999:QSS

[1172] M. Sgroi, L. Lavagno, Y. Watanabe, and A. Sangiovanni-Vincentelli. Quasistatic scheduling of embedded software using equal conflict nets. Lecture Notes in Computer Science, 1639:208– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddad:1999:TAR

[1173] S. Haddad and D. Poitrenaud. Theoretical aspects of recursive Petri nets. Lecture Notes in Computer Science, 1639: 228–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1999:PNT

[1174] C. Schneider and J. Wehler. Petri net theory — problems solved by commutative algebra. Lecture Notes in Computer Science, 1639:248–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrzastowski-Wachtel:1999:TUR

[1175] P. Chrzastowski-Wachtel. Testing undecidability of the reachability in Petri nets with the help of 10th Hilbert problem. Lecture Notes in Computer Science, 1639:268–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMichelis:1999:NTW

[1176] G. De Michelis. Net theory and workflow models. Lecture Notes in Computer Science, 1639:282–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1999:CIA

[1177] W. Vogler. Concurrent implementation of asynchronous transition systems. Lecture Notes in Computer Science, 1639:284–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fanchon:1999:TCN

[1178] J. Fanchon. Trace channel nets. Lecture Notes in Computer Science, 1639: 304–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juhas:1999:RAA

[1179] G. Juhas. Reasoning about algebraic generalisation of Petri nets. Lecture Notes in Computer Science, 1639:324– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1999:BAM

[1180] E. Best, R. Devillers, and M. Koutny.

The box algebra — A model of nets and process expressions. Lecture Notes in Computer Science, 1639:344–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couvreur:1999:DIB

[1181] J.-M. Couvreur and D. Poitrenaud. Detection of illegal behaviours based on unfoldings. Lecture Notes in Computer Science, 1639:364–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheung:1999:FCI

[1182] T.-Y. Cheung and Y. Lu. Five classes of invariant-preserving transformations on colored Petri nets. Lecture Notes in Computer Science, 1639:384–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baar:1999:VII

[1183] T. Baar, E. Kindler, and H. Voelzer. Verifying intuition — ILF checks DAWN proofs. Lecture Notes in Computer Science, 1639:404–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sowa:1999:CGD

[1184] J. Sowa. Conceptual graphs: Draft Proposed American National Standard. Lecture Notes in Computer Science, 1640:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mineau:1999:CPE

[1185] G. W. Mineau. Constraints on processes: Essential elements for the validation and execution of processes. Lecture Notes in Computer Science, 1640: 66–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nenkova:1999:UMA

[1186] A. Nenkova and G. Angelova. User modeling as an application of actors. Lecture Notes in Computer Science, 1640:83–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brennan:1999:SUH

[1187] J. Brennan. Spatial universals as the human spatial notion. Lecture Notes in Computer Science, 1640:90–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Messaadia:1999:KES

[1188] K. Messaadia and M. Oussalah. Knowledge engineering with semantic and transfer links. Lecture Notes in Computer Science, 1640:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farkas:1999:PFS

[1189] J. Farkas and J. Sarbo. A peircean framework of syntactic structure. Lecture Notes in Computer Science, 1640: 112–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hess:1999:CBB

[1190] J. Hess and W. R. Cyre. A CG-based behavior extraction system. Lecture Notes in Computer Science, 1640:127– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1999:ECG

[1191] B. Moulin and H. Irandoust. Extending the conceptual graph approach to represent evaluative attitudes in discourse. Lecture Notes in Computer Science, 1640:140–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dogru:1999:ISL

[1192] S. Dogru and J. R. Slagle. Implementing a semantic lexicon. Lecture Notes in Computer Science, 1640:154–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1999:ATO

[1193] B. Moulin, M. Gouiaa, and S. Delisle. Analysis of task-oriented conversations into conceptual graph structures. Lecture Notes in Computer Science, 1640: 168–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Racky:1999:UCG

[1194] J. Racky and M. Pandit. Using conceptual graphs as a common representation for data and configuration in an active image processing system. Lecture Notes in Computer Science, 1640:176–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braeuner:1999:SSL

[1195] T. Braeuner, C. Donner, and P. Oehrstroem. A software system for learning peircean graphs. Lecture Notes in Computer Science, 1640:184–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabbaj:1999:SCG

[1196] A. Kabbaj. Synergy: A conceptual graph activation-based language. Lecture Notes in Computer Science, 1640: 198–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fun:1999:DCB

[1197] P.-W. Fun and R. H. Kemp. On developing case-based tutorial systems with conceptual graphs. Lecture Notes in Computer Science, 1640:214–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1999:EKW

[1198] P. Martin and P. Eklund. Embedding knowledge in Web documents: CGs versus XML-based metadata languages. Lecture Notes in Computer Science, 1640:230–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabbaj:1999:SHO

[1199] A. Kabbaj. Synergy as an hybrid object-oriented conceptual graph language. Lecture Notes in Computer Science, 1640:247–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Southey:1999:NJA

[1200] F. Southey and J. G. Linders. NO-TIO — A Java API for developing CG tools. Lecture Notes in Computer Science, 1640:262–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thanitsukkarn:1999:MAS

[1201] T. Thanitsukkarn and A. Finkelstein. Multiperspective analysis of the Sisyphus-I room allocation task modeled in a CG meta-representation language. Lecture Notes in Computer Science, 1640:272–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damianova:1999:UCG

[1202] S. Damianova and K. Toutanova. Using conceptual graphs to solve a resource allocation task. Lecture Notes in Computer Science, 1640:297–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1999:WSP

[1203] P. Martin and P. Eklund. WebKB and the Sisyphus-I problem. Lecture Notes in Computer Science, 1640:315– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mineau:1999:CGU

[1204] G. W. Mineau. Constraints and goals under the conceptual graph formalism: One way to solve SCG-1 problem. Lecture Notes in Computer Science, 1640:334–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baget:1999:PGB

[1205] J. F. Baget, D. Genest, and M. L. Mugnier. A pure graph-based solution to the SCG-1 initiative. Lecture Notes in Computer Science, 1640:355–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganter:1999:CAL

[1206] B. Ganter and R. Wille. Contextual attribute logic. Lecture Notes in Computer Science, 1640:377–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groh:1999:ACR

[1207] B. Groh and P. Eklund. Algorithms for creating relational power context families from conceptual graphs. Lecture Notes in Computer Science, 1640:389– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prediger:1999:LCG

[1208] S. Prediger and R. Wille. The lattice of concept graphs of a relationally scaled context. Lecture Notes in Computer Science, 1640:401–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raban:1999:CIS

[1209] R. Raban and B. Garner. Contexts in information systems development. Lecture Notes in Computer Science, 1640:415–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mineau:1999:CSR

[1210] G. W. Mineau, G. Stumme, and R. Wille. Conceptual structures represented by conceptual graphs and formal concept analysis. Lecture Notes in Computer Science, 1640:423–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baget:1999:SCI

[1211] J.-F. Baget. A simulation of coidentity with rules in simple and nested graphs. Lecture Notes in Computer Science, 1640:442–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braeuner:1999:CGA

[1212] T. Braeuner, J. F. Nilsson, and A. Rasmussen. Conceptual graphs as algebras — with an application to analogical reasoning. Lecture Notes in Computer Science, 1640:456–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbett:1999:UCC

[1213] D. Corbett and R. Woodbury. Unification over constraints in conceptual structures. Lecture Notes in Computer Science, 1640:470–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1999:TDF

[1214] F. Baader, R. Molitor, and S. Tobies. Tractable and decidable fragments of conceptual graphs. Lecture Notes in Computer Science, 1640:480–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerdiles:1999:DSC

[1215] G. Kerdiles. Dynamic semantics for conceptual graphs. Lecture Notes in Computer Science, 1640:494–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbett:1999:CVA

[1216] D. Corbett. A case for variablearity relations: Definitions and domains. Lecture Notes in Computer Science, 1640:508–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solovyev:1999:GSP

[1217] V. Solovyev. Graph structures in parametric spaces for representation of verbs. Lecture Notes in Computer Science, 1640:510–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keeler:1999:PPO

[1218] M. Keeler and W. Tepfenhart. PORT:

Peirce online resource testbed. Lecture Notes in Computer Science, 1640:511– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tepfenhart:1999:ACA

[1219] W. Tepfenhart and M. Holder. Assuring computer agent communications. Lecture Notes in Computer Science, 1640:513–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebban:1999:TLS

[1220] Marc Sebban and Gilles Richard. From theoretical learnability to statistical measures of the learnable. Lecture Notes in Computer Science, 1642:3– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420003.
pdf.

Cordon:1999:AMD

[1221] Oscar Cordón and Francisco Herrera. ALM: A methodology for designing accurate linguistic models for intelligent data analysis. Lecture Notes in Computer Science, 1642:15– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420015.
pdf.

Nock:1999:TPI

[1222] Richard Nock and Pascal Jappy. A “top-down and prune” induction scheme for constrained decision committees. Lecture Notes in Computer Science, 1642:27–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420027.
pdf.

Kosters:1999:MCA

[1223] Walter A. Kosters, Elena Marchiori, and Ard A. J. Oerlemans. Mining clusters with association rules. Lecture Notes in Computer Science, 1642:39– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420039.
pdf.

Swift:1999:ECS

[1224] Stephen Swift, Allan Tucker, and Xiaohui Liu. Evolutionary computation to search for strongly correlated variables in high-dimensional time-series. Lecture Notes in Computer Science, 1642: 51–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420051.
pdf.

Elomaa:1999:BDT

[1225] Tapio Elomaa. The biases of decision tree pruning strategies. Lecture Notes in Computer Science, 1642:63– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420063.
pdf.

Talavera:1999:FSR

[1226] Luis Talavera. Feature selection as retrospective pruning in hierarchical clustering. Lecture Notes in Computer Science, 1642:75–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420075.
pdf.

Silipo:1999:DPI

[1227] Rosaria Silipo and Michael R. Berthold.

Discriminative power of input features in a fuzzy model. Lecture Notes in Computer Science, 1642:87– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420087.
                                        pdf.

Firoiu:1999:LER

[1228] Laura Firoiu and Paul Cohen. Learning elements of representations for redescribing robot experiences. Lecture Notes in Computer Science, 1642:99– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420099.
                                        pdf.

Huang:1999:SOS

[1229] Xingang Huang and Feng Zhao. “seeing” objects in spatial datasets. Lecture Notes in Computer Science, 1642: 111–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420111.
                                        pdf.

Matsumoto:1999:IMM

[1230] Kazunori Matsumoto and Kazuo Hashimoto. Intelligent monitoring method using time varying binomial distribution models for pseudo-periodic communication traffic. Lecture Notes in Computer Science, 1642:123–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420123.
pdf.

Flexer:1999:MHI

[1231] Arthur Flexer and Herbert Bauer.

Monitoring human information processing via intelligent data analysis of EEG recordings. Lecture Notes in Computer Science, 1642:137–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420137.
pdf.

Andrienko:1999:KBV

[1232] Gennady Andrienko and Natalia Andrienko. Knowledge-based visualization to support spatial data mining. Lecture Notes in Computer Science, 1642:149–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1642/16420149.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1642/16420149.
pdf.

Hofmann:1999:PTM

[1233] Thomas Hofmann. Probabilistic topic maps: Navigating through large text collections. Lecture Notes in Computer Science, 1642:161–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420161.
                                        pdf.

Yang:1999:GTM

[1234] Li Yang. 3D grand tour for multidimensional data and clusters. Lecture Notes in Computer Science, 1642:173– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420173.
                                        pdf.

Potharst:1999:DTA

[1235] Rob Potharst and Jan C. Bioch. A decision tree algorithm for ordinal classification. Lecture Notes in Computer Science, 1642:187–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420187.
                                        pdf.

Sebastiani:1999:DDU

[1236] Paola Sebastiani, Marco Ramoni, Paul Cohen, John Warwick, and James Davis. Discovering dynamics using Bayesian clustering. Lecture Notes in Computer Science, 1642:199–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420199.
pdf.

Talavera:1999:IDK

[1237] Luis Talavera and Javier Béjar. Integrating declarative knowledge in hierarchical clustering tasks. Lecture Notes in Computer Science, 1642:211– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420211.
pdf.

Aladjem:1999:NLD

[1238] Mayer Aladjem. Nonparametric linear discriminant analysis by recursive optimization with random initialization. Lecture Notes in Computer Science, 1642:223–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420223.
pdf.

Kelly:1999:SCP

[1239] Mark G. Kelly, David J. Hand, and Niall M. Adams. Supervised classification problems: How to be both judge and jury. Lecture Notes in Computer Science, 1642:235–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420235.
                                        pdf.

Li:1999:TPG

[1240] Cen Li and Gautam Biswas. Temporal pattern generation using hidden Markov model based unsupervised classification. Lecture Notes in Computer Science, 1642:245–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420245.
                                        pdf.

Huellermeier:1999:ESS

[1241] E. Huellermeier. Exploiting similarity for supporting data analysis and problem solving. Lecture Notes in Computer Science, 1642:257–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hullermeier:1999:ESS

[1242] Eyke Hüllermeier. Exploiting similarity for supporting data analysis and problem solving. Lecture Notes in Computer Science, 1642:257–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420257.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1642/16420257.
pdf.

Nascimento:1999:MPM

[1243] Susana Nascimento, Boris Mirkin, and Fernando Moura-Pires. Multiple prototype model for fuzzy clustering. Lecture Notes in Computer Science, 1642: 269–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420269.
pdf.

Eggermont:1999:CGP

[1244] Jeroen Eggermont, Agoston E. Eiben, and Jano I. van Hemert. A comparison of genetic programming variants for data classification. Lecture Notes in Computer Science, 1642:281– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420281.
pdf.

Klawonn:1999:FCB

[1245] Frank Klawonn and Annette Keller.

Fuzzy clustering based on modified distance measures. Lecture Notes in Computer Science, 1642:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420291.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420291.
                                        pdf.

Valtchev:1999:BCO

[1246] Petko Valtchev. Building classes in object-based languages by automatic clustering. Lecture Notes in Computer Science, 1642:303–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420303.
                                        pdf.

Mertens:1999:AEC

[1247] Bart J. A. Mertens and David J. Hand. Adjusted estimation for the combination of classifiers. Lecture Notes in Computer Science, 1642:317– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420317.
                                        pdf.

Yang:1999:DDT

[1248] Jihoon Yang, Rajesh Parekh, Vasant Honavar, and Drena Dobbs. Data-driven theory refinement using KBDistAl. Lecture Notes in Computer Science, 1642:331–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420331.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1642/16420331.
pdf.

Easley:1999:RAI

[1249] Matthew Easley and Elizabeth Bradley.

Reasoning about input-output modeling of dynamical systems. Lecture Notes in Computer Science, 1642:343– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420343.
pdf.

Almond:1999:USA

[1250] Russell Almond. Undoing statistical advice. Lecture Notes in Computer Science, 1642:357–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420357.
pdf.

Guimaraes:1999:MTK

[1251] Gabriela Guimarães and Alfred Ultsch.

A method for temporal knowledge conversion. Lecture Notes in Computer Science, 1642:369–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420369.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1642/16420369.
pdf.

Gunetti:1999:IDT

[1252] Daniele Gunetti and Giancarlo Ruffo. Intrusion detection through behavioral data. Lecture Notes in Computer Science, 1642:383–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1642/16420383.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1642/16420383.pdf.

Macrossan:1999:BNN

[1253] Paula E. Macrossan, Hussein A. Abbass, Kerry Mengersen, Michael Towsey, and Gerard Finn. Bayesian neural network learning for prediction in the Australian dairy industry. Lecture Notes in Computer Science, 1642: 395–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420395.
pdf.

Talbert:1999:ESD

[1254] Doug Talbert and Doug Fisher. Exploiting sample-data distributions to reduce the cost of nearest-neighbor searches with Kd-trees. Lecture Notes in Computer Science, 1642:407–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420407.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1642/16420407.
pdf.

Tax:1999:PFD

[1255] David M. J. Tax, Alexander Ypma, and Robert P. W. Duin. Pump failure detection using support vector data descriptions. Lecture Notes in Computer Science, 1642:415–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420415.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420415.
pdf.

Poddig:1999:DMD

[1256] Thorsten Poddig and Claus Huber.

Data mining for the detection of turning points in financial time series. Lecture Notes in Computer Science, 1642: 427–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420427.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420427.
pdf.

Yang-Stephens:1999:CAC

[1257] Bokyung Yang-Stephens, M. Charles Swope, Jeffrey Locke, and Isabelle Moulinier. Computer-assisted classification of legal abstracts. Lecture Notes in Computer Science, 1642:437– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1642/16420437.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420437.
                                        pdf.

Ikkai:1999:SCL

[1258] Yoshitomo Ikkai, Kazuhisa Ikeda, Norihisa Komoda, Akira Yamane, and Isao Tone. Sequential control logic inferring method from observed plant I/O data. Lecture Notes in Computer Science, 1642:449–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420449.
                                        pdf.

Cheng:1999:EES

[1259] Gongxian Cheng, Kwawen Cho, Xiaohui Liu, George Loizou, and John X. Wu. Evaluating an eye screening test. Lecture Notes in Computer Science, 1642:461–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420461.
                                        pdf.

Pena:1999:ARS

[1260] José M. Peña, Sylvain Létourneau, and Fazel Famili. Application of rough sets algorithms to prediction of aircraft component failure. Lecture Notes in Computer Science, 1642:473– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420473.
pdf.

Fuernkranz:1999:ESI

[1261] J. Fuernkranz. Exploiting structural information for text classification on the WWW. Lecture Notes in Computer Science, 1642:487–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furnkranz:1999:ESI

[1262] Johannes Fürnkranz. Exploiting structural information for text classification on the WWW. Lecture Notes in Computer Science, 1642:487–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420487.
pdf.

Choi:1999:MAW

[1263] Yong S. Choi and Suk I. Yoo. Multiagent Web information retrieval: Neural network based approach. Lecture Notes in Computer Science, 1642:499– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1642/16420499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1642/16420499.
pdf.

Tauritz:1999:AIF

[1264] Daniel R. Tauritz and Ida G. Sprinkhuizen-Kuyper. Adaptive information filtering algorithms. Lecture Notes in Computer Science, 1642:513– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420513.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420513.
                                        pdf.

Fatemi:1999:CGA

[1265] Nastaran Fatemi and Philippe Mulhem. A conceptual graph approach for video data representation and retrieval. Lecture Notes in Computer Science, 1642:525–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1642/16420525.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1642/16420525.
                                        pdf.

Frankel:1999:ASD

[1266] Y. Frankel, M. Yung, and P. MacKenzie. Adaptively-secure distributed public-key systems. Lecture Notes in Computer Science, 1643:4–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korte:1999:HLD

[1267] B. Korte. How long does a bit live in a computer? Lecture Notes in Computer Science, 1643:28–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravi:1999:AAT

[1268] R. Ravi and F. S. Salman. Approximation algorithms for the traveling purchaser problem and its variants in network design. Lecture Notes in Computer Science, 1643:29–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diks:1999:IKB

[1269] K. Diks, D. Krizanc, E. Kranakis, and A. Pelc. The impact of knowledge on broadcasting time in radio networks (extended abstract). Lecture Notes in Computer Science, 1643:41–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwama:1999:MRD

[1270] K. Iwama and E. Miyano. Multipacket routing on 2-D meshes and its application to fault-tolerant routing. Lecture Notes in Computer Science, 1643:53–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1999:IAL

[1271] P. Crescenzi, L. Dardini, and R. Grossi.

IP address lookup made fast and simple. Lecture Notes in Computer Science, 1643:65–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bar-Noy:1999:LLB

[1272] A. Bar-Noy, A. Freund, and J. Naor.

On-line load balancing in a hierarchical server topology. Lecture Notes in Computer Science, 1643:77–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AufderHeide:1999:PGP

[1273] F. M. Auf der Heide, M. Westermann, and B. Voecking. Provably good and practical strategies for nonuniform data management in networks. Lecture Notes in Computer Science, 1643:89–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phillips:1999:AAR

[1274] S. J. Phillips and J. R. Westbrook. Approximation algorithms for restoration capacity planning. Lecture Notes in Computer Science, 1643:101–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bar-Yehuda:1999:EAI

[1275] R. Bar-Yehuda and D. Rawitz. Efficient algorithms for integer programs with two variables per constraint (extended abstract). Lecture Notes in Computer Science, 1643:116–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skutella:1999:CQP

[1276] M. Skutella. Convex quadratic programming relaxations for network scheduling problems. Lecture Notes in Computer Science, 1643:127–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moehring:1999:RCP

[1277] R. H. Moehring, F. Stork, M. Uetz, and A. S. Schulz. Resource-constrained project scheduling: Computing lower bounds by solving minimum cut problems. Lecture Notes in Computer Science, 1643:139–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epstein:1999:ASS

[1278] L. Epstein and J. Sgall. Approximation schemes for scheduling on uniformly related and identical parallel machines. Lecture Notes in Computer Science, 1643:151–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azar:1999:LTT

[1279] Y. Azar and O. Regev. Off-line temporary tasks assignment. Lecture Notes in Computer Science, 1643:163–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bischof:1999:LBU

[1280] S. Bischof, T. Schickinger, and A. Steger. Load balancing using bisectors — A tight average-case analysis. Lecture Notes in Computer Science, 1643:172–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1999:AEA

[1281] T. Jansen and I. Wegener. On the analysis of evolutionary algorithms — A proof that crossover really can help. Lecture Notes in Computer Science, 1643:184–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicodeme:1999:MS

[1282] P. Nicodeme, B. Salvy, and P. Flajolet. Motif statistics. Lecture Notes in Computer Science, 1643:194–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heun:1999:APF

[1283] V. Heun. Approximate protein folding in the HP side chain model on extended cubic lattices (extended abstract). Lecture Notes in Computer Science, 1643:212–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crauser:1999:CSA

[1284] A. Crauser and P. Ferragina. On constructing suffix arrays in external memory. Lecture Notes in Computer Science, 1643:224–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laber:1999:SSD

[1285] E. S. Laber, R. L. Milidiu, and A. A. Pessoa. Strategies for searching with different access costs. Lecture Notes in Computer Science, 1643:236–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ting:1999:IAB

[1286] C. Ting and M.-Y. Kao. On the informational asymmetry between upper and lower bounds for ultrametric evolutionary trees. Lecture Notes in Computer Science, 1643:248–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cicalese:1999:OBS

[1287] F. Cicalese and D. Mundici. Optimal binary search with two unreliable tests and minimum adaptiveness. Lecture Notes in Computer Science, 1643:257–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roura:1999:IML

[1288] S. Roura. Improving mergesort for linked lists. Lecture Notes in Computer Science, 1643:267–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manzini:1999:EAL

[1289] G. Manzini. Efficient algorithms for online symbol ranking compression (extended abstract). Lecture Notes in Computer Science, 1643:277–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1999:LUW

[1290] E. J. Anderson, A. R. Karlin, K. Hildrum, and A. Rasala. On list update and work function algorithms. Lecture Notes in Computer Science, 1643:289– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bein:1999:SPP

[1291] W. W. Bein, L. L. Larmore, and M. Chrobak. The 3-Server problem in the plane (extended abstract). Lecture Notes in Computer Science, 1643:301– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1999:QCI

[1292] V. Berry, T. Jiang, P. Kearney, and M. Li. Quartet cleaning: Improved algorithms and simulations. Lecture Notes in Computer Science, 1643:313– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaertner:1999:FRS

[1293] B. Gaertner. Fast and robust smallest enclosing balls. Lecture Notes in Computer Science, 1643:325–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nardelli:1999:ESM

[1294] E. Nardelli, M. Talamo, and P. Vocca. Efficient searching for multi-dimensional data made simple (extended abstract). Lecture Notes in Computer Science, 1643:339–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazelle:1999:GSR

[1295] B. Chazelle. Geometric searching over the rationals. Lecture Notes in Computer Science, 1643:354–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finocchiaro:1999:CDP

[1296] D. V. Finocchiaro and M. Pellegrini. On computing the diameter of a point set in high dimensional Euclidean space. Lecture Notes in Computer Science, 1643:366–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolliopoulos:1999:NLT

[1297] S. G. Kolliopoulos and S. Rao. A nearly linear-time approximation scheme for the Euclidean k-median problem. Lecture Notes in Computer Science, 1643: 378–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bar-Noy:1999:SMC

[1298] A. Bar-Noy, M. M. Halldorsson, G. Kortsarz, and R. Salman. Sum multicoloring of graphs. Lecture Notes in Computer Science, 1643:390–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krysta:1999:EAA

[1299] P. Krysta and K. Lorys. Efficient approximation algorithms for the achromatic number. Lecture Notes in Computer Science, 1643:402–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishii:1999:AVC

[1300] T. Ishii, H. Nagamochi, and T. Ibaraki.

Augmenting a (k −1)-vertex-connected multigraph to an 1-edge-connected and k-vertex-connected multigraph. Lecture Notes in Computer Science, 1643: 414–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verweij:1999:OAM

[1301] B. Verweij and K. Aardal. An optimisation algorithm for maximum independent set with applications in map labelling. Lecture Notes in Computer Science, 1643:426–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1999:DTM

[1302] M.-Y. Kao, T.-W. Lam, W.-K. Sung, and H.-F Ting. A decomposition theorem for maximum weight bipartite matchings with applications to evolutionary trees. Lecture Notes in Computer Science, 1643:438–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drori:1999:FES

[1303] L. Drori and D. Peleg. Faster exact solutions for some NP-Hard problems (extended abstract). Lecture Notes in Computer Science, 1643:450–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moura:1999:PAP

[1304] L. Moura. A polyhedral algorithm for packings and designs. Lecture Notes in Computer Science, 1643:462–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akhavi:1999:TPR

[1305] A. Akhavi. Threshold phenomena in random lattices and efficient reduction algorithms. Lecture Notes in Computer Science, 1643:476–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ageev:1999:FMN

[1306] A. A. Ageev. On finding the maximum number of disjoint cuts in Seymour graphs. Lecture Notes in Computer Science, 1643:490–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheriyan:1999:CPL

[1307] J. Cheriyan, T. Jordan, and R. Ravi. On 2-Coverings and 2-Packings of laminar families. Lecture Notes in Computer Science, 1643:510–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pak:1999:RCG

[1308] I. Pak. Random Cayley graphs with O(log |G|) generators are expanders. Lecture Notes in Computer Science, 1643:521–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hell:1999:FDA

[1309] P. Hell, R. Shamir, and R. Sharan.

A fully dynamic algorithm for recognizing and representing proper interval graphs. Lecture Notes in Computer Science, 1643:527–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1999:FGM

[1310] X. He, M.-Y. Kao, and H.-I. Lu. A fast general methodology for information theoretically optimal encodings of graphs. Lecture Notes in Computer Science, 1643:540–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ajtai:1999:GHI

[1311] Miklós Ajtai. Generating hard instances of the short basis problem. Lecture Notes in Computer Science, 1644: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440001.
pdf.

Cardelli:1999:WAC

[1312] Luca Cardelli. Wide area computation.

Lecture Notes in Computer Science, 1644:10–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1644/16440010.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1644/16440010.
pdf.

Compton:1999:PTC

[1313] Kevin J. Compton and Scott Dexter. Proof techniques for cryptographic protocols. Lecture Notes in Computer Science, 1644:25–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440025.
pdf.

Crary:1999:TSL

[1314] Karl Crary and Greg Morrisett. Type structure for low-level programming languages. Lecture Notes in Computer Science, 1644:40–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440040.
pdf.

Cucker:1999:RCF

[1315] Felipe Cucker. Real computations with fake numbers. Lecture Notes in Computer Science, 1644:55–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440055.
pdf.

Bruijn:1999:MAM

[1316] N. G. de Bruijn. A model for associative memory, a basis for thinking and consciousness. Lecture Notes in Computer Science, 1644:74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440074.
pdf.

DeBruijn:1999:MAM

[1317] N. G. De Bruijn. A model for associative memory, a basis for thinking and consciousness. Lecture Notes in Computer Science, 1644:74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440074.
pdf.

Edalat:1999:NIE

[1318] Abbas Edalat and Marko Krznarić.

Numerical integration with exact real arithmetic. Lecture Notes in Computer Science, 1644:90–104, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440090.
                                        pdf.

Hartmanis:1999:OAN

[1319] Juris Hartmanis. Observations about the nature and state of computer science (keynote address). Lecture Notes in Computer Science, 1644:105– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440105.
                                        pdf.

Rozenberg:1999:DCN

[1320] Grzegorz Rozenberg and Arto Salomaa. DNA computing: New ideas and paradigms. Lecture Notes in Computer Science, 1644:106–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440106.
                                        pdf.

Vitter:1999:ODS

[1321] Jeffrey Scott Vitter. Online data structures in external memory. Lecture Notes in Computer Science, 1644:119– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440119.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1644/16440119.
pdf.

Watanabe:1999:CLT

[1322] Osamu Watanabe. From computational learning theory to discovery science. Lecture Notes in Computer Science, 1644:134–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440134.
pdf.

Allender:1999:BDA

[1323] Eric Allender, Andris Ambainis, David A. Mix Barrington, Samir Datta, and Huong LêThanh. Bounded depth arithmetic circuits: Counting and closure. Lecture Notes in Computer Science, 1644:149–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440149.
pdf.

Alur:1999:PTL

[1324] Rajeev Alur, Kousha Etessami, Salvatore La Torre, and Doron Peled. Parametric temporal logic for “model measuring“. Lecture Notes in Computer Science, 1644:159–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440159.
                                        pdf.

Alur:1999:CHS

[1325] Rajeev Alur, Sampath Kannan, and Mihalis Yannakakis. Communicating hierarchical state machines. Lecture Notes in Computer Science, 1644:169– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440169.
                                        pdf.

Andreev:1999:SPR

[1326] Alexander E. Andreev, Juri L. Baskakov, Andrea E. F. Clementi, and José D. P. Rolim. Small pseudo-random sets yield hard functions: New tight explicit lower bounds for branching programs (extended abstract). Lecture Notes in Computer Science, 1644:179– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440179.
                                        pdf.

Bednarczyk:1999:GMP

[1327] Marek A. Bednarczyk and Andrzej M. Borzyszkowski. General morphisms of Petri nets (extended abstract). Lecture Notes in Computer Science, 1644: 190–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440190.
pdf.

Berman:1999:STI

[1328] Piotr Berman and Marek Karpinski.

On some tighter inapproximability results (extended abstract). Lecture Notes in Computer Science, 1644:200– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440200.
pdf.

Bouyer:1999:DCT

[1329] Patricia Bouyer and Antoine Petit. Decomposition and composition of timed automata. Lecture Notes in Computer Science, 1644:210–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440210.
pdf.

Buhrman:1999:NAI

[1330] Harry Buhrman, Tao Jiang, Ming Li, and Paul Vitányi. New applications of the incompressibility method (extended abstract). Lecture Notes in Computer Science, 1644:220–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440220.
                                        pdf.

Cardelli:1999:MTM

[1331] Luca Cardelli, Giorgio Ghelli, and Andrew D. Gordon. Mobility types for mobile ambients. Lecture Notes in Computer Science, 1644:230–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440230.
                                        pdf.

Clote:1999:PFL

[1332] Peter Clote. Protein folding, the levinthal paradox and rapidly mixing Markov chains. Lecture Notes in Computer Science, 1644:240–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440240.
                                        pdf.

Cortier:1999:DFS

[1333] Veronique Cortier, Harald Ganzinger, Florent Jacquemard, and Margus Veanes. Decidable fragments of simultaneous rigid reachability. Lecture Notes in Computer Science, 1644:250– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440250.
pdf.

Crochemore:1999:TCU

[1334] M. Crochemore, F. Mignosi, A. Restivo, and S. Salemi. Text compression using antidictionaries. Lecture Notes in Computer Science, 1644:261–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440261.
pdf.

DeSantis:1999:NIZ

[1335] A. De Santis, G. Di Crescenzo, and G. Persiano. Non-interactive zeroknowledge: A low-randomness characterization of NP (extended abstract). Lecture Notes in Computer Science, 1644:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santis:1999:NIZ

[1336] A. De Santis, G. Di Crescenzo, and G. Persiano. Non-interactive zeroknowledge: A low-randomness characterization of NP. Lecture Notes in Computer Science, 1644:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440271.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440271.
                                        pdf.

Dickhoefer:1999:TAT

[1337] M. Dickhoefer and T. Wilke. Timed alternating tree automata: The automata-theoretic solution to the TCTL model checking problem. Lecture Notes in Computer Science, 1644: 281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dickhofer:1999:TAT

[1338] Martin Dickhöfer and Thomas Wilke. Timed alternating tree automata: The automata-theoretic solution to the TCTL model checking problem. Lecture Notes in Computer Science, 1644: 281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440281.
                                        pdf.

Dodis:1999:STT

[1339] Yevgeniy Dodis and Sanjeev Khanna. Space-time tradeoffs for graph properties. Lecture Notes in Computer Science, 1644:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1644/16440291.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1644/16440291.pdf.

Dufourd:1999:BRP

[1340] C. Dufourd, P. Jančar, and Ph.

Schnoebelen. Boundedness of reset P/T nets. Lecture Notes in Computer Science, 1644:301–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440301.
pdf.

Engelfriet:1999:TWF

[1341] Joost Engelfriet and Hendrik Jan Hoogeboom. Two-way finite state transducers and monadic second-order logic. Lecture Notes in Computer Science, 1644:311–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1644/16440311.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1644/16440311.pdf.

Flesca:1999:POR

[1342] Sergio Flesca and Sergio Greco. Partially ordered regular languages for graph queries (extended abstract). Lecture Notes in Computer Science, 1644: 321–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440321.
pdf.

Frick:1999:DFO

[1343] Markus Frick and Martin Grohe. Deciding first-order properties of locally tree-decomposable graphs. Lecture Notes in Computer Science, 1644:331– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440331.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440331.
                                        pdf.

Galpin:1999:CPA

[1344] Vashti Galpin. Comparison of process algebra equivalences using formats. Lecture Notes in Computer Science, 1644:341–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1644/16440341.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440341.
                                        pdf.

Gavoille:1999:CRT

[1345] Cyril Gavoille and Nicolas Hanusse. Compact routing tables for graphs of bounded genus (extended abstract). Lecture Notes in Computer Science, 1644:351–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1644/16440351.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440351.
                                        pdf.

Gottlob:1999:CLC

[1346] Georg Gottlob, Nicola Leone, and Francesco Scarcello. Computing LOGCFL certificates. Lecture Notes in Computer Science, 1644:361–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440361.
pdf.

Grossi:1999:ETM

[1347] Roberto Grossi and Giuseppe F. Italiano. Efficient techniques for maintaining multidimensional keys in linked data structures (extended abstract). Lecture Notes in Computer Science, 1644:372–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1644/16440372.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1644/16440372.
pdf.

Gupta:1999:CPT

[1348] Arvind Gupta, Damon Kaller, and Thomas Shermer. On the complements of partial k-trees. Lecture Notes in Computer Science, 1644:382– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440382.
pdf.

Hammar:1999:ARK

[1349] Mikael Hammar and Bengt J. Nilsson. Approximation results for kinetic variants of TSP. Lecture Notes in Computer Science, 1644:392–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440392.
                                        pdf.

Hassin:1999:DPP

[1350] Yehuda Hassin and David Peleg. Distributed probabilistic polling and applications to proportionate agreement. Lecture Notes in Computer Science, 1644:402–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1644/16440402.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440402.
                                        pdf.

Zhang:1999:MLCb

[1351] Yu Zhang and Alan K. Mackworth. A multi-level constraintbased controller for the Dynamo98 robot soccer team. Lecture Notes in Computer Science, 1644:402–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1604/16040402.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1604/16040402.
                                        pdf.

Hirshfeld:1999:BED

[1352] Yoram Hirshfeld and Mark Jerrum.

Bisimulation equivalence is decidable for normed process algebra (extended abstract). Lecture Notes in Computer Science, 1644:412–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440412.
pdf.

Hirshfeld:1999:FDM

[1353] Yoram Hirshfeld and Alexander Rabinovich. A framework for decidable metrical logics. Lecture Notes in Computer Science, 1644:422–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440422.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440422.
pdf.

Hromkovic:1999:PVI

[1354] Juraj Hromkovič and Georg Schnitger. On the power of Las Vegas. II. two-way finite automata. Lecture Notes in Computer Science, 1644:433– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440433.
pdf.

Iwama:1999:SMI

[1355] Kazuo Iwama, David Manlove, Shuichi Miyazaki, and Yasufumi Morita. Stable marriage with incomplete lists and ties. Lecture Notes in Computer Science, 1644:443–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440443.
pdf.

Jiang:1999:ACC

[1356] Tao Jiang, Ming Li, and Paul Vitányi. Average-case complexity of shellsort (preliminary version). Lecture Notes in Computer Science, 1644:453–462, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440453.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440453.

pdf. See [1] for the final version of this paper.

Kim:1999:LTC

[1357] Dong Kyue Kim and Kunsoo Park. Linear-time construction of twodimensional suffix trees (extended abstract). Lecture Notes in Computer Science, 1644:463–472, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1644/16440463.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1644/16440463.pdf.

Kirsten:1999:CBS

[1358] Daniel Kirsten. A connection between the star problem and the finite power property in trace monoids (extended abstract). Lecture Notes in Computer Science, 1644:473–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440473.
pdf.

Kirsten:1999:TTA

[1359] Daniel Kirsten and Jerzy Marcinkowski.

Two techniques in the area of the star problem. Lecture Notes in Computer Science, 1644:483–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440483.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440483.
pdf.

Krause:1999:AOV

[1360] Matthias Krause, Petr Savický, and Ingo Wegener. Approximations by OBDDs and the variable ordering problem. Lecture Notes in Computer Science, 1644:493–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440493.
pdf.

Kucera:1999:SPS

[1361] Antonín Kučera and Richard Mayr. Simulation preorder on simple process algebras. Lecture Notes in Computer Science, 1644:503–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1644/16440503.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1644/16440503.pdf.

Laneve:1999:SC

[1362] Cosimo Laneve and Björn Victor. Solos in concert. Lecture Notes in Computer Science, 1644:513–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440513.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440513.
                                        pdf.

Lanthier:1999:SAP

[1363] Mark Lanthier, Anil Maheshwari, and Jörg-Rüdiger Sack. Shortest anisotropic paths on terrains. Lecture Notes in Computer Science, 1644:524– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440524.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440524.
                                        pdf.

Lepisto:1999:RBL

[1364] A. Lepistö. Relations between local and global periodicity of words. Lecture Notes in Computer Science, 1644: 534–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440534.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440534.
pdf.

Lepistoe:1999:RBL

[1365] A. Lepistoe. Relations between local and global periodicity of words (extended abstract). Lecture Notes in Computer Science, 1644:534–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lingas:1999:EMC

[1366] Andrzej Lingas, Hans Olsson, and Anna Östlin. Efficient merging, construction, and maintenance of evolutionary trees. Lecture Notes in Computer Science, 1644:544–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440544.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440544.
pdf.

Miculan:1999:FLS

[1367] Marino Miculan. Formalizing a lazy substitution proof system for μ-calculus in the calculus of inductive constructions. Lecture Notes in Computer Science, 1644:554–564, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1644/16440554.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440554.
                                        pdf.

Nichitiu:1999:LED

[1368] Codrin Nichitiu and Eric Rémila. Leader election by d dimensional cellular automata. Lecture Notes in Computer Science, 1644:565–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440565.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440565.
                                        pdf.

Niedermeier:1999:NUB

[1369] Rolf Niedermeier and Peter Rossmanith. New upper bounds for MaxSat. Lecture Notes in Computer Science, 1644:575–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1644/16440575.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1644/16440575.pdf.

Olshevsky:1999:PRE

[1370] Vadim Olshevsky and Victor Y. Pan. Polynomial and rational evaluation and interpolation (with structured matrices). Lecture Notes in Computer Science, 1644:585–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440585.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1644/16440585.
pdf.

Pagh:1999:LRS

[1371] Rasmus Pagh. Low redundancy in static dictionaries with O(1) worst case lookup time. Lecture Notes in Computer Science, 1644:595–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440595.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440595.
pdf.

Peichl:1999:FAG

[1372] Timo Peichl and Heribert Vollmer.

Finite automata with generalized acceptance criteria. Lecture Notes in Computer Science, 1644:605–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440605.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440605.
pdf.

Peleg:1999:VAD

[1373] David Peleg and Eilon Reshef. A variant of the arrow distributed directory with low average complexity (extended abstract). Lecture Notes in Computer Science, 1644:615–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440615.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440615.
                                        pdf.

Power:1999:CFC

[1374] John Power and Hayo Thielecke. Closed Freyd- and κ-categories. Lecture Notes in Computer Science, 1644: 625–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440625.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440625.
                                        pdf.

Riecke:1999:TEC

[1375] Jon G. Riecke and Hayo Thielecke. Typed exceptions and continuations cannot macro-express each other. Lecture Notes in Computer Science, 1644: 635–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440635.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440635.
                                        pdf.

Rutten:1999:APS

[1376] J. J. M. M. Rutten. Automata, power series, and coinduction: Taking input derivatives seriously (extended abstract). Lecture Notes in Computer Science, 1644:645–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440645.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1644/16440645.
pdf.

Sanders:1999:AMS

[1377] Peter Sanders. Accessing multiple sequences through set associative caches. Lecture Notes in Computer Science, 1644:655–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440655.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440655.
pdf.

Senizergues:1999:XXX

[1378] Géraud Sénizergues. T (A) = T (B)?

Lecture Notes in Computer Science, 1644:665–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1644/16440665.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1644/16440665.
pdf.

Szegedy:1999:MVL

[1379] Mario Szegedy. Many-valued logics and holographic proofs. Lecture Notes in Computer Science, 1644:676– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1644/16440676.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1644/16440676.
pdf.

Umans:1999:CIS

[1380] Christopher Umans. On the complexity and inapproximability of shortest implicant problems. Lecture Notes in Computer Science, 1644:687–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440687.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440687.
                                        pdf.

Weihrauch:1999:WPT

[1381] Klaus Weihrauch and Ning Zhong. The wave propagator is Turing computable. Lecture Notes in Computer Science, 1644:697–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1644/16440697.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440697.
                                        pdf.

Woeginger:1999:FAW

[1382] Gerhard J. Woeginger. An FPTAS for agreeably weighted variance on a single machine (extended abstract). Lecture Notes in Computer Science, 1644: 707–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1644/16440707.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1644/16440707.
                                        pdf.

Tiskin:1999:EBS

[1383] A. Tiskin. Erratum: Bulk-synchronous parallel multiplication of Boolean matrices. Lecture Notes in Computer Science, 1644:717–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1644/16440717.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1644/16440717.pdf. See [2].

Anonymous:1999:AIi

[1384] Anonymous. Author index. Lecture Notes in Computer Science, 1644:719– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1644/1644auth.pdf.

Kida:1999:SAP

[1385] Takuya Kida, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Shift-and approach to pattern matching in LZW compressed text. Lecture Notes in Computer Science, 1645:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450001.
pdf.

Navarro:1999:GPA

[1386] Gonzalo Navarro and Mathieu Raffinot. A general practical approach to pattern matching over Ziv–Lempel compressed text. Lecture Notes in Computer Science, 1645:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450014.
pdf.

Shibata:1999:PMT

[1387] Yusuke Shibata, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Pattern matching in text compressed by using antidictionaries. Lecture Notes in Computer Science, 1645:37– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450037.
pdf.

Liben-Nowell:1999:SSD

[1388] David Liben-Nowell. On the structure of syntenic distance. Lecture Notes in Computer Science, 1645:50– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450050.
pdf.

Batzoglou:1999:PMR

[1389] Serafim Batzoglou and Sorin Istrail. Physical mapping with repeated probes: The hypergraph superstring problem. Lecture Notes in Computer Science, 1645:66–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450066.
pdf.

El-Mabrouk:1999:HGR

[1390] Nadia El-Mabrouk and David Sankoff.

Hybridization and genome rearrangement. Lecture Notes in Computer Science, 1645:78–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1645/16450078.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1645/16450078.pdf.

Ben-Dor:1999:CPS

[1391] Amir Ben-Dor, Itsik Pe’er, Ron Shamir, and Roded Sharan. On the complexity of positional sequencing by hybridization. Lecture Notes in Computer Science, 1645:88–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450088.
pdf.

Lancia:1999:GGS

[1392] Giuseppe Lancia and R. Ravi.

GESTALT: Genomic Steiner alignments. Lecture Notes in Computer Science, 1645:101–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1645/16450101.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1645/16450101.pdf.

Hirschberg:1999:BNS

[1393] Daniel S. Hirschberg. Bounds on the number of string subsequences. Lecture Notes in Computer Science, 1645: 115–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1645/16450115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1645/16450115.
                                        pdf.

Sim:1999:APS

[1394] Jeong Seop Sim, Costas S. Iliopoulos, Kunsoo Park, and William F. Smyth. Approximate periods of strings. Lecture Notes in Computer Science, 1645: 123–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1645/16450123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1645/16450123.
                                        pdf.

Brodal:1999:FMP

[1395] Gerth Stølting Brodal, Rune B. Lyngsø, Christian N. S. Pedersen, and Jens Stoye. Finding maximal pairs with bounded gap. Lecture Notes in Computer Science, 1645:134– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450134.
pdf.

Yakoo:1999:DDS

[1396] H. Yakoo. A dynamic data structure for reverse lexicographically sorted prefixes. Lecture Notes in Computer Science, 1645:150–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokoo:1999:DDS

[1397] Hidetoshi Yokoo. A dynamic data structure for reverse lexicographically sorted prefixes. Lecture Notes in Computer Science, 1645:150–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450150.
pdf.

Navarro:1999:NIM

[1398] Gonzalo Navarro and Ricardo Baeza-Yates. A new indexing method for approximate string matching. Lecture Notes in Computer Science, 1645:163– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450163.
pdf.

Karhumaeki:1999:CSI

[1399] J. Karhumaeki, W. Plandowski, and W. Rytter. The compression of subsegments of images described by finite automata. Lecture Notes in Computer Science, 1645:186–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaki:1999:CSI

[1400] Juhani Karhumäki, Wojciech Plandowski, and Wojciech Rytter. The compression of subsegments of images described by finite automata. Lecture Notes in Computer Science, 1645:186– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1645/16450186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1645/16450186.
                                        pdf.

Ristov:1999:ZLC

[1401] Strahil Ristov and Eric Laporte. Ziv lempel compression of huge natural language data tries using suffix arrays. Lecture Notes in Computer Science, 1645:196–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1645/16450196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1645/16450196.
                                        pdf.

Akutsu:1999:MSE

[1402] Tatsuya Akutsu, Kyotetsu Kanaya, Akira Ohyama, and Asao Fujiyama.

Matching of spots in 2D electrophoresis images. point matching under nonuniform distortions. Lecture Notes in Computer Science, 1645:212–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450212.
pdf.

Wenk:1999:AED

[1403] Carola Wenk. Applying an edit distance to the matching of tree ring sequences in dendrochronology. Lecture Notes in Computer Science, 1645:223– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450223.
pdf.

Navarro:1999:FMD

[1404] Gonzalo Navarro and Ricardo Baeza-Yates. Fast multi-dimensional approximate pattern matching. Lecture Notes in Computer Science, 1645:243– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450243.
pdf.

Wang:1999:FCR

[1405] Zhuozhi Wang and Kaizhong Zhang.

Finding common RNA secondary structures from RNA sequences. Lecture Notes in Computer Science, 1645: 258–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1645/16450258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1645/16450258.
                                        pdf.

Zhuozhi:1999:FCR

[1406] W. Zhuozhi and Z. Kaizhong. Finding common RNA secondary structures from RNA sequences. Lecture Notes in Computer Science, 1645:258–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1999:FCS

[1407] Patricia A. Evans. Finding common subsequences with arcs and pseudoknots. Lecture Notes in Computer Science, 1645:270–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1645/16450270.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1645/16450270.pdf.

Kaizhong:1999:CSB

[1408] Z. Kaizhong, W. Lusheng, and M. Bin. Computing similarity between RNA structures. Lecture Notes in Computer Science, 1645:281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1999:CSB

[1409] Kaizhong Zhang, Lusheng Wang, and Bin Ma. Computing similarity between RNA structures. Lecture Notes in Computer Science, 1645:281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1645/16450281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1645/16450281.
pdf.

Jennings:1999:AOSb

[1410] N. R. Jennings. Agent-oriented software engineering. Lecture Notes in Computer Science, 1647:1–7, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herlea:1999:SBR

[1411] D. E. Herlea, C. M. Jonker, J. Treur, and N. J. E. Wijngaards. Specification of behavioural requirements within compositional multi-agent system design. Lecture Notes in Computer Science, 1647:8–27, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barber:1999:AOD

[1412] K. S. Barber, T. H. Liu, and D. C.

Han. Agent-oriented design. Lecture Notes in Computer Science, 1647:28– 40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charlton:1999:DPM

[1413] P. Charlton and E. Mamdani. A developer’s perspective on mult-agent system design. Lecture Notes in Computer Science, 1647:41–51, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergenti:1999:DER

[1414] F. Bergenti and A. Poggi. A development environment for the realization of open and scalable multi-agent systems. Lecture Notes in Computer Science, 1647:52–62, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botti:1999:MAH

[1415] V. Botti, C. Carrascosa, V. Julian, and J. Soler. Modelling agents in hard realtime environments. Lecture Notes in Computer Science, 1647:63–76, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cremonini:1999:MAS

[1416] M. Cremonini, A. Omicini, and F. Zambonelli. Multi-agent systems on the Internet. extending the scope of coordination towards security and topology. Lecture Notes in Computer Science, 1647:77–88, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ElFallah-Seghrouchni:1999:PEM

[1417] Amal El Fallah-Seghrouchni, S. Haddad, and H. Mazouzi. Protocol engineering for multi-agent interaction. Lecture Notes in Computer Science, 1647:89–101, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pitt:1999:DAC

[1418] J. Pitt and A. Mamdani. Designing agent communication languages for multi-agent systems. Lecture Notes in Computer Science, 1647:102–114, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carron:1999:TAC

[1419] T. Carron, H. Proton, and O. Boissier.

A temporal agent communication language for dynamic multi-agent systems. Lecture Notes in Computer Science, 1647:115–127, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amandi:1999:MPL

[1420] A. Amandi, A. Zunino, and R. Iturregui. Multi-paradigm languages supporting multi-agent development. Lecture Notes in Computer Science, 1647: 128–139, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroeder:1999:EAF

[1421] M. Schroeder. An efficient argumentation framework for negotiating autonomous agents. Lecture Notes in Computer Science, 1647:140–149, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbuceanu:1999:NSP

[1422] M. Barbuceanu. Negotiating service provisioning. Lecture Notes in Computer Science, 1647:150–161, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffiths:1999:CPS

[1423] N. Griffiths and M. Luck. Cooperative plan selection through trust. Lecture Notes in Computer Science, 1647:162– 174, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1999:ESR

[1424] N. David, J. S. Sichman, and H. Coelho. Extending social reasoning to cope with multiple partner coalitions. Lecture Notes in Computer Science, 1647:175–187, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1999:BMA

[1425] C. Castelfranchi and R. Falcone. Basic mental attitudes of a collaborating agent: Cognitive primitives for MAS. Lecture Notes in Computer Science, 1647:188–209, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1999:SS

[1426] A. Moreno, U. Cortes, and T. Sales. Subjective situations. Lecture Notes in Computer Science, 1647:210–220, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1999:FAM

[1427] C. M. Jonker and J. Treur. Formal analysis of models for the dynamics of trust based on experiences. Lecture Notes in Computer Science, 1647:221– 232, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daswani:1999:EEC

[1428] Neil Daswani and Dan Boneh. Experimenting with electronic commerce on the PalmPilot. Lecture Notes in Computer Science, 1648:1– 16, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480001.
pdf.

Krawczyk:1999:BCC

[1429] Hugo Krawczyk. Blinding of credit card numbers in the SET protocol. Lecture Notes in Computer Science, 1648:17– 28, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480017.
pdf.

Juels:1999:TTS

[1430] Ari Juels. Trustee tokens: Simple and practical anonymous digital coin tracing. Lecture Notes in Computer Science, 1648:29–45, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480029.
pdf.

Sander:1999:FCN

[1431] Tomas Sander and Amnon Ta-Shma. Flow control: A new approach for anonymity control in electronic cash systems. Lecture Notes in Computer Science, 1648:46–61, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480046.
                                        pdf.

Yacobi:1999:RMC

[1432] Yacov Yacobi. Risk management for E-cash systems with partial real-time audit. Lecture Notes in Computer Science, 1648:62–71, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480062.
                                        pdf.

Ezawa:1999:AEC

[1433] Kazuo J. Ezawa, Gregory Napiorkowski, and Mariusz Kossarski. Assessment of effectiveness of counterfeit transaction detection systems for smart card based electronic cash. Lecture Notes in Computer Science, 1648:72– 85, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480072.
                                        pdf.

Kohlas:1999:RAP

[1434] Reto Kohlas and Ueli Maurer. Reasoning about public-key certification: On bindings between entities and public keys. Lecture Notes in Computer Science, 1648:86–103, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480086.
pdf.

Fox:1999:OCS

[1435] Barbara Fox and Brian LaMacchia. Online certificate status checking in financial transactions: The case for re-issuance. Lecture Notes in Computer Science, 1648:104–117, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480104.
pdf.

Shamir:1999:PSS

[1436] Adi Shamir and Nicko van Someren.

Playing “hide and seek” with stored keys. Lecture Notes in Computer Science, 1648:118–124, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1648/16480118.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1648/16480118.pdf.

Breitbach:1999:CCM

[1437] Markus Breitbach and Hideki Imai.

On channel capacity and modulation of watermarks in digital still images. Lecture Notes in Computer Science, 1648:125–139, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480125.
                                        pdf.

Abdalla:1999:TMB

[1438] Michel Abdalla, Yuval Shavitt, and Avishai Wool. Towards making broadcast encryption practical. Lecture Notes in Computer Science, 1648:140– 157, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480140.
                                        pdf.

Kravitz:1999:CAC

[1439] David W. Kravitz and David M. Goldschlag. Conditional access concepts and principles. Lecture Notes in Computer Science, 1648:158–172, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480158.
                                        pdf.

Callas:1999:FUI

[1440] Jon Callas, Joan Feigenbaum, David Goldschlag, and Erin Sawyer. Fair use, intellectual property, and the information economy (panel session summary). Lecture Notes in Computer Science, 1648:173–183, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480173.
pdf.

Schechter:1999:AAM

[1441] Stuart Schechter, Todd Parnell, and Alexander Hartemink. Anonymous authentication of membership in dynamic groups. Lecture Notes in Computer Science, 1648:184–195, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480184.
pdf.

Ateniese:1999:SOI

[1442] Giuseppe Ateniese and Gene Tsudik.

Some open issues and new directions in group signatures. Lecture Notes in Computer Science, 1648:196–211, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1648/16480196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1648/16480196.
pdf.

MacKenzie:1999:AIH

[1443] Philip MacKenzie and Jeffrey Sorensen.

Anonymous investing: Hiding the identities of stockholders. Lecture Notes in Computer Science, 1648:212–229, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480212.
                                        pdf.

Stubblebine:1999:FLA

[1444] Stuart G. Stubblebine and Paul F. Syverson. Fair on-line auctions without special trusted parties. Lecture Notes in Computer Science, 1648:230– 240, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480230.
                                        pdf.

Frankel:1999:CRA

[1445] Yair Frankel and Moti Yung. Cryptosystems robust against “dynamic faults” meet enterprise needs for organizational “change control”. Lecture Notes in Computer Science, 1648:241– 252, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1648/16480241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1648/16480241.
                                        pdf.

Jakobsson:1999:IMI

[1446] Markus Jakobsson and Joy Müller. Improved magic ink signatures using hints. Lecture Notes in Computer Science, 1648:253–267, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1648/16480253.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1648/16480253.pdf.

Ullman:1999:SAD

[1447] J. D. Ullman. Some advances in datamining techniques. Lecture Notes in Computer Science, 1649:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bar-Yossef:1999:QST

[1448] Z. Bar-Yossef, Y. Kanza, Y. Kogan, and W. Nutt. Querying semantically tagged documents on the World-WideWeb. Lecture Notes in Computer Science, 1649:2–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konopnicki:1999:WEQ

[1449] D. Konopnicki and O. Shmueli. WWW exploration queries. Lecture Notes in Computer Science, 1649:20–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shoval:1999:SFM

[1450] P. Shoval, B. Shapira, and U. Hanani.

Strategies for filtering E-mail messages combining content-based and sociological filtering with user-stereotypes. Lecture Notes in Computer Science, 1649: 40–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliveira:1999:IQE

[1451] C. Oliveira, L. G. V. Resende, and R. Lehmann. Interactive query expansion in a meta-search engine. Lecture Notes in Computer Science, 1649:50– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henrich:1999:OQC

[1452] A. Henrich and S. Jamin. On the optimization of queries containing regular path expressions. Lecture Notes in Computer Science, 1649:58–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumann:1999:DAA

[1453] P. Baumann. A database array algebra for spatio-temporal data and beyond. Lecture Notes in Computer Science, 1649:76–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sapir:1999:DRT

[1454] A. Sapir and E. Gudes. Dynamic relationships and their propagation and concurrency semantics in objectoriented databases. Lecture Notes in Computer Science, 1649:94–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfson:1999:TMO

[1455] O. Wolfson, P. Sistla, B. Xu, and J. Zhou. Tracking moving objects using database technology in DOMINO. Lecture Notes in Computer Science, 1649: 112–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mengchi:1999:ODO

[1456] L. Mengchi. OLOG: A deductive object database language. Lecture Notes in Computer Science, 1649:120–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motro:1999:MFM

[1457] A. Motro. Multiplex: A formal model for multidatabases and its implementation. Lecture Notes in Computer Science, 1649:138–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golani:1999:TAR

[1458] M. Golani and O. Etzion. Temporal active rules. Lecture Notes in Computer Science, 1649:159–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birk:1999:CEJ

[1459] Y. Birk and M. Mokryn. Cost-effective jukebox storage via hybrid file-block caching. Lecture Notes in Computer Science, 1649:173–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christophides:1999:ODI

[1460] V. Christophides, C. Houstis, S. Lalis, and H. Tsalapata. Onthology-driven integration of scientific repositories. Lecture Notes in Computer Science, 1649:190–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenhoff:1999:GVM

[1461] M. Schoenhoff and M. Straessler.

Global version management for a federated turbine design environment. Lecture Notes in Computer Science, 1649: 203–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peleg:1999:OPD

[1462] M. Peleg and D. Dori. From objectprocess diagrams to a natural objectprocess language. Lecture Notes in Computer Science, 1649:221–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelstein:1999:CEC

[1463] O. Edelstein, A. Yaeli, and G. Zodik. e-Colabra: An enterprise collaboration and reuse environment. Lecture Notes in Computer Science, 1649:229– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohan:1999:WMI

[1464] C. Mohan. Workflow management in the Internet age. Lecture Notes in Computer Science, 1649:237–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koksal:1999:CBW

[1465] P. Koksal, I. Cingil, and A. Dogac. A component-based workflow system with dynamic modifications. Lecture Notes in Computer Science, 1649:238– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cicekli:1999:TRA

[1466] N. K. Cicekli. A temporal reasoning approach to model workflow activities. Lecture Notes in Computer Science, 1649:256–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gidron:1999:DCE

[1467] Y. Gidron, I. Ben-Shaul, and Y. Aridor. Dynamic configuration and enforcement of access control for mobile components. Lecture Notes in Computer Science, 1649:267–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolff:1999:MVP

[1468] J. E. Wolff and A. B. Cremers. The my VIEW project: A data warehousing approach to personalized digital libraries. Lecture Notes in Computer Science, 1649:277–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nestorov:1999:IDM

[1469] S. Nestorov and S. Tsur. Integrating data mining with relational DBMS: A tightly-coupled approach. Lecture Notes in Computer Science, 1649:295– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carmel:1999:MDH

[1470] D. Carmel and Y. S. Maarek. Morphological disambiguation for Hebrew search systems. Lecture Notes in Computer Science, 1649:312–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arcos:1999:ADC

[1471] Josep Lluís Arcos, Dolores Cañamero, and Ramon López de Mántaras.

Affect-driven CBR to generate expressive music. Lecture Notes in Computer Science, 1650:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500001.
pdf.

Blanzieri:1999:PBM

[1472] Enrico Blanzieri and Francesco Ricci. Probability based metrics for nearest neighbor classification and casebased reasoning. Lecture Notes in Computer Science, 1650:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500014.
pdf.

Branting:1999:AEI

[1473] L. Karl Branting. Active exploration in instance-based preference modeling. Lecture Notes in Computer Science, 1650:29–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500029.
pdf.

Branting:1999:MDE

[1474] L. Karl Branting and Yi Tao. A multiple-domain evaluation of stratified case-based reasoning. Lecture Notes in Computer Science, 1650:44– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500044.
pdf.

Brueninghaus:1999:BCB

[1475] S. Brueninghaus and K. D. Ashley.

Bootstrapping case base development with annotated case summaries. Lecture Notes in Computer Science, 1650: 59–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruninghaus:1999:BCB

[1476] Stefanie Brüninghaus and Kevin D.

Ashley. Bootstrapping case base development with annotated case summaries. Lecture Notes in Computer Science, 1650:59–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1650/16500059.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1650/16500059.pdf.

Carrick:1999:ACS

[1477] Christina Carrick, Qiang Yang, Irene Abi-Zeid, and Luc Lamontagne. Activating CBR systems through autonomous information gathering. Lecture Notes in Computer Science, 1650: 74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500074.
                                        pdf.

Coello:1999:ICH

[1478] Juan Manuel Adán Coello and Ronaldo Camilo dos Santos. Integrating CBR and heuristic search for learning and reusing solutions in real-time task scheduling. Lecture Notes in Computer Science, 1650:89– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500089.
                                        pdf.

Fuchs:1999:TUT

[1479] Béatrice Fuchs, Jean Lieber, Alain Mille, and Amedeo Napoli. Towards a unified theory of adaptation in case-based reasoning. Lecture Notes in Computer Science, 1650:104– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500104.
                                        pdf.

Fuchs:1999:KLT

[1480] Béatrice Fuchs and Alain Mille. A knowledge-level task model of adaptation in case-based reasoning. Lecture Notes in Computer Science, 1650:118– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500118.
pdf.

Goeker:1999:DUC

[1481] M. Goeker and T. Roth-Berghofer. Development and utilization of a casebased help-desk support system in a corporate environment. Lecture Notes in Computer Science, 1650:132–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goker:1999:DUC

[1482] Mehmet Göker and Thomas Roth-Berghofer. Development and utilization of a case-based help-desk support system in a corporate environment. Lecture Notes in Computer Science, 1650:132–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1650/16500132.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1650/16500132.
pdf.

Gomez-Albarran:1999:MCL

[1483] Mercedes Gómez-Albarrán, Pedro A.

González-Calero, Belén Díaz-Agudo, and Carlos Fernández-Conde. Modelling the CBR life cycle using description logics. Lecture Notes in Computer Science, 1650:147–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500147.
                                        pdf.

Garza:1999:EAC

[1484] Andrés Gómez de Silva Garza and Mary Lou Maher. An evolutionary approach to case adaptation. Lecture Notes in Computer Science, 1650:162– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500162.
                                        pdf.

GomezdeSilvaGarza:1999:EAC

[1485] A. Gomez de Silva Garza and M. L. Maher. An evolutionary approach to case adaptation. Lecture Notes in Computer Science, 1650:162–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonWangenheim:1999:RCB

[1486] C. Gresse von Wangenheim. RE-MEX — A case-based approach for reusing software measurement experienceware. Lecture Notes in Computer Science, 1650:173–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wangenheim:1999:RCB

[1487] Christiane Gresse von Wangenheim. REMEX — A case-based approach for reusing software measurement experienceware. Lecture Notes in Computer Science, 1650:173–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500173.
pdf.

Lawton:1999:ULT

[1488] James H. Lawton, Roy M. Turner, and Elise H. Turner. A unified long-term memory system. Lecture Notes in Computer Science, 1650:188– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500188.
pdf.

Leake:1999:CCI

[1489] David B. Leake and David C. Wilson. Combining CBR with interactive knowledge acquisition, manipulation, and reuse. Lecture Notes in Computer Science, 1650:203–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500203.
pdf.

Leake:1999:WEW

[1490] David B. Leake and David C. Wilson. When experience is wrong: Examining CBR for changing tasks and environments. Lecture Notes in Computer Science, 1650:218–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500218.
                                        pdf.

Lei:1999:CLR

[1491] Celestino Lei, Otakar Babka, and Laurinda A. G. Garanito. Case library reduction applied to pile foundations. Lecture Notes in Computer Science, 1650:233–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1650/16500233.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500233.
                                        pdf.

McLaren:1999:CRA

[1492] Bruce McLaren and Kevin D. Ashley. Case representation, acquisition, and retrieval in SIROCCO. Lecture Notes in Computer Science, 1650:248– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500248.
                                        pdf.

Melis:1999:FIP

[1493] Erica Melis and Carsten Ullrich. Flexibly interleaving processes. Lecture Notes in Computer Science, 1650:263– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500263.
pdf.

Munoz-Avila:1999:CRP

[1494] Héctor Muñoz-Avila. A case retention policy based on detrimental retrieval. Lecture Notes in Computer Science, 1650:276–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1650/16500276.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1650/16500276.
pdf.

Munoz-Avila:1999:UGC

[1495] Héctor Muñoz-Avila, Daniel C. Mc-Farlane, David W. Aha, Len Breslow, James A. Ballas, and Dana S. Nau. Using guidelines to constrain interactive case-based HTN planning. Lecture Notes in Computer Science, 1650:288– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500288.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500288.
pdf.

Portinale:1999:SQC

[1496] Luigi Portinale, Pietro Torasso, and Paolo Tavano. Speed-up, quality, and competence in multi-modal case-based reasoning. Lecture Notes in Computer Science, 1650:303–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500303.
                                        pdf.

Seitz:1999:CBM

[1497] Alexander Seitz. A case-based methodology for planning individualized case oriented tutoring. Lecture Notes in Computer Science, 1650:318–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500318.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500318.
                                        pdf.

Smyth:1999:BCC

[1498] Barry Smyth and Elizabeth McKenna. Building compact competent casebases. Lecture Notes in Computer Science, 1650:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1650/16500329.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1650/16500329.pdf.

Smyth:1999:FBR

[1499] Barry Smyth and Elizabeth McKenna. Footprint-based retrieval. Lecture Notes in Computer Science, 1650:343– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1650/16500343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500343.
pdf.

Abi-Zeid:1999:CAC

[1500] Iréne Abi-Zeid, Qiang Yang, and Luc Lamontagne. Is CBR applicable to the coordination of search and rescue operations? A feasibility study. Lecture Notes in Computer Science, 1650:358– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500358.
pdf.

Auriol:1999:ICB

[1501] Eric Auriol, Richard M. Crowder, Rob MacKendrick, Roger Rowe, and Thomas Knudsen. Integrating casebased reasoning and hypermedia documentation: An application for the diagnosis of a welding robot at odense steel shipyard. Lecture Notes in Computer Science, 1650:372–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500372.
pdf.

Bellazzi:1999:IRB

[1502] Riccardo Bellazzi, Stefania Montani, Luigi Portinale, and Alberto Riva. Integrating rule-based and case-based decision making in diabetic patient management. Lecture Notes in Computer Science, 1650:386– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500386.
                                        pdf.

Conruy:1999:MCK

[1503] Noël Conruy and David Grosser. Managing complex knowledge in natural sciences. Lecture Notes in Computer Science, 1650:401–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500401.
                                        pdf.

Conruyt:1999:MCK

[1504] N. Conruyt and D. Grosser. Managing complex knowledge in natural sciences. Lecture Notes in Computer Science, 1650:401–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuddihy:1999:EME

[1505] Paul Cuddihy and William Cheetham. ELSI: A medical equipment diagnostic system. Lecture Notes in Computer Science, 1650:415–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1650/16500415.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500415.
pdf.

Fagan:1999:CBR

[1506] Michael Fagan and Konrad Bloor.

Case-based reasoning for candidate list extraction in a marketing domain. Lecture Notes in Computer Science, 1650: 426–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500426.
pdf.

Ficet-Cauchard:1999:CRI

[1507] Valérie Ficet-Cauchard, Christine Porquet, and Marinette Revenu. CBR for the reuse of image processing knowledge: A recursive retrieval/ adaptation strategy. Lecture Notes in Computer Science, 1650:438–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500438.
pdf.

Han:1999:VFG

[1508] Younghyun Han and Kunwoo Lee. Virtual function generators: Representing and reusing underlying design concepts in conceptual synthesis of mechanisms for function generation. Lecture Notes in Computer Science, 1650: 453–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500453.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500453.
                                        pdf.

Hayes:1999:SCV

[1509] Conor Hayes and Padraig Cunningham. Shaping a CBR view with XML. Lecture Notes in Computer Science, 1650:468–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1650/16500468.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500468.
                                        pdf.

Leake:1999:IIR

[1510] David B. Leake, Larry Birnbaum, Kristian Hammond, Cameron Marlow, and Hao Yang. Integrating information resources: A case study of engineering design support. Lecture Notes in Computer Science, 1650:482– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500482.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500482.
                                        pdf.

Main:1999:HCB

[1511] Julie Main and Tharam S. Dillon. A hybrid case-based reasoner for footwear design. Lecture Notes in Computer Science, 1650:497–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500497.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500497.
pdf.

Melchiors:1999:FMC

[1512] Cristina Melchiors and Liane M. R.

Tarouco. Fault management in computer networks using case-based reasoning: DUMBO system. Lecture Notes in Computer Science, 1650:510– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500510.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500510.
pdf.

Perner:1999:ACI

[1513] Petra Perner. An architecture for a CBR image segmentation system. Lecture Notes in Computer Science, 1650: 525–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500525.
pdf.

Praehofer:1999:SRS

[1514] Herbert Praehofer and Josef Kerschbaummayr. Supporting reusability in a system design environment by case-based reasoning techniques. Lecture Notes in Computer Science, 1650: 535–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500535.
                                        pdf.

Schmidt:1999:CBR

[1515] Rainer Schmidt, Bernhard Pollwein, and Lothar Gierl. Case-based reasoning for antibiotics therapy advice. Lecture Notes in Computer Science, 1650: 550–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500550.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500550.
                                        pdf.

Smyth:1999:SDW

[1516] Barry Smyth and Paul Cotter. Surfing the digital wave: Generating personalised TV listings using collaborative, case-based recommendation. Lecture Notes in Computer Science, 1650: 561–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1650/16500561.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1650/16500561.
                                        pdf.

Taki:1999:CBQ

[1517] Hirokazu Taki, Satoshi Hori, and Norihjiro Abe. Case-based quality management system using expectation values. Lecture Notes in Computer Science, 1650:572–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500572.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500572.
pdf.

Varma:1999:IDD

[1518] Anil Varma. ICARUS: Design and deployment of a case-based reasoning system for locomotive diagnostics. Lecture Notes in Computer Science, 1650:581– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1650/16500581.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1650/16500581.
pdf.

Papadimitriou:1999:TQ

[1519] Christos H. Papadimitriou. Topological queries. Lecture Notes in Computer Science, 1651:3–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510003.
pdf.

Sellis:1999:RIS

[1520] Timos Sellis. Research issues in spatiotemporal database systems. Lecture Notes in Computer Science, 1651:5– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510005.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510005.
                                        pdf.

Guenther:1999:LBW

[1521] O. Guenther. Looking both ways: SSD 1999 10. Lecture Notes in Computer Science, 1651:12–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunther:1999:LBW

[1522] Oliver Günther. Looking both ways: SSD 1999±10. Lecture Notes in Computer Science, 1651:12–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510012.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510012.
                                        pdf.

Stell:1999:GGU

[1523] John G. Stell and Michael F. Worboys. Generalizing graphs using amalgamation and selection. Lecture Notes in Computer Science, 1651:19– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510019.
                                        pdf.

DeFloriani:1999:DSS

[1524] L. De Floriani, P. Magillo, and E. Puppo. Data structures for simplicial multi-complexes. Lecture Notes in Computer Science, 1651:33–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Floriani:1999:DSS

[1525] Leila De Floriani, Paola Magillo, and Enrico Puppo. Data structures for simplicial multi-complexes. Lecture Notes in Computer Science, 1651:33– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510033.
pdf.

Hoerhammer:1999:SIS

[1526] M. Hoerhammer and M. Freeston.

Spatial indexing with a scale dimension. Lecture Notes in Computer Science, 1651:52–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horhammer:1999:SIS

[1527] Mike Hörhammer and Michael Freeston. Spatial indexing with a scale dimension. Lecture Notes in Computer Science, 1651:52–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510052.
pdf.

Boehm:1999:XOS

[1528] C. Boehm, G. Klump, and H.-P. Kriegel. XZ-ordering: A space-filling curve for objects with spatial extension. Lecture Notes in Computer Science, 1651:75–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bohm:1999:XOS

[1529] Christian Böhm, Gerald Klump, and Hans-Peter Kriegel. XZ-Ordering: A space-filling curve for objects with spatial extension. Lecture Notes in Computer Science, 1651:75–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510075.
                                        pdf.

Choubey:1999:GGT

[1530] Rupesh Choubey, Li Chen, and Elke A. Rundensteiner. GBI: A generalized Rtree bulk-insertion strategy. Lecture Notes in Computer Science, 1651:91– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510091.
                                        pdf.

Pfoser:1999:CUM

[1531] Dieter Pfoser and Christian S. Jensen. Capturing the uncertainty of movingobject representations. Lecture Notes in Computer Science, 1651:111–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510111.
pdf.

Shek:1999:DSC

[1532] Eddie C. Shek, Giovanni Giuffrida, Suhas Joshi, and Son K. Dao. Dynamic spatial clustering for intelligent mobile information sharing and dissemination. Lecture Notes in Computer Science, 1651:132–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510132.
pdf.

Theodoridis:1999:GSD

[1533] Yannis Theodoridis, Jefferson R. O.

Silva, and Mario A. Nascimento. On the generation of spatiotemporal datasets. Lecture Notes in Computer Science, 1651:147–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510147.
pdf.

Zhou:1999:EPA

[1534] Xiaofang Zhou, David Truffet, and Jiawei Han. Efficient polygon amalgamation methods for spatial OLAP and spatial data mining. Lecture Notes in Computer Science, 1651:167– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510167.
                                        pdf.

Lin:1999:EMP

[1535] Xuemin Lin, Xiaomei Zhou, and Chengfei Liu. Efficiently matching proximity relationships in spatial databases. Lecture Notes in Computer Science, 1651:188–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510188.
                                        pdf.

Xuemin:1999:EMP

[1536] L. Xuemin, Z. Xiaomei, and L. Chengfei. Efficiently matching proximity relationships in spatial databases. Lecture Notes in Computer Science, 1651:188– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ankerst:1999:SHS

[1537] Mihael Ankerst, Gabi Kastenmüller, Hans-Peter Kriegel, and Thomas Seidl. 3D shape histograms for similarity search and classification in spatial databases. Lecture Notes in Computer Science, 1651:207–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510207.
pdf.

Park:1999:MWS

[1538] Ho-Hyun Park, Guang-Ho Cha, and Chin-Wan Chung. Multi-way spatial joins using R-trees: Methodology and performance evaluation. Lecture Notes in Computer Science, 1651:229– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510229.
pdf.

Corral:1999:AJT

[1539] Antonio Corral, Michael Vassilakopoulos, and Yannis Manolopoulos. Algorithms for joining R-trees and linear region quadtrees. Lecture Notes in Computer Science, 1651:251–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510251.
pdf.

Becker:1999:APP

[1540] Ludger Becker, André Giesen, Klaus H.

Hinrichs, and Jan Vahrenhold. Algorithms for performing polygonal map overlay and spatial join on massive data sets. Lecture Notes in Computer Science, 1651:270–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510270.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510270.
                                        pdf.

Papadopoulos:1999:PES

[1541] Apostolos Papadopoulos, Philippe Rigaux, and Michel Scholl. A performance evaluation of spatial join processing strategies. Lecture Notes in Computer Science, 1651:286–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510286.
                                        pdf.

Voisard:1999:ADG

[1542] Agnès Voisard. Abduction and deduction in geologic hypermaps. Lecture Notes in Computer Science, 1651:311– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1651/16510311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510311.
                                        pdf.

Schneider:1999:UMS

[1543] Markus Schneider. Uncertainty management for spatial data in databases: Fuzzy spatial data types. Lecture Notes in Computer Science, 1651:330– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510330.
pdf.

Ravada:1999:OSE

[1544] Siva Ravada and Jayant Sharma. Oracle8i spatial: Experiences with extensible databases. Lecture Notes in Computer Science, 1651:355–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510355.
pdf.

Talamo:1999:SES

[1545] Maurizio Talamo, Franco Arcieri, Giancarlo Conia, and Enrico Nardelli. SICC: An exchange system for cadastral information. Lecture Notes in Computer Science, 1651:360–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1651/16510360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1651/16510360.
pdf.

Brinkhoff:1999:RTT

[1546] Thomas Brinkhoff. Requirements of traffic telematics to spatial databases. Lecture Notes in Computer Science, 1651:365–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1651/16510365.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1651/16510365.
                                        pdf.

Truszkowski:1999:ATN

[1547] W. Truszkowski, H. Hallock, and J. Kurien. Agent technology from a NASA perspective. Lecture Notes in Computer Science, 1652:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishida:1999:DCK

[1548] T. Ishida, J.-I. Akahani, K. Hiramatsu, and K. Isbister. Digital city kyoto: Towards a social information infrastructure. Lecture Notes in Computer Science, 1652:34–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelenbe:1999:ASI

[1549] E. Gelenbe. Autonomous search for information in an unknown environment. Lecture Notes in Computer Science, 1652:47–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodsky:1999:RMA

[1550] A. Brodsky, L. Kerschberg, and S. Varas. Resource management in agent-based distributed environments. Lecture Notes in Computer Science, 1652:61–??, 1999. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1999:MAA

[1551] C. M. Jonker, R. A. Lam, and J. Treur.

A multi-agent architecture for an intelligent Website in insurance. Lecture Notes in Computer Science, 1652:86– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saeyor:1999:FCB

[1552] S. Saeyor and M. Ishizuka. Formation of cooperative behavior among information agents in Web repository change monitoring service. Lecture Notes in Computer Science, 1652:101– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staab:1999:GSW

[1553] S. Staab, C. Braun, I. Bruder, and A. Duesterhoeft. GETESS — searching the Web exploiting German texts. Lecture Notes in Computer Science, 1652: 113–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ORiordan:1999:ABS

[1554] C. O’Riordan and H. Sorensen. An agent-based system for intelligent collaborative filtering. Lecture Notes in Computer Science, 1652:125–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomaa:1999:IAC

[1555] H. Gomaa. Inter-agent communication in cooperative information agent-based systems. Lecture Notes in Computer Science, 1652:137–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sullivan:1999:IRC

[1556] D. G. Sullivan, A. Glass, B. J. Grosz, and S. Kraus. Intention reconciliation in the context of teamwork: An initial empirical investigation. Lecture Notes in Computer Science, 1652:149– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puuronen:1999:SET

[1557] S. Puuronen and V. Terziyan. A similarity evaluation technique for cooperative problem solving with a group of agents. Lecture Notes in Computer Science, 1652:163–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deen:1999:CMC

[1558] S. M. Deen. A computational model for a cooperating agent system. Lecture Notes in Computer Science, 1652: 175–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanachi:1999:MAB

[1559] C. Hanachi, N. Hameurlain, and C. Sibertin-Blanc. Mobile agents behaviours: From declarative specifications to implementation. Lecture Notes in Computer Science, 1652:196–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Theilmann:1999:MSS

[1560] W. Theilmann and K. Rothermel. Maintaining specialized search engines through mobile filter agents. Lecture Notes in Computer Science, 1652:208– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vieira:1999:EMA

[1561] W. Vieira and L. M. Camarinha-Matos.

Execution monitoring in adaptive mobile agents. Lecture Notes in Computer Science, 1652:220–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitamura:1999:MAM

[1562] Y. Kitamura, Y. Mawarimichi, and S. Tatsumi. Mobile-agent mediated place oriented communication. Lecture Notes in Computer Science, 1652:232– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wellman:1999:AEC

[1563] M. P. Wellman. Agents and electronic commerce: Mechanisms and protocols. Lecture Notes in Computer Science, 1652:243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papazoglou:1999:RAT

[1564] M. P. Papazoglou. The role of agent technology in business to business electronic commerce. Lecture Notes in Computer Science, 1652:245–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerschberg:1999:ABF

[1565] L. Kerschberg and S. Banerjee. An agency-based framework for electronic business. Lecture Notes in Computer Science, 1652:265–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:SAM

[1566] X. F. Wang, X. Yi, K. Y. Lam, and C. Q. Zhang. Secure agent-mediated auctionlike negotiation protocol for Internet retail commerce. Lecture Notes in Computer Science, 1652:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheth:1999:SIB

[1567] A. Sheth, V. Kashyap, and T. Lima. Semantic information brokering: How can a multi-agent approach help? Lecture Notes in Computer Science, 1652: 303–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tesch:1999:AMA

[1568] T. Tesch and P. Fankhauser. Arbitration and matchmaking for agents with conflicting interests. Lecture Notes in Computer Science, 1652:323–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langley:1999:ACI

[1569] P. Langley, C. Thompson, R. Elio, and A. Haddadi. An adaptive conversational interface for destination advice. Lecture Notes in Computer Science, 1652:347–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewis:1999:ADD

[1570] M. Lewis. Anticipation, delegation, and demonstration: Why talking to agents is hard. Lecture Notes in Computer Science, 1652:365–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanElst:1999:CCA

[1571] L. Van Elst and F. Schmalhofer. A cooperative comprehension assistant for intranet-based information environments. Lecture Notes in Computer Science, 1652:390–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexander:1999:AA

[1572] D. S. Alexander and J. M. Smith. The architecture of ALIEN. Lecture Notes in Computer Science, 1653:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denazis:1999:DIO

[1573] S. Denazis, K. Miki, J. Vicente, and A. Campbell. Designing interfaces for open programmable routers. Lecture Notes in Computer Science, 1653:13– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menage:1999:RRC

[1574] P. Menage. Rcane: A resource controlled framework for active network services. Lecture Notes in Computer Science, 1653:25–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sivakumar:1999:PPN

[1575] R. Sivakumar, N. Venkitaraman, and V. Bharghavan. The Protean Programmable Network Architecture: Design and initial experience. Lecture Notes in Computer Science, 1653:37– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rizzo:1999:DPF

[1576] M. Rizzo, B. Briscoe, J. Tassel, and K. Damianakis. A dynamic pricing framework to support a scalable, usage-based charging model for packetswitched networks. Lecture Notes in Computer Science, 1653:48–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marshall:1999:AIN

[1577] I. Marshall, M. Fry, L. Velasco, and A. Ghosh. Active information networks and XML. Lecture Notes in Computer Science, 1653:60–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloman:1999:PSP

[1578] M. Sloman and E. Lupu. Policy specification for programmable networks. Lecture Notes in Computer Science, 1653:73–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vanet:1999:SCD

[1579] G. Vanet and Y. Kiriha. A selfconfiguring data caching architecture based on active networking techniques. Lecture Notes in Computer Science, 1653:85–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delgrossi:1999:ICA

[1580] L. Delgrossi, G. Di Fatta, D. Ferrari, and G. Lo Re. Interference and communications among active network applications. Lecture Notes in Computer Science, 1653:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeumer:1999:GMA

[1581] C. Baeumer and T. Magedanz. The grasshopper mobile agent platform enabling short-term active broadband intelligent network implementation. Lecture Notes in Computer Science, 1653: 109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardoe:1999:LPS

[1582] R. Cardoe, J. Finney, A. C. Scott, and W. D. Shepherd. LARA: A prototype system for supporting high performance active networking. Lecture Notes in Computer Science, 1653:117– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1999:PIS

[1583] A. Cohen and S. Rangarajan. A programming interface for supporting IP traffic processing. Lecture Notes in Computer Science, 1653:132–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihai:1999:NGC

[1584] N. Mihai and G. Vanecek. New generation of control planes in emerging data networks. Lecture Notes in Computer Science, 1653:144–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tschudin:1999:ANO

[1585] C. Tschudin. An active networks overlay network (ANON). Lecture Notes in Computer Science, 1653:156–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busse:1999:ADA

[1586] I. Busse, S. Covaci, and A. Leichsenring. Autonomy and decentralization in active networks: A case study for mobile agents. Lecture Notes in Computer Science, 1653:165–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1999:TAH

[1587] D. C. Lee, M. T. Jones, S. F. Midkiff, and P. M. Athanas. Towards active hardware. Lecture Notes in Computer Science, 1653:180–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juhola:1999:IEN

[1588] A. Juhola, I. Marshall, S. Covaci, and T. Velte. The impact of AN on established network operators. Lecture Notes in Computer Science, 1653:188– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenwood:1999:UAP

[1589] D. P. A. Greenwood and D. Gavalas. Using active processes as the basis for an integrated distributed network management architecture. Lecture Notes in Computer Science, 1653:199–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norden:1999:AAF

[1590] S. Norden and K. F. Wong. AN-MAC: An architectural framework for network management and control using active networks. Lecture Notes in Computer Science, 1653:212–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raz:1999:ANA

[1591] D. Raz and Y. Shavitt. An active network approach to efficient network management. Lecture Notes in Computer Science, 1653:220–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Redlich:1999:VNC

[1592] J.-P. Redlich, M. Suzuki, and S. Weinstein. Virtual networks for customizable traffic treatments. Lecture Notes in Computer Science, 1653:232– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugauchi:1999:FNM

[1593] K. Sugauchi, S. Miyazaki, K. Yoshida, and K. Nakane. Flexible network management using active network framework. Lecture Notes in Computer Science, 1653:241–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campbell:1999:MSV

[1594] A. T. Campbell, J. Vicente, and D. A.

Villela. Managing spawned virtual networks. Lecture Notes in Computer Science, 1653:249–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willmott:1999:AOR

[1595] S. Willmott and B. Faltings. Active organisations for routing. Lecture Notes in Computer Science, 1653:262– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1999:DIC

[1596] J.-H. Yeh, R. Chow, and R. Newman. A dynamic interdomain communication path setup in active network. Lecture Notes in Computer Science, 1653:274–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhushan:1999:ANC

[1597] B. Bhushan and J. Hall. Active network challenges to TMN. Lecture Notes in Computer Science, 1653:285–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kulkarni:1999:SAN

[1598] A. Kulkarni, G. Minden, V. Frost, and J. Evans. Survivability of active networking services. Lecture Notes in Computer Science, 1653:299–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hjalmtysson:1999:CD

[1599] G. Hjalmtysson and S. Bhattacharjee. Control on demand. Lecture Notes in Computer Science, 1653:315–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karnouskos:1999:ABS

[1600] S. Karnouskos, I. Busse, and S. Covaci. Agent based security for the active network infrastructure. Lecture Notes in Computer Science, 1653:330–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siddiqi:1999:HAE

[1601] K. Siddiqi, A. Tannenbaum, and S. W.

Zucker. A Hamiltonian approach to the eikonal equation. Lecture Notes in Computer Science, 1654:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Worthington:1999:TSS

[1602] P. L. Worthington and E. R. Hancock. Topographic surface structure from 2D images using shape-from-shading. Lecture Notes in Computer Science, 1654:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1999:HSI

[1603] D. Zhang and M. Hebert. Harmonic shape images: A representation for 3D free-form surfaces based on energy minimization. Lecture Notes in Computer Science, 1654:30–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donatini:1999:DES

[1604] P. Donatini, P. Frosini, and C. Landi.

Deformation energy for size functions. Lecture Notes in Computer Science, 1654:44–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Figueiredo:1999:FMM

[1605] M. A. T. Figueiredo, J. M. N. Leitao, and A. K. Jain. On fitting mixture models. Lecture Notes in Computer Science, 1654:54–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cazorla:1999:BMF

[1606] M. A. Cazorla, F. Escolano, D. Gallardo, and R. Rizo. Bayesian models for finding and grouping junctions. Lecture Notes in Computer Science, 1654:70–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chardin:1999:SII

[1607] A. Chardin and P. Perez. Semiiterative inferences with hierarchical energy-based models for image analysis. Lecture Notes in Computer Science, 1654:83–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Descombes:1999:MVK

[1608] X. Descombes and E. Pechersky. Metropolis vs Kawasaki dynamic for image segmentation based on Gibbs models. Lecture Notes in Computer Science, 1654:99–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jalobeanu:1999:HES

[1609] A. Jalobeanu, L. Blanc-Feraud, and J. Zerubia. Hyperparameter estimation for satellite image restoration by a MCMCML method. Lecture Notes in Computer Science, 1654:115–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morris:1999:AVM

[1610] R. D. Morris. Auxiliary variables for Markov random fields with higher order interactions. Lecture Notes in Computer Science, 1654:131–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rostaing:1999:UMI

[1611] P. Rostaing, J.-N. Provost, and C. Collet. Unsupervised multispectral image segmentation using generalized Gaussian noise model. Lecture Notes in Computer Science, 1654:142–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dias:1999:ABC

[1612] J. M. B. Dias. Adaptive Bayesian contour estimation: A vector space representation approach. Lecture Notes in Computer Science, 1654:157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubota:1999:APB

[1613] T. Kubota. Adaptive pixel-based data fusion for boundary detection. Lecture Notes in Computer Science, 1654:173– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coughlan:1999:BTS

[1614] J. M. Coughlan and A. L. Yuille.

Bayesian tree search with expected O(N) convergence rates for road tracking. Lecture Notes in Computer Science, 1654:189–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boykov:1999:NAE

[1615] Y. Boykov, O. Veksler, and R. Zabih. A new algorithm for energy minimization with discontinuities. Lecture Notes in Computer Science, 1654:205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cross:1999:CHC

[1616] A. D. J. Cross and E. R. Hancock. Convergence of a hill climbing genetic algorithm for graph matching. Lecture Notes in Computer Science, 1654:221– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rangarajan:1999:NDM

[1617] A. Rangarajan, H. Chui, and E. Mjolsness. A new distance measure for nonrigid image matching. Lecture Notes in Computer Science, 1654:237–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torsello:1999:CTR

[1618] A. Torsello and M. Pelillo. Continuoustime relaxation labeling processes. Lecture Notes in Computer Science, 1654: 253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yin:1999:RAU

[1619] L. Yin and A. Basu. Realistic animation using extended adaptive mesh for model based coding. Lecture Notes in Computer Science, 1654:269–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguiar:1999:MLI

[1620] P. M. Q. Aguiar and J. M. F. Moura. Maximum likelihood inference of 3D structure from image sequences. Lecture Notes in Computer Science, 1654: 285–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oakley:1999:MRI

[1621] J. Oakley, J. Missimer, and G. Szekely.

Magnetic resonance imaging based correction and reconstruction of positron emission tomography images. Lecture Notes in Computer Science, 1654:301– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Svensen:1999:MRF

[1622] M. Svensen, F. Kruggel, and D. Y.

Von Cramon. Markov random field modelling of fMRI data using a mean field EM-algorithm. Lecture Notes in Computer Science, 1654:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aczel:1999:RTT

[1623] Peter Aczel. On relating type theories and set theories. Lecture Notes in Computer Science, 1657:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570001.
pdf.

Ahn:1999:CMC

[1624] René Ahn and Tijn Borghuis. Communication modelling and contextdependent interpretation: An integrated approach. Lecture Notes in Computer Science, 1657:19–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570019.
pdf.

Coquand:1999:GBT

[1625] Thierry Coquand and Henrik Persson. Gröbner bases in type theory. Lecture Notes in Computer Science, 1657:33– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570033.
pdf.

Despeyroux:1999:MLC

[1626] Joëlle Despeyroux and Pierre Leleu. A modal lambda calculus with iteration and case constructs. Lecture Notes in Computer Science, 1657:47– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570047.
pdf.

Dowek:1999:PNM

[1627] Gilles Dowek and Benjamin Werner. Proof normalization modulo. Lecture Notes in Computer Science, 1657:62– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1657/16570062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570062.
pdf.

Filliatre:1999:PIP

[1628] Jean-Christophe Filliâtre. Proof of imperative programs in type theory. Lecture Notes in Computer Science, 1657: 78–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570078.
pdf.

Fridlender:1999:IFT

[1629] Daniel Fridlender. An interpretation of the fan theorem in type theory. Lecture Notes in Computer Science, 1657: 93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570093.
pdf.

Goubault-Larrecq:1999:CTS

[1630] Jean Goubault-Larrecq. Conjunctive types and SKInT. Lecture Notes in Computer Science, 1657:106–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570106.
pdf.

Kammuller:1999:MSD

[1631] Florian Kammüller. Modular structures as dependent types in isabelle. Lecture Notes in Computer Science, 1657:121–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1657/16570121.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1657/16570121.
                                        pdf.

Kleymann:1999:MVC

[1632] Thomas Kleymann. Metatheory of verification calculi in LEGO. Lecture Notes in Computer Science, 1657:133– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1657/16570133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1657/16570133.
                                        pdf.

Liquori:1999:BPE

[1633] Luigi Liquori. Bounded polymorphism for extensible objects. Lecture Notes in Computer Science, 1657:149– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1657/16570149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1657/16570149.
                                        pdf.

Maietti:1999:AEQ

[1634] Maria Emilia Maietti. About effective quotients in constructive type theory.

Lecture Notes in Computer Science, 1657:164–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1657/16570164.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1657/16570164.
pdf.

Pfenning:1999:AEU

[1635] Frank Pfenning and Carsten Schürmann.

Algorithms for equality and unification in the presence of notational definitions. Lecture Notes in Computer Science, 1657:179–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570179.
pdf.

Sambin:1999:PBP

[1636] Giovanni Sambin and Silvia Gebellato.

A preview of the basic picture: A new perspective on formal topology. Lecture Notes in Computer Science, 1657: 194–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1657/16570194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1657/16570194.
pdf.

Wehner:1999:LSN

[1637] Rüdiger Wehner. Large-scale navigation: The insect case. Lecture Notes in Computer Science, 1661:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1661/16610001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1661/16610001.
                                        pdf.

Roefer:1999:RNU

[1638] T. Roefer. Route navigation using motion analysis. Lecture Notes in Computer Science, 1661:21–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rofer:1999:RNU

[1639] Thomas Röfer. Route navigation using motion analysis. Lecture Notes in Computer Science, 1661:21–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1661/16610021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1661/16610021.
                                        pdf.

Sorrows:1999:NLR

[1640] Molly E. Sorrows and Stephen C. Hirtle. The nature of landmarks for real and electronic spaces. Lecture Notes in Computer Science, 1661:37– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1661/16610037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1661/16610037.
                                        pdf.

Tversky:1999:PVT

[1641] Barbara Tversky and Paul U. Lee. Pictorial and verbal tools for conveying routes. Lecture Notes in Computer Science, 1661:51–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1661/16610051.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1661/16610051.pdf.

Lovelace:1999:EGR

[1642] Kristin L. Lovelace, Mary Hegarty, and Daniel R. Montello. Elements of good route directions in familiar and unfamiliar environments. Lecture Notes in Computer Science, 1661:65– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610065.
pdf.

Fontaine:1999:PRI

[1643] Sylvie Fontaine and Michel Denis. The production of route instructions in underground and urban environments. Lecture Notes in Computer Science, 1661:83–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610083.
pdf.

Frank:1999:OSA

[1644] Andrew U. Frank. One step up the abstraction ladder: Combining algebras — from functional pieces to a whole. Lecture Notes in Computer Science, 1661:95–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610095.
pdf.

Remolina:1999:FRS

[1645] Emilio Remolina, Juan A. Fernandez, Benjamin Kuipers, and Javier Gonzalez. Formalizing regions in the spatial semantic hierarchy: An AHgraphs implementation approach. Lecture Notes in Computer Science, 1661: 109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610109.
pdf.

Timpf:1999:ALD

[1646] Sabine Timpf. Abstraction, levels of detail, and hierarchies in map series. Lecture Notes in Computer Science, 1661: 125–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610125.
pdf.

Winter:1999:TRH

[1647] Stephan Winter. Topological relations in hierarchical partitions. Lecture Notes in Computer Science, 1661: 141–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610141.
pdf.

Eschenbach:1999:PCQ

[1648] Carola Eschenbach. A predication calculus for qualitative spatial representations. Lecture Notes in Computer Science, 1661:157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610157.
pdf.

Ligozat:1999:SMS

[1649] Gérard Ligozat. Simple models for simple calculi. Lecture Notes in Computer Science, 1661:173–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610173.
pdf.

Moeller:1999:TDR

[1650] R. Moeller and M. Wessel. Terminological default reasoning about spatial information: A first step. Lecture Notes in Computer Science, 1661:189– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:1999:TDR

[1651] Ralf Möller and Michael Wessel. Terminological default reasoning about spatial information: A first step. Lecture Notes in Computer Science, 1661:189– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610189.
pdf.

Kulik:1999:RAC

[1652] Lars Kulik and Alexander Klippel. Reasoning about cardinal directions using grids as qualitative geographic coordinates. Lecture Notes in Computer Science, 1661:205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610205.
pdf.

Guarino:1999:RIC

[1653] Nicola Guarino. The role of identity conditions in ontology design. Lecture Notes in Computer Science, 1661:221– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1661/16610221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610221.
pdf.

Masolo:1999:AVI

[1654] Claudio Masolo and Laure Vieu. Atomicity vs. infinite divisibility of space. Lecture Notes in Computer Science, 1661:235–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1661/16610235.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1661/16610235.
pdf.

Galton:1999:MDS

[1655] Antony Galton. The mereotopology of discrete space. Lecture Notes in Computer Science, 1661:251–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610251.
pdf.

Smith:1999:A

[1656] Barry Smith. Agglomerations. Lecture Notes in Computer Science, 1661: 267–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610267.
pdf.

Mark:1999:OGO

[1657] David M. Mark, Barry Smith, and Barbara Tversky. Ontology and geographic objects: An empirical study of cognitive categorization. Lecture Notes in Computer Science, 1661:283– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610283.
pdf.

Cohn:1999:MC

[1658] A. G. Cohn and A. C. Varzi. Modes of connection. Lecture Notes in Computer Science, 1661:299–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610299.
pdf.

Gero:1999:RRA

[1659] John S. Gero. Representation and reasoning about shapes: Cognitive and computational studies in visual reasoning in design. Lecture Notes in Computer Science, 1661:315–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610315.
pdf.

Kuhn:1999:AIS

[1660] Werner Kuhn. An algebraic interpretation of semantic networks. Lecture Notes in Computer Science, 1661:331– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610331.
pdf.

Andrienko:1999:DCS

[1661] Gennady Andrienko and Natalia Andrienko. Data characterization schema for intelligent support in visual data analysis. Lecture Notes in Computer Science, 1661:349–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610349.
pdf.

Mallot:1999:RTR

[1662] Hanspeter A. Mallot, Sabine Gillner, Sibylle D. Steck, and Matthias O. Franz. Recognition-triggered response and the view-graph approach to spatial cognition. Lecture Notes in Computer Science, 1661:367–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610367.
pdf.

Raubal:1999:FMP

[1663] Martin Raubal and Michael Worboys. A formal model of the process of wayfinding in built environments. Lecture Notes in Computer Science, 1661: 381–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1661/16610381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1661/16610381.
                                        pdf.

Kettani:1999:SMB

[1664] Driss Kettani and Bernard Moulin. A spatial model based on the notions of spatial conceptual map and of object’s influence areas. Lecture Notes in Computer Science, 1661:401– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1661/16610401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1661/16610401.
                                        pdf.

Stell:1999:GG

[1665] John G. Stell. Granulation for graphs. Lecture Notes in Computer Science, 1661:417–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1661/16610417.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1661/16610417.
                                        pdf.

Bittner:1999:OER

[1666] Thomas Bittner. On ontology and epistemology of rough location. Lecture Notes in Computer Science, 1661:433– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610433.
pdf.

Ellis:1999:QSR

[1667] Christopher D. Ellis and Douglas M.

Johnston. Qualitative spatial representation for situational awareness and spatial decision support. Lecture Notes in Computer Science, 1661:449– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610449.
pdf.

Musto:1999:QMR

[1668] Alexandra Musto, Klaus Stein, Kerstin Schill, Andreas Eisenkolb, and Wilfried Brauer. Qualitative motion representation in egocentric and allocentric frames of reference. Lecture Notes in Computer Science, 1661:461– 476, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1661/16610461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1661/16610461.
pdf.

Aversa:1999:AMP

[1669] R. Aversa, N. Mazzocca, and U. Villano. Analytical modeling of parallel application in heterogeneous computing environments: A study of Cholesky factorization. Lecture Notes in Computer Science, 1662:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacci:1999:STI

[1670] B. Bacci, S. Gorlatch, C. Lengauer, and S. Pelagatti. Skeletons and transformations in an integrated parallel programming environment. Lecture Notes in Computer Science, 1662:13–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chi:1999:SUA

[1671] C.-H. Chi and J.-L. Yuan. Sequential unification and aggressive lookahead mechanisms for data memory accesses. Lecture Notes in Computer Science, 1662:28–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doroshenko:1999:CMF

[1672] A. E. Doroshenko, I. S. Kononenko, and A. Y. Korotun. A coordination model and facilities for efficient parallel computation. Lecture Notes in Computer Science, 1662:42–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godlevsky:1999:PSP

[1673] A. Godlevsky, M. Gazak, and L. Hluchy. Parallelizing of sequential programs on the basis of pipeline and speculative features of the operators. Lecture Notes in Computer Science, 1662:48–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorbunova:1999:KMP

[1674] K. O. Gorbunova. Kinetic model of parallel data processing. Lecture Notes in Computer Science, 1662:54–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1999:PAP

[1675] P. Hartmann. PSA approach to population models for parallel genetic algorithms. Lecture Notes in Computer Science, 1662:60–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konshin:1999:HAN

[1676] V. N. Konshin and V. A. Garanzha.

Highly accurate numerical methods for incompressible 3D fluid flows on parallel architectures. Lecture Notes in Computer Science, 1662:68–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maric:1999:DTS

[1677] S. Maric and Z. Jovanovic. Dynamic task scheduling with precedence constraints and communication delays. Lecture Notes in Computer Science, 1662:77–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Likhoded:1999:TDS

[1678] N. A. Likhoded. Two-dimensional scheduling of algorithms with uniform dependencies. Lecture Notes in Computer Science, 1662:92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mostefaoui:1999:CLC

[1679] A. Mostefaoui, M. Raynal, and M. Takizawa. Consistent lamport clocks for asynchronous groups with process crashes. Lecture Notes in Computer Science, 1662:98–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudov:1999:CAL

[1680] S. Pudov. Comparative analysis of learning methods of cellular-neural associative memory. Lecture Notes in Computer Science, 1662:108–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Selikhov:1999:EPR

[1681] A. V. Selikhov. Emergence and propagation of round autowave in cellular neural network. Lecture Notes in Computer Science, 1662:120–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Virbitskaite:1999:PBA

[1682] I. B. Virbitskaite and E. A. Pokozy. Parametric behaviour analysis for time Petri nets. Lecture Notes in Computer Science, 1662:134–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wanek:1999:BAA

[1683] H. Wanek and E. Schikuta. A blackboard approach for the automatic optimization of parallel I/O operations. Lecture Notes in Computer Science, 1662:141–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1999:RES

[1684] C.-H. Yeh, E. A. Varvarigos, and H. Lee. Routing and embeddings in super Cayley graphs. Lecture Notes in Computer Science, 1662:151–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandini:1999:ICA

[1685] S. Bandini, G. Erbacci, and G. Mauri.

Implementing cellular automata based models on parallel architectures: The CAPP project. Lecture Notes in Computer Science, 1662:167–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bessonov:1999:MAP

[1686] O. Bessonov, D. Fougere, K. D. Quoc, and B. Roux. Methods for achieving peak computational rates for linear algebra operations on superscalar RISC processors. Lecture Notes in Computer Science, 1662:180–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birov:1999:PML

[1687] L. Birov, A. Prokofiev, Y. Bartenev, and A. Vargin. The parallel mathematical libraries project (PMLP): Overview, design innovations, and preliminary results. Lecture Notes in Computer Science, 1662:186–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bystrov:1999:IMC

[1688] A. V. Bystrov and I. B. Virbitskaite. Implementing model checking and equivalence checking for time Petri nets by the RT-MEC tool. Lecture Notes in Computer Science, 1662:194– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capretti:1999:LCP

[1689] G. Capretti, M. R. Lagana, and L. Ricci. Learning concurrent programming: A constructionist approach. Lecture Notes in Computer Science, 1662: 200–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalton:1999:SPA

[1690] D. Dalton. The speedup performance of an associative memory based logic simulator. Lecture Notes in Computer Science, 1662:207–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giloi:1999:HLP

[1691] W. K. Giloi, H. W. Pohl, and A. Schramm. A high-level programming environment for distributed memory architectures. Lecture Notes in Computer Science, 1662:217–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konovalov:1999:VSF

[1692] A. Konovalov, V. Samofalov, and S. Scharf. Virtual shared files: Towards user-friendly inter-process communications. Lecture Notes in Computer Science, 1662:223–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krajecki:1999:OOE

[1693] M. Krajecki. An object oriented environment to manage the parallelism of the FIIT applications. Lecture Notes in Computer Science, 1662:229–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1999:PSS

[1694] J. Li, J. Wang, and H. Kameda. Performance studies of shared-nothing parallel transaction processing systems. Lecture Notes in Computer Science, 1662: 235–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1999:STE

[1695] T. Ludwig, J. Trinitis, and R. Wismueller. Synergetic tool environments. Lecture Notes in Computer Science, 1662:248–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mostefaoui:1999:LIC

[1696] A. Mostefaoui, M. Raynal, and P. Verissimo. Logically instantaneous communication on top of distributed memory parallel machines. Lecture Notes in Computer Science, 1662:258– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedretti:1999:TCA

[1697] K. T. Pedretti, T. L. Casavant, R. C.

Braun, and T. E. Scheetz. Three complementary approaches to parallelization of local BLAST service on workstation clusters. Lecture Notes in Computer Science, 1662:271–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peter:1999:ILS

[1698] Y. Peter and H. Guyennet. An implementation of the lifecycle service object mobility on CORBA. Lecture Notes in Computer Science, 1662:283–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serot:1999:SSB

[1699] J. Serot, D. Ginhac, and J.-P Derutin. SKIPPER: A skeleton-based parallel programming environment for realtime image processing applications. Lecture Notes in Computer Science, 1662:296–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlassov:1999:QMM

[1700] V. Vlassov and A. Kraynikov. A queuing model of a multi-threaded architecture: A case study. Lecture Notes in Computer Science, 1662:306–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1999:BPA

[1701] W. Zheng, S. Khan, and H. Xie. BSP performance analysis and prediction: Tools and application. Lecture Notes in Computer Science, 1662:313–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sancese:1999:MPV

[1702] S. Sancese, P. Ciancarini, and A. Messina. Message passing vs tuple space coordination in an aerodynamics application. Lecture Notes in Computer Science, 1662:320–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achasova:1999:TED

[1703] S. M. Achasova. Two examples of distributed architecture for solving combinatorial optimization problems. Lecture Notes in Computer Science, 1662: 335–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cappello:1999:PNB

[1704] F. Cappello, O. Richard, and D. Etiemble. Performance of the NAS benchmarks on a cluster of SMP PCs using a parallelization of the MPI programs with OpenMP. Lecture Notes in Computer Science, 1662:339–350, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorojevets:1999:CAP

[1705] M. Dorojevets. COOL approach to petaflops computing. Lecture Notes in Computer Science, 1662:351–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goossens:1999:HSO

[1706] B. Goossens, H. Essafi, and M. Pic.

Hardware and software optimizations for multimedia databases. Lecture Notes in Computer Science, 1662:365– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markova:1999:CRA

[1707] V. Markova. Cellular recursive algorithm architecture for long integers multiplication in arrays of restricted size. Lecture Notes in Computer Science, 1662:371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandini:1999:PMB

[1708] S. Bandini, G. Mauri, G. Pavesi, and C. Simone. A parallel model based on cellular automata for the simulation of pesticide percolation in the soil. Lecture Notes in Computer Science, 1662:383–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandman:1999:CSC

[1709] O. L. Bandman. Comparative study of cellular-automata diffusion models. Lecture Notes in Computer Science, 1662:395–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolshakov:1999:CRM

[1710] K. Bolshakov, A. Borshchev, A. Filippoff, and Y. Karpov. Creating and running mobile agents with XJ DOME. Lecture Notes in Computer Science, 1662:410–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haefner:1999:PIA

[1711] H. Haefner, W. Schoenauer, and R. Weiss. Parallelization and integration of the LU and ILU algorithm in the LINSOL program package. Lecture Notes in Computer Science, 1662:417–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hochberger:1999:CDM

[1712] C. Hochberger, R. Hoffman, and S. Waldschmidt. CDL++ for the description of moving objects in cellular automata. Lecture Notes in Computer Science, 1662:428–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaporin:1999:PSL

[1713] I. E. Kaporin and I. N. Konshin. Parallel solution of large sparse SPD linear systems based on overlapping domain decomposition. Lecture Notes in Computer Science, 1662:436–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1999:RPP

[1714] Y.-C. Kim and Y.-K. Jun. Restructuring parallel programs for on-the-fly race detection. Lecture Notes in Computer Science, 1662:446–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petcu:1999:SIV

[1715] D. Petcu. Solving initial value problems with a multiprocessor code. Lecture Notes in Computer Science, 1662: 452–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz-Andino:1999:PIC

[1716] A. Ruiz-Andino, L. Araujo, F. Saenz, and J. Ruz. Parallel implementation of constraint solving. Lecture Notes in Computer Science, 1662:466–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aversa:1999:EPD

[1717] R. Aversa and B. Di Martino. Experiences on parallelization of divide and conquer algorithms with parallel paradigms. Lecture Notes in Computer Science, 1662:473–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerin:1999:DMP

[1718] C. Cerin. Differentiating message passing interface and bulk synchronous parallel computation models. Lecture Notes in Computer Science, 1662:477– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farid:1999:RCA

[1719] T. Farid and D. Zerbino. Realization of complex arithmetic on cellular automata. Lecture Notes in Computer Science, 1662:479–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katkov:1999:PPS

[1720] S. I. Katkov and E. Y. Ruban. Parallel programming system based on Super Pascal language. Lecture Notes in Computer Science, 1662:481–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klimova:1999:GTM

[1721] O. V. Klimova. Group-theoretic methods for parallel computation of convolution. Lecture Notes in Computer Science, 1662:483–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCollum:1999:KEI

[1722] B. McCollum, P. Milligan, and P. H.

Corr. Knowledge exploitation for improved data distribution. Lecture Notes in Computer Science, 1662:485–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McMullan:1999:ESA

[1723] P. J. P. McMullan, P. Milligan, and P. H. Corr. An expert system approach to data distribution and distribution analysis. Lecture Notes in Computer Science, 1662:487–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1999:ACA

[1724] P. J. Parsons. Automated communication analysis and code generation for non-contiguous partitions of twodimensional arrays. Lecture Notes in Computer Science, 1662:489–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beletkov:1999:WST

[1725] D. Beletkov, M. Ostapkevich, S. Piskunov, and I. Zhileev. WinALT, a software tool for fine-grain algorithms and structures synthesis and simulation. Lecture Notes in Computer Science, 1662:491– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolshakov:1999:DTS

[1726] K. Bolshakov, Y. Karpov, and A. Sintotski. DEALed — A tool suite for distributed real-time systems development. Lecture Notes in Computer Science, 1662:497–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1999:PPT

[1727] C. Lang, M. Trehel, and P. Baptiste. PLATINUM: A placement tool based on process initiative. Lecture Notes in Computer Science, 1662:503–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrzejak:1999:OSP

[1728] A. Andrzejak and K. Fukuda. Optimization over k-set polytopes and efficient k-set enumeration. Lecture Notes in Computer Science, 1663:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neyer:1999:LSR

[1729] G. Neyer. Line simplification with restricted orientations. Lecture Notes in Computer Science, 1663:13–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1999:JPS

[1730] P. Berman, A. B. Kahng, D. Vidhani, and A. Zelikovsky. The T -join problem in sparse graphs: Applications to phase assignment problem in VLSI mask layout. Lecture Notes in Computer Science, 1663:25–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodnik:1999:RAO

[1731] A. Brodnik, S. Carlsson, E. D. Demaine, and J. I. Munro. Resizable arrays in optimal time and space. Lecture Notes in Computer Science, 1663: 37–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagh:1999:HDE

[1732] R. Pagh. Hash and displace: Efficient evaluation of minimal perfect hash functions. Lecture Notes in Computer Science, 1663:49–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leiserson:1999:ITD

[1733] C. E. Leiserson. Invited talk: Design and analysis of algorithms for shared-memory multiprocessors. Lecture Notes in Computer Science, 1663: 55–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patrignani:1999:COC

[1734] M. Patrignani. On the complexity of orthogonal compaction. Lecture Notes in Computer Science, 1663:56– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barequet:1999:OCO

[1735] G. Barequet, P. Bose, and M. T. Dickerson. Optimizing constrained offset and scaled polygonal annuli. Lecture Notes in Computer Science, 1663:62– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyar:1999:AFG

[1736] J. Boyar, K. S. Larsen, and M. N.

Nielsen. The accommodating function — A generalization of the competitive ratio. Lecture Notes in Computer Science, 1663:74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bender:1999:PGT

[1737] M. A. Bender and C. Chekuri. Performance guarantees for the TSP with a parameterized triangle inequality. Lecture Notes in Computer Science, 1663: 80–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:1999:RMV

[1738] X. Deng, E. Milios, and A. Mirzaian. Robot map verification of a graph world. Lecture Notes in Computer Science, 1663:86–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broecker:1999:SRS

[1739] C. A. Broecker and S. Schuierer. Searching rectilinear streets completely. Lecture Notes in Computer Science, 1663:98–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1999:GMT

[1740] K. Jansen and L. Porkolab. General multiprocessor task scheduling: Approximate solutions in linear time. Lecture Notes in Computer Science, 1663: 110–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arkin:1999:LBS

[1741] E. M. Arkin, M. A. Bender, J. S. B. Mitchell, and S. S. Skiena. The lazy bureaucrat scheduling problem. Lecture Notes in Computer Science, 1663: 122–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magnenat-Thalmann:1999:ITG

[1742] N. Magnenat-Thalmann, W. Lee, P. Beylot, and D. Sankoff. Invited talk: Generating 3D virtual populations from pictures of a few individuals. Lecture Notes in Computer Science, 1663:134–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1999:TQM

[1743] P. Bose and P. Morin. Testing the quality of manufactured balls. Lecture Notes in Computer Science, 1663:145– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosaraju:1999:OST

[1744] S. R. Kosaraju, T. M. Przytycka, and R. Borgstrom. On an optimal split tree problems. Lecture Notes in Computer Science, 1663:157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benoit:1999:RTH

[1745] D. Benoit, E. D. Demaine, J. I. Munro, and V. Raman. Representing trees of higher degree. Lecture Notes in Computer Science, 1663:169–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1999:IDM

[1746] A. Amir, D. Keselman, G. M. Landau, and M. Lewenstein. Indexing and dictionary matching with one error. Lecture Notes in Computer Science, 1663:181–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukovszki:1999:NRF

[1747] T. Lukovszki. New results on fault tolerant geometric spanners. Lecture Notes in Computer Science, 1663:193– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goodrich:1999:TVE

[1748] M. T. Goodrich and J. G. Kloss. Tiered vectors: Efficient dynamic arrays for rank-based sequences. Lecture Notes in Computer Science, 1663:205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazirani:1999:ITG

[1749] U. V. Vazirani. Invited talk: Go-with-the-winner heuristic. Lecture Notes in Computer Science, 1663:217–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barequet:1999:PSV

[1750] G. Barequet, M. T. Dickerson, and R. L. S. Drysdale. 2-Point site Voronoi diagrams. Lecture Notes in Computer Science, 1663:219–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chin:1999:PAF

[1751] F. Chin, D. T. Lee, and C. A. Wang. A parallel algorithm for finding the constrained Voronoi diagram of line segments in the plane. Lecture Notes in Computer Science, 1663:231–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broecker:1999:PIS

[1752] C. A. Broecker and A. Lopez-Ortiz.

Position-independent street searching. Lecture Notes in Computer Science, 1663:241–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chakraborty:1999:AAD

[1753] S. Chakraborty and S. Biswas. Approximation algorithms for 3-D common substructure identification in drug and protein molecules. Lecture Notes in Computer Science, 1663:253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:TBB

[1754] C. A. Wang and B. Yang. A tight bound for beta-skeleton of minimum weight triangulations. Lecture Notes in Computer Science, 1663:265–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bespamyatnikh:1999:RSD

[1755] S. Bespamyatnikh and M. Segal. Rectilinear static and dynamic discrete 2-Center problems. Lecture Notes in Computer Science, 1663:276–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stege:1999:GTS

[1756] U. Stege. Gene trees and species trees:

The gene-duplication problem is fixedparameter tractable. Lecture Notes in Computer Science, 1663:288–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1999:EWS

[1757] A. Czumaj, I. Finch, L. Gasieniec, and A. Gibbons. Efficient Web searching using temporal factors. Lecture Notes in Computer Science, 1663:294– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iturriaga:1999:ELA

[1758] C. Iturriaga and A. Lubiw. Elastic labels around the perimeter of a map. Lecture Notes in Computer Science, 1663:306–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bespamyatnikh:1999:OFL

[1759] S. Bespamyatnikh, K. Kedem, and M. Segal. Optimal facility location under various distance functions. Lecture Notes in Computer Science, 1663:318– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusfield:1999:TSE

[1760] D. Gusfield and C. Martel. Thresholds for sports elimination numbers: Algorithms and complexity. Lecture Notes in Computer Science, 1663:330– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodal:1999:DRS

[1761] G. S. Brodal and R. Fagerberg. Dynamic representations of sparse graphs. Lecture Notes in Computer Science, 1663:342–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitter:1999:ITO

[1762] J. S. Vitter. Invited talk: Online data structures in external memory. Lecture Notes in Computer Science, 1663:352–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coron:1999:SRP

[1763] Jean-Sébastien Coron, David Naccache, and Julien P. Stern. On the security of RSA padding. Lecture Notes in Computer Science, 1666:1– 18, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660001.
pdf.

Kipnis:1999:CHP

[1764] Aviad Kipnis and Adi Shamir. Cryptanalysis of the HFE public key cryptosystem by relinearization. Lecture Notes in Computer Science, 1666:19– 30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660019.
pdf.

Nguyen:1999:HHS

[1765] Phong Nguyen and Jacques Stern.

The hardness of the hidden subset sum problem and its cryptographic implications. Lecture Notes in Computer Science, 1666:31–46, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660031.
pdf.

Maurer:1999:ITC

[1766] Ueli Maurer. Information-theoretic cryptography (extended abstract). Lecture Notes in Computer Science, 1666: 47–64, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660047.
pdf.

Aumann:1999:ITS

[1767] Yonatan Aumann and Michael O. Rabin. Information theoretically secure communication in the limited storage space model. Lecture Notes in Computer Science, 1666:65–79, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660065.
pdf.

Beimel:1999:ANN

[1768] Amos Beimel, Tal Malkin, and Silvio Micali. The all-or-nothing nature of two-party secure computation. Lecture Notes in Computer Science, 1666:80– 97, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660080.
pdf.

Canetti:1999:AST

[1769] Ran Canetti, Rosario Gennaro, Stanisław Jarecki, Hugo Krawczyk, and Tal Rabin. Adaptive security for threshold cryptosystems. Lecture Notes in Computer Science, 1666:98–115, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660098.
pdf.

Gilboa:1999:TPR

[1770] Niv Gilboa. Two party RSA key generation (extended abstract). Lecture Notes in Computer Science, 1666:116– 129, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660116.
pdf.

Abe:1999:RDM

[1771] Masayuki Abe. Robust distributed multiplication without interaction. Lecture Notes in Computer Science, 1666:130–147, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1666/16660130.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1666/16660130.pdf.

Schoenmakers:1999:SPV

[1772] Berry Schoenmakers. A simple publicly verifiable secret sharing scheme and its application to electronic voting. Lecture Notes in Computer Science, 1666:148–164, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660148.
                                        pdf.

Knudsen:1999:TDS

[1773] Lars R. Knudsen, M. J. B. Robshaw, and David Wagner. Truncated differentials and Skipjack. Lecture Notes in Computer Science, 1666:165– 180, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660165.
                                        pdf.

Johansson:1999:FCA

[1774] Thomas Johansson and Fredrik Jönsson. Fast correlation attacks based on turbo code techniques. Lecture Notes in Computer Science, 1666:181–197, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1666/16660181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660181.
pdf.

Maitra:1999:HNR

[1775] Subhamoy Maitra and Palash Sarkar.

Highly nonlinear resilient functions optimizing Siegenthaler’s inequality. Lecture Notes in Computer Science, 1666: 198–215, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1666/16660198.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1666/16660198.
pdf.

Black:1999:UFS

[1776] J. Black, S. Halevi, H. Krawczyk, T. Krovetz, and P. Rogaway. UMAC: Fast and secure message authentication. Lecture Notes in Computer Science, 1666:216–233, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1666/16660216.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1666/16660216.pdf.

Etzel:1999:QHA

[1777] Mark Etzel, Sarvar Patel, and Zulfikar Ramzan. Square Hash: Fast message authentication via optimized universal hash functions. Lecture Notes in Computer Science, 1666:234– 251, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660234.
                                        pdf.

An:1999:CVM

[1778] Jee Hea An and Mihir Bellare. Constructing VIL-MACs from FIL-MACs: Message authentication under weakened assumptions. Lecture Notes in Computer Science, 1666:252–269, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660252.
                                        pdf.

Bellare:1999:SEP

[1779] Mihir Bellare, Oded Goldreich, and Hugo Krawczyk. Stateless evaluation of pseudorandom functions: Security beyond the birthday barrier. Lecture Notes in Computer Science, 1666:270– 287, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660270.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660270.
                                        pdf.

Nguyen:1999:CGG

[1780] Phong Nguyen. Cryptanalysis of the Goldreich-Goldwasser-Halevi cryptosystem from Crypto ’97. Lecture Notes in Computer Science, 1666:288– 304, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660288.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660288.
pdf.

Coppersmith:1999:WQS

[1781] Don Coppersmith. Weakness in quaternion signatures. Lecture Notes in Computer Science, 1666:305–314, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660305.
pdf.

Ye:1999:CS

[1782] D.-F. Ye, K.-Y. Lam, and Z.-D. Dai.

Cryptanalysis of “2R” schemes. Lecture Notes in Computer Science, 1666:315– 325, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding-Feng:1999:CS

[1783] Ye Ding-Feng, Lam Kwok-Yan, and Dai Zong-Duo. Cryptanalysis of “2R” schemes. Lecture Notes in Computer Science, 1666:315–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660315.
pdf.

Boneh:1999:FL

[1784] Dan Boneh, Glenn Durfee, and Nick Howgrave-Graham. Factoring N = p^{r}q for large r. Lecture Notes in Computer Science, 1666:326–337, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660326.
                                        pdf.

Boneh:1999:EPK

[1785] Dan Boneh and Matthew Franklin. An efficient public key traitor scheme (extended abstract). Lecture Notes in Computer Science, 1666:338–353, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660338.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660338.
                                        pdf.

Fiat:1999:DTT

[1786] Amos Fiat and Tamir Tassa. Dynamic traitor tracing. Lecture Notes in Computer Science, 1666:354–371, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660354.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660354.
                                        pdf.

Gafni:1999:EMI

[1787] Eli Gafni, Jessica Staddon, and Yiqun Lisa Yin. Efficient methods for integrating traceability and broadcast encryption. Lecture Notes in Computer Science, 1666:372–387, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660372.
pdf.

Kocher:1999:DPA

[1788] Paul Kocher, Joshua Jaffe, and Benjamin Jun. Differential power analysis. Lecture Notes in Computer Science, 1666:388–397, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1666/16660388.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1666/16660388.pdf.

Chari:1999:TSA

[1789] Suresh Chari, Charanjit S. Jutla, Josyula R. Rao, and Pankaj Rohatgi. Towards sound approaches to counteract power-analysis attacks. Lecture Notes in Computer Science, 1666:398– 412, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660398.
pdf.

Camenisch:1999:SEG

[1790] Jan Camenisch and Markus Michels. Separability and efficiency for generic group signature schemes (extended abstract). Lecture Notes in Computer Science, 1666:413–430, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1666/16660413.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1666/16660413.pdf.

Bellare:1999:FSD

[1791] Mihir Bellare and Sara K. Miner. A forward-secure digital signature scheme. Lecture Notes in Computer Science, 1666:431–448, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1666/16660431.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1666/16660431.pdf.

Garay:1999:AFO

[1792] Juan A. Garay, Markus Jakobsson, and Philip MacKenzie. Abuse-free optimistic contract signing. Lecture Notes in Computer Science, 1666:449– 466, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660449.
                                        pdf.

Goldreich:1999:CSZ

[1793] Oded Goldreich, Amit Sahai, and Salil Vadhan. Can statistical zero knowledge be made noninteractive? or on the relationship of SZK and NISZK. Lecture Notes in Computer Science, 1666:467– 484, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660467.
pdf.

DiCrescenzo:1999:CZK

[1794] G. Di Crescenzo and R. Ostrovsky. On concurrent zero-knowledge with preprocessing. Lecture Notes in Computer Science, 1666:485–502, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzo:1999:CZK

[1795] Giovanni Di Crescenzo and Rafail Ostrovsky. On concurrent zeroknowledge with pre-processing (extended abstract). Lecture Notes in Computer Science, 1666:485–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660485.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660485.
pdf.

Boyko:1999:SPO

[1796] Victor Boyko. On the security properties of OAEP as an all-or-nothing transform. Lecture Notes in Computer Science, 1666:503–518, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660503.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660503.
pdf.

Bellare:1999:NME

[1797] Mihir Bellare and Amit Sahai. Non-malleable encryption: Equivalence between two notions, and an indistinguishability-based characterization. Lecture Notes in Computer Science, 1666:519–536, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660519.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660519.
pdf.

Fujisaki:1999:SIA

[1798] Eiichiro Fujisaki and Tatsuaki Okamoto. Secure integration of asymmetric and symmetric encryption schemes. Lecture Notes in Computer Science, 1666:537– 554, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660537.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660537.
pdf.

Sander:1999:AAE

[1799] Tomas Sander and Amnon Ta-Shma. Auditable, anonymous electronic cash (extended abstract). Lecture Notes in Computer Science, 1666:555–572, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660555.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660555.
pdf.

Naor:1999:OTA

[1800] Moni Naor and Benny Pinkas. Oblivious transfer with adaptive queries. Lecture Notes in Computer Science, 1666: 573–590, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1666/16660573.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1666/16660573.
pdf.

Gilboa:1999:CCR

[1801] Niv Gilboa and Yuval Ishai. Compressing cryptographic resources (extended abstract). Lecture Notes in Computer Science, 1666:591–608, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1666/16660591.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1666/16660591.
pdf.

Kumar:1999:CCB

[1802] Ravi Kumar, Sridhar Rajagopalan, and Amit Sahai. Coding constructions for blacklisting problems without computational assumptions. Lecture Notes in Computer Science, 1666:609– 623, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660609.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660609.
                                        pdf.

Poovendran:1999:ITA

[1803] R. Poovendran and J. S. Baras. An information theoretic analysis of rootedtree based secure multicast key distribution schemes. Lecture Notes in Computer Science, 1666:624–638, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1666/16660624.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1666/16660624.
                                        pdf.

Anonymous:1999:AIj

[1804] Anonymous. Author index. Lecture Notes in Computer Science, 1666:639– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1666/16660639.pdf.

Goldberg:1999:SPA

[1805] Andrew V. Goldberg. Selecting problems for algorithm evaluation. Lecture Notes in Computer Science, 1668:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680001.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1668/16680001.
pdf.

McColl:1999:BAW

[1806] Bill McColl. BSP algorithms — “write once, run anywhere”. Lecture Notes in Computer Science, 1668:12– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680012.
pdf.

Mehlhorn:1999:TYL

[1807] Kurt Mehlhorn. Ten years of LEDA: Some thoughts. Lecture Notes in Computer Science, 1668:14– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680014.
pdf.

Jimenez:1999:CSP

[1808] Victor M. Jiménez and Andrés Marzal.

Computing the K shortest paths: A new algorithm and an experimental comparison. Lecture Notes in Computer Science, 1668:15–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680015.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1668/16680015.
pdf.

Giegerich:1999:EIL

[1809] Robert Giegerich, Stefan Kurtz, and Jens Stoye. Efficient implementation of lazy suffix trees. Lecture Notes in Computer Science, 1668:30– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680030.
pdf.

Dascal:1999:ELR

[1810] Shlomit Dascal and Uzi Vishkin. Experiments with list ranking for explicit multi-threaded (XMT) instruction parallelism (extended abstract). Lecture Notes in Computer Science, 1668:43– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680043.
pdf.

Werneck:1999:FMC

[1811] Renato Fonseca F. Werneck, João Carlos Setubal, and Arlindo F. da Conceição. Finding minimum congestion spanning trees. Lecture Notes in Computer Science, 1668:60–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680060.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1668/16680060.
pdf.

Kavvadias:1999:EAT

[1812] Dimitris J. Kavvadias and Elias C.

Stavropoulos. Evaluation of an algorithm for the transversal hypergraph problem. Lecture Notes in Computer Science, 1668:72–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680072.
pdf.

Glover:1999:CHD

[1813] Fred Glover, Gregory Gutin, Anders Yeo, and Alexey Zverovich. Construction heuristics and domination analysis for the asymmetric TSP. Lecture Notes in Computer Science, 1668:85– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680085.
pdf.

Hatzis:1999:CMN

[1814] K. Hatzis, G. Pentaris, P. Spirakis, and B. Tampakas. Counting in mobile networks: Theory and experimentation. Lecture Notes in Computer Science, 1668:95–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680095.
                                        pdf.

Schulz:1999:DAL

[1815] Frank Schulz, Dorothea Wagner, and Karsten Weihe. Dijkstra’s algorithm on-line: An empirical case study from public railroad transport. Lecture Notes in Computer Science, 1668:110– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680110.
                                        pdf.

Fatourou:1999:IEE

[1816] Panagiota Fatourou, Paul Spirakis, Panagiotis Zarafidis, and Anna Zoura. Implementation and experimental evaluation of graph connectivity algorithms using LEDA. Lecture Notes in Computer Science, 1668:124–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680124.
                                        pdf.

Aharoni:1999:LZC

[1817] Yuval Aharoni, Dan Halperin, Iddo Hanniel, Sariel Har-Peled, and Chaim Linhart. On-line zone construction in arrangements of lines in the plane. Lecture Notes in Computer Science, 1668:

139–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680139.
pdf.

Flato:1999:DIP

[1818] Eyal Flato, Dan Halperin, Iddo Hanniel, and Oren Nechushtan. The design and implementation of planar maps in C GAL. Lecture Notes in Computer Science, 1668:154–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680154.
pdf.

Comes:1999:EUI

[1819] Jochen Comes and Mark Ziegelmann.

An easy to use implementation of linear perturbations within C GAL. Lecture Notes in Computer Science, 1668: 169–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680169.
pdf.

Rahman:1999:ACE

[1820] Naila Rahman and Rajeev Raman. Analysing cache effects in distribution sorting. Lecture Notes in Computer Science, 1668:183–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680183.
                                        pdf.

Navarro:1999:FRE

[1821] Gonzalo Navarro and Mathieu Raffinot. Fast regular expression search. Lecture Notes in Computer Science, 1668:198–212, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1668/16680198.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1668/16680198.pdf.

Korda:1999:EEH

[1822] Maureen Korda and Rajeev Raman. An experimental evaluation of hybrid data structures for searching. Lecture Notes in Computer Science, 1668:213– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680213.
                                        pdf.

Crauser:1999:LSE

[1823] Andreas Crauser and Kurt Mehlhorn. LEDA-SM: Extending LEDA to secondary memory. Lecture Notes in Computer Science, 1668:228–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680228.
pdf.

Fredman:1999:PQT

[1824] Michael L. Fredman. A priority queue transform. Lecture Notes in Computer Science, 1668:243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680243.
pdf.

Bouganis:1999:IIE

[1825] Athanasios Bouganis, Ioannis Caragiannis, and Christos Kaklamanis. Implementation issues and experimental study of a wavelength routing algorithm for irregular alloptical networks. Lecture Notes in Computer Science, 1668:258–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680258.
pdf.

Huson:1999:ELD

[1826] Daniel H. Huson, Kelly Ann Smith, and Tandy J. Warnow. Estimating large distances in phylogenetic reconstruction. Lecture Notes in Computer Science, 1668:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680271.
                                        pdf.

Hanke:1999:PCR

[1827] Sabine Hanke. The performance of concurrent red-black tree algorithms. Lecture Notes in Computer Science, 1668: 286–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680286.
                                        pdf.

Bojesen:1999:PEC

[1828] Jesper Bojesen, Jyrki Katajainen, and Maz Spork. Performance engineering case study: Heap construction. Lecture Notes in Computer Science, 1668: 301–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1668/16680301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1668/16680301.
                                        pdf.

Caramia:1999:FSL

[1829] Massimiliano Caramia and Paolo Dell’Olmo. A fast and simple local search for graph coloring. Lecture Notes in Computer Science, 1668:316– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680316.
pdf.

Boghossian:1999:BBA

[1830] Nicolas Boghossian, Oliver Kohlbacher, and Hans-Peter Lenhof. BALL: Biochemical algorithms library. Lecture Notes in Computer Science, 1668:330– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680330.
pdf.

Brengel:1999:ESP

[1831] Klaus Brengel, Andreas Crauser, Paolo Ferragina, and Ulrich Meyer. An experimental study of priority queues in external memory. Lecture Notes in Computer Science, 1668:345–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1668/16680345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1668/16680345.
pdf.

Anonymous:1999:AIk

[1832] Anonymous. Author index. Lecture Notes in Computer Science, 1668:361– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/1668/1668auth.pdf.

Broder:1999:CRP

[1833] A. Z. Broder and M. Mitzenmacher. Completeness and robustness properties of min-wise independent permutations. Lecture Notes in Computer Science, 1671:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saks:1999:LDS

[1834] M. Saks, A. Srinivasan, S. Zhou, and D. Zuckerman. Low discrepancy sets yield approximate min-wise independent permutation families. Lecture Notes in Computer Science, 1671:11– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alon:1999:ISH

[1835] N. Alon, U. Arad, and Y. Azar. Independent sets in hypergraphs with applications to routing via fixed paths. Lecture Notes in Computer Science, 1671: 16–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gudmundsson:1999:AMM

[1836] J. Gudmundsson, C. Levcopoulos, and G. Narasimhan. Approximating minimum Manhattan networks. Lecture Notes in Computer Science, 1671:28– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doerr:1999:AMC

[1837] B. Doerr and A. Srivastav. Approximation of multi-color discrepancy. Lecture Notes in Computer Science, 1671: 39–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kellerer:1999:PTA

[1838] H. Kellerer. A polynomial time approximation scheme for the multiple knapsack problem. Lecture Notes in Computer Science, 1671:51–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihail:1999:SCR

[1839] M. Mihail. Set cover with requirements and costs evolving over time. Lecture Notes in Computer Science, 1671:63– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halldorsson:1999:MPG

[1840] M. M. Halldorsson and G. Kortsarz.

Multicoloring planar graphs and partial k-trees. Lecture Notes in Computer Science, 1671:73–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parnas:1999:TDG

[1841] M. Parnas and D. Ron. Testing the diameter of graphs. Lecture Notes in Computer Science, 1671:85–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dodis:1999:ITA

[1842] Y. Dodis, O. Goldreich, E. Lehman, and S. Raskhodnikova. Improved testing algorithms for monotonicity. Lecture Notes in Computer Science, 1671: 97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aumann:1999:LCT

[1843] Y. Aumann, J. Haastad, M. O. Rabin, and M. Sudan. Linear consistency testing. Lecture Notes in Computer Science, 1671:109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morris:1999:IBS

[1844] B. J. Morris. Improved bounds for sampling contingency tables. Lecture Notes in Computer Science, 1671:121– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wigderson:1999:PDA

[1845] A. Wigderson. Probabilistic and deterministic approximations of the permanent. Lecture Notes in Computer Science, 1671:130–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldreich:1999:IDB

[1846] O. Goldreich and A. Wigderson. Improved derandomization of BPP using a hitting set generator. Lecture Notes in Computer Science, 1671:131– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baltz:1999:PCS

[1847] A. Baltz, T. Schoen, and A. Srivastav. Probabilistic construction of small strongly sum-free sets via large sidon sets. Lecture Notes in Computer Science, 1671:138–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moehring:1999:SMS

[1848] R. H. Moehring, A. S. Schulz, and M. Uetz. Stochastic machine scheduling: Performance guarantees for LPbased priority policies. Lecture Notes in Computer Science, 1671:144–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fotakis:1999:ERA

[1849] D. A. Fotakis and P. G. Spirakis. Efficient redundant assignments under fault-tolerance constraints. Lecture Notes in Computer Science, 1671:156– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imreh:1999:SMC

[1850] C. Imreh and J. Noga. Scheduling with machine cost. Lecture Notes in Computer Science, 1671:168–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1999:LTA

[1851] K. Jansen, R. Solis-Oba, and M. Sviridenko. A linear time approximation scheme for the job shop scheduling problem. Lecture Notes in Computer Science, 1671:177–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feige:1999:RRS

[1852] U. Feige. Randomized rounding for semidefinite programs — variations on the MAX CUT example. Lecture Notes in Computer Science, 1671:189– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clementi:1999:HRP

[1853] A. E. F. Clementi, P. Penna, and R. Silvestri. Hardness results for the power range assignment problem in packet radio networks. Lecture Notes in Computer Science, 1671:197–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1999:NAA

[1854] C. T. Cheng. A new approximation algorithm for the demand routing and slotting problem with unit demands on rings. Lecture Notes in Computer Science, 1671:209–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Condon:1999:AGP

[1855] A. E. Condon and R. M. Karp. Algorithms for graph partitioning on the planted partition model. Lecture Notes in Computer Science, 1671:221–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettie:1999:RTW

[1856] S. Pettie and V. Ramachandran. A randomized time-work optimal parallel algorithm for finding a minimum spanning forest. Lecture Notes in Computer Science, 1671:233–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Batu:1999:FAP

[1857] T. Batu, R. Rubinfeld, and P. White. Fast approximate PCPs for multidimensional bin-packing problems. Lecture Notes in Computer Science, 1671:

245–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1999:PAS

[1858] R. A. Martin and D. Randall. Pfaffian algorithms for sampling routings on regions with free boundary conditions. Lecture Notes in Computer Science, 1671:257–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1999:SUM

[1859] S. Albers and G. Schmidt. Scheduling with unexpected machine breakdowns. Lecture Notes in Computer Science, 1671:269–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afrati:1999:SCN

[1860] F. Afrati, E. Bampis, C. Kenyon, and I. Mills. Scheduling on a constant number of machines. Lecture Notes in Computer Science, 1671:281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andou:1999:ARTb

[1861] Tomohito Andou. Andhill-98: A RoboCup team which reinforces positioning with observation. Lecture Notes in Computer Science, 1674:338– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040338.
pdf.

Sewell:1999:LIC

[1862] Peter Sewell, Paweł T. Wojciechowski, and Benjamin C. Pierce. Locationindependent communication for mobile agents: A two-level architecture. Lecture Notes in Computer Science, 1686: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1686/16860001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1686/16860001.
pdf.

VanRoy:1999:LRO

[1863] Peter Van Roy, Per Brand, Seif Haridi, and Raphaël Collet. A lightweight reliable object migration protocol. Lecture Notes in Computer Science, 1686: 32–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1686/16860032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1686/16860032.
pdf.

Vitek:1999:SFS

[1864] Jan Vitek and Giuseppe Castagna. Seal: A framework for secure mobile computations. Lecture Notes in Computer Science, 1686:47–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1686/16860047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1686/16860047.
pdf.

Rowstron:1999:RTS

[1865] Antony Rowstron and Stuart Wray. A run-time system for WCL. Lecture Notes in Computer Science, 1686:78– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1686/16860078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1686/16860078.
pdf.

Ramming:1999:PLI

[1866] J. Christopher Ramming. PML:

A language interface to distributed voice-response units. Lecture Notes in Computer Science, 1686:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1686/16860097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1686/16860097.
pdf.

Duggan:1999:DCI

[1867] Dominic Duggan. Derivatives: A construct for Internet programming. Lecture Notes in Computer Science, 1686: 113–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1686/16860113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1686/16860113.
pdf.

Hicks:1999:NPU

[1868] Michael Hicks, Pankaj Kakkar, Jonathan T.

Moore, Carl A. Gunter, and Scott Nettles. Network programming using PLAN. Lecture Notes in Computer Science, 1686:127–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1686/16860127.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1686/16860127.pdf.

Costa:1999:UTC

[1869] Augusto Cesar Pinto Loureiro da Costa and Guilherme Bittencourt. UFSC-Team: A cognitive multi-agent approach to the RoboCup ’98 simulator league. Lecture Notes in Computer Science, 1694:371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1604/16040371.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1604/16040371.pdf.

PintoLoureirodaCosta:1999:UTC

[1870] Augusto Cesar Pinto Loureiro da Costa and Guilherme Bittencourt. UFSC-Team: A cognitive multiagent approach to the RoboCup ’98 simulator league. Lecture Notes in Computer Science, 1694:371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040371.
pdf.

Thimbleby:1999:DPH

[1871] H. Thimbleby. Design probes for handheld and ubiquitous computing. Lecture Notes in Computer Science, 1707: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070001.
pdf.

Stenton:1999:MCC

[1872] S. P. Stenton. Maintaining context and control in a digital world. Lecture Notes in Computer Science, 1707: 20–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070020.
pdf.

Schmandt:1999:EM

[1873] Ch. Schmandt. Everywhere messaging.

Lecture Notes in Computer Science, 1707:22–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1707/17070022.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1707/17070022.
pdf.

Susani:1999:MUS

[1874] M. Susani. Mobile, ubiquitous and the sense of space. Lecture Notes in Computer Science, 1707:28–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070028.
pdf.

Mikhak:1999:CMH

[1875] B. Mikhak, F. Martin, M. Resnick, R. Berg, and B. Silverman. The children’s machines: Handheld and wearable computers too. Lecture Notes in Computer Science, 1707:31– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070031.
pdf.

Brody:1999:PBH

[1876] A. B. Brody and E. J. Gottsman. Pocket BargainFinder: A handheld device for augmented commerce. Lecture Notes in Computer Science, 1707:44– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070044.
pdf.

Jose:1999:SFL

[1877] R. José and N. Davies. Scalable and flexible location-based services for ubiquitous information access. Lecture Notes in Computer Science, 1707:52– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070052.
pdf.

Couderc:1999:ECA

[1878] P. Couderc and A.-M. Kermarrec. Enabling context-awareness from network-level location tracking. Lecture Notes in Computer Science, 1707: 67–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070067.
pdf.

Pentland:1999:PI

[1879] A. Pentland. Perceptual intelligence.

Lecture Notes in Computer Science, 1707:74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1707/17070074.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1707/17070074.
pdf.

Schmidt:1999:AIC

[1880] A. Schmidt, K. A. Aidoo, A. Takaluoma, U. Tuomela, K. Van Laerhoven, and W. Van de Velde. Advanced interaction in context. Lecture Notes in Computer Science, 1707:89– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070089.
                                        pdf.

Fjeld:1999:EBB

[1881] M. Fjeld, F. Voorhorst, M. Bichsel, K. Lauche, M. Rauterberg, and H. Krueger. Exploring brick-based navigation and composition in an augmented reality. Lecture Notes in Computer Science, 1707:102–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070102.
                                        pdf.

Allen:1999:HCP

[1882] J. P. Allen. Handheld computing predictions: What went wrong? Lecture Notes in Computer Science, 1707:117– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070117.
                                        pdf.

Stabell-Kulo:1999:OEA

[1883] T. Stabell-Kulø, F. Dillema, and T. Fallmyr. The open-end argument for private computing. Lecture Notes in Computer Science, 1707:124– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070124.
pdf.

Roman:1999:IPD

[1884] M. Román, A. Singhai, D. Carvalho, Ch. Hess, and R. H. Campbell. Integrating PDAs into distributed systems: 2K and PalmORB. Lecture Notes in Computer Science, 1707:137– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070137.
pdf.

Silva:1999:DIA

[1885] M. J. Silva and A. P. Afonso. Designing information appliances using a resource replication model. Lecture Notes in Computer Science, 1707:150– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070150.
pdf.

McCarthy:1999:CMA

[1886] J. F. McCarthy and E. S. Meidel.

A CTIVE M AP : A visualization tool for location awareness to support informal interactions. Lecture Notes in Computer Science, 1707:158– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070158.
                                        pdf.

Kortuem:1999:CES

[1887] G. Kortuem, Z. Segall, and Th.G. Cowan Thompson. Close encounters: Supporting mobile collaboration through interchange of user profiles. Lecture Notes in Computer Science, 1707:171– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070171.
                                        pdf.

Liechti:1999:DPF

[1888] O. Liechti and T. Ichikawa. A digital photography framework supporting social interaction and affective awareness. Lecture Notes in Computer Science, 1707:186–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070186.
                                        pdf.

Cheverst:1999:RCS

[1889] K. Cheverst, N. Davies, K. Mitchell, and A. Friday. The role of connectivity in supporting contextsensitive applications. Lecture Notes in Computer Science, 1707:193–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070193.
pdf.

Pascoe:1999:IDC

[1890] J. Pascoe, N. Ryan, and D. Morse. Issues in developing context-aware computing. Lecture Notes in Computer Science, 1707:208–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1707/17070208.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1707/17070208.pdf.

Ancona:1999:RMC

[1891] M. Ancona, G. Dodero, and V. Gianuzzi. RAMSES: A mobile computing system for field archaeology. Lecture Notes in Computer Science, 1707: 222–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070222.
pdf.

Holmquist:1999:TBA

[1892] L. E. Holmquist, J. Redström, and P. Ljungstrand. Token-based access to digital information. Lecture Notes in Computer Science, 1707:234– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070234.
                                        pdf.

Kohtake:1999:IID

[1893] N. Kohtake, J. Rekimoto, and Y. Anzai. InfoStick: An interaction device for inter-appliance computing. Lecture Notes in Computer Science, 1707:246– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070246.
                                        pdf.

Beigl:1999:USC

[1894] M. Beigl. Using spatial co-location for coordination in ubiquitous computing environments. Lecture Notes in Computer Science, 1707:259–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070259.
                                        pdf.

Falk:1999:AR

[1895] J. Falk, J. Redström, and S. Björk. Amplifying reality. Lecture Notes in Computer Science, 1707:274–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1707/17070274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070274.
pdf.

Kristoffersen:1999:DIS

[1896] S. Kristoffersen and F. Ljungberg. Designing interaction styles for a mobile use context. Lecture Notes in Computer Science, 1707:281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070281.
pdf.

Masui:1999:PET

[1897] T. Masui. POBox: An efficient text input method for handheld and ubiquitous computers. Lecture Notes in Computer Science, 1707:289–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070289.
pdf.

Aiken:1999:MUC

[1898] R. J. Aiken, A. Abramski, J. Bates, and Th. Blackadar. Middleware for ubiquitous computing. Lecture Notes in Computer Science, 1707:301–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1707/17070301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070301.
                                        pdf.

Abowd:1999:TBU

[1899] G. D. Abowd, A. K. Dey, P. J. Brown, N. Davies, M. Smith, and P. Steggles. Towards a better understanding of context and context-awareness. Lecture Notes in Computer Science, 1707:304– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070304.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070304.
                                        pdf.

Gellersen:1999:MAT

[1900] H.-W. Gellersen, M. Beigl, and H. Krull. The MediaCup: Awareness technology embedded in an everyday object. Lecture Notes in Computer Science, 1707:308–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070308.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070308.
                                        pdf.

Beigl:1999:PCI

[1901] M. Beigl. Point & click — interaction in smart environments. Lecture Notes in Computer Science, 1707:311– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1707/17070311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070311.
pdf.

Baber:1999:WIA

[1902] Ch. Baber, D. J. Haniff, and R. Buckley. Wearable information appliances for the emergency services: HotHelmet. Lecture Notes in Computer Science, 1707:314–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070314.
pdf.

Hoisko:1999:UWC

[1903] J. Hoisko. Using wearable computer as an audiovisual memory prosthesis. Lecture Notes in Computer Science, 1707: 317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070317.
pdf.

Panayi:1999:TS

[1904] M. Panayi, W. Van de Velde, D. Roy, O. Cakmakci, K. De Paepe, and N. O. Bernsen. Today’s stories. Lecture Notes in Computer Science, 1707:320– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070320.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070320.
                                        pdf.

VandeVelde:1999:SEW

[1905] W. Van de Velde. On the self evaluation of a wearable assistant. Lecture Notes in Computer Science, 1707:324– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070324.
                                        pdf.

Persa:1999:PAR

[1906] S. Persa and P. Jonker. On positioning for augmented reality systems. Lecture Notes in Computer Science, 1707: 327–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070327.
                                        pdf.

Oppermann:1999:HNI

[1907] R. Oppermann, M. Specht, and I. Jaceniak. Hippie: A nomadic information system. Lecture Notes in Computer Science, 1707:330–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070330.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070330.
                                        pdf.

ORafferty:1999:RCL

[1908] R. P. O’Rafferty, M. J. O’Grady, and G. M. P. O’Hare. A rapidly configurable location-aware information system for an exterior environment. Lecture Notes in Computer Science, 1707: 334–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070334.
pdf.

Brachtl:1999:MCM

[1909] M. Brachtl, L. Miksícek, and P. Slavík.

Mobile computing in machine engineering applications. Lecture Notes in Computer Science, 1707:337–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070337.
pdf.

Smit:1999:CRH

[1910] G. J. M. Smit, T. Bos, P. J. M.

Havinga, S. J. Mullender, and J. Smit. Chameleon — reconfigurability in hand-held multimedia computers. Lecture Notes in Computer Science, 1707: 340–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070340.
pdf.

Jones:1999:EWS

[1911] M. Jones, G. Buchanan, and N. Mohd-Nasir. An evaluation of WebTwig — A site outliner for handheld Web access. Lecture Notes in Computer Science, 1707:343–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1707/17070343.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070343.
                                        pdf.

Haniff:1999:HFM

[1912] D. J. Haniff, Ch. Baber, and W. Edmondson. Human factors of multimodal ubiquitous computing. Lecture Notes in Computer Science, 1707:346– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070346.
                                        pdf.

Donnelly:1999:UES

[1913] A. Donnelly and E. Barnstedt. URCP: Experimental support for multimodal interfaces. Lecture Notes in Computer Science, 1707:349–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070349.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070349.
                                        pdf.

Wan:1999:MMC

[1914] D. Wan. Magic medicine cabinet: A situated portal for consumer healthcare. Lecture Notes in Computer Science, 1707:352–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070352.
pdf.

Dempski:1999:AWW

[1915] K. L. Dempski. Augmented workspace:

The world as your desktop. Lecture Notes in Computer Science, 1707:356– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070356.
pdf.

Redstrom:1999:C

[1916] J. Redström, P. Jaksetic, and P. Ljungstrand. The ChatterBox. Lecture Notes in Computer Science, 1707: 359–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070359.
pdf.

Glance:1999:PVN

[1917] N. Glance and D. Snowdon. Pollen:

Virtual networks that use people as carriers. Lecture Notes in Computer Science, 1707:362–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070362.
pdf.

Kanter:1999:VCA

[1918] Th. Kanter and H. Gustafsson. VoIP in context-aware communication spaces. Lecture Notes in Computer Science, 1707:365–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1707/17070365.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1707/17070365.
pdf.

vanderMeer:1999:PEA

[1919] S. van der Meer, S. Arbanowski, and R. Popescu-Zeletin. A platform for environment-aware applications. Lecture Notes in Computer Science, 1707: 368–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070368.
pdf.

Afonso:1999:DIU

[1920] A. P. Afonso, M. J. Silva, J. P. Campos, and F. S. Regateiro. The design and implementation of the ubidata information dissemination framework. Lecture Notes in Computer Science, 1707: 371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070371.
pdf.

Burger:1999:CAD

[1921] C. Burger and O. Schramm. Coauthoring in dynamic teams with mobile individuals. Lecture Notes in Computer Science, 1707:374–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070374.
pdf.

Kubach:1999:ULA

[1922] U. Kubach and K. Rothermel.

A universal, location-aware hoarding mechanism. Lecture Notes in Computer Science, 1707:377–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1707/17070377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1707/17070377.
pdf.

Chalmers:1999:QCA

[1923] D. Chalmers and M. Sloman. QoS and context awareness for mobile computing. Lecture Notes in Computer Science, 1707:380–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070380.
                                        pdf.

Molina-Jimenez:1999:ACC

[1924] C. Molina-Jiménez and L. Marshall. Anonymous and confidential communications from an IP addressless computer. Lecture Notes in Computer Science, 1707:383–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070383.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070383.
                                        pdf.

Lim:1999:AHN

[1925] H. Lim and Ch. Kim. Ad-hoc network routing for centralized information sharing systems. Lecture Notes in Computer Science, 1707:386–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1707/17070386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1707/17070386.
                                        pdf.

Anonymous:1999:AIl

[1926] Anonymous. Author index. Lecture Notes in Computer Science, 1707:389– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1707/1707auth.pdf.

Hertfordshire:1999:SBM

[1927] Winand H. Dittrich. Seeing biological motion — is there a role for cognitive strategies? Lecture Notes in Computer Science, 1739:3– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390003.
pdf.

Hummels:1999:EPG

[1928] Caroline Hummels and Kees Overbeeke. The expressive power of gestures: Capturing scent in a spatial shape. Lecture Notes in Computer Science, 1739:23–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1739/17390023.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1739/17390023.pdf.

Wanderley:1999:NOP

[1929] Marcelo M. Wanderley. Nonobvious performer gestures in instrumental music. Lecture Notes in Computer Science, 1739:37–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390037.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1739/17390037.
                                        pdf.

Angeli:1999:EAM

[1930] Antonella De Angeli, Fréderic Wolff, Laurent Romary, and Walter Gerbino. The ecological approach to multimodal system design. Lecture Notes in Computer Science, 1739:49–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390049.
                                        pdf.

Roby-Brami:1999:ATU

[1931] Agnès Roby-Brami, Mounir Mokhtari, Isabelle Laffont, Nezha Bennis, and Elena Biryukova. Analysis of trunk and upper limb synergies. Lecture Notes in Computer Science, 1739:53– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390053.
                                        pdf.

Nolker:1999:GVR

[1932] Claudia Nölker and Helge Ritter. GREFIT: Visual recognition of hand postures. Lecture Notes in Computer Science, 1739:61–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390061.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1739/17390061.
pdf.

Triesch:1999:TIL

[1933] Jochen Triesch, Jan Wieghardt, Eric Maël, and Christoph von der Malsburg. Towards imitation learning of grasping movements by an autonomous robot. Lecture Notes in Computer Science, 1739:73–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390073.
pdf.

Lyons:1999:LSC

[1934] Damian M. Lyons and Daniel L.

Pelletier. A line-scan computer vision algorithm for identifying human body features. Lecture Notes in Computer Science, 1739:85–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390085.
pdf.

Marcel:1999:HPR

[1935] Sébastien Marcel and Olivier Bernier.

Hand posture recognition in a bodyface centered space. Lecture Notes in Computer Science, 1739:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390097.
                                        pdf.

Wu:1999:VBG

[1936] Ying Wu and Thomas S. Huang. Vision-based gesture recognition: A review. Lecture Notes in Computer Science, 1739:103–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1739/17390103.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1739/17390103.pdf.

Boehme:1999:PLP

[1937] Hans-Joachim Boehme, Ulf-Dietrich Braumann, Andrea Corradini, and Horst-Michael Gross. Person localization and posture recognition for human-robot interaction. Lecture Notes in Computer Science, 1739:117– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390117.
                                        pdf.

Martin:1999:SGR

[1938] J. Martin, D. Hall, and J. L. Crowley. Statistical gesture recognition through modelling of parameter trajectories. Lecture Notes in Computer Science, 1739:129–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390129.
pdf.

Howell:1999:GRV

[1939] A. Jonathan Howell and Hilary Buxton. Gesture recognition for visually mediated interaction. Lecture Notes in Computer Science, 1739:141– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390141.
pdf.

Gherbi:1999:IPG

[1940] Rachid Gherbi and Annelies Braffort. Interpretation of pointing gesture: The PoG system. Lecture Notes in Computer Science, 1739:153– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390153.
pdf.

Kamp:1999:CVS

[1941] Jean-François Kamp, Franck Poirier, and Philippe Doignon. Control of invehicle systems by gestures. Lecture Notes in Computer Science, 1739:159– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390159.
                                        pdf.

University:1999:FSL

[1942] Christian Cuxac. French sign language: Proposition of a structural explanation by iconicity. Lecture Notes in Computer Science, 1739:165– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390165.
                                        pdf.

Hienz:1999:HBC

[1943] Hermann Hienz, Britta Bauer, and Karl-Friedrich Kraiss. HMM-based continuous sign language recognition using stochastic grammars. Lecture Notes in Computer Science, 1739:185– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390185.
                                        pdf.

Sagawa:1999:MAS

[1944] Hirohiko Sagawa and Masaru Takeuchi. A method for analyzing spatial relationships between words in sign language recognition. Lecture Notes in Computer Science, 1739:197–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390197.
pdf.

Vogler:1999:TSA

[1945] Christian Vogler and Dimitris Metaxas.

Toward scalability in ASL recognition: Breaking down signs into phonemes. Lecture Notes in Computer Science, 1739:211–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1739/17390211.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1739/17390211.
pdf.

Lebourque:1999:CSS

[1946] Thierry Lebourque and Sylvie Gibet.

A complete system for the specification and the generation of sign language gestures. Lecture Notes in Computer Science, 1739:227–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390227.
pdf.

Losson:1999:SSS

[1947] Olivier Losson and Jean-Marc Vannobel. Sign specification and synthesis. Lecture Notes in Computer Science, 1739:239–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1739/17390239.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390239.
                                        pdf.

Lu:1999:ACD

[1948] Shan Lu and Seiji Igi. Active character: Dynamic reaction to the user. Lecture Notes in Computer Science, 1739: 253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390253.
                                        pdf.

Julliard:1999:RPM

[1949] Frédéric Julliard and Sylvie Gibet. Reactiva’motion project: Motion synthesis based on a reactive representation. Lecture Notes in Computer Science, 1739:265–268, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390265.
                                        pdf.

Fabri:1999:EAN

[1950] Marc Fabri, David J. Moore, and Dave J. Hobbs. The emotional avatar: Non-verbal communication between inhabitants of collaborative virtual environments. Lecture Notes in Computer Science, 1739:269–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390269.
pdf.

Technology:1999:CRG

[1951] Ipke Wachsmuth. Communicative rhythm in gesture and speech. Lecture Notes in Computer Science, 1739: 277–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390277.
pdf.

Sowa:1999:TSI

[1952] Timo Sowa, Martin Fröhlich, and Marc Erich Latoschik. Temporal symbolic integration applied to a multimodal system using gestures and speech. Lecture Notes in Computer Science, 1739:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1739/17390291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1739/17390291.
pdf.

LaViola:1999:MIF

[1953] Joseph J. LaViola Jr. A multimodal interface framework for using hand gestures and speech in virtual environment applications. Lecture Notes in Computer Science, 1739:303–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390303.
                                        pdf.

Panayi:1999:SRG

[1954] Marilyn Panayi, David Roy, and James Richardson. Stimulating research into gestural human machine interaction. Lecture Notes in Computer Science, 1739:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1739/17390317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1739/17390317.
                                        pdf.

Anonymous:1999:AIm

[1955] Anonymous. Author index. Lecture Notes in Computer Science, 1739:333– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1739/1739auth.pdf.

Vallecillo:1999:OI

[1956] Antonio Vallecillo, Juan Hernández, and José M. Troya. Object interoperability. Lecture Notes in Computer Science, 1743:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1743/17430001.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1743/17430001.pdf.

Wohlrab:1999:OOO

[1957] Lutz Wohlrab, Francisco Ballesteros, Frank Schubert, Henning Schmidt, and Ashish Singhai. Object-orientation and operating systems. Lecture Notes in Computer Science, 1743:22–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430022.
pdf.

Kent:1999:USF

[1958] Stuart Kent, Andy Evans, and Bernhard Rumpe. UML semantics FAQ. Lecture Notes in Computer Science, 1743:33–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1743/17430033.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1743/17430033.
pdf.

Borne:1999:OOA

[1959] Isabelle Borne, Serge Demeyer, and Galal Hassan Galal. Object-oriented architectural evolution. Lecture Notes in Computer Science, 1743:57–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430057.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1743/17430057.
pdf.

Tomek:1999:MUO

[1960] Ivan Tomek, Alicia Diaz, Ronald Melster, António Rito Silva, Miguel Antunes, Vinny Cahill, Zièd Choukair, Alexandro Fernandez, Mads Haahr, Eric Jul, Atsushi Ohnishi, Luís Rodrigues, Erich R. Schmidt, Cristian Ţăpuş, Waldemar Wieczerzycki, and Daniel M. Zimmerman. Multi-user object-oriented environments. Lecture Notes in Computer Science, 1743:80– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430080.
pdf.

Jacobs:1999:FTJ

[1961] Bart Jacobs, Gary T. Leavens, Peter Müller, and Arnd Poetzsch-Heffter. Formal techniques for Java programs. Lecture Notes in Computer Science, 1743:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430097.
pdf.

Rashid:1999:WPS

[1962] Awais Rashid, David Parsons, and Alexandru Telea. 9th workshop for PhD students in object oriented systems. Lecture Notes in Computer Science, 1743:116–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430116.
pdf.

Noble:1999:AOO

[1963] James Noble, Jan Vitek, Doug Lea, and Paulo Sergio Almeida. Aliasing in object oriented systems. Lecture Notes in Computer Science, 1743:136– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430136.
pdf.

Ducasse:1999:EOO

[1964] Stéphane Ducasse and Oliver Ciupke.

Experiences in object-oriented reengineering. Lecture Notes in Computer Science, 1743:164–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/1743/17430164.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1743/17430164.pdf.

Szyperski:1999:COP

[1965] Clemens Szyperski, Jan Bosch, and Wolfgang Weck. Component-oriented programming. Lecture Notes in Computer Science, 1743:184–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/1743/17430184.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1743/17430184.pdf.

Lalanda:1999:OTP

[1966] Philippe Lalanda, Jan Bosch, Ramon Lerchundi, and Sophie Cherki. Object technology for product-line architectures. Lecture Notes in Computer Science, 1743:193–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1743/17430193.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1743/17430193.pdf.

Guerrini:1999:OOD

[1967] Giovanna Guerrini, Isabella Merlo, Elena Ferrari, Gerti Kappel, and Adoracion de Miguel. Objectoriented databases. Lecture Notes in Computer Science, 1743:207–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1743/17430207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1743/17430207.
                                        pdf.

Mohr:1999:PHP

[1968] Bernd Mohr, Federico Bassetti, and Kei Davis. Parallel/highperformance object-oriented scientific computing. Lecture Notes in Computer Science, 1743:222–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430222.
pdf.

Seffah:1999:IHF

[1969] Ahmed Seffah and Cameron Hayne. Integrating human factors into use cases and object-oriented methods. Lecture Notes in Computer Science, 1743:240– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430240.
pdf.

Boer:1999:OOC

[1970] Frank S. de Boer. Object-oriented and constraint programming for time critical applications. Lecture Notes in Computer Science, 1743:255–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430255.
pdf.

JardimNunes:1999:ISD

[1971] Nuno Jardim Nunes, Marco Toranzo, João Falcão e Cunha, Jaelson Castro, Srdjan Kovacevic, Dave Roberts, Jean-Claude Tarby, Mark Collins-Cope, and Mark van Harmelen. Interactive system design and object models. Lecture Notes in Computer Science, 1743: 267–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1743/17430267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1743/17430267.
                                        pdf.

Bergmans:1999:AOP

[1972] Lodewijk Bergmans and Cristina Videira Lopes (Editors). Aspect-oriented programming. Lecture Notes in Computer Science, 1743:288–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1743/17430288.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1743/17430288.pdf.

Nestmann:1999:SOP

[1973] Uwe Nestmann and António Ravara. Semantics of objects as processes (SOAP). Lecture Notes in Computer Science, 1743:314–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1743/17430314.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1743/17430314.pdf.

Abreu:1999:QAO

[1974] Fernando Brito e Abreu, Horst Zuse, Houari Sahraoui, and Walcelio Melo. Quantitative approaches in objectoriented software engineering. Lecture Notes in Computer Science, 1743:326– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430326.
pdf.

Angster:1999:IOD

[1975] Erzsébet Angster, Joseph Bergin, and László Böszörményi. Introducing OO design and programming with special emphasis on concrete examples. Lecture Notes in Computer Science, 1743: 338–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430338.
pdf.

Cohen:1999:PSO

[1976] Sholom Cohen. Panel session: Object technology and systematic reuse. Lecture Notes in Computer Science, 1743: 362–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1743/17430362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1743/17430362.
pdf.

Baquero:1999:PS

[1977] Carlos Baquero. Poster session. Lecture Notes in Computer Science, 1743: 371–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1743/17430371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1743/17430371.
                                        pdf.

Anonymous:1999:AI

[1978] Anonymous. Author index. Lecture Notes in Computer Science, 1743:385– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1743/1743auth.pdf.

Yang:1999:WVO

[1979] Jin Yang and Ioannis Kriaras. Wireless VoIP: Opportunities and challenges. Lecture Notes in Computer Science, 1748:3–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1748/17480003.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480003.
                                        pdf.

Moon:1999:GMP

[1980] Y. S. Moon, H. C. Ho, and Kenneth Wong. GSM mobile phone based communication of multimedia information: A case study. Lecture Notes in Computer Science, 1748:14– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480014.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480014.
                                        pdf.

Aron:1999:WAR

[1981] Ionut D. Aron and Sandeep K. S.

Gupta. A witness-aided routing protocol for mobile ad-hoc networks with unidirectional links. Lecture Notes in Computer Science, 1748:24– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480024.
pdf.

Mutafungwa:1999:AOB

[1982] Edward Mutafungwa and Kamugisha Kazaura. Assessing opportunities for broadband optical wireless local loops in an unbundled access network. Lecture Notes in Computer Science, 1748: 34–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480034.
pdf.

Lee:1999:SMI

[1983] Wang-Chien Lee, Johnson Lee, and Karen Huff. On simulation modeling of information dissemination systems in mobile environments. Lecture Notes in Computer Science, 1748:45– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480045.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1748/17480045.
                                        pdf.

Lee:1999:PTP

[1984] Victor C. S. Lee, Sang H. Son, and Kwok-Wa Lam. On the performance of transaction processing in broadcast environments. Lecture Notes in Computer Science, 1748:61– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480061.
                                        pdf.

Mok:1999:TPA

[1985] Eppie Mok, Hong Va Leong, and Antonio Si. Transaction processing in an asymmetric mobile environment. Lecture Notes in Computer Science, 1748: 71–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480071.
                                        pdf.

Reddy:1999:SLM

[1986] P. Krishna Reddy and Masaru Kitsuregawa. Speculative lock management to increase concurrency in mobile environments. Lecture Notes in Computer Science, 1748:82–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480082.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1748/17480082.
pdf.

Lee:1999:OCC

[1987] Victor C. S. Lee and Kwok wa Lam.

Optimistic concurrency control in broadcast environments: Looking forward at the server and backward at the clients. Lecture Notes in Computer Science, 1748:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480097.
pdf.

Gordon:1999:MWT

[1988] Steven Gordon and Jonathan Billington. Modelling the WAP transaction service using coloured Petri nets. Lecture Notes in Computer Science, 1748: 109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480109.
pdf.

Lei:1999:EUD

[1989] Hui Lei, Kang-Woo Lee, Marion Blount, and Carl Tait. Enabling ubiquitous database access with XML. Lecture Notes in Computer Science, 1748: 119–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1748/17480119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480119.
                                        pdf.

Tagashira:1999:DEI

[1990] Shigeaki Tagashira, Fumitake Inada, Keizo Saisho, and Akira Fukuda. Design and evaluation of an information announcement mechanism for mobile computers. Lecture Notes in Computer Science, 1748:135–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480135.
                                        pdf.

Wu:1999:ARS

[1991] Shiow yang Wu and Shih-Hsun Ho. Active rule system for adaptive mobile data access. Lecture Notes in Computer Science, 1748:146–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480146.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480146.
                                        pdf.

Hwang:1999:PST

[1992] Se-Jin Hwang and Myong-Soon Park. Preserving smooth traffic and high presentation QoS for VBR-encoded video. Lecture Notes in Computer Science, 1748:156–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1748/17480156.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1748/17480156.
pdf.

Ratner:1999:PRS

[1993] David H. Ratner, Peter L. Reiher, Gerald J. Popek, and Richard G. Guy. Peer replication with selective control. Lecture Notes in Computer Science, 1748: 169–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480169.
pdf.

Xu:1999:CCL

[1994] Jianliang Xu, Xueyan Tang, Dik Lun Lee, and Qinglong Hu. Cache coherency in location-dependent information services for mobile environment. Lecture Notes in Computer Science, 1748:182–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1748/17480182.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1748/17480182.
pdf.

Kuroda:1999:ADC

[1995] Masahiro Kuroda, Takashi Sakakura, Takashi Watanabe, and Tadanori Mizuno. An autonomous data coherency protocol for mobile devices. Lecture Notes in Computer Science, 1748:194–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1748/17480194.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480194.
                                        pdf.

Chen:1999:LMS

[1996] Chao-Chun Chen and Chiang Lee. Location management strategies for reducing both movement cost and locating cost simultaneously. Lecture Notes in Computer Science, 1748:209– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480209.
                                        pdf.

Sufatrio:1999:IMS

[1997] Sufatrio and Kwok-Yan Lam. Internet mobility support optimized for client access and its scalable authentication framework. Lecture Notes in Computer Science, 1748:220–229, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1748/17480220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1748/17480220.
                                        pdf.

Chang:1999:PAC

[1998] Kyung-Ah Chang, Byung-Rae Lee, and Tai-Yun Kim. Proxy agent consistency service based on CORBA. Lecture Notes in Computer Science, 1748: 230–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480230.
pdf.

Chlamtac:1999:IMO

[1999] Imrich Chlamtac. Integrating mobile objects into the wirelessly wired world: The need for energy efficient algorithms. Lecture Notes in Computer Science, 1748:243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480243.
pdf.

Franklin:1999:DUC

[2000] Michael Franklin. Databases unplugged: Challenges in ubiquitous data management. Lecture Notes in Computer Science, 1748:244–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1748/17480244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1748/17480244.
pdf.

Anonymous:1999:AIn

[2001] Anonymous. Author index. Lecture Notes in Computer Science, 1748:245– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/1748/1748auth.pdf.